A Complete Bibliography of Publications in Biographical Memoirs of Fellows of the Royal

Society

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

04 October 2024 Version 1.09

Title word cross-reference

10th [Hen90]. 13 [Ano68]. 1911- [KP78]. 1913-5 [Doc98]. 1915-14 [Dom98]. 1918-22 [MS98]. 1940 [Her55]. 1945 [Fil56]. 1950 [Jon03]. 1952 [Sch55]. 1953 [Bus60, Fle55b]. 1954 [BC55, Cla55b, CA56, Fle55a, Mot55, New55, Tho55, Tur55a, Tur55b, Wat55, Wat56, Wil55]. 1955 [And57, Ark55, Bed57, Cha55, Chi56, Cla55a, Col56, Dal55, Die56, Dub56, Gad56, God57, Hau56, Haw56, Joh03, Kin56, New57, Red56, Sco56, Ses62, Smi55, Whi55, Wig56, Wil56]. 1956 [Baw58, Boy56, Bul57, Chr56, Cow57, Cra57, Fle57, Fos57, Fre57, Gad57, Har56, Hod56, Kot60, Kur58, Pan57, Pet58, Pug58, Sal57, Sta58, Str56, Str57b, Tem56, Tho56, Til57, Woo57]. 1957 [And58, Bur03, Cha58, Cox57, Dal57, Dod57, DE58, Hil58, Hin58, Kar58, Rob60, Rub58, Smi58, Str57a, Tay58, Tho58, Tit58, Whi58, vE57]. 1958 [AT60, AC60, Bar60, Bla60b, Bow58, BO58, Bri58, BCB61, Coo58, CH58, Cox58, Dal58, Dar58, Eme60, Fre60, HP60, Jon60b, Kei60, McC60, Pei60, Tay62, Tod58, Tur58]. 1959 [Abe61, BC60, Bla60a, Cam60, Che62, Coo60, Far61, Fel60, Fra60, Haw60, Hin60, Man60, Men60, New60, Poo60, Pug61, Rat60, Ril60, Rox07, Tho60b, Til61, Tis60, Wil60]. 1960 [Baw61, Cha61, Cla61, Cor60, Dar61, DL03, Ewa60, Har61, Jon60a, Mit61, New61, Pfe60, Rel61, Rus61, Sal61, Sau61, Sko61, Tho60a, Tho61, Wil61, Woo61]. 1961 [Ait62, Baw63, Ber63, CC61, Dal61, Dal62, Flo63, GY62, MDHW03, Hei61, Hub71, LB67, Mar66, Med61, MP62, New62, Oak62, PR63, Sav63, Tur62, Wal61, Whi62, Yon62]. 1962 [Adr63, Bas63, Boy63, Coc63, Dar62, Dea63, Har63, HL63, Lov64, Mar64, Ram64, Rob64, Tho63, YM63]. 1963 [Adr64, Car64, Car65, Chi64, Cow64, Cun64, Dou64, Gal03, Gol66, Hir63, Hol66, Ing64, Man64, Men64, Pro64, Rat64, Rub64, Sal64, Sch64, Sho63, Smi63, Tay64, Tem65]. 1964 [Bel66, Bra65, Cla71, Col64, GF65, Hay65, Kuh65, McC65, MF66, Pan65, Pir66, SE65, Sti65, Tho65, Wes64]. 1965 [Adr66, BR73, Dee67, Dor66, Dun66, Eva66, Fel67, Gol65, Har66, Jon66, LS66, Mor66, Oak66, Por66, Rat66, RWN66, Sal66, SK66, Stu65, Stu66, Sut65, Syn67, TF67, Tho66, Wal66, Whi66b]. 1966 [BRE68, Bow67, Boy66, BS67, Bul67, BT68, Bul66, Coc67, DM68, Dav70, Dea66, Hun03, Ing68, Jam67, KN69, Lew67, Oak68, Pen67, Pic68, Rat71, Sch67, Sho67, Ste67, Tol67, Whi66a]. 1967 [And67, Bet68, Bom67, Bri68, Bur68, BZ67, CP69, Cre67, DP70, Kay68, KS69, Men70, Oak71, OP68, Pea68, Pon68, Por68, Pri68, Pug67, Rus68, Smi68, Sta69, Tho73a, WB68]. 1968 [Abr71, All70, Bat69, BR71, Bra70, Bul69, DG70, Dun68, Fel70, Fri70, Gra69, HCH73, Hol71, KL69, KM69, KM70, Mai69, Pro68, Ray69, Spe70, Tab69]. 1969 [All72, Bro70, Cla70, Dow70, Fal70, FP71, Gar70a, Gar70b, Hab71, HB71, HG70, HT70, Roc70, Sil03, Ske70, Smi69, Stu70, Win70]. 1970 [All76, And71, Baw71, BC72, Bha71, Bla72, Bur71, CH71, Chr72, Col71, Coo72, Cow71, Dav71, Fea71, Fle71, GB71, Goo72, HW71, Hil72, Kay71, KS71, Kre72, Kre73, Mac03, Mas71, Oak73b, Par72, Rot72, Sho71, Sho72, Sut70, TH71, TG71, Wig71, Woo71]. 1971 [AC72, BF72, Bir72, CP72, Cot72, Cra72, Dar72, GKW73, Hod75, Hod80, Isl78, Kay72, KP74, MP74, Phi79, Ruz72, Vog72, Wil72, YF74, Zuc73]. 1972 [All73, Bar73, Cas73, Cou73, Den74, Gib73, Har73, HPR72, Hod73, Hub75, Hux03, Lip73, MY73, Neu73, Oak73a, Ogs73, Pic74, Pir73, Rao73, Smi73, Syk73, Tho73b, Tod73]. 1973 [Baw75, BM74, CHW74, DN74, Dic75, DR74, Gar74, Hog74, HH74, Jon74, PW74, Phi74, Ran75, Rat75, Ros74, Stu74, TK74, Tho74, WP75, Whe74]. 1974 [AB74, Bar75a, Bar75b, Bul75, Ele76, Haw76, Kem75, Lov75, MF76, Meh75, MP75, SR76, Stu75, TP75, Tod75, Wil75]. 1975 [Adr76, Ati76, Bak76, BR79, Bat76, Bro77, Buc77, CU76, CF76, Cru76, Eva76, For77, Gar75, GW76, HM77, Lis78, McR76, Moo77, Ray76, RG76, Rob76, Rob77, SP76, Stø76, Syn76, TC76, Wel78]. 1976 [Bak77, Ber77, Bla77, Cla77, EF78, Glu77, HW77, Jef76, Jef77, Joh77, LHF78, Lwo77, Mac77, MT78, MP77, Neu78, OM80, Pig77, PJ80, Rus77, Stu77, SB77, WH77]. 1977 [Bur78, Col78, Dun78, GPPG78, Hod79, Jon79, Kat78, KP78, May79, Nut77, PB79, PSS78, PR78, Pri78, Pug78, Rus78, Sal78, Sta79, SP79, WB03, YPS78]. 1978 [And79, Att91, Bro80, CK80, CH79, Cla80, CD81, DT81, DS79, Gar80, Ker79, KW79, Kre80, Kur83, LMA80, Mar79, Pat79, SG79, Tod79, Whi79, WZ79, Wis83, YH82]. 1979 [Abr83, All80, BM81, CB81, Den81, Els82, Fry80, Gar81a, Her81, JP81, KD82, LE80, Med80, NS82, Pat82, Pei81, PR81, Sut80, TC81, WM84, Zuc81]. 1980 [Bar86, Bar81, Bel81, Bla81, Ble82, CR83, CM81, Gar81b, GP83, Har81, Har82, HLH82, HGM81, Kat82, Kek81, Lea91, MD81, McC81, McK87, MP82, Pay81, Ric81, She82, Sho83, Smi82, Ubb81, Whi83]. 1981 [Ban83, Bea82, Bil90, BW84, Cha90, CRST83, CL82, Dun82, Eve84, Goo82, Gra82, Hay82, Her82, JG82, Jon83, Kas82, Kem83, KBP82, Ken82, KW84, Lew83, Mot82, Per82, Phi83, Wig82, Zuc83]. 1982 [All83, Cla86, Dai85, Dal83, Dun83a, Dun83b, FW83, GRT84, Hut84, JRS84, JB84, KH83, LH84, Mil84, NW84, Sho85b, TO83, Tho83, Wig83]. 1983 [All84, BBD84, Bla85, CMO87, CV85, Cha85a, Cot84, Cur84, Gow90, GR85, Jef85, Lan85, Lig85, LS85, ML85, Mof85, Neu84, Pas87, PM94, PB84, Ric85, Ste84, Thi84]. 1984 [Ada85, Bla94, Cha85b, Cla85, DP86, Den86, DS86, DST85, GBC85, GS86, Hay85, Hor86, Hux92, Leg85, Lew86, Mon86, Pat90, YHP^{+}86, QT86, Ran86, Sho85a, Sou87, Sta86, Ste86, Sto94, Stu86, Whi86, Wil87, WB85]. 1985 [Alb87, Alp88, Bel86, BH88b, Dac88, Fen87, FP95, Fog86, GS90, Hey87, HH88, MM87, Mar86, Mat90b, MM96, NJ86, Per88, Per87, Pow87, Rob86, SL88, Stu85, Tyr87, Wes88, Whi94b, ZK90]. 1986 [AB88, BH88a, Bis87, Dai90, Eab90, FD87, Fox87, Gra90, Har87, Hil87, Hil88, Hin87, HW87, Hir90, Hog87, JW00, LH90, Man87, McC88, Mor92, Mot87, Sau87, SV91, SF91, Tat90, TC87, Vau87, WW86, Woo97a]. 1987 [Abr88, AW88, Ask90, Bai88, BW90, Bra92, Bud88, CD88, Dat88, Gar88, Gin94, Gre01, Ken91a, Kib88, KH90, LH88, MS90, Mar88, Mat88a, Mat88b, Mit90, Mor88, Nor91, Och90, PM88, Pol92, Ser88, SQ88, Tei03, Tod88, Tyr91, VT88, Whi90]. 1988 [Arn91b, BS90, Cla95, CV94, Dow94, Duc91, Edw96, Elv91, Hen90, Hey91, Hil90b, Hin90, Key90, LP91, MH90, Mad90, Mat90a, Meh02, Neu90, Nut90, Pir90, Pre90, Qua94, Ran90, Spi91, SL90, SW90, TM90, Wei90]. 1989 [Arn94, BF91, Coo90, Cor90, Cul90, Gab02, Har91, Hil90a, Hor95, Jam90, KS92, Lig90, MR22, NA95, PG92, Pol00, Pri94, Rob92, Rog91, Roo96, Rub91, Sac94, Swa96, Tho91, Woo90]. 1990 [All91, Arn91a, BS04, BC92, BB92, BN92, BP99, Coo94, DE95, Doc98, Edw94, FS92, GV92, Gre00a, HD92, Jon91, Joy92, Ken91b, Lew92, Mes91, Mit91, Mui94, Mul94, Pol06, Ree94, SL94, Tat94, Wil94b, fWT94, YL92]. 1991 [Ben94, Bey96, BMW92, Bul94, Cam94, Can14, Dol94, Gri95, HNW94, Hol94, Lov11, Mak94, McL00, Mot92, Pea95, Pet92, Pip94, Saw92, SC99, Sta94, WA94, Woo95a, Woo94]. 1992 [All94, Bai94, Bun94, Bur94, Cox95, Cru04, CP94, Fed94, Fei95, HH95, Hil97, Hon95, Hul94, Imu95, Kil94, LM94, Oli94, Rob95, Sau94, Sla94, Spe94, Spr95, Whi94a, Wil94a, Woo95b, Wun97]. 1993 [Big02, BB97, BG95, Bol97, BB96, Bra95, Cul95, DAB97, Dol95, EE95, For95, Gar95, Gre98, Gre96, GM97, Jon23a, Kri95, Kro95, Kus01, NG96, Owe95, Pit94, RP96, RE04, Wad97, Wei94, Wes95, Wid95, WR95]. 1994 [Akh96, AH97, Ati96, Bar98, Bea95, Ble96b, BH96, BI96, Cha97, Cro97, Dal97, Dod02, Dul95, Dun96, EL96, FJ95, Fin95, GI96, Gor96, Gre00b, GS13, Gut96, HD98, Hol96, Hop96, JG98, Jel99, KO99, LKK97, LJ97, Loc96, Man03, MS96, MS98, Mil97, MR96, MBW99, Mor97, Ove97, Phi95, Por96, ST99, Von97, Wai97]. 1995 [AA98, Ale96, Ble96a, Ble97, BH97, Col02, CE96, Dom98, Flo08, GR00b, Har99, Hew99, Hux97, JW97, KH97, KE04, Lee07, Oxl23, Pat97, PF99, Pea98, PS97, RC07, Row96, SVW00, Son99, Str96, Tay96, Whi96, Woo97b]. 1996 [All99, AM01, Ano98, BB99b, BM98, CJ01, Cru00, Far98, FH98, FR98, GG00, Gui97, Hal99, Jac00, Kib98, Lig98, LBG98, Mab99, NH13, Pip98, Ree98, Rot99, Scr99, Lor98, Sle97, Smi98, Smi99, Tho98, Wol98]. 1997 [BB99a, Ble99, Boy98, BK00, CJ99, Cap99, Car99, CCG^{+}00, Coo98, Coo99, CA01, GI00, Hol01, Kin00, Man98, Mes98, Mes99, MS08, Mur98, New98, Ost07, Pie99, SB08a, Vin98, WW99, Whi99, vC99]. 1998 [Akh00, All00, And06, ADE00, Bel01, BN01, CS98, Chi99, Cla00, Cro00, Dys99, Dys11, Eva99, Fle99, For00, For99, GWM11, GR00a, Gun00, Han00, Hay00, Hux00, LM02, Mac00, Mac12, Mit00, NN00, Ped01, QRJ12, Rit00, Rob00, San00, Sel07, Ser99, Wel00, Wol01]. 1999 [Ale01, AS02, Ave08, BS01, Bur01, But01, Cha01, Cla08, Coc03, EP10, Fea04, Fla03, Gui02, Hof07, How01, Joh00, JJ14, Kev05, KM15, Law01, Lyo01, MW01, MP07, Pag01, Pat01, Qui01, Ruf02, Sid02, Smi05a, Sto03, Whe01, Whi03, Whi01, Win01]. 1st [Lor98].

2000 [AA02, Ash02, Ast01, AS03, BG02, Ble01, Bod04, BS05, Bry01, Bun02, Cha02, Coo01, Cox03, CVW02, Cus05, DWS02, Dom21, DH04, Fei05, Fow02, FA03, Gra04, Har04, HM03, Hay02, Hew02, HP01, JGP06, Lar09, Lin01, LL05, Mas02, McC03, Mof02, Nee02, RG06, Wäs02, Wea02, WL02]. 2001 [App07, AL23, BIR12, BT05, Cha03, Cia02, DL04, EFP17, EB08, Fit22, Hit06, HT03, Jam09, Lew02, MT08, Mor05, O’H06, Pep03, QG03, RR05, RB20, She03, SB08b, Stu09, VJ05, WK03, Wol03]. 2002 [Atk04, Ber04, Blo04, Bre05, Con09, CW11, Cow04, FP04, Jor04, Lov04, McL04, ML03, MR19, NA05, Ott04, Pra10, Pre10, Rai08, Sin03, Whi04, You12]. 2003 [Bar04, Bat03, Cow05, Cut05, DM06, ES05, Eva05, HW04, KH04, Law08, Pat05, Pud21, RW06, Sak07, Sti05, WB05, Woo05]. 2004 [AKS06, AR09, Ber05, Bon06, Bow05, BM05, BM17, BE05, Buc06, CK06, CH05, Cro19, GLR07, GM06, GW07, HS07, HF09, Hit14, Igg06, Ive06, Mad06, McC08, Mon06, MS06, Pip05, Qui05, RS07b, Row04, Smi05b, Spa14, Wat06, Wei05, Wel05, YN06]. 2005 [Aca08, Ati07, BS08, BN06, BH08, Boc06, BC07b, CR07, Dou07, FGOS15, Fie08, FM21, Gar06, Gre06, Har18, HF07, HF06, HF08, KTR07, LB07, LMW06, MH08, PB10, RS07a, Rey07, SS06, SLO10, Wal06, Yat06, vEG06]. 2006 [AS16, Ant15, AD21, BS14, Bro09, Bro10, CL14, Ell09, Gia16, GCJM18, Gur12, HP08, HP09, Hun08, KC22, LW09a, Lau07, ME11, Moo15, Ree08, RE08, RK10]. 2007 [Bar11, Bar21, BC07a, BH18, BC18, Ben08, BS13, Car14, CL08, Cla09, CJ23, DM09, DJ13, Edm20, Ell08, Fer17, Gre12, Gre13, Gri11, Hil11, HKB23, JC19, Kin09, KR11, Leh12, Mah24, MR13, Pap19, Pic16, CBE09, Sma10, Vin19, Vog08, Wal08, Wil23c]. 2008 [Ada09, AB15, AM16, Aut19, Bar23, BG11, Bos12, Buc10, CEP15, Con22, CT14, DGV11, Fit10, FAL21, FE12, HW10, IG21, LW09b, MT21, Mar12, Pal09, PO21, PC10, PC11, Ran09, Ser09, SH24, Stu13a, Ter10, Vin09, Web10, Yud10]. 2009 [ASS16, BM13, CJ12, CC10, CDS12, Cro10, FS20, Gra10, GTK11, Hut23, Jon10, Jon21, JS10, Lon11b, Owe09, Phi13, PT11, Sou11, SBG10, SBG11, TW21, Tre11]. 2010 [All13, Ano19, Big10, BBM13, Bra12, Bra15, BLT11, BLT12, BM22, BK17, Bur21, CK14, CEW18, DGS18, For12, Gar21, Glo20, GB13, Gly11, HG11, Hir15, HJ11, Hum11, KT18, LH18, Lon11a, LR17, LS12, MB21, Mil18, Pet11, Rae12, Ram22, Rei14, SR17, SM24, Sen19, Tay17, TO12, Wah23, Wil13, van19]. 2011 [Bre14, BSD13, Bur19, Cle20, Cra20, DSD20, Edw18, FM19, FGW13, GB14, Gil20, Gol17, GN19, JG13, Mak18, Mor16, OG14, PB17, Ric14, Sca18, Sew15]. 2012 [AM13, Ati14, BB22, Bro13, BM23, Cer19, CKT15, CG20, DGL16, Dod20, DR22, Eyr13, Gai19, GM21, Gre19, Gut21, Har20, HT14, Hop23, Hor14, JWW^{+}20, Kel17, Lyn15, ML19, McK22, Pas19, PD18, Sco14, Sim18, SK13, Tho13, TB13, Uns14, Wil23a]. 2013 [Aga23, BFO^{+}19, BY16, BG14, BR14, Bro15, BS16, Dav18, Dor18, Gar15, GR16, Gra20, Gre14, GSL19, Her14, HB22, HW17, JH20, MS19, McC22, MO18, MF18, Mor17, RS15, Ree15, Sha15, SA15, Smi15, Wei24, Wur16]. 2014 [Ash16, Bou15, Bro23, CR16a, CR16b, CBPC15, DE19, FR24, GS18, Has16, How15, Hyn15, Joy16, KL19, LN16, MIH24, May15, Min22, MS15, MA21, Mor19, Nai21, Nic16, PSW16, RH18, Ric15, RSD16, SW21a, ST19, SD15, VC18, Wri24]. 2015 [Ben19, BR18, Bir17, Boy23, Bro16, BFIT18, CV17, DG16, DFSV23, Fea18, GV18, Goo16, GLD18, GCK17, Gri18, HV20, HHP22, HT16, HCB18, Hun17, HR18, Kin16, MJR24, MM21, Pea18, Pud17, Sha21, Smi16, TH16, TC18, War17, Web17]. 2016 [Ale22, BSHC18, Bor17, BE24, Cal21, CST17, Col18, Cow21, CG21, CSB22, DT20, DS21, Edw20, Han18, MH24, HL18, Hua18, Lai20, LPP19, Leg17, LC17, Lyo22, MT19, OKC21, PB20, Ran22, RRS20, Rea19, RW18, SH18, SL18, SO20, Tho20, Vin23, Wat18]. 2017 [ACW20, Bal24, Bro19, BB18, Cla21, CH19, Faw18, HU19, LR19, LG20, MH19, Mor21, Nag19, PSCV21, Pau18, RP20, Riv20, She18, SW21b, SDV19, Tak22, Tho19, Tur23, VL24, Vel19]. 2018 [CEG^{+}19, CB22, Cre24, Cro20, DM20, Dou19, Edw21, Eva20, Flo20, GVS19, Goo20, HL23, HB20, Hig24, KT20, LW20, MW20, Mas23, Mid21, Ogd20, RH22, RC24, SC23, Ste20, Tho24, WF19, Wat19]. 2019 [Ber20, BW22, BT20, FSC20, Gad21, GW20, Gor24, Hib21, Hit20, Kin22, Leg22, Mil20, Mui23, NL22, RSG21, RR24, San21, Sea20, SH21b, WB20, WD21, Woo21]. 2020 [AS21, BL24, Bra22, CGZ22, Cra21, CNO23, Cur22, Dys22, Ell22, EW22, Eva22, GT23, KHG21, Joh23, JKL^{+}22, KC23, Lau23, MO23, NC21, PR22, Pet24, Pro23, Rei21, Rob23, SL21, SOD22, SH21a, WPL21, Whi22, Wil22, WBF23]. 2021 [AGJ^{+}22, Ash22, Azz24, BF24, Bya22, CO22, Dav23, DG23, GMP24, Irv23, Jon23b, LL23, Lon21c, Lon22b, Mac23, Mac22, MP23, Mül22, NS24, GN24, PK24, PF21, Sad23, SSGP23, SB22, TW22, TJ24, TS24, TG23, Wil23b]. 2022 [Bha24, Bro24, Cra23, DT24, Hai24, KM24, LY24, LK23, Lon22a, Lon23b, MB24, MP24, Rei24, BS24, Sow24, AS24, SSEF23]. 2023 [BM24, Caw24, Cra24, Pud24, Rus24, ŠR24]. 2nd [Hux97, Wal08].

3rd [Lin01].

562 [Ano71a]. 5th [Tod73].

68 [Lon20a]. 69 [Lon20b].

70 [Lon21b].

A.C [Ble01, CJ99, CA01]. Aaron [Cro20]. Abdus [Kib98]. Abel [Rus68]. Abercrombie [Med80]. Abragam [Gol17]. Abraham [JJ14]. Abram [Bur71]. Abrikosov [VL24]. AC [BY16, FE12, FGW13, Nic16, Pra10]. Adair [JP81, Row96]. Adam [CHW74]. Adams [Jam90, Sta86, Str56]. Addenda [Ano87]. Addendum [Ano60, Ano66, Ano88, Ano90a]. Addison [LJ97, Mor05]. Adelaide [Abr71]. Adler [Sho67]. Adolf [Eme60, KW84, Whe74, AA98]. Adrian [Hod79, Hux97, Kin00, BH88a]. Adrien [DP86]. Agnes [DM06, Tho60a]. Ahmed [Tho20]. Ainsley [Cer19]. Ainslie [KH83]. Aitken [WB68]. Akers [Fle55a]. Alan [BC92, BM13, Fle55a, GPPG78, Han00, Hor14, Hux00, Pol06, Qui05, SM24, SK13, Son99, Wah23, Yon62, All00, Bos12, Bou15, BS16, BC07b, Cru04, Edw20, GCK17, Hil90a, HU19, HKB23, Hum11, JB84, JKL^{+}22, KL19, Kek81, LP91, MJR24, Mas23, New55, Pas19, Ray76, Rub64]. Alastair [Oxl23, AA02]. Albert [CL08, HD98, Hog74, JW00, SP79, TO83, Tho65, Wat56, AM01, Bas63, Bur68, Cha01, Hay02, Lip73, Mar66, Mat90a, SM24, SW90, Whi55, Woo57]. Alberto [YF74]. Albery [HB22]. Albin [Thi84]. Albrecht [LB07, BT05]. Alec [Cha03, Gab02, CR07, HG11]. Aleksandr [CR83]. Alekseevich [VL24]. Alers [Sch64]. Alexander [Ada85, Baw71, Chr56, Cro00, DE19, Gab02, Gre14, Hor86, Jel99, Lew67, Mat88b, NJ86, Oxl23, Pol92, Pug61, Rat75, Tho58, Wid95, Wil23c, Ber77, BK00, Col56, Dav18, Dod57, Dow94, Gar80, GS90, HP60, HV20, HW71, Her82, Hol71, Oak66, Smi99, Tho56, WB68]. Alexei [VL24]. Alfred [AB74, Chi99, DGL16, Doc98, Hey91, LW09b, ML19, May79, Med61, Pfe60, Sho83, Wig56, Baw71, Gre98, JRS84, New60, Sch67, Ske70, Wei90, Wei05, WP75, Woo97a]. Alfven [Pea98]. Alger [Ken91b]. Algernon [Hen90]. Alice [MO18]. Alick [And67]. Alistair [GMP24]. Alister [Sav63, Mar86]. Allan [DM09, LL23, Smi82, Ble97, Can14, LS12, TM90]. Allen [AL23, Rus24, Sad23, Syk73, Wil55, WBF23, Gar81b]. Allibone [Law08]. Alwyn [BE05, Pud24]. Alwyne [DS79, Sle97]. Alymer [BF72]. AM [BSD13]. Amaldi [Rub91]. Amazaspovich [LBG98]. Ambartsumian [LBG98]. Ambrose [CVW02, FGOS15]. Amoroso [Sho85b]. Amos [LW09a, Tho13]. Amyand [CO22]. Anastasios [SL18]. Anatole [Gol17]. Andersen [BL24]. Anderson [Bri58, HCB18, HD92]. Andrade [Cot72]. Andre [JG98, Ser99]. Andrei [Ken91a]. Andrew [DSD20, She03, Sim18, AGJ^{+}22, Aut19, Kin22, PB84, Pug78, AS24, SSEF23, Win01]. Andrewes [Tyr91]. Andrews [JG13]. Andries [BS67]. Angus [Bur01]. Ann [HT03, GV92]. Anne [CJ23, Gut21, KT20]. Anrep [Gad56]. Anstruther [Tho60b]. Anthony [BS90, Buc06, Lin01, RG06, Sea20, CEP15, Che62, Cle20, Fea18, Fit10, Gur12, How15, Hut23, Nag19, Sau94]. Antony [Lon22b]. AO [Hum11, Kin16]. Appleton [Rat66]. April [AM13, Aga23, Akh96, Ale96, AM01, ACW20, AW88, AL23, Arn91a, AB88, Ast01, Bai94, BB99a, Bar75a, BH88a, BF24, Bel81, Bel01, Ben94, BC18, Ben08, Ben19, BS14, BR18, BG02, BB96, Bra92, BG14, BE05, BW22, Bro80, Bro09, BM22, BFIT18, Bur01, BT20, CJ99, CF76, CGZ22, Cha90, CH05, CL08, Cla80, Cla85, Cla95, CJ23, CRST83, Con22, Coo90, Cor90, CG20, CM81, Cox03, Cra21, Cro00, CNO23, CL14, CBPC15, Cru00, Cru04, Cul90, Cur84, Cur22, Dai90, Dav18, DE95, DG16, DAB97, Dol94, DE19, Dun83a, DR22, EF78, EB08, Ele76, Eva20, Eve84, Eyr13, Fei05, FSC20, Fit22, Flo20, FS92, FGW13, Gai19, Gar15, Gar95, Gar88, Gra10, GRT84, Gre98, GM06, Hal99, HW10, Har81, HV20, KHG21, Hay00]. April [HD98, Hen90, Her82, Hil90a, Hin87, Hin90, HW87, Hit20, HB22, MH24, Hod75, Hof07, HL18, HKB23, Hor86, How01, Hub75, Hul94, Hun03, Hun17, HP01, Hux92, Hux97, Isl78, Jam09, Joh03, Jon21, Jor04, JS10, Kem75, KBP82, Ken91a, Ken91b, KD82, Kre80, Kro95, Lai20, Lau23, LL23, Leg22, Leg17, LJ97, Loc96, LC17, Lon21c, Lon22a, LG20, LMA80, MH90, Mac12, MM87, Mak18, MIH24, McC88, McC22, MO23, McK87, MO18, Mes99, MP07, MBW99, Mon86, Mor17, Mor05, Mor19, Mot87, NN00, Nai21, Neu84, NH13, Oli94, Ove97, Pal09, PB79, Pas19, Pat79, Pat97, Pay81, Pea95, Pea98, PM88, Per88, PR22, PF21, Pit94, PR78, Pre90, Pri78, QRJ12, Rat75, RRS20, RS07a, Ree15, Rei14, Ric81, RG76, Riv20, Rob95, Rog91]. April [Rus78, Sak07, San21, SL21, Scr99, SB08a, Sea20, Ser88, She82, She18, Sho85a, SLO10, Sla94, Sle97, SB08b, SH21a, Spa14, Sta79, Sta94, Ste86, Ste84, Stu13a, Tho19, TO83, Tho24, Tod75, TC81, TC18, Ubb81, Vau87, VJ05, Von97, Wad97, Wal08, Wal06, Wäs02, WA94, Wei05, Wei94, WB20, Whi22, Wil23b, Wil22, Woo90, Woo94, YPS78, van19]. Arber [Tho60a]. Archer [Lov04]. Archibald [Bro77, Str96, Kat78, Tur55a]. Argyris [Spa14].

Aristide [LY24]. Arkell [Cox58]. Armand [TG71]. Armstrong [Mak18]. Arndt [CL14]. Arne [KP74]. Arnold [CF76, Dow94, EW22, KT18, Rub64, Neu90]. Arnott [Ree15]. Arrow [Vel19]. Arthur [All76, Ati96, Bar21, Bir17, BH97, BC07b, DG70, GBC85, Har66, JW97, Jel99, KTR07, Key90, Kre73, Mad06, Mad90, Mof85, RH22, Sca18, SQ88, Tod73, Bis87, Bra70, Cha55, Cla55a, Cla08, Col78, Cra57, Dal58, Dun66, Elv91, God57, Hau56, Hit06, Jon74, Jon79, Leh12, MBW99, MF66, MP62, Pri78, Pro68, Pug58, RC07, Rub58, Ruz72, Wil94b, Woo97b]. Ashby [HH95]. Ashesh [Mah24]. Ashley [Neu90]. Ashton [LL05]. Ashworth [Bud88, Sen19]. Askonas [MO18]. Astbury [Ber63, KL19]. Astley [Wea02]. Aston [Pol00]. Athelstan [Rae12]. Athlone [Dod57]. Atiyah [Hit20]. Atkins [Poo60]. Attlee [Bri68]. Aubrey [BSHC18]. Auerbach [Bea95]. August [Abr88, Abr83, Aca08, Adr76, AM01, All94, AB88, Ati96, Bai94, Bar21, BFO^{+}19, BS04, Bar81, Baw75, Bel86, Bel01, BH88b, Bla81, Ble96b, Ble99, Bou15, Boy23, Bra12, Bra22, BW22, BMW92, Bro13, Bro16, Bro15, Buc77, BS16, Bur03, BM13, CR16a, Car14, Cer19, Cha01, CEP15, Cha03, CC10, Cla00, CR07, Cla08, CD81, Con09, Coo98, Cow21, CT14, Cro20, CE96, CP94, DP86, Dal83, DGL16, Dav23, DL03, DGS18, Dom21, DGV11, Dun82, Dun96, Edw21, Edw20, Fea18, FH98, Fei05, Fen87, Fit10, Fit22, Fla03, FS20, FP04, FW83, FGW13, Gar81a, Gar06, GB13, Gly11, Goo16, GBC85, Gor96, Gra04, GLR07, Gui02, Gut96, Gut21, Han00, Har18, Har82, HLH82, Has16, HB20, HH95, Hil87, HF07, HP08, HP09]. August [HU19, Hod79, HF06, HF08, Hol01, Hor14, HCB18, How01, Hua18, Hum11, Hun03, Hun08, HGM81, Hux03, JGP06, Jon21, JKL^{+}22, Kat82, Kek81, Kel17, KBP82, KP78, KC22, KO99, KE04, Kin16, KR11, KW84, KD82, Kus01, LR19, Lau07, Lau23, LBG98, LS85, MT21, Mac23, Mac22, MM87, Mar12, MIH24, Mas23, Mat90a, MP23, McL00, McL04, Med80, MT08, MM21, MF18, MM96, MR96, MBW99, MR22, MR19, NA95, NS24, Neu78, NG96, NW84, Ogd20, Pat01, Pau18, PM94, Per82, Pet92, Phi83, Pip98, Pir90, Pri78, QG03, RH22, Ran75, Ree08, RC07, Ric14, Rit00, Rot99, Ruf02, Sal78, SL21, Sau87, SM24, Sco14, BS24, Sen19, Ser99, Ser09, Sha15, She18, Sin03, ST99, SW21b, Spa14, Spi91, Sti05, Stø76, SL88]. August [SD15, Syn76, Tay96, TP75, Tho20, Vel19, Vin19, Von97, Wat18, Wea02, Wei24, Wel78, Wel05, Wes88, Whi03, WK03, WB03, Wil23a, Wil23c, Win01, Woo05, Zuc81, vEG06]. Augustine [Dea63]. Augustus [Per82]. Aulis [SO20]. Austin [Dol94]. Autar [Igg06]. Avery [Dub56]. Avrion [MB24]. Axel [Dal83]. Axelrod [Ive06]. Axford [All13]. Aylmer [YM63].

B [Akh96, BB97, BG95, CS98, Coo98, CE96, Doc98, Edw96, Far98, FH98, FR98, Gre96, HD98, KH97, Kib98, Kro95, Lig98, Loc96, Mak94, Mes98, Mil97, MR96, New98, Ove97, Owe95, RP96, Ree94, Roo96, Sle97, Smi98, Str96, Tho98, Wad97]. Bach [CC61]. Bacon [Wil94a]. Baddiley [Buc10]. Bagnold [Ken91b]. Bagshawe [BS24]. Bailey [Chi64, Smi58, Stu65]. Bairstow [Tem65]. Baker [Hey87, Hod56, KH90, Mas23, ML03, WB85].

Baldwin [Lon11a]. Balls [Har61]. Bamford [Fea04]. Bangham [HG11]. Banks [Jon60a]. Bankside [Gow90]. Bannerman [DS21]. Baptiste [Oak62]. Barbara [Fed94]. Barber [Bey96, Dar72]. Barcroft [GR00a]. Bardeen [Pip94]. Barenblatt [CB22]. Barker [BC72, Row96, Tit58]. Barlow [Cul90]. Barnard [Ste20]. Barnes [PR81]. Baron [Abr71, BK00, CC61, CR16b, GB71, HH95, Hey87, Hod79, Hol71, Hux97, Kay72, KH04, Kro95, Lov75, Mak94, Mit91, Pic68, Pic74, Ree94, Sch64, Lor98, Tho13, Win70, Woo95b]. Barr [PS97]. Barrer [Ree98]. Barrington [Mat90b, TK74]. Barrow [Ell22]. Barry [HD98, Pre10, Sin03]. Bartlett [Bro70, Whi04]. Barton [LM02]. Basil [Bro16, DP70, All73, HH88, Jon21, Pat05]. Basinski [How01]. Batchelor [Mof02]. Bates [Dal97, Kur83]. Bateson [LR19]. Battersby [Stu65]. Bawden [Pir73]. Bawn [Pud21]. Baylor [LY24]. Beadle [Hor95]. Beale [Pow87, PT11]. Beals [Her81]. Beament [Gre06]. Beckwith [Rae12]. Beddington [RR05]. Bedford [Bro80]. Bedson [Dow70]. Beer [Bar73]. Bell [BP99, CJ01, Eva20, RE04, Rob95, SH21a]. Bellamy [Car14]. Belle [Ave08]. Ben [Edw94]. Benedict [RB20]. Benjamin [Hun03, ML85, JW97, Swa96]. Bennet [Bro77]. Bennet-Clark [Bro77]. Bennett [Haw60, LH88, HM77]. Benson [Hil58]. Benton [CR16a]. Benzer [Gre12, Gre13]. Beresford [MS15, Pan65, Pet58]. Bergel [Tod88]. Bergersen [BSD13]. Bernal [Hod80]. Bernard [Bur19, DGL16, Sho63, Edm20, FGW13, Per82, Sak07]. Bernardo [YF74]. Bernhard [Pra10]. Berridge [Pet24]. Berrill [Scr99]. Bertram [BR73, BM98, BM05, Hub71, Sta86, Cow05]. Bertrand [Kre73]. Besag [DGS18]. Besicovitch [Bur71]. Bessemer [All72]. Best [YH82]. Bethe [LB07]. Bethune [Oxl23]. Beverton [CE96]. Beynon [Boy23, Tho98]. Bhabha [Pen67]. Bhatnagar [Ses62]. Biggs [NS24]. Bijvoet [GP83]. Bilby [Gre14]. Billingham [Bre05]. Binnie [Woo90, Woo97a]. Biographical [Lon20a, Lon20b, Lon21b, Ano71a, Lin56, McR76, Wil95]. Birch [RC07]. Birchall [KH97]. Bisat [Stu74]. Bishop [GV92, PD18, Pri94]. Black [MB21, She82]. Blackett [Lov75]. Blackman [Har81, Pas87, Por68]. Blackmore [For12]. Blakemore [MP24]. Blaschko [BB96]. Blaxter [WA94]. Bleaney [Ell09]. Blin [Ell08]. Blin-Stoyle [Ell08]. Bloch [WL02]. Blow [HF09]. Blundell [Wäs02]. Bogdan [Dai85]. Bohm [Hil97]. Bohr [Coc63]. Bolton [WR95]. Bond [Nut90]. Bondi [Rox07]. Bone [Mac22]. Bonsall [Gil20]. Bonsor [Wig82]. Book [Ano71a, McR76, Wil95]. Boon [Wai97]. Boone [Stu75]. Bordet [Oak62]. Boris [Abr83, Wig71, van19]. Borissovich [Gin94]. Borlaug [Phi13]. Born [Fle57, Flo20, KS71]. Bose [Meh75]. Bosnich [Ben19]. Boswell [Mit61]. Bott [Ati07, Wat19]. Bourne [HW77]. Bovet [Oli94]. Bowden [Tab69]. Bowen [Bel81, BMW92, Til57]. Box [Smi15]. Boycott [Wäs02]. Boyd [Goo82, Kay72]. Boys [Cou73]. Boyse [Bar21]. Brachet [Pir90]. Bradford [Dol94]. Bradley [Lip73, ST19, Tay17]. Bradshaw [Fit10]. Bragg [Phi79]. Brain [Pic68]. Brambell [Oak73b]. Brandon [HH95]. Brebis [Ell09]. Brechin [Kay72].

Brenner [WB20]. Breslow [Bro19]. Brewer [Ogs73]. Brian [For00, GWM11, Gar81a, GY62, Loc96, LW09b, Mit90, MP24, MR13, PSW16, Wäs02]. Brice [Ben19]. Bridges [Win70]. Bridget [Vau87]. Bridgman [New62]. Briggs [Rob86]. Brighton [Fle57]. Brigitte [MO18]. Brims [She82]. Brink [SB22]. Brisco [Cla95]. Broadbent [Wei94]. Brockhouse [Cow05]. Broglie [Abr88, Wil61]. Bronk [Adr76]. Brooke [Hun03]. Brooksby [Sel07]. Brouwer [KN69]. Brown [Adr66, Coc03, DL03, Gai19, Gar75, MP74, RS07b, Vin98]. Browning [Oak73a]. Bruce [Men70, Gre14, PR22]. Brugt [Har04]. Bruno [Fel60]. Brunt [Sut65]. Bruyne [Kin00]. Bryan [Bro23, Cap99, Gra90]. Bryant [KW79]. Bryce [Has16]. BT [Bul94]. Buckingham [CO22]. Bucks [Woo95b]. Budden [Dou07]. Bülbring [BB92]. Bullard [McK87]. Bullen [Jef77]. Bullerwell [Dun78]. Bullough [Wil22]. Bulman [Stu75]. Bunn [Arn91a]. Burbidge [LR17, SL21]. Burch [All84]. Burcham [DGV11]. Burdon [Pir66]. Burgh [Bar75a]. Burhop [MD81]. Burke [Hib21]. Burkill [Pit94, Pit98]. Burkitt [EE95]. Burma [Zuc81]. Burn [BW84]. Burnet [Fen87]. Burnham [Kro95]. Burnley [GB71]. Burns [RB20, Sel07, TC81]. Burnstock [NC21]. Burton [BI96, Pet11]. Burtt [Arn91b]. Butenandt [AA98]. Butler [But01, May79]. Butterworth [Dor18]. Buxton [Wig56].

C [BB97, BG95, CS98, CE96, Doc98, Edw96, FR98, Gre96, HD98, Kro95, LKK97, Lig98, Loc96, LG20, Mil97, MR96, New98, Ove97, Pip98, PS97, RP96, Roo96, Str96, Tho98]. C.B [Coo99]. C.B.E [All99, CJ99, Car99, Coo99, Cro00, Fow02, Mab99, Nee02, Pag01, Rob00, San00, Wea02, Whe01]. C.H [Ash02, Coo99, GR00b]. C.M.G [Bur01, DWS02]. Cabot [Gra90]. Cain [Cla08]. Caldwell [Den81, ML19, Pau18]. Caldwell-Moore [ML19]. Calestous [Cla21]. Callaghan [Kel17]. Callan [Gal03]. Calland [KP78]. Callow [Neu84]. Calvin [SB08a]. Cambridge [Hod79, Hux97]. Camden [Pic74]. Cameron [LC17, Oak68]. Campbell [Bro23, Bur94, Nai21, Sav63, Wes95]. Campion [Kri95]. Cannon [Smi63]. Carew [CA01]. Carl [Cha58, Dun96, Oak73a, Ran86, Rus68, Sal64]. Carrington [BS16, Wig82]. Cartan [Ser09]. Cartwright [Hay00, Web17]. Caryle [Her81]. Casey [RRS20]. Casida [HL23]. Casimir [Har04]. Cassels [DFSV23, Hol96]. Catcheside [FJ95]. Cattanach [PR22]. Caullery [Kei60]. Cavalli [Edw21]. Cavalli-Sforza [Edw21]. Cavendish [TP75]. CB [Bro16]. CBE [Bar04, BSHC18, BY16, Blo04, BM05, Bro15, Bur19, CEG^{+}19, CC10, Cla09, Cow04, Cra20, Dor18, DGV11, Edw18, Eva05, Fea18, FE12, Gar15, Gly11, GR16, Goo16, GLD18, HW10, Har18, HW04, Hun17, JJ14, Jor04, Lov04, LL05, Mac12, MW20, Moo15, Mor19, Owe09, Pic16, PSW16, QG03, Ric14, Riv20, Sou11, TH16, Wal06, WB05, Wel05, WK03, Yat06]. CC [Sto03]. Cecil [Dou19, Kay68, Owe09, Tod73, All84, DN74, FP71, GBC85, McC60, Pud21]. César [NA05]. CH [Blo04, Bro15, CEG^{+}19, MR22, NA05, PB10, WF19]. Chadwick [MF76, Ogd20]. Chain [Abr83]. Chalmers [Men64, Leg85].

Chaloner [CSB22]. Chambers [San21]. Chandler [ACW20, Wad97]. Chandra [Rao73, Lan85]. Chandrasekhar [GLR07, Tay96]. Chandrasekhara [Bha71]. Chaplin [Irv23]. Chapman [Bow58, Cow71, Qui01]. Charles [Bor17, Bro70, Bro19, But01, Can14, Car64, CKT15, DGL16, Dic75, DAB97, Fal70, FR98, Joh03, Jon74, Joy92, Joy16, Mar79, McC08, Mid21, NN00, New60, Pat05, Pir73, Pit94, Pit98, Qua94, RRS20, Smi98, SW90, Tho55, All83, AB74, And58, Arn91a, Ben08, Bla60a, Chi56, Cla61, DG70, GWM11, HPR72, HGM81, Lai20, Lyo01, Man03, Mat88b, MW01, MS19, Moo15, Mor92, Neu78, Pfe60, Por96, QRJ12, Ric81, Rob00, SLO10, SC99, Tho63, WM84, fWT94, Woo71, Woo21, YH82]. Charlotte [Bea95]. Charlton [ST19, TW22]. Charnley [NW84]. Charnock [Car99]. Chart [Sch64]. Chatt [EL96]. Chelsea [Lov75]. Chern [Hit14]. Cherry [BT68]. Cherwell [Tho58]. Chesters [Gre00b]. Chibnall [SW90]. Chilton [Joh00]. Chilver [Tho13]. Chisholm [Pud17]. Christian [Atk04, SB08b]. Christie [MT78]. Christoffel [Coo01]. Christopher [Den81, Edw96, GM06, Hun08, Ran22, Tur58, Whe01, CC10, DS79, HB20, RR24, Sho72, Tyr91, Woo95b]. Christophers [SG79]. Christopherson [DH04]. Churchill [Jon66]. Ciprian [Sho85b]. Clack [AS21]. Clapham [Wil94b]. Clark [Bro77, DM20, Zuc73]. Clarke [Bra15, BR14, Bro23, Gol65, Her14, Wea02]. Clarricoats [SH21a]. Claude [LKK97, Cun64, FGOS15, Hin60, Jam09, NG96, TP75]. Clavering [Mar86]. Clayton [CRST83]. Clement [Bri68, Fea04]. Clemo [LS85]. Clennel [Kin22]. Clifford [But01, KE04, LJ97, BLT11, BLT12]. Clifton [Bat69]. Clive [Dom21]. CM [DM09, Hyn15]. CMG [SL18]. Coales [Mac03]. Cochran [Woo97b, Woo05]. Cockayne [Hir15]. Cockcroft [OP68, Saw92]. Cockerell [Whe01]. Cohn [BS14]. Cole [Fra60, Hux92]. Colebrook [Oak71]. Coles [Cap99]. Colin [Mad06, MP24, RH18, For99, Smi05b, SH21b, Wri24]. Collar [Bis87]. Collier [Hew99]. Collingwood [CH71, GB13]. Collip [BR73]. Coln [Mit91]. Coluthur [RSG21]. Commander [Har04]. Compton [BM17]. Conant [KW79]. Conrad [Ano19, Mil18, Pat82, Rob77]. Constant [HCH73]. Constantine [Dow94, New61]. Conway [Cur22, Mai69]. Conyngham [Bul57]. Cook [Elv91, PM88, Qui05, Rob76]. Cooke [Bun94]. Coombs [LW09a]. Cooper [Edw18, Whi94b]. Copner [New98]. Copp [Mac00]. Cori [Ran86]. Corneille [Kei60]. Corner [Mab99, Zuc83]. Cornforth [BY16]. Correction [Ano13, Ano19, BLT12, Gre13, HP09, HF08, PC11, Sma10, SBG11]. Corrigenda [Ano66, Ano73, Ano74, Ano78, Ano79, Ano80, Ano81, Ano82, Ano84, Ano90b]. Corrigendum [Ano75, Ano83, Ano85, Ano86]. Cosslett [Mul94]. Costa [Cot72]. Cotton [CL08]. Cottrell [SK13]. Coulson [AB74]. Cour [Lew86]. Courrier [Tat90]. Cowan [Rai08]. Cowdery [Hol01]. Cowey [Pas19]. Cowley [Bir17, JKL^{+}22, MS06]. Cowling [Mes91]. Cox [Bel01, Cru00, Rei24, RWN66]. Coxeter [RW06]. Craig [HR18, WB68].

Craigie [And79, Sac94]. Crane [ML85]. Cranfield [Tho13]. Crew [Hog74]. Crichton [Ruf02]. Crick [BM17]. Crighton [HP01]. Cripps [Sch55]. Crisp [FS92, McK87]. Crombie [Pat01]. Cronin [Wat18]. Cross [HD98, Hut84]. Crossland [Bur19]. Crossley [Cha97]. Crow [HT14]. Crowe [Fle99]. Crowfoot [Dod02]. Cruickshank [BH18]. Crundall [Cre67]. Crutzen [Mül22]. Cudmore [Baw58]. Cullen [Dav18]. Cunningham [Ram64]. Curran [Fle99]. Curtis [RP20, Roc70]. Cushing [SH24]. Cuthbert [Edw20]. Cyril [Hay02, HU19, LKK97, Sco56, Wea02, Whi01, Eva76, Lew83, LJ97, Tho73a].

D [BFIT18, Hux97, Owe95]. D.B.E [Hay00]. D.L [Gab02, Lor98]. D.S.C [BB99a]. Dacie [LMW06]. Dai [LL23]. Dainton [GI00]. Dale [Fel70]. Dalgarno [HV20]. Dalitz [AS16]. Dalton [AM16]. Daly [Bar75a]. Dalziel [Gut96, Bow67]. Dame [BB22, Hay00, Owe95, Pep03, vEG06]. Damian [Lau23]. Dan [CJ12]. Danckwerts [Den86]. Daniel [Dys99, Gai19, HHP22, KM15, Oli94, Tay17]. Danielli [Ste86]. Daniels [Cox03]. Darashaw [Stu70]. Darlington [Lew83]. Darwin [Gly11, Tho63]. Datta [Bar23]. Davenport [Bla77, HB71]. David [BF24, BS01, Bre14, Bro19, CR07, Coc63, Dal97, Dou19, Ell22, Fit10, Fit22, GCJM18, Hig24, Ing64, LL23, LC17, Mar66, QRJ12, Rei24, Sau94, AS24, SSEF23, ACW20, BM81, BFO^{+}19, BC07a, BF72, Bow58, CK06, Caw24, CG20, DST85, Fei95, Fie08, FJ95, GI96, HF09, Hil97, Hir15, Hor14, HP01, HR18, Hux03, Joh00, KM70, KTR07, Kin09, Leg22, Man64, Mas71, MT78, McL04, Men64, Mil20, PW74, Pip05, PB17, RP20, Rei21, RH18, Sau61, Sau87, Sha15, SH24, ŠR24, Sti65, SB22, Sut65, SDV19, Tho74, Web17, Wis83, Wur16]. Davidson [Neu73, Woo71]. DAVIES [Kus01, GLD18, Nee02, Pud24]. Davis [Vin19]. Davison [Cor60, GCK17]. Davydovitch [KL69]. Dawes [Lig98]. Dawson [Cus05, Irv23, BS24]. DBE [CJ23]. Deacon [Cha85b]. Dean [Lew83]. Deane [GLD18]. Dearnaley [SBG10, SBG11]. Debye [Dav70]. December [Akh00, All00, All99, All91, Alp88, AS02, AA02, Arn94, AKS06, Ash22, Ask90, AH97, Ati96, Ati07, Att91, BB99a, Bar11, Bar86, Bat03, BSHC18, BY16, Bha24, Big02, Bil90, BS01, Ble97, BB92, Bos12, Boy23, BBM13, BG14, Bro80, Bro13, BM23, Bry01, Bud88, BB18, Bur94, CK06, CMO87, CU76, Car99, CJ12, Cha02, Cha97, Cla21, Col18, Col02, CW11, Coo99, Cro19, CNO23, CH19, Cur22, Dai85, Dav18, DM20, DWS02, DL04, Dic75, Doc98, Dod20, Dom98, Dor18, Duc91, DS21, Dul95, DR22, Dys11, Dys22, EFP17, Edw18, Ell08, EP10, EW22, Eva22, FGOS15, FSC20, FR24, FP04, For77, For00, Fow02, FAL21, Gar80, GT23, GCJM18, Gin94, Glo20, Gol17, GV18, GV92, GM21, GR85, GI00, GRT84, Gre96, Gre00b]. December [Gre19, Gri11, GSL19, GR00b, Gur12, HL23, Han00, Har91, Has16, HS07, Hay00, HB20, Hig24, Hil97, Hil90b, Hin90, Hir15, Hit14, HB22, Hog87, Hon95, HR18, Hut23, Hux00, HD92, Igg06, Ive06, Jon79, JB84, Jon23a, Joy16, KH97, Ken82, Kin22, LKK97, Lar09, LN16, Lew83, Lig90, Lis78, Lon11a, Mac23, Mad06, MR13, Mat90b, ML19, MS98, McC88, McC08, MO23, McL04, MP07, MS15, MB24, Mof85, Mor88, Mor92, MP77, Mot87, Mot92, Mül22, NS24, NS82, Nic16, Nor91, O’H06, Och90, OG14, Ott04, PSCV21, Pas87, Pas19, PM94, Pep03, Per88, PR78, Pow87, Pro23, QT86, Qui05, Ran86, Rat75, RP20, RS07a, Ric85, RC07, Ric14, Rob76, Rog91, Rub91, Sac94, SH18, BS24, Sel07, SOD22, SL94, Sou87, SH21b]. December [Sto03, Stu77, ST19, SF91, SP79, Tei03, TJ24, Ter10, TC87, Tho24, Tre11, TO12, Tyr91, VC18, Wäs02, Web17, Wei90, Wel78, WP75, Whi79, Whi22, Whi90, Wid95, Wil94a, Wil94b, WD21, Wol03, Woo97a, Woo21, Yud10, Zuc83]. Dee [Cur84]. Delbrück [Hay82]. Delisle [RB20]. Demetrios [SL18]. Denbigh [Row04]. Denham [PM94]. Denis [FS20, Har91, EE95, Key90, LY24, Phi95]. Dennis [Mor16, Sla94, All80, EP10, FS92, Gre00a, Gre19, Qui01]. Dent [HH74, Neu78]. Denton [Cla09]. Denys [Jef85, Mit91]. Derek [KH97, LM02, PK24, CW11, KH83, Stu13a]. Derman [DH04]. Desch [McC60]. Desmond [Hod80]. Desty [Phi95]. Detlev [Adr76]. Devendra [GM21]. Devereux [GF65]. Dewar [Mur98]. DFC [Jon10]. Diarmuid [Mor19]. Dickens [TC87]. Dickinson [San21]. Dickson [Whi66a]. Died [Fle57]. Dietrich [OM80]. Digby [HS07]. Dirac [DP86]. Ditchburn [Ser88]. Dixey [Dun83a]. Dixon [Ash22, Bra12, Cha55, Hin60, OKC21, Per88, Wol03]. Dixon-Lewis [Bra12]. DL [How15]. Dmitrievich [Tak22]. Dobson [HW77, RR24]. Dobzhansky [For77]. Dodd [SF91]. Dodds [Dic75]. Dodson [Dod20]. Doll [PB10]. Domagk [Col64]. Donald [Gra04, Joh77, Nai21, Pri78, CD88, Cro19, DM09, Eva20, GF65, JS10, MM96, Nee02, Nic16, ST19, Wei94]. Donnan [Fre57]. Donohue [Pri94]. Doodson [Pro68]. Dorey [Smi73]. Dorinthea [CJ23]. Dorothy [CJ99, Dod02, Tei03]. Douglas [Ati76, Bil90, Cla00, Cun64, HG11, Her82, HHP22, Hod79, OP68, RE04, Rit00, AT60, Bow05, Dar58, DWS02, Mac00, Ott04, RS07a, SA15, Wol01, YPS78]. Downie [TM90]. Downs [Jon91, Mor19]. Dowson [Bra22]. Drummond [HGM81, RP96, Whi99]. Drury [Kek81]. Duc [Wil61]. Dudley [Boy56, LPP19, SR17, Ubb81]. Duke [Hen90, Lon21c, LMA80]. Duke-Elder [LMA80]. Duncan [Rel61, Bra22, Bur21]. Dunham [Joh03]. Durward [BH18]. Dwyer [Mor88]. Dyson [All91].

E. [CS98, Doc98, Far98, FH98, Kib98, Mes98, New98, Tho98]. Eaborn [Smi05b]. Eamon [Syn76]. Earl [Bri68, Dod57, Hog87, Kay68, Kre73, Lin01, GN24, Wal08, Zuc81]. East [Mak94]. Eastbourne [Fle57]. Eastwood [Jon83]. Ebenezer [Ing68]. Eberhard [Sko61]. Eccles [CA01, Rat71]. Eckersley [Rat60]. Edensor [Bur78]. Edgar [Cow04, DN74, GV18, Her82, MH90, Web17, Cru76, Hod79, Mar64]. Edgell [Dea63]. Edinburgh [Lon21c]. Edith [BB92]. Editorial [Ano13, Lon16, Lon17, Lon18a, Lon18b, Lon19a, Lon19b, Lon21a, Lon22c, Lon22d, Lon23a, Mea08, Mea09, Mea10, Mea11, Stu12, Stu13b, Stu14, Stu15, Lon20a, Lon20b, Lon21b, Lon23c, Lon24a, Lon24b]. Edman [PB79]. Edmond [EB08, LE80]. Edmund [BH88a, Lyo01, MH08, Whi90, Bat69, Bel81, Cla95, SP76, Tem56]. Edoardo [Rub91]. Edred [Mab99]. Edward [Baw61, Bur68, BT20, Cha01, Cha85b, Cla61, Eyr13, FD87, Gad21, Gar88, Gar21, Goo72, Goo16, Gra69, Gre98, HL23, Jef77, Jel99, JJ14, Kay68, Law08, Mot55, Pol00, Rob86, Rob95, Roc70, Sin03, Smi15, Str96, WF19, Zuc73, vE57, AR09, Bar21, Ble99, Bor17, BMW92, Car64, CH71, Cla80, Cot72, Dal55, Dic75, Gar74, Gui02, Hyn15, Ing64, Mai69, McK87, Pal09, Pay81, Phi83, Rat66, Riv20, She03, Stu65, Tho66, TG71, Win70]. Edwards [Bar04, Fer17, Gar15, New98, Pfe60, War17]. Edwin [Cro10, Lon23b, Pud21, WPL21, BK17, Cam94, Hil87, Ter10]. Egbertus [KN69]. Egerton [New60]. Eglinton [BE24]. Egon [Ale01, Bar81, FW83, NA95]. Einstein [Whi55]. Ekins [Vin23]. Elder [LMA80]. Elderfield [MH24]. Eleanor [SL21]. Eleazer [DJ13]. Elected [Bur03, DP86, DL03, Gal03, MDHW03, Hun03, Hux03, Joh03, Jon03, Mac03, Pep03, Rox07, Sil03, WB03]. Eley [HHP22, Hog74]. Elijah [Pug58]. Elion [Ave08]. Elizabeth [GB14, Gut21, Lon22a]. Ellesmere [Joh00]. Ellington [Woo21]. Elliot [BO58, Sou11]. Elliott [Dal61, IG21, Pic16, Tho24, Wal06]. Ellis [Cox03, Har99, HGM81, Mul94]. Elmer [CP69]. Elsie [Ash02]. Elton [SC99]. Elwin [Ble01]. Elwood [Jam09]. Emanuel [Gia16]. Emeléus [Gre96]. Emett [Har81]. Emmanuel [Fle55b, Sho85b, Wil72]. Emmeline [Ran75]. Emperor [Cor90]. Endel [Cra24]. Enders [Tyr87]. Eng [Roo96, Woo97a]. Engledow [Bel86]. Enrico [BC55]. Enrique [Neu78]. Ephraim [Edm20, ASS16]. Erdélyi [Jon79]. Erdõs [BB99b]. Eric [Ada09, BN92, Cla09, Pig77, Wei94, Ber04, Bla81, Cal21, Ele76, HH95, Jon83, MD81, Ste20]. Erich [HLH82]. Erik [Ban83, Ran22, Pat90]. Ernest [Cru00, Dou64, DGV11, KM69, Kus01, Law01, MBW99, Phi13, PC10, PC11, SSGP23, Ter10, Whi94a, Whi03, BCB61, Col71, Coo58, CP94, DP70, Fle71, Hun08, Mak18, Mat90b, Rey07, SB77]. Ernst [Ati14, DGS18, Lee07, Ril60, Abr83, Boc06]. Errata [Ano71b, Ano90c, Ano91, Ano92, Ano94]. Erratum [Ano68, Ano77]. Errol [Stu85]. Ervin [BS08]. Erwin [Hei61, PB20]. Eshelby [Bil90]. Esmor [HW17]. Essen [Coo98]. Esslemont [Bar98]. Ettingdean [Win70]. Eugen [Kur58, Eve84]. Eugène [FA03, SVW00]. Euler [Bla85]. Eustace [Ing68]. Evan [Lon11a]. Evans [AC72, BM22, CM81, DG70, DST85, Gre00a, Hut23, Kin16, Qua94]. Evelyn [Lov11, Pri94, Pow87]. Everard [Cul90]. Everest [Sal78]. Everett [Bre05, Ott04]. Evershed [Str57b]. Evgenii [ZK90]. Ewald [BH88b]. Ewart [Jon03, Whi96]. Ewing [Bul75]. Ewins [Caw24, Dal58]. Eynsham [Pic68]. Eyre [PSW16]. Ezer [Dar62].

F [Mil97, Roo96, Woo97a]. F.R.Ae.S [Pag01]. F.R.Eng [Chi99, Hew99, Pag01]. FAA [Bos12, BBM13, BSD13, FGW13, Hum11, HR18, Kin16, Rae12]. Fabes [Smi73]. Faddeev [Tak22]. Fage [Col78]. Fairley [Boy66]. Falcon [Ano98]. Falconer [Bow05]. Falkow [RC24]. Fallowfield [CC61]. Fano [BIR12]. Farley [Cre24]. Farren [TH71]. Fasham [AB15]. Fatt [Ash16]. Fauré [Wil72]. Fauré-Fremiet [Wil72]. FBA [Mac12]. Fearnsides [Bul69]. Feather [CD81]. Featherstone [Gil20]. February [Ada85, Ada09, AS16, All80, All83, Ant15, AM16, AR09, Ave08, Bak76, Bak77, BR79, Bar75a, BM81, BH88a, Bat03, Ben08, Ber04, BIR12, BR18, Ber20, Bha24, Big10, Bir17, Bis87, Bla85, Ble96a, BL24, Blo04, Boc06, BG11, Bol97, BN92, BM98, Bou15, Bow05, Bra95, BM05, Bre14, Bro23, BLT11, BLT12, Bro24, Bun94, BM13, Bur21, CK06, CS98, CR16a, Cap99, CMO87, Cha90, CL08, CR16b, CO22, Cle20, CV94, CDS12, Cro10, CH19, Cul95, DM20, Dav23, DM06, DE19, Dul95, Dun96, Dys99, Dys22, EE95, Eva20, Fei95, FR98, Fle99, Gab02, Gai19, GB14, Gar75, GW20, Gar81b, Gia16, Gil20, Goo16, GM21, Gra20, GTK11, Gri18, Gui97, Gun00, Har81, Har99, HLH82, Hew02, Hil90a, HT16, Hil11, Hir90, HW04]. February [Hof07, Hog87, HF06, HF08, HW17, HKB23, Hon95, Hor14, HW77, Hua18, Hux00, Hyn15, Imu95, Jac00, Jam09, Jef76, JRS84, Joh00, JG82, KT20, Kem75, Ker79, Key90, Kib88, KW79, KS92, KH83, Law01, Leg85, LMW06, Loc96, Lwo77, Lyn15, LM94, Lyo22, Mac77, Mad90, Mah24, Man03, Mar88, Mar86, MW20, Mas23, Mat90b, May79, Meh75, Meh02, Mit90, MS06, Mor88, Mor21, Mor19, MP77, Mui94, GN24, NS82, Neu84, Neu90, Nic16, Nut77, Nut90, OM80, PSS78, Pau18, Pea18, PK24, Pet92, Pet24, Pit94, PB17, Rae12, Ran22, RS15, Ric85, RB20, Ric81, Ric15, Riv20, RW06, Rob86, RS07b, Sac94, SH18, Sau94, SL18, Lor98, Sew15, Sha21, SG79, SO20, SK13, Smi98, Smi05b, SB08b, ŠR24, Spe94, Spi91, SQ88]. February [Stu75, SD15, Tak22, Tay17, Thi84, Tho20, TC76, Tod88, Tyr87, Vel19, Vin09, VC18, Wal08, War17, WB05, Wel00, Whi99, WH77, Wil23c, Win01, Wol98, Wol03, Woo97a, Woo95b, Wur16, Yat06, YH82, ZK90]. Federico [Och90]. Feilden [BM05]. Feldberg [BB97]. Felix [Cra57, FA03, BS67, DR22]. Fell [Vau87]. Fellows [Ano71a, Lin56, McR76, Wil95]. Fenner [BBM13]. Feodor [KD82]. Ferdinand [All73, Blo04, Ran86]. Fergus [Wes95]. Fermi [BC55]. Fermor [CA56]. Fèvre [AB88]. Feynman [Meh02]. FIBiol [RSD16]. FICE [BM13]. Field [WPL21]. Fielding [Sim18]. Fildes [GKW73]. Finch [Bla72, CH19]. Fincham [HF06, HF08]. Finney [Gri11]. Finniston [Pet92]. FInstP [Joy16]. First [Bri68, Hog87, Hol66, Pic68, Sch64]. Fisher [YM63]. Flavell [Mac03]. Fleck [Hol71]. Fleetwood [Hey87, Kur83]. Fleischmann [Wil23a]. Fleming [Col56, Mat88b]. Fleure [Gar70a]. Florence [Mor97]. Florey [Abr71]. Floyd [CGZ22]. Fogg [Wal06]. Folley [Par72]. Forbes [Mas71]. Ford [Cla95, Har99, Lyo01, Wil13]. Forrest [All94]. Forster [SC23]. Foster [Bel66]. Fotis [LG20]. fourth [GN24]. Fowler [Wol98]. Fox [Smi68]. Foyle [CH71]. Fozard [Far98]. Fraenkel [BT20]. FRAeS [Mor16]. Frances [FR24]. Francis [Boy56, Cou73, CNO23, Lew86, MH19, Mor05, Pip98, VC18, WW99, Wil23b, All76, And71, Bar75b, BM17, Cha90, Cre24, Fra60, Gib73, Hog74, MH90, Mar79, Oak73b, Phi74, Por66, SL90, Wil94a]. Franck [Kuh65]. Francois [Mor17]. Frank [AL23, Cox57, FH98, Hal99, Jam90, KM15, LP91, NN00, Whi96, Woo95b, AS03, Bel86, BC18, BBM13, Bro10, Bur94, CL08, Dun83a, Fen87, Fin95, For00, Gil20, Goo72, Har20, MS96, Men60, MS08, PC10, PC11, Ran90, Tab69, TC87, Whi58]. Frankel [Eva99]. Franklin [DM68, HT14, Hil88, Tyr87, MP75]. Frantisek [SV91]. Franz [LN16, Dys99, JG82, Kur58, Tod88]. Fraser [Mar79, Pol92, BSD13]. Frazer [Pug61]. Fred [Bur03, RS07b]. Freddie [TG23]. Frederic [Gra82, Ste86, Bla60b, Bro70, Cla55b, KP78]. Frederick [AKS06, Ber20, BR14, Bul66, Col71, Cru04, Gar21, Gre00a, Hod56, Kil94, KH90, MR96, NN00, O’H06, Rey07, Tho55, Yat06, Bro15, Cha61, Coo94, DS86, EFP17, Fle57, Fre57, GI00, HH74, Jef85, Joh23, KE04, Kri95, Mil84, Pir73, PR63, Sta69, Tho58, Whi03]. Fredrik [Sal64, Cha58]. Freeman [KM24, LM94, Dys22]. Freeth [All76]. Fremiet [Wil72]. FREng [BM13, DG16, Eva05, Goo16, HW04, HJ11, PSW16, SK13, Tho13, Yat06]. Freudenberg [Ste84]. Frew [DWS02]. Frey [Mat90a]. Frey-Wyssling [Mat90a]. Friedel [SD15]. Friedlander [EFP17]. Friedrich [AA98, CV17, Fil56, Ati14, Eme60]. Frisch [Pei81, Tho83]. Frits [Tol67]. Fritz [AM13, JW00]. Fröhlich [BT05, Mot92]. FRS [Bur03, DL03, Eva05, Gal03, MDHW03, Hun03, Hux03, Joh03, Jon03, LR19, Pep03, Rox07, SL18, Sil03, Vin09, WB03]. FRSC [Hyn15, RS15]. FRSE [GR16]. FRSNZ [KM15]. FTSE [FE12]. Fuchs [Law01]. Fukui [BN01].

G [Cor90, GMP24, Hut23, LKK97, Ree94, Whe74]. G.C.B [Lor98]. G.C.M.G [Lor98]. Gabor [All80]. Gabriel [Bro13]. Gaddum [Fel67]. Gadsby [Mil20]. Gale [FS20, Rey07]. Galton [Tho63]. Gareth [Hil11]. Garner [Baw61]. Garnet [Gal03]. Garnham [LKK97]. Garrett [Har91, Por96]. Garton [Con09, Wah23]. Gascoyne [Tod73]. Gascoyne-Cecil [Tod73]. Gaskell [Bra92]. Gass [Cox95]. Gasser [Adr64]. Gaston [Kei60]. Gatenby [WR95]. Gates [Rob64, TK74]. Gavin [RW18, Bar73]. Gaydon [Wei05]. Gaylord [Whi86]. Gaze [Hop23]. Gee [All99]. Gehring [MA21]. Gell [Sil03]. Gelli [ST99]. Gelston [Poo60]. Gennes [JC19]. Geoffrey [CT14, Gar15, GG00, Hit06, MF18, RS07a, Rus24, All99, Bat76, Bha24, BE24, BM05, Cot84, Faw18, Har81, Lig98, LR17, NC21, PT11, SBG10, SBG11, Vog72, Wal61, Woo90]. Georg [Att91, Sho83]. George [AR09, Bar04, Bra70, Bul69, Che62, CCG^{+}00, Cul95, DT81, Dou07, Fre57, God57, Gui02, Hib21, HP01, Kin09, Lea91, Mac23, Mak94, Mes91, Mil84, MR96, PO21, YHP^{+}86, Phi83, Pud24, Ran90, RS15, Row04, Rus61, Sal66, Sil03, Smi99, Spe94, Stu66, Vin98, Wat56, Wei24, Adr63, Akh00, Ark55, Bat03, Bla72, Ble96a, BS05, Bun94, Car14, Cha85b, Cla70, Coc67, Dai85, Dom98, FP04, Fog86, Gol65, GR16, Haw60, Hor95, Kil94, Lov11, LS85, MP74, MP82, Mof02, Moo77, Nut90, Pro64, RE04, Rob86, SS06, Smi15, Stu77, Tho55, Tit58, Tod79, Wah23, Wal08, WW99, Wel00, Whi66b, Whi86, Wol03, Zuc83].

Gerald [Bar75b, HW87, Hol94, Bul57, MS15]. Gerard [EFP17, Nut77]. Gerardus [PG92]. Gerhard [Ble96b, Har04, Col64, Sto03]. Germain [Bas63]. Gertrude [Ave08]. Gherbury [Mar66]. Gibb [HP60, Hin60]. Gibson [LP91, OG14, PC10, PC11]. Giffard [Lin01]. Gilbert [CSB22, GT23, JP81, Ker79, Tay62]. Gill [BH88a]. Gilles [JC19]. Gillespie [Dav23]. Gilman [Eab90]. Ginzburg [Lon11b]. Giuseppe [Big02]. Glanville [Bak77]. Glass [Wis83]. Gleb [Gad56]. Glenn [Hof07]. Glenny [Oak66]. Glueckauf [Eve84]. Gluecksohn [Pap19]. Gluecksohn-Waelsch [Pap19]. Glyn [New60]. Glynn [LK23]. Gödel [Kre80]. Godfrey [BC60, SDV19, Wel05, McC22]. Godwin [Wes88]. Gold [Alb87, Bon06, SB77]. Goldsbrough [Pro64]. Goldstein [Lig90]. Goodeve [Ric81]. Goodrich [Gar81b]. Goodwin [PO21]. Gopalasamudram [VJ05]. Gordon [Bel01, Cru00, Cun64, GKW73, Gra20, Hew99, Pau18, Wei05, Whi79, Aca08, CB81, HW77, Oak68, OKC21, Ram64, She82, Sti05, Wal06, Wal66]. Gorer [Med61]. Goring [FR98]. Gorrie [MW01]. Gosta [Pea98]. Gough [Dor66]. Govind [Cow21, KC23]. Gowing [Mac12]. Graaff [Bom67]. Graaff-Hunter [Bom67]. Grace [BS90, Tod58]. Graeme [Web10]. Graham [Adr66, AA02, BC07a, Cox95, GB13, Hin58, HKB23, MO23, PSS78, Smi63, AW88, Bra12, Con22, Lyo22]. Granit [Gri95]. Grant [Tho91]. Granville [Abe61, Tho98]. Graves [EF78]. Gray [CH58, GR16, Gui02, Lew67, Lis78, LS66, Sco14]. Greaves [Red56]. Green [Cha01, Cra21, DE19]. Greenland [Gre19]. Greenwood [Bha24, Bur21, JWW^{+}20, Pat97]. Gregory [Doc98, Hir90, LH18, PR63]. Grenville [Chi99]. Griffith [Rub64]. Griffiths [Dar62, Mit00]. Grigg [SSGP23]. Grigorievich [For77]. Grigory [CB22]. Grime [PF21]. Gros [Zuc73]. Grüneberg [LH84]. Guerrier [TH16]. Guest [VT88]. Gugenheim [PR63]. Guggenheim [TG71]. Guido [Sid02]. Guillery [She18]. Guinness [Kay68]. Gulland [BC92]. Gustaf [Whe74]. Gustav [Flo20]. Gutfreund [TG23]. Guthrie [FJ95]. Guttmann [Whi83]. Guy [DH04, Sta58, Dod20, Gra82, Tho60b]. Gwendolen [Mor97]. Gwyn [BW90, Hil11]. Gwynne [DST85].

H [Cra21, Mil97, Pip98]. Hadamard [Car65]. Haddow [Ber77, Gar80]. Hadji [Spa14]. Hague [GTK11]. Hahn [PB20, Spe70]. Hailsham [Lew02]. Hal [Rob77]. Haldan [GR85]. Haldane [Pir66]. Hales [SLO10]. Hall [GV18, GRT84, Hau56, MJR24, RW18]. Hallett [Fel70]. Hallpike [WM84]. Halsbury [Lin01]. Hamilton [Boy66, Gra04, Oak73a, AS24, Rus77]. Hämmerling [Har82]. Hammersley [GW07]. Hammick [Bow67]. Hammond [SE65]. Hamnall [Mit61]. Hamshaw [Har63]. Hanbury [DL03]. Handy [CKT15]. Hanes [fWT94]. Hankey [Sch64]. Hannah [Bra15]. Hannes [Pea98]. Hans [Ale96, CF76, Dea66, KW84, Lar09, LB07, LH84, NH13, Rot72]. Hanski [SO20]. Hanson [Ran75]. Harborne [Pre10]. Hardinge [Lin01]. Hardy [Fox87, Mar86, McL04]. Harington [HPR72]. Harish [Lan85]. Harish-Chandra [Lan85]. Harland [Hut84]. Harley [SL94]. Harold [App07, BG95, BW84, Cal21, CMO87, Cla71, Coo90, DG16, Gra90, Hog87, Leg17, LM02, LS66, Mac00, Pet92, Pud17, Sut70, Wat19, And57, Bir72, CRST83, Cul90, Gal03, HB71, Har56, MS98, Ogs73, Ric85, RW06, Smi68, Smi69, Tho65, Woo61]. Harper [TW21]. Harrie [BBD84]. Harris [Cha85a, Cro10, GS18, Gra69, Hop96, Vog72]. Harrison [Abe61, Gun00, Kev05, Pea68]. Harry [Ati07, BG02, BH08, BM17, Glo20, McC22, PK24, Sow24, Vin09, For12, Gre96, Gri18, Haw76, Hop96, McC81, Mot87, Ric14, SR76, Sou11, Sto94, Wes88]. Hart [MIH24]. Hartley [Dal57, Ogs73]. Hartline [GR85]. Hartree [Dar58]. Hartridge [Rus77]. Harvey [Coo72]. Hasbrouck [Ble82]. Hassan [Tho20]. Haszeldine [MT19]. Hatton [Sal66]. Haughton [Dun83b]. Havelock [BR71]. Haward [Dun68]. Hawes [DM20]. Hawkes [Dun82]. Hawking [CEG^{+}19]. Hawkins [All70]. Haworth [Fei05, Jon91]. Hawthorne [GN19]. Haydon [Key90, Smi69]. Hayes [BH96, BM24]. Hayne [HCH73]. Head [Hum11]. Heaphy [CE96]. Heaps [BF24]. Heath [Man98]. Hebb [MM96]. Hedley [Syn67]. Heidelberger [Sta94]. Heilbron [Coo60]. Heilbronn [CF76]. Heinrich [GW20, Kre72, Mot82, Ble96b, Her55, Kar58, Ril60]. Heinz [Bla94, Sho71]. Heisenberg [MP77]. Heitler [Mot82]. Helen [Har18, Nor91]. Helge [Pat90]. Helio [ST99]. Hemley [McC81]. Hems [JW97]. Henderson [Bod04, JH20, KM70]. Hendred [Mak94]. Hendrik [Coo01, Har04]. Henning [Hay82]. Henri [Ser09]. Henricus [PG92]. Henrik [Coc63]. Henriksen [Lar09]. Henry [AS16, Bak77, BC18, BR71, CS98, Cro19, Dun83b, ES05, Fea04, Fel67, GS18, GWM11, JG13, LL05, Mab99, Man87, MD81, McC60, New61, O’H06, OKC21, Phi95, Pud21, Rat71, RE08, Sal61, SH24, Tho13, Car99, CCG^{+}00, Cox03, Eab90, Far61, Fel70, FM21, Gar88, GV18, GR00a, MH24, Hod56, Nut77, Sal57, Sta58, Str57a, Wat56, Woo94, Wun97]. Henton [Eva05]. Her [Lon22a]. Herbert [Ask90, Boy23, Cam94, DE58, Elv91, Jon03, JKL^{+}22, McC08, Por68, PT11, Red56, YH82, Adr64, Ait62, All70, AC72, Bai88, Bla77, Dat88, Dor66, Gar70a, GY62, Jon60a, McL00, Mot92, Smi63, Sut70, TG23, Wil55]. Herman [Ada85, Rei21]. Hermann [Pra10, Rox07, Dac88, New57, Pon68]. Herries [Hir90, Son99]. Herriot [Ram64]. Herzberg [Eva99, Sto03]. Heslop [Gun00, Pea68]. Heslop-Harrison [Gun00]. Hevesy [Coc67]. Hewish [Lon22b]. Hewitt [Ber04, Tur55b]. Hey [CD88, Hew02]. Heyrovský [BZ67]. Hickling [Wat56]. Hide [Rea19]. Hideki [Kem83]. Higgins [GM06, SH18]. Highness [Lon21c]. Higman [Con22]. Hilda [Aga23]. Hildebrand [Coo72]. Hilditch [Mor66]. Hiley [BLT11, BLT12]. Hill [Ben94, Bul67, CJ99, Dol94, Hux03, Kat78, Mac23, Sad23, Sew15, Wat55]. Hills [Hil87, Lon23b]. Hilton [Fer17, Tod58]. Himsworth [BG95]. Hinde [BSHC18]. Hindle [CB81, Gar74]. Hinshelwood [Tho73a]. Hinton [Gow90, Sal78, Sav63]. Hirihito [Cor90]. Hirsch [MS90]. Hirst [BB99a, SP76]. Hirzebruch [Ati14]. Hoare [GBC85]. Hobson [Man60].

Hodge [Ati76]. Hodgkin [Dod02, Hux00]. Hogben [Wel78]. Hogg [Lew02]. Holder [YPS78]. Holliday [HF08]. Holmes [Dun66, Wal66]. Holroyd [Ros74]. Holroyde [CD88]. Holt [CDS12]. Holtedahl [Stø76]. Homan [Hin87]. Homi [Pen67]. Hondros [SL18]. Honeycombe [CBE09, Sma10]. HonFRSC [CR16b]. Honor [Vau87]. Hooker [YHP^{+}86]. Hoole [Stu77]. Hooley [HB20]. Hopkins [MS98]. Horace [MS98, Dar72]. Horlock [DG16]. Horn [Bro13]. Horne [Bry01]. Hörstadius [Jac00]. Horton [Cur22, Whi58]. Houghton [Whi22]. Hounsfield [Wel05]. Houssay [YF74]. Houthem [Sil03]. Howard [AM16, Mit00, SR17, Tyr91, Wol98, Abr71, Dul95, Glo20, Mon86, Sal78]. Howarth [Stu09]. Howieson [OG14]. Hoyle [Bur03, Whi99]. Hubel [Wur16]. Hubert [Sac94, Bed57, Jon10, MM21]. Hückel [HLH82]. Hudson [Pat79, Stu66]. Huggett [Bra70]. Hugh [AKS06, Gre00b, Hir15, May15, Ott04, Qui05, SOD22, Whi94b, Wil13, BB96, GM06, Har63, Hen90, HW17, Kem75, Sad23, SC23, Tho61]. Hughes [Ing64]. Hugo [Dal83]. Hulst [Coo01]. Hume [Hux97, Ray69]. Hume-Rothery [Ray69]. Humphrey [Ask90]. Humphry [Pat97]. Hunt [BM23]. Hunter [Bom67, Lau07, MP07, Wur16]. Hutchinson [Arn91b, Hub75, Lov11]. Hutchison [Roo96]. Hutton [RE04]. Huxley [Bak76, HW17, Sim18]. Hynek [FW83].

Iain [Mar12, Nai21]. Ian [All13, CG20, Den74, DFSV23, Wil23c, Cha02, Coo60, Cox95, Dor18, GVS19, Han18, LK23]. Ieuan [Duc91]. Iggo [Cer19]. Igorevich [KT18]. II [Lon22a]. Ilkka [SO20]. Illing [Fal70]. Ing [SR76]. Inge [Bol97]. Ingham [Bur68]. Ingle [Bla72]. Inglis [TP75]. Ingold [Sho72]. Ingram [Bat76, KC22]. Irene [Pre90]. Irvine [HL63]. Irving [Hyn15, Tay58]. Irwin [Wel00]. Isaac [Bul94]. Isaacs [And67]. Isaak [Bya22]. Isaakovich [CB22]. Isabella [Har18]. Issigonis [Dow94]. Ivan [Bar75a, CV85]. Iveagh [Kay68]. Iversen [Rob23]. Ivor [Cur84, Stu85].

J [BH18, BFIT18, Caw24, GMP24]. J. [Bet68, FM19]. Jack [Bre14, DS79, Leg85, CR16b]. Jackson [GB71, Jon60b, KH83]. Jacob [Mor17]. Jacobus [PG92]. Jacques [Car65, Cia02, Lwo77, SD15]. Jacques-Louis [Cia02]. Jacquin [Fle55b]. Jaeger [Pat82]. Jake [SW21a]. Jakob [MA21]. James [AGJ^{+}22, AB88, BFO^{+}19, Bra65, Bro24, Buc10, CEP15, Chi56, CH79, Cla80, Cla08, Cla09, Cre24, Dal58, DM68, Dav23, Gre06, Gre19, Gur12, HP08, HP09, Hon95, HH74, Kil94, Leg22, Lip73, MW20, Mat90b, MB21, Mur98, Ped01, SH21b, Tho24, Til61, Tod73, WB05, Whi01, Woo97b, All91, And79, BR73, Bom67, BN92, CK80, CGZ22, DE19, Dun68, Fox87, Fre60, HT14, HL63, Hew02, Hol96, IG21, KH97, KM69, KW79, Kuh65, Lau07, Lis78, Mad90, MF76, McC65, MO23, Mit00, Mor88, Neu73, Rob76, Ste86, SQ88, SF91, Tod75, VC18, Wat55, Wat18, WZ79]. Jamieson [All73]. Jamison [HNW94]. Jan [KN69, Woo57, Hul94]. Janet [Owe95, FP95]. János [GS13]. January [Aca08, Ada09, AS16, AA98, Ale01, All13, All76, All83, All91, AM16, ADE00, Arn91a, Arn91b, Ash16, Atk04, Ave08, Bal24, Bar98, Bea82, BC18, Ber77, Ber04, BS14, BN06, Ber20, BH88b, Bir17, Ble96a, Ble97, BL24, BB96, Bor17, Bra22, BS05, Bra15, BH96, Bro77, BMW92, BK00, Bro09, BM24, Bro16, Bry01, BN01, BS16, Bun94, Bur19, BS13, Bya22, CD88, CS98, CEG^{+}19, Caw24, CEP15, Cha85a, CH05, CR83, Cla86, Cla95, CR07, Cla09, CST17, CO22, CRST83, Con22, Cor90, CDS12, CJ01, Cra23, CVW02, CEW18, Cro00, Cro19, CA01, Dal97, DL03, DS79, Den81, DM06, DG23, Dod20, Dom98, DJ13, EF78, Edm20, Edw21, ES05, Ell08, Els82, Eyr13, Far98, Faw18, Fea18, Fit10, FS92, For77, Fry80, GS18, Gar75, Gar06]. January [Gia16, GVS19, Goo20, GLD18, GR00a, GP83, Gui02, Gut96, Hai24, HM77, HT14, HV20, KHG21, Haw76, Hay85, Hew99, Hit20, Hod75, HJ11, How15, Hum11, HGM81, HD92, JH20, Jam90, Joh23, Joh77, Joh03, JWW^{+}20, JKL^{+}22, Joy92, KM24, Kem83, Ker79, Kib98, Kil94, Kin09, KR11, Kre80, Kur83, LW09a, LKK97, LY24, Lau07, LPP19, LW20, Lig98, Lig90, Lin01, LR17, Lov11, LH88, LL05, Lyo01, Mab99, MH90, Mac00, Mak18, MD81, ME11, McC81, MO18, ML03, Meh75, Mes98, MM21, MS15, Mot82, Mot92, Mül22, New98, NJ86, Nut77, Nut90, O’H06, Owe95, Owe09, Oxl23, Pat90, Ped01, Phi95, Pip94, Pip05, Pol92, Pri94, Pug78, QT86, RC24, Rei24, RW18, RS07b, Row04, Rus77, SM24, SB08a, SVW00, SS06, Ser88]. January [Sho83, SA15, Smi05a, Smi99, SH21a, SSGP23, Sta79, AS24, Ste84, SSEF23, Stu86, Stu85, SB77, SW90, Tat94, Tay17, TJ24, TO83, TS24, Tod88, TM90, Vog08, Wal06, Wei90, WW86, Wes95, WL02, WB20, Whi04, WK03, WZ79, WD21, Wis83, Wun97, YL92, vEG06]. Japan [Cor90]. Jardine [Hun17]. Jaroslav [BZ67]. Jarvis [Gra20]. Jean [Oak62, Ran75, Wei90, And06, Bla60b, Pir90]. Jefferson [Wal61]. Jeffery [Tit58]. Jeffrey [Pre10]. Jeffreys [Coo90]. Jehangir [Pen67]. Jellicoe [Wal08]. Jencks [KR11]. Jenkinson [PB17]. Jennifer [AS21]. Jeremy [BF24, HW10]. Jerne [AH97]. Jerzy [KBP82]. Jessell [BW22]. Jim [MM21]. Jinks [Mat88a]. Jiri [DR22]. Joachim [Har82]. Joan [HCB18]. Joel [Cas73, Yud10]. Johan [Sal64]. Johann [AA98, KS69, Ste84]. Johannes [Fil56, GP83]. John [AB15, AD21, BY16, Bel81, Ber04, Bla81, BBM13, Bro16, Buc06, BS13, CK06, Cha61, CC10, Cow04, CA01, CE96, DG16, Dol95, Dor66, Dys22, Ell08, Eyr13, FS92, Gar70a, Gar80, Gib73, Gra10, Hig24, Hir15, HB22, HF08, Hol01, Hun08, Jon23b, KTR07, Kev05, LMW06, LC17, Lov04, Mab99, ME11, Mes98, Mon06, Pag01, Par72, Pau18, Pet24, Por66, Ric15, Rub58, SC23, Scr99, Ske70, SH21a, TB13, Tho66, Tur23, Wal08, WF19, Whi22, Wil75, All94, Ano19, Ant15, AL23, Arn94, Ask90, Ati96, BB99a, Bar11, Bel66, BC72, BN06, Ber20, BC92, Bil90, Ble82, Boy98, Boy23, Bra95, BR14, Bri58, BM24, BM23, Buc77, Bud88, BB18, BP99, Bur78, CH05, Cha97, Chr56, CC61, CST17, Cor60, CDS12, Cro10]. John [CH19, Cur22, Dea63, DG23, DFSV23, Dom21, DE58, Edw18, EB08, Ell22, Far98, Fel67, Fer17, Flo08, FE12, Gad57, Gar95, Goo82, Gra69, Gre00b, GW07, Gun00, HL23, Has16, Hay85, Hew99, Hey87, Hin58, Hir63, Hod80, HF06, Hub75, HD92, Jam90, Joh77, Jon60b, Kay72, Lin01, Lon11a, LM94, MH19, MT21, Mac03, Mas02, Mat88a, May15, May79, McC03, Mil18, Mof85, MS06, Mot55, GN24, New61, NW84, OP68, Pat82, Pea68, Pet58, PF21, Pip94, Pir66, Pit94, Pit98, Pol92, Rob60, RSD16, Row96, Sac94, Sca18, Sel07, Sen19, SW21a, SE65, Sle97, Smi05a, SL94, SB08b, Spa14, Spr95, Sta79, Sta86, Str57b, TH16, TW22, Tod58, TW21, Tur55b, Tur58, Tyr87, Uns14, Vin19, VT88, WPL21, Whi62, Wig83, WR95]. John [Wil87, WB85, Wil56, Woo95a, WBF23, vE57]. Johns [HS07]. Johnson [BB22, JB84, Rae12, Rei14, Sin03, Smi16]. Joliot [Bla60b]. Jolly [Rel61]. Jon [DM20]. Jones [Cla55b, Coo99, Eva22, HM77, MH90, Mot55, Mot87, Pug67, SA15, Sta79, Woo61]. Jordan [Ril60]. Joscelyn [Cox58]. Jose [KO99]. Joseph [AM13, Akh00, Azz24, Bra70, BM24, Bur21, Dav70, Fra60, Hil97, HT16, Mai69, PM88, Pon68, Sca18, Tay17, Vel19, Arn91b, CU76, EL96, GR00b, HF07, Lew67, Lov64, Mad06, Mar88, Ran09, Tur62]. Joshua [BG11, BW84]. Jozef [Mül22, HF07]. Jr [Cor60]. Juda [MS90]. Jules [Kei60, Oak62]. Julian [Bak76, DGS18, MIH24]. Julius [Edm20, Gre96, Pri68, Ive06]. July [Ale22, All76, All84, App07, ADE00, AGJ^{+}22, Arn94, AR09, AD21, Ati76, BS08, Bai88, Ban83, Bar11, Bar21, Bar04, Bar75b, BH18, Bel86, BIR12, Big10, Ble01, BF91, Boc06, Bod04, BB92, Boy98, Bra12, Bra15, BM17, Bro10, Bro24, BFIT18, BW84, Bun02, BK17, BP99, CK80, Can14, CB22, CR16b, Cla77, Cla08, CJ23, Cle20, CV94, Coc03, Coo01, Coo94, Cow05, Cra21, Cre24, Cru76, CE96, Cus05, Dac88, Dal83, DM09, DE95, DFSV23, Dod02, Dol94, Dol95, Edm20, ES05, Elv91, FD87, Far98, FS20, Flo20, Fog86, FW83, Gad21, GWM11, Gar21, Gar81b, Gra82, Gra90, GG00, GW76, Gri11, Gui97, Har04, HNW94, Her81, Her82, Hey91, HP08, HP09, Hir90, Hit06, HW17, Hop96, HR18, Hut84, Hux92, Hux03, IG21, JW97]. July [Jel99, JW00, Joy16, KL19, KH04, KO99, LH18, LB07, Leg85, LE80, LK23, Lew86, LMW06, Lov75, Lov04, Mac22, Man87, Man98, Mar88, MF76, May79, MS19, McC81, McC22, McC03, MF18, Mil97, MM96, Mor97, Nag19, New98, OKC21, Pat82, Pat05, PF99, PO21, Pea18, Ped01, PB10, Phi79, Pie99, Pig77, PC10, PC11, Pol00, Pol06, PSW16, Pre10, Qua94, Qui05, Ran09, Ree08, Rei24, Rei21, Rey07, Rit00, Rob23, Rot99, Ruf02, Sad23, SR76, Ser09, Sho85a, SK13, Son99, Sou11, Spr95, SP76, Ste20, Sto94, SSEF23, Stu13a, SB22, SB77, Tho13, TB13, TC81, Tre11, Uns14, Vin23, War17, Wat19, Wei24, Whi99, Whi83, WM84, Wig82, Wig83, WR95, Wil75, Wil13, fWT94, Woo95a, Woo95b, Woo97b, Woo05, Woo21]. July [YN06]. Juma [Cla21]. JUNE [Ano19, Abr83, Alb87, All99, All80, Alp88, AB15, Ash02, Ati76, AS03, Aut19, Azz24, Bak76, Bak77, Bal24, Bar98, BFO^{+}19, Bar86, BC07a, Bar81, Bat76, BF24, Ben19, BC92, BG02, BH08, Bla94, BS01, Bon06, BW90, BR14, BM17, BE05, Bro23, Bul94, Bun02, Bur03, Bur78, But01, CK14, CK80, Cap99, CU76, Cha03, CB22, Cla21, CKT15, CB81, CG21, Cro10, Cru00, DT81, Dat88, DWS02, Den81, DS86, DG23, Dou07, Dou19, DST85, DS21, Edw96, Ell09, EW22, Eva76, FM19, Faw18, Fed94, FH98, FJ95, Fin95, For12, GB13, Gly11, Gol17, Goo82, GI96, Gow90, Gra90, Gre96, Gre19, GS90, HL23, Han18, Har20, Har87, Hay02, HF09, Her81, Hib21, Hop23, Hor86, Hor95, How15, HT03, Irv23, Isl78]. June [JW97, Jac00, JGP06, Jef77, JW00, JP81, JJ14, Joy92, Kas82, Kat78, KH04, KTR07, KT18, KP78, Kri95, KH83, LL23, Lea91, Lee07, LK23, Lew92, Lin01, Lon21c, Lon23b, LH90, LH88, LS12, LM94, MH19, Mac03, Mak94, Man98, MR13, Mar79, Mas02, Mat88a, MH08, ME11, McC03, MB21, McK22, MP82, ML03, Mes91, Mil84, Mil18, Mor17, Mor16, Mul94, Nee02, NL22, NC21, Ost07, Ott04, Pag01, MP24, Pas87, Pat01, PB84, Pet11, PD18, PB20, PS97, PT11, Pro23, Rai08, Ran09, RS15, RP20, Ree98, Rei14, Rei21, RE04, Rob86, RSD16, RG06, Sau87, Sco14, Sew15, She82, She03, Smi05a, Smi99, Sou11, SH21b, Sta94, SBG10, SBG11, Stø76, SL88, Stu85, SL90, TC87, Tho83, Tod75, Tod79, Tyr91]. June [VL24, Vin09, VT88, Wad97, Wah23, WA94, Web10, Whe01, Whi94a, Whi96, Whi94b, Whi86, Whi04, WR95, Wil87, WB85, Woo95a, Zuc81, vC99, van19].

K.B.E [Ble01, Cru00, Hux00, Joh00]. K.C.B [BG02]. K.G [Zuc81]. Kafatos [LG20]. Kahn [Dys99]. Kaj [AH97, Tis60]. Kao [Mid21]. Kapitza [Sho85a]. Karel [SV91]. Kariamanikkam [LB67]. Karl [Mil97, Ril60, Att91, Bar60, Baw75, Ste84, Tho83]. Kármán [Gol66]. Karrer [Isl78]. Kassanis [HH88]. Katchalski [ASS16]. Katchalski-Katzir [ASS16]. Kathleen [Hod75]. Katritzky [Bou15]. Katz [Sak07]. Katzir [ASS16]. Kay [Bla77, MO23]. Kayser [Her55]. KBE [CBE09, TW22]. KCB [Rox07]. Kearton [Woo95b]. Keen [Per82]. Keffer [GR85]. Keightley [Ele76]. Keilin [Man64]. Keith [Cla55a, Col02, Men64, Mof02, Bel01, CG21, Gut96, Jef77, Lig85]. Kekwick [CVW02]. Kelk [Sho72]. Keller [Win01]. Kelly [How15]. Kemball [Rob00]. Kemmer [Dys11]. Kemp [Nor91]. Kendall [CK80, Kin09]. Kendrew [Hol01]. Kenichi [BN01]. Kennaway [Coo58]. Kennedy [Bra95, Sut80]. Kenner [Tod75, Tod79]. Kenneth [BG14, Cla61, Hum11, Neu84, Roo96, Chi64, DM68, DT24, Dou07, Hux92, JH20, Jon23b, Kas82, Lew92, Min22, Pet11, Row04, BS24, Smi16, Vel19, WA94]. Kensington [CHW74]. Kent [FD87]. Kenyon [Sta79, Tur62]. Kerensky [Hor86]. Kermack [Dav71]. Kerr [Hin58, CBE09, Sma10]. Kershaw [Kay71]. Keynes [Gly11, Hux03]. Khalatnikov [Bya22]. Kibble [DS21]. Kidd [MP75]. Kilburn [WK03]. Kimura [Cro97]. King [Fos57, Har56, Sho63, Whi03, Whe74]. Kingdon [Haw56]. Kingsley [Joh03]. Kistiakowsky [Dai85]. Kitching [Sle97]. Kiyoshi [NL22]. Klaus [CV17]. Kleczkowski [Baw71]. Klug [Cro20]. Kluyver [Woo57]. Knight [Har04]. Knott [BB18]. Knowles [Bar75b, HW10]. Knox [Goo82, Tho60b]. Knut [Vog08]. Kodicek [FW83]. Kohn [Col18, HL18]. Koiter [vC99]. Kolmogorov [Ken91a]. Konrad [KS92, WL02]. Kornberg [Leh12]. Kosterlitz [NH13]. Krebs [KW84]. Krishnan [LB67]. Kronberger [Rot72]. Kroto [Leg17]. Kruskal [GCJM18]. KT [FP04, KH04, Wel05, HH95, Woo95b]. Kücherman [OM80].

Kuen [Mid21]. Kuffler [Kat82]. Kuhn [Ble96b]. Kumar [Cow21, Rat64]. Kurt [CV94, Kre80, Sho83, Thi84]. Kurti [San00]. Kuypers [PG92].

L [Cra21, Hux97]. Lack [Tho74]. Lade [Lov64]. Laing [KH04]. Lainson [Sha21]. Laker [SL94]. Lal [GM21]. Lamb [CC10, Dav18]. Lambert [Sto94, Tho74]. Lancelot [Wel78]. Land [Cam94, CNO23]. Landau [KL69]. Lander [TW21]. Lane [Cle20]. Lang [Aut19, Sal61, Tis60, Whi66a]. Langley [SP76]. Langmuir [Tay58]. Langstreth [Smi16]. Langton [LH18]. Lansdowne [Pow87]. Lappert [LN16]. Lars [Rob23, LHF78]. Lashley [Bar60]. Latimer [Mon86]. Laue [Ewa60]. Laughton [Sea20]. Laura [CJ23]. Laurence [Ble01, Coo58, Gor96, Hil88, Rae12, Dee67]. Laver [Web10]. Lawrence [Har61, Phi79, WBF23, Goo20]. Lawrenson [Tur23]. Laws [CBPC15]. Layton [Sco56]. Lazarevich [Lon11b]. Leach [FSC20]. Leader [All70]. Leathes [Pet58]. Leblond [Ben08]. Lecomber [Spe94]. Lecorney [ES05]. Lederberg [BG11]. Ledwith [Fea18]. Ledyard [BS05]. Lee [Cha55, Kay68]. Lees [Ark55, Sau94]. Leeuw [Coo01, Har04]. Lefschetz [Hod73]. Lehmann [Bol97, Dac88]. Leicester [Pat05]. Leigh [CA56]. Leiper [Gar70b]. Leloir [Och90]. Lemberg [RG76]. Lemieux [Bun02]. Lennard [Mot55]. Lennard-Jones [Mot55]. Lennox [Uns14]. Lenox [Bul57]. Lenox-Conyngham [Bul57]. Leo [GSL19, LE80]. Leon [SW21b]. Leonard [Bel86, Bow58, Jon66, Mat88a, Pic74, SS06, All72, Boy63, Cox57, Dun82, Har87, Lew86, Oak71, PO21, QG03, Tem65, Til61]. Leonidovich [Sho85a]. Leopold [PJ80]. Leray [And06]. Leslie [Ano98, AS03, Eva76, FAL21, Lyn15, SL90, CMO87, DJ13, Hil72, Kur83, Mor16, Pat01, Rob23, RWN66, Stu09, Whi94a, Whi94b]. Lester [CP94]. Lev [KL69]. Levi [Til57]. Lewis [Baw58, Bra12, CJ12, CR16b, Hau56, LH88, MIH24, Sta58, CA56, Eyr13, TS24]. Liddell [Phi83]. Lifshitz [ZK90]. Liggins [GB13]. Lighthill [Ped01]. Lighton [Flo08]. Lillian [MR19]. Lilly [Cla00]. Lindemann [Tho58]. Linderstrøm [Tis60]. Linderstrøm-Lang [Tis60]. Lindor [MP74]. Linnane [Nag19]. Linnett [Buc77]. Linstead [BRE68]. Linus [Dun96]. Lionel [Rob60, Fei05, Har73, Oxl23, Pol00]. Lions [Cia02]. Lipmann [JW00]. Lipson [Woo94]. Lisa [Hun17]. Lise [Fri70]. Lissmann [Ale96]. Littlewood [Bur78]. Llewellyn [Bow67, MP23, PS97]. Lloyd [App07, BC07a, Hux00, Kin16]. Lockspeiser [Edw94]. Loewi [Dal62]. London [Ano71a, Sho71]. Longman [Gre06]. Longuet [GM06, SH18]. Longuet-Higgins [GM06, SH18]. Lonsdale [Hod75]. Lord [FR98, FP04, Joh00, Lew02, Gow90, Jel99, Ram22]. Lorenz [KS92, Pal09]. Louis [Bro09, Fow02, Gre01, PB20, Abr88, Cas73, Cia02, Coo98, Fos57, FA03, LS66, Stu86]. Louisa [vEG06]. Louise [BB22, Cut05]. Loutit [LM94]. Lovatt [DG70]. Lovell [DGL16]. Low [MF18]. Lowe [Sti05]. Lowenstein [Ale01]. Luca [Edw21]. Lucien [Lwo77]. Lucy [Hay00]. Luddenham [FP04]. Ludwig [Hay82, BT20, Bus60, Tak22, Whi83]. Luigi [Edw21]. Luis [Och90]. Luitzen [KN69]. Lund [TH16]. Lwoff [JG98]. Lyall [Bur01]. Lydwell [Rat60]. Lyman [Ost07]. Lynden [Eva20]. Lynden-Bell [Eva20].

Lynen [KD82]. Lyon [FR24, WA94]. Lythgoe [Jon21]. Lyttleton [BH97].

M [Akh96, Cra23, Dul95, FH98, Kro95, LKK97, Mak94, Ree94]. M.B.E [CCG^{+}00, Hew02]. Maass [Flo63]. Macalister [Rus78]. Macaulay [Adr63]. MacDiarmid [HKB23]. MacDonald [Men64, RW06, Cla77, Cre24, Haw60, Hol96, KH04]. MacFarland [BT68]. MacFarlane [BW90, GMP24, Fen87]. Macgregor [Bod04]. MacIntosh [Bur94, PR22]. MacIntyre [Mar12]. Mackaness [Car14]. MacKay [LC17]. Mackenzie [Leg85]. Mackintosh [Ble97, Pea18]. MacLeish [Sau87]. MacLennan [Rei21]. Macmillan [Hog87, SW21a]. Macnaghten [Hay85]. Maddock [Duc91]. Maddox [Gra10]. Maddrell [SOD22]. Mahalanobis [Rao73]. Maheshwari [Ste67]. Mahler [CV94]. Maitland [Bed57, CBPC15]. Maitlis [Cra23]. Maizels [Mac77]. Majesty [Cor90, Lon22a]. Makins [Lor98]. Malcolm [BB99a, Irv23, Vin98, Cla00, Cra21, Her14, Pea95, Per88, Pud17]. Maldon [Che62]. Maling [Ree98]. Mambillikalathil [Cow21]. Mandelstam [Yud10]. Manley [Kas82]. Mann [Jon23a, Mil84]. Manne [Att91]. Mannering [MP62]. Mansfield [Cal21, Mor21]. Manton [Fry80, Pre90]. Mapson [Hil72]. Marcel [Bas63]. March [Abr88, AS21, AA98, Ale22, All13, AGJ^{+}22, Arn91b, AS03, BB99b, Bar04, Bar75b, Bat76, BH18, Bea95, Ben94, Bla85, Ble82, Ble96b, Ble99, BF91, BW90, Bor17, Bow05, Boy98, BE24, Bro10, Bro19, Buc06, BW84, Cam94, CEG^{+}19, Car14, CV85, Caw24, Cer19, Cha85b, Chi99, CC10, CCG^{+}00, CST17, Con09, Coo90, CM81, Cox95, CG21, Cra20, CT14, CL14, DT24, DS79, DGS18, DAB97, Dol95, Duc91, Edw94, Edw96, EE95, Eva76, FD87, FM19, Fer17, FP95, FR98, Flo08, For99, For12, Gab02, Gal03, Gar21, Gil20, Gin94, Glu77, Goo20, GBC85, Gor24, Gra04, GR85, Gra82, GLR07, Gre00a, Gre06, GCK17, GM06, GS90, GW76, Gri95, GW07, GP83, GR00b, Gur12, Hal99, Har82, Har20, MDHW03, HNW94, Hay82, Hay85]. March [HD98, HG11, Hey87, Hig24, Hil90b, HT16, Hir15, HH88, Hit06, HW04, HL18, Hol01, HW77, Irv23, Joh00, Jon83, Jon91, Jon03, Kel17, KC23, Ken82, Kin16, Kin00, KW79, KH90, Kur83, Kus01, LR19, LY24, LB07, Leh12, LN16, LE80, Lew83, Lew92, LM02, LP91, Lov04, LMA80, LS85, MH19, Mac00, Mak94, ML85, ML19, MB21, McL00, Mil84, Mil20, Min22, MS08, Mit91, Mof02, Mon06, MA21, Mor05, Mor92, NN00, Nai21, GN24, Neu90, NA05, Ogd20, OKC21, Oli94, OG14, Ost07, PB79, Pat97, PK24, Pep03, PB84, Phi79, Phi83, Phi13, Pie99, Pip05, Pir90, QG03, Ran90, RR05, Ray76, Ree94, Rey07, RH18, Rob00, RW06, RE08, Row96, Rus24, Sak07, San21, SS06, SH24, Smi98, Smi05b, Smi15, Smi16, SC99]. March [Sta86, Sto94, Sto03, Stu77, SL90, SB22, Sut80, SDV19, Tak22, TH16, Tat90, Tat94, TW22, TP75, Tho98, TW21, TG23, Tur23, Ubb81, VL24, Vin98, VT88, Wai97, WF19, WW86, Whi90, Whi83, Wid95, WH77, Wil87, Wil75, Wil23a, WZ79, Woo94, YH82]. Marcus [Ble01, McL00]. Margaret [SL21, Aga23, Mac12]. Margerie [Fle55b]. Maria [Owe95, PG92]. Marie [Fle55b, KT20]. Marion [LE80, MT21]. Mark [Ble01, WD21, TJ24]. Markham [Els82, MR22]. Markl [MM21]. Markovich [Bya22]. Marmaduke [All00]. Marmion [EB08]. Marquess [Tod73]. Marrian [Cha61, Gra82]. Marsden [Fle71, QRJ12]. Marshall [BM98, BM13, FR98, Rus78, Tho60b]. Marsham [Rob92]. Marston [Abr71, Syn67]. Marthe [Cut05]. Martin [Ark55, Chi56, Fle55b, GP83, KC22, Lov04, MT78, Phi74, Pow87, RR24, RK10, TO12, Wil23b, AS02, GCJM18, GS86, Mes99, MF18, Sav63, Wat19, Wil23a]. Martyn [Mas71, Atk04]. Mary [Dod02, Har18, Hay00, HCB18, Mac12, MR19, Tei03, BF91, FR24]. Marylebone [Lew02, Ram22]. Maskell [BCB61]. Mason [BC60, Bro16, Gri11]. Massey [BBD84, WB03]. Masson [HL63]. Mather [Lew92]. Mathison [New55]. Matthews [AS03, Gra90, Har87, Har99, Kib88, Whi99]. Mattinson [Ken82]. Matveevich [CV85]. Maurice [Bul75, CK14, DP86, Mor92, SB22, Ubb81, Woo97a, AKS06, Bul67, ES05, Hog87, Kei60, Ove97, Sch64, Whi04, Wil61]. Mawson [AT60]. Max [Blo04, Ewa60, Hay82, KS71, Pic74, RG76]. Maxwell [Cha85a, MS06, Rai08, Ada85]. May [AM13, Akh00, All84, AB15, Ano98, App07, AL23, ASS16, Ash22, Ash02, Ati14, BC07a, BBD84, Bea95, Ber05, BN06, Bey96, BT05, Bla94, BG95, BS90, Blo04, BG11, Bol97, BH97, Bon06, Bra95, BM05, BR14, BSD13, BLT11, BLT12, Buc10, Bul75, BI96, Bur01, BC07b, BS13, BT20, But01, Cal21, Cam94, CGZ22, Cha85a, Chi99, Cia02, Cla00, CW11, Cra23, Cra24, Cru76, Cru04, CA01, Cus05, Dat88, DSD20, DG16, DL04, DT20, Dod02, Dom21, Dou19, Dun83b, Dys99, Eab90, EL96, EFP17, Edw20, Fei95, FJ95, Fin95, FR24, Fle99, Fow02, FE12, Fry80, GB14, GWM11, GT23, GMP24, GR16, GV92, Gow90, Gra20, GN19, GW07, Gun00, Gut21, Har04, KHG21, Haw76, Her14, Hew02, Hil87, Hil11, Hod80, HCB18, Hyn15]. May [Ive06, JG98, Jef76, JC19, JG82, Jon91, Jon03, JJ14, Jon23b, JS10, KM24, Ken91b, KTR07, KC22, Kin22, KM15, Kro95, LH18, Lea91, LH84, Lig98, Lig85, Lon22b, LP91, Lov11, Lwo77, MJR24, Mar86, MW01, McC08, Med80, Meh02, MT08, Mes99, MR96, MB24, MP75, MS06, Moo77, MA21, Mor16, Mui94, MT19, Nee02, NJ86, NC21, Owe09, Oxl23, Pal09, PSS78, Pat82, YHP^{+}86, Pea95, Pea98, Pip94, Pol00, PS97, Pre90, Pud24, Qui01, Rae12, RC24, Ram22, RP96, RR05, Rea19, Row04, Rus77, Sad23, San00, SR17, Saw92, Sca18, Ser99, Sha15, Sha21, She03, Sil03, SO20, Sim18, SW21a, Sin03, CBE09, Sma10, Smi82, SC99, Sta86, AS24, Ste20, Sti05, SBG10, SBG11, Str96, Stu09, Stu75, SF91, SP79, Swa96]. May [Tho98, Tho19, TW21, TC18, Vau87, Wah23, Wei94, Wes88, Wes95, Whi96, Whi79, Whi03, Wil94a, Wil13, Wil94b, Wol01, fWT94, Wri24, Yat06, You12]. Mayer [Ree94]. Maynard [CH05, Lov75]. Mayneord [Spi91]. Mayo [KO99]. Mayr [Boc06]. MBE [Bro10, HT16, HW17, Ott04, PT11, SA15]. McCance [PB84, Wid95]. McCarthy [Bra70]. McClintock [Fed94]. McCollum [CP69]. McCrea [MP07]. McCullagh [Baw58]. McCulloch [Mak18]. McFadyen [McC65]. Mcgarel [Lew02]. McGee [Mor88].

McGillivray [Gai19]. McGregor [Wil23c]. McIvor [She82]. McLachlan [SSEF23]. McLaren [CJ23, HS07]. McLean [AC72]. McLeod [Bos12, Has16, WZ79]. McMichael [Dol95]. McNeill [Ale22, McC65]. Meade [Hai24]. Mearns [Kay72]. Medawar [Mit90]. Mees [Cla61]. Meghnad [Kot60]. Meinesz [BS67]. Meirion [PR78]. Meitner [Fri70]. Melbourne [Men70]. Mellanby [Dal55]. Mellor [Lor98]. Melson [Wun97]. Melvill [HM77]. Melville [BG02]. Melvin [SB08a]. Memoirs [Ano71a, Lin56, Lon20a, Lon20b, Lon21b, McR76, Wil95]. Mendel [Fel60]. Mendelssohn [Sho83]. Menon [Cow21]. Menter [HP08, HP09]. Menzies [Whi79]. Mercer [For00]. Meredith [Mit91, PW74, Stu75]. Merrill [CEP15]. Merrison [Gab02]. Merton [HG70, RG06]. Mervyn [HF09]. Mestel [SW21b]. Metcalf [Nic16]. Michael [BW22, GW07, Hit20, HW04, Hop23, LK23, MP23, Mor05, Ped01, Pet24, Red56, Wei24, AB15, Ast01, Bey96, BS90, Bro24, Bry01, BFIT18, BS13, CNO23, FS20, JGP06, LN16, Med80, Mit91, Mur98, Pau18, PM94, Pic16, Ric15, SH18, Sta94, WH77, WD21]. Michel [JG98]. Michell [Che62]. Mieczyslaw [DE95]. Mikhailovich [ZK90]. Milana [Fry80]. Miledi [PSCV21]. Miles [Bai94, Neu90]. Miller [BS08, HW77, Ruf02, Sta69]. Millington [Gor96]. Millner [DL04]. Mills [Ble99, FGW13, Man60]. Milne [Pea95]. Milner [AC60, Cle20]. Milstein [NA05]. Miriam [vEG06]. Mitchell [Bar11, Cow04, Mar88, Sla94, Stu77, WW99]. Mitchison [MB24]. Mitra [Mah24, Rat64]. Mollison [Cra20]. Monod [Lwo77]. Mont [GB13]. Montague [Pet92, Ken82, Mac77]. Monteagle [Cul90]. Monteath [Arn94]. Monteith [Uns14]. Moon [BI96]. Moorbath [DT20]. Moore [ML19, Stu13a]. Moran [Hey91]. Mordell [Cas73]. Morgan [Bro80, Wal66, WB05]. Morien [Bro80]. Moritz [BS14]. Morley [Mor16, ML85]. Morris [Coo60, Hol66, QT86, Baw63]. Morrogh [Eva05]. Morse [Woo90]. Mortimer [BLT11, BLT12, Pig77]. Morton [Bea82, GPPG78]. Moses [Pas87]. Mothes [Thi84]. Motoo [Cro97]. Mott [Pip98]. Mountbatten [Zuc81]. Mourant [MBW99]. Moyle [Tei03]. MRIA [GR16]. Muggleton [Bai88]. Muhammad [Kib98]. Muir [BM13, Cam60, Har18]. Mülertz [Cha58]. Muller [Pon68]. Mulliken [LH90]. Munk [GW20]. Munro [Smi68, SF91]. Murdoch [FM19, Ran09]. Muriel [BM74]. Murison [Han18]. Murray [BG14, GB14, PS97]. Murrell [CST17].

Nabarro [Bro10]. Nagai [NL22]. Naismith [Cha02]. Naomi [Bar23]. Napier [BB22]. Narayana [VJ05]. Nassau [Har04]. Nath [Meh75]. Nathan [Smi55]. Nathaniel [Ree94]. Nederlandse [Coo01, Har04]. Needham [GR00b, HW04, Tei03]. Néel [FA03]. Neil [Tho19, Boy66, CHW74]. Neill [JWW^{+}20]. Neish [Jon74]. Nelder [Sen19]. Nelson [BR18, Fei95, Rob92]. Nesmeyanov [CR83]. Ness [MDHW03]. Netherton [Sta79]. Neuberger [AM01]. Neumann [Pra10]. Nevill [Pip98]. Neville [Bul67, Cot72, Cow05, HJ11, Lea91, MT19, PR81, Whi66b].

Newbold [Wel05]. Newitt [Ubb81]. Newland [Ash22]. Newman [Ada85, Joy16]. Newnham [CR16b]. Newsom [Vin19]. Newsom-Davis [Vin19]. Newton [Dar72, Sch67]. Neyman [KBP82]. Nicholas [Sal57, SLO10, App07, CKT15, Dys11, MB24, Pea18, San00]. Nicholson [Wil56]. Nicol [Mad06]. Niels [AH97, Ban83, Coc63]. Nielsen [Vog08]. Nigel [Kek81, Pro23]. Nikolaas [Hin90]. Nikolaevich [CR83, Dai90, Ken91a]. Nikolai [Dai90]. Nils [Sko61]. Nishizuka [YN06]. Nockolds [Mui94]. Noel [Hil58]. Noreen [GB14]. Nørlund [Ban83]. Norman [CST17, Neu73, Tho73a, Wad97, Whi94b, Ano98, BM98, CD81, Hil88, Hon95, JWW^{+}20, Kin00, LL05, Phi13, Pie99, Scr99, Syk73, Til57, TC18]. Norris [Str57a]. Norrish [DT81]. Northcote [Cro19]. Northcott [RS07a]. Northumberland [Hen90]. Norton [Pal09]. Nosherwan [Stu70]. November [Adr76, AS21, All94, And06, AA02, Ant15, Atk04, Azz24, BS08, Bar23, BBD84, Baw75, Bel81, Bey96, BT05, BB97, BG95, Bla77, Bod04, BBM13, BE24, Bre14, Bre05, Bro77, Bro15, Buc10, Buc77, BC07b, CK14, Car99, Cha01, Cha02, Cha85b, CV17, Cla80, Cla85, CD81, Col18, Col02, Coo01, CG20, Cow21, CJ01, Cra20, CSB22, CVW02, CEW18, Cro97, Cro20, Cul90, Cul95, CP94, DT81, Dai85, Dal97, DSD20, DS86, Dor18, DGV11, Dow94, DH04, Dun78, Dun83a, Eab90, EL96, Edw18, Ell09, EP10, Ell22, Els82, Eva99, FGOS15, Fea04, Fie08, For00, FM21, FA03, Gal03, Gar88, GMP24, GLD18, Gra10, GI00, Gre98, Gre00a, Gre06, GCK17, Gre12, Gre13, Gre14, GM97, Han18, Har18, Har91, Har87, Har99, HM03, HG11]. November [Hew99, Hil88, HF07, HW87, HU19, Hod79, Hol94, HJ11, Hul94, HP01, Hut84, Jam90, Joh23, JWW^{+}20, JG13, Jon23a, Kas82, KL19, Kek81, Key90, Kib88, Kib98, KE04, Kin00, KW84, KS92, LPP19, Law08, Law01, Lew86, LJ97, Lon11b, LHF78, LG20, Lov75, LS12, Lyo22, Mac12, MW20, Mat90a, May15, Mid21, Mon06, Mot82, Mul94, Nag19, NG96, Nor91, Pap19, Pay81, PO21, PR22, Pet11, PSW16, Pud17, Pud21, RSG21, Ram22, Ran75, Ric15, Rob92, Rob77, RW18, Rob95, Roo96, Rox07, RK10, SC23, San00, SR17, Saw92, SV91, SVW00, Lor98, SG79, Sid02, Sim18, SA15, SL94, Sou87, Sow24, ŠR24, Ste86, Stu86, ST19, Sut80, TH16, Ter10, Thi84, Tho91, Tho83, Tod79, Wat06, WW99, Wea02, WP75]. November [Wig83, WB03, Wil23b, Wil22, Wol98, Woo90, Woo97b, YN06, YL92, Yud10]. Nuffield [Hol66]. Numa [Imu95]. Nunes [Bro10]. Nutman [Ber05]. Nye [Ber20, GTK11]. Nyholm [Cra72].

O [Coo98, Far98, FH98, KH97, Kro95, Mak94, Mes98, PS97, Sle97, Smi98]. O.B.E [Bry01, Cha02, Chi99, For00, Gre00b, NN00, Pol00, Woo95b]. O.C [How01]. O.M [BK00, Dod02, Hux00, Zuc81]. O.N.Z [Har99]. OA [Rae12]. Oakley [Eva76]. Oatley [Smi98]. OBE [DGL16, Dav18, HJ11, KH04, PB10, Stu09, Vin09]. O’Brien [LW20]. OC [RS15]. Occhialini [Big02]. Ochoa [GM97]. OCTOBER [Ano19, Aga23, Akh96, Ale01, ACW20, AKS06, Ash02, AH97, Ast01, Ati14, Ban83, BSHC18, Bea82, Ber05, BH08, BB97, Ble82, Ble01, Bos12, Bra92, Bre05, BSD13, BK00, Bro19, BM22, BN01, Buc06, Bud88, BI96, BB18, Bur19, BP99, Bur21, Bya22, Cal21, Can14, CF76, Cha97, CV17, Cla77, Cla86, CKT15, Coo94, Cot84, Cow04, Cow05, Cox95, Cox03, CSB22, Cre24, CBPC15, DP86, DM09, Den86, DT20, Dic75, Dow94, Dun82, DJ13, Dys11, Edw94, Fea04, Fer17, Fie08, Fla03, For95, For99, FAL21, Fox87, Gad21, GS18, Gar95, GW20, GVS19, Glo20, Glu77, Gor96, GS86, GR00a, Gre12, Gre13, Gre00b, Gre01, Gri95, GSL19, GS13, Hai24, HM77, HM03, Hen90, Her14, HH95, HHP22, Hib21, Hil97, HH88, Hit14, Hol94]. October [Hol96, Hor95, Hun17, Hux97, IG21, JH20, JC19, Joh77, Jon79, Jon83, Jon10, Kat82, KH97, Ken91a, Kev05, Kin09, Lai20, Lan85, Leg17, Leh12, LH84, Lew02, LW20, Lis78, Lon11b, LHF78, LH90, Lyn15, Lyo01, MS90, Man03, Man87, Mar79, MS96, MF76, Mat88a, MP23, MW01, MH08, May15, MS19, MS98, McK22, Mes98, Mil18, Min22, Mit90, Mit00, Mon86, MP75, Moo15, Mor97, Mor21, Mui23, MT19, Mur98, NA05, NH13, Ove97, Owe95, Pap19, Pat05, Pat90, Pei81, Per87, Pet24, PB10, Phi95, Pic16, Por96, Pra10, PT11, Pud17, Pug78, PR81, Qua94, Qui01, RSG21, RH22, Ran86, RP96, RRS20, Ray76, Ree94, RH18, RE04, RG76, Rob23, Rob92, Rob76, RR24, RSD16, RE08, Roo96, RG06, Row96, Rus24]. October [Sau94, Scr99, Sen19, Sho85b, Sil03, Sle97, Smi15, SP76, SQ88, SDV19, Syn76, Tat90, Tay96, TW22, Tho13, TB13, TS24, TG23, TO12, Tur23, VJ05, Vin98, Vin19, Wai97, WPL21, WB05, Wat06, WW99, Wat19, Web17, Wel00, WL02, Whi94a, Whi86, Whi01, Wig82, WB85, Wol01, WBF23, ZK90]. Offord [Hay02]. Ogilvy [Dav71]. Ogston [Smi99]. O’Hara [Ric15]. Øjvind [Wes64]. Olaf [Stø76]. Olding [MM96]. Oleg [Hor86]. Olek [Owe09]. Olgierd [Owe09]. Oliphant [Ble01]. Olive [CG20]. Oliver [GW76, Sad23, Tho56, Bal24, Gre01, PSS78, Stu75]. Ollis [BS01]. Olof [Pea98]. OM [Bar04, Blo04, BM17, Bro15, Cro20, FP04, KHG21, Hit20, MB21, Sim18]. Onsager [LHF78]. Oort [Hul94]. Oppenheimer [Bet68]. Orange [Har04]. Orde [Coo01]. Order [Har04]. Orgel [DJ13]. Orlebar [PD18]. Orme [HL63]. Orowan [NA95]. Orr [Kay72]. Orsenhelm [Kin56]. Oscar [Pro23, Man98, Pan57]. Osmund [Riv20]. Osvald [Pat90]. Oswald [Wad97, Dub56]. Ottewill [Vin09]. Otto [Bla94, Eva99, Jac00, Kin56, Ale01, Cow64, Dal62, Dar61, Flo63, Kre72, Pei81, Spe70]. Owain [Sou87]. Owen [BM24, CH79, For95, YL92, Pug67, TO12, Wil60]. Oxford [KHG21].

Packer [Jon23b]. Page [Yat06]. Pagel [Edm20]. Paget [Moo77]. Paige [AR09]. Paintal [Igg06]. Pal [Swa96]. Palmer [Kin22, GN24]. Panchanan [Ste67]. Paneth [Eme60]. Pantin [Rus68]. Paolo [Big02]. Parke [BF91, Wei24]. Parker [HR18]. Parkes [Pol06]. Parrington [Cha90]. Parry [Ree08]. Parsons [DE58, EE95]. Partridge [Bai94]. Pascal [Sch64]. Pashley [JS10]. Pasquill [MS96]. Pateman [Sca18]. Paton [Bur01, HT16, RP96]. Patricia [Pep03, Bra15]. Patrick [BRE68, LR19, ML19, SS06, Wal08, Cra20, Edw96, Fit22, Hey91, Lov75, RG06, Wig56]. Patterson [For99]. Paul [FE12, Kel17, SVW00, Ser09, Wei90, Ash16, BB99b, BS14, BH88b, DP86, Gia16, GKW73, Gra20, Isl78, Kib88, KO99, LW20, Mül22, Sow24, YL92]. Pauli [Pei60]. Pauling [Dun96]. Pawsey [Lov64]. PC [Sto03]. PCNZM [Kel17]. Pearsall [Cla71]. Pearson [Bar81]. Peart [Gor24]. Peary [Swa96]. Pease [Wat06]. Peat [HT70]. Pehr [PB79]. Peierls [Lee07]. Pelham [Smi15]. Penelope [RR05]. Penfield [EF78]. Penley [JJ14]. Penman [Mon86]. Pennell [Ken82]. Penney [Mak94]. Pennycuick [SH21b]. Penrose [Har73]. Percival [BG95, Dal57, SP79]. Percy [CJ01, DSD20, Hen90, Mor66, Syk73, FD87, Gar81a, LKK97, Mit61, New62]. Pereira [McC08, ST99]. Perham [BR18]. Perkins [Sco56]. Perry [KH04, Tre11]. Perutz [Blo04]. Petch [Hon95]. Peter [Ati14, Baw71, BH88b, FD87, MR22, Mor21, Wat55, van19, Cra23, Dav70, Den86, Den81, GTK11, HJ11, JG13, KH90, Lau23, Med61, Mit90, NS24, Ogd20, Pat97, PD18, RW18, Sco14, Sla94, SH21a, Spe94, Tho19, Tur23, Wol98]. Peters [TO83]. Petrovitch [Wig71]. Petrus [KS69]. Pettersson [Dea66]. Pettifor [SDV19]. Peyton [And71]. Pfeiffer [Fil56]. Pfeil [All72]. Philip [AD21, Edw18, Fog86, Gui97, IG21, Lon21c, PF21, Smi05a, Tab69, Vin23, BI96, Cla77, CT14, Cur84, GRT84, Hal99, Hib21, Hir90, Sil03, TB13]. Philippe [Ben08]. Phillip [Ber05, McK22]. Phillips [Dow70, Joh00, Meh02, Por96, TO12, VT88, Wat19]. Pickavance [Hol94]. Pickering [CK80, MP82]. Pickett [BF24]. Pickett-Heaps [BF24]. Pickford [MR19]. Pierce [Hey91]. Pierre [Abr88, Fle55b, JC19, SL88]. Pierre-Gilles [JC19]. Pilkington [Oxl23]. Pillinger [Wri24]. Piotr [Sho85a]. Piper [SOD22]. Pippard [LW09b, Ske70]. Pirie [Pie99]. Pitt [BH08, Tat94]. Pitt-Rivers [Tat94]. Plaskett [McC81]. Platt [KR11]. Plowright [Big10]. Polani [Gia16]. Polányi [WH77]. Polge [Hun08]. Polkinghorne [TW22]. Pollock [AS02]. Ponsonby [Bul57]. Pontecorvo [Sid02]. Popjàk [Akh00]. Pople [Buc06]. Popper [Mil97]. Porter [FP04, Lov04, Nor91, Per87, Woo21]. Portevin [Bas63]. Postgate [RSD16]. Powell [BFIT18, FP71, Gui97, McL00, Ray76]. Pp [Ano71a, McR76]. Prandtl [Bus60]. Prasantha [Rao73]. Prelog [ADE00]. Prescott [SQ88]. Preston [Cus05]. Price [DAB97]. Priestley [GWM11]. Prince [Lon21c]. Pringle [Wig83]. Prokofievitch [MY73]. Prosad [Mah24]. Proudman [CU76]. Pryce [ES05]. Pugh [Wil75]. Pugsley [Chi99]. Pumphrey [Pri68]. Punnett [Cre67]. Purcell [Ble99]. Purslove [BFO^{+}19]. Pye [Sau61].

Quarrier [Sut80]. Quastel [MS90]. Quayle [Ant15]. Queen [Lon22a]. Quentin [Mac22, OG14]. Quintin [Lew02].

R [Rei24]. R.D.I [Whe01]. Race [Cla85]. Rado [Rog91]. Ragnar [Gri95]. Raimund [Mil97]. Rainer [She18]. Rainey [Joy92]. Raisman [Faw18]. Raistrick [Bir72]. Rajendra [BR79]. Ralph [Col18, Fla03, HG70, HNW94, Haw76, Syn67, CVW02, Cro00, Fow02, Ken91b, Sha21]. Ramachandran [VJ05]. Ramage [Mui23]. Ramalingaswami [MT08]. Raman [Bha71]. Ramsay [DM09, DG23, Mad90]. Randal [WB85]. Randall [HW10, Sau61, Wil87]. Randell [Cap99]. Randle [AD21]. Rankin [Wel00]. Rankine [Tho56]. Raoul [Ati07]. Raphael [Cro00]. Ratcliffe [Bud88]. Raven [Cha85b]. Rawcliffe [CB81]. Ray [Jon03, KM24]. Raymond [Abr88, Atk04, BH08, RSD16, SR76, She03, AB88, BH97, Bun02, CE96, Goo16, Hop23, RRS20, Rea19]. Rayner [Dar58]. Raynor [Cot84]. Read [GVS19, Hir63, Sut70, Tur55a]. Reason [Whi90]. Redcliffe [Smi55]. Rede [GN19]. Redman [GW76]. Rees [Bla60a, Jon10, LL23, Moo15, Mor97, Sha15]. Reginald [All84, Bro10, Con09, RWN66, BRE68, Bra65, Coo99, Cre67, Cus05, Joy92, Rob64, Rus61, Saw92]. Reichstein [Rot99]. Relf [Col71]. Rendel [Wat06]. Rene [Wei90, Bas63]. Renner [Dar61]. Renton [Dal61]. Review [Ano71a, Lin56, McR76, Wil95]. Rex [Bry01, Cha90, Gad21, LS12, Irv23]. Reynolds [Riv20]. Ricardo [Haw76, PSCV21]. Richard [Aca08, Aut19, Bri68, Hol66, Kev05, LM02, Mar64, MH08, PM88, Ray76, Wad97, AS16, Ash22, BR18, Chr72, CBPC15, Fea71, Fil56, GT23, GPPG78, Gly11, Gor96, Har99, Hux97, LH18, Lon23b, McC88, Meh02, NS82, Pri94, Pri68, Ree98, Rog91, San21, Sch55, Smi58, WW86, Whi90]. Richards [Gad21, Por66, Sch67, Sou87]. Richardson [CM81, Jef85, Lau23, Tur55a, Wil60]. Richey [KM69]. Rickard [Dee67, SG79]. Ridder [Coo01]. Rideal [Ele76]. Ridley [App07, Sal57]. Ridsdale [Pro64]. Riet [McC88]. Riley [CDS12, Fla03]. Rimington [NG96]. Ringrose [Poo60]. Ringwood [Gre98]. Rishbeth [Woo95a]. Ritchie [Ran09]. Rivers [AS02, Tat94]. Rivett [Mar66]. Robert [AB15, Bar04, Bet68, BM05, Cla70, Dal97, Gar15, KHG21, HPR72, HF06, HF08, Joh23, Jon23a, Jor04, Mon06, Mui94, Pei81, Per87, CBE09, Sma10, Smi05a, Sta69, Wai97, Wol01, WBF23, YL92, Ale22, BSHC18, Ben94, BW90, Cam60, CS98, Cla85, Coc03, CH58, DL03, DL04, Die56, Gar70b, Gar06, Glu77, HT16, Jef76, Jon91, Kus01, LW09a, LH90, Mui23, MT19, Neu84, Pan65, Pig77, Pug61, Rat75, Rob95, RE08, Sco56, Ser88, Smi82, Stu66, Tat90, Tod73, TC76, TC81, Whi79, Wid95]. Roberts [Eyr13, Hil11, Ker79, PK24, Pol92, Sow24]. Robertson [Arn94, BM74, MDHW03, HW71, Hil90a, Pug78]. Robertus [BK00]. Robin [LW09a, Tho73b, DM20, HNW94]. Robinson [And57, CW11, Sho63, TC76]. Robson [HJ11]. Rochester [Wol03]. Rockefeller [Cor60]. Roderic [Doc98]. Roderick [Bis87, RP20, GW76]. Rodney [Ant15, GLD18, Per87, Sew15]. Roger [Lai20, LS85, Tho24, BS08, Bir17, Cla86, Ell08, HW04, Hua18, Mes98, Lor98, Vin23, Von97]. Rogers [Boy63, FGOS15, Oak73b]. Rogosinski [Hay65]. Roman [DE95]. Romer [WP75]. Ronald [LR17, Bro19, CJ01, Cra72, DT81, Dav23, Joy16, Ros74, Sal66, SSGP23, Tho91, Vin09, Wil22, YM63]. Room [HW87]. Roper [And57]. Rosa [RR05]. Rosalind [Tat94]. Rose [Joh77, SL90]. Rosenbrock [Glo20]. Rosenhead [Stu86]. Rosenheim [Pic74]. Roslington [AC60]. Ross [Abe61]. Rotblat [HF07]. Roth [CV17, RK10]. Rotherham [QG03]. Rothery [Ray69]. Rothschild [Ree94, vEG06]. Roughton [Gib73]. Roundell [GN24]. Rous [And71]. Rouse [Phi74]. Rowland [Har20]. Rowledge [Rub58]. Rowson [Pol00]. Roy [Ble97, Bou15, Her14, Oak68, Wil94b, Els82]. Royal [Ano71a, Lin56, Lon21c, McR76, Wil95]. Royden [Gra10]. Royle [Riv20]. Royston [LW09a]. Rudolf [RG76, Lee07]. Rudolph [Flo20, Jon23a, TO83]. Ruggles [Rob64]. Runcorn [Col02]. Rundle [Woo71]. Rupert [Bre05, Kay68]. Rushbrooke [Dom98]. Rushton [Bar86]. Rushworth [Wal08]. Russel [Cla85]. Russell [DS86, Kre73, Pic68, Str96, Str57a, Tho66]. Ruth [HT03]. Rutherford [MDHW03]. Rutter [MP23]. Ruzicka [PJ80]. Ryde [Whi62]. Rylands [Bar73]. Ryle [GS86].

Sabine [Cla70]. Sadler [Ber05]. Saffman [CT14]. Saha [Kot60]. Salam [Kib98]. Salaman [Smi55]. Salimuzzaman [Akh96]. Salisbury [Cla80, Tod73]. Salome [Pap19]. Salpeter [Ter10]. Salt [Bat03]. Saltcoats [Hol71]. Sam [War17]. Samoilovitch [Bur71]. Samuel [All94, Fle99, SA15, fWT94, AC60, Cou73, DR74, Dow70, SG79, Tre11]. Sandage [LS12]. Sanderson [LH90, Pir66]. Sanger [Bro15, HT03]. Sargent [Lyn15]. Sargeson [Bos12]. Satyendra [Meh75]. Saul [Sho67]. Saunders [For95]. Sawicki [DE95]. Sawney [Stu74]. Sawyer [Mas02]. Sayer [Man98]. Schild [Bla94]. Schmidt [Vog08]. Schmidt-Nielsen [Vog08]. Schneider [RS15]. Schonland [All73]. Schrödinger [Hei61]. Schwarzschild [Mes99]. Sciama [EP10]. Scott [BF72, Bow05, CGZ22, DFSV23, MR22, RW06, Tho61, TH71]. Seaborg [Hof07]. Sealy [vE57]. Seares [PW74]. Searle [Tho55]. Seaton [BS13]. Sebastian [CC61, Wat06]. Second [Kay68]. Selborne [GN24]. Selby [AW88]. Selwyn [SH18]. Semenov [Dai90]. September [Alb87, All00, AW88, AD21, Ash16, Ati07, Att91, Aut19, BR79, BB99b, BM81, BB22, BS04, Bar23, BY16, BC92, Bla77, Bla81, BN92, Bul94, Bur94, Bur78, CD88, CJ99, CV85, CJ12, CR83, CCG^{+}00, Cla09, CB81, Coo98, Coo99, Cow04, Cra24, CL82, Cut05, Dai90, DT24, Doc98, Dou07, DST85, DH04, Dun78, Ele76, Ell22, Eva05, Eva22, Eve84, Fed94, Fen87, FP95, Fle57, Fog86, For95, Fox87, Gar15, Gar81a, GCJM18, GV18, GR16, Goo82, GS86, GG00, GTK11, Gre14, GN19, Gri18, GM97, GS13, MDHW03, Hay82, HHP22, Hey91, Hey87, Hil88, Hod80, Hol96, Hop23, Hop96, Hub75, Hut23, Igg06, JG98, Jef85, Jef77, Jel99, JP81, JB84, Jon10, JG13, Jon23b, Jor04, Kat78, KC23, Kem83, Kil94, KM15, Kri95]. September [Law08, Lee07, Leg22, LM02, Lig85, LW09b, LR17, Lon22b, Lon22a, Lon23b, LL05, LBG98, Mab99, Mac77, Mac03, Mad90, Mah24, Mar12, MS96, Mas02, ML85, Mat88b, McK87, MP82, Mid21, Mil97, MS08, Mit91, Mit00, Mof85, Moo15, Moo77, MR22, Mur98, Neu78, NL22, Och90, OM80, Pag01, PSCV21, Pat79, PF99, YHP^{+}86, Pei81, PM88, Per82, Per87, PG92, Phi13, Pic16, Pig77, PB20, Pip98, Pol92, Pol06, Por96, Pow87, Pre10, PJ80, Pri94, Pud21, Pud24, PR81, QRJ12, Rai08, Ran90, Rea19, Ree98, Ree15, RB20, Rob00, Rob77, RR24, Rox07, Rub91, RK10, SR76, SL18, Sea20, Sho83, Sho85b, Sid02, SLO10, SW21a, SOD22, ST99, Sla94, CBE09, Sma10, Smi16, SW21b, Son99, Sow24, Spe94, Spr95, SSGP23, Stu09]. September [Swa96, Tei03, TC76, Tyr87, TM90, Uns14, Vin23, Vog08, WPL21, Wat18, Web10, Wei05, Whi94b, WM84, Whi01, Wis83, Wun97, Wur16, Zuc83, vC99]. Series [Ble96a]. Seshadri [BR79]. Severo [GM97]. Sewall [Hil90b]. Sewell [Pan65]. Seymour [CR16a, Pan65, Pea18, Sea20, Gre12, Gre13]. Sforza [Edw21]. Shackleton [DL04, Jel99, ME11]. Shannon [Jam09]. Shanti [Ses62]. Sharman [Lig98]. Sharpe [Bar81]. Sharples [Har73]. Shaw [Den74]. Sheila [Pep03]. Sheina [Rus78]. Sheldon [Boy56]. Shen [Hit14]. Shenstone [Gar81b]. Sheppard [Cla77, SP79, TC18]. Sherbon [Hil87]. Shercliff [Mof85]. Sherfield [Lor98]. Sherlock [Pep03]. Sherrington [RH18]. Sherry [Har20]. Sherwood [Har20, WP75]. Shiing [Hit14]. Shiing-Shen [Hit14]. Shoenberg [Pip05]. Shoppee [Man03]. Shortt [Gar88]. Shosaku [Imu95]. Shotton [Coo94]. Siddiqui [Akh96]. Sidney [Dun83b, Tay64, TK74]. Sidnie [Fry80]. Siegbahn [Att91]. Siegmund [BB97]. Sigmungd [Kin56]. Simon [Kur58, MM21, SOD22]. Simonsen [Rob60]. Simpson [Gol65, Whi86]. Singh [Igg06]. Sinton [Chr56]. Sir [All13, AM16, AD21, Bar04, BG02, BH08, BG95, Ble01, Bod04, BG14, BE05, Bro13, Bro16, Buc10, Buc06, Bur03, Bur01, Bur19, BM13, But01, CK14, Cha03, Chi99, Cla09, Col18, Coo90, Cow04, Cro20, Cru00, Cul95, CA01, Dal97, DG16, Dol95, Dou19, DH04, DS21, EW22, Eva99, Eva22, FH98, Fla03, Fle99, For95, FAL21, Gab02, Gad21, Gar21, GMP24, Gor24, Gra10, GG00, Gre06, GN19, Han00, MDHW03, HD98, Hig24, Hil11, HP08, HP09, Hit20, Hol01, Joh03, Jon03, JJ14, Jor04, Kel17, Kev05, LR19, Law01, LL23, Leg17, LMW06, LM02, Loc96, LW09b, LC17, MP23, ML19, ME11, McC08, MB21, Mid21, Mil97, MR96, Mon06, MR22, Mor05, Mor21, Mor19, NN00, O’H06, MP24, Ped01, PK24, Pet92, Pet24]. Sir [PB10, Pip98, Pol06, Qui05, Ran22, RP96, Rei24, Roo96, Rox07, Rus24, RK10, Sea20, Sim18, CBE09, Sma10, SK13, Smi98, Son99, Str96, Tho98, Tho24, Vin98, Wad97, War17, WF19, Wea02, Wei24, Wel05, Whe01, Whi96, Whi22, Whi01, Wil13, Wil23b, Wil23c, Yat06]. Sisir [Rat64]. Sivaramakrishna [GLR07]. Skempton [Cha03]. Skinner [Jon60a, WM84]. Skottsberg [Sal64]. Slack [Lai20]. Slater [Bor17, Kay71]. Smallman [Goo16]. Smith [Ast01, BN92, BK17, CH05, Cow57, CP94, Dat88, DSD20, Dou19, Eva66, Goo72, Goo82, Gri18, Hab71, Han18, Her81, Kas82, Ric14, Sau87, ŠR24, Whi96]. Smithies [Bal24]. Smithson [JP81]. Smyth [BM81]. Sneath [JG13]. Sneddon [Cha02]. Snow [Cla70]. Society [Ano71a, McR76, Ano71a, McR76, Lin56, Wil95]. Soddy [Fle57]. Solly [Kro95]. Solomon [Woo94, Hod73]. Sondheimer [JG82]. Sonneborn [Bea82]. Sophia [DM06]. Sorell [Bak76]. Southwell [Chr72]. Southwood [MH08]. Spalding [LPP19]. Spath [Cox57]. Spear [Ada09]. Spence [Glu77]. Spencer [Adr64, Bar60, CEP15, Haw56, Jon66, Stu66, Til61, Woo61].

Sperry [Von97]. Spinks [JRS84]. Spitzer [Ost07]. Squire [GY62]. Srinivasa [LB67]. St [Bra70, Lew02, Mit91]. Stacey [Ove97]. Stafford [McC22, Sch55]. Stanier [Cla86, Har66]. Stanislao [Big02]. Stanislaw [How01]. Stanley [Alp88, Bai88, Dom98, Gor24, Hew02, HF06, HF08, Mar88, Mas02, PF99, Wil55, Bai94, Col02, HT70, Men70, YHP^{+}86, RC24, Smi73]. Stapledon [Rus61]. Steacie [Mar64]. Stebbins [BS05]. Stedman [Cru76]. Steitz [RH22]. Stensiö [Pat90]. Stephen [Con09, Dou64, Bar98, CEG^{+}19, DT20, Gri11, Har91, Kat82, MY73, Mui94]. Stephens [TB13]. Stephenson [Yon62]. Steptoe [Edw96]. Sterling [Pol06]. Steuart [Mur98]. Stevens [HM03]. Stevenson [Whi04]. Steward [Kri95]. Stewart [BBD84, BP99, CS98, Gar06, Hux92, LMA80, MR13, O’H06, PB17, Ruf02]. Stewartson [Lig85]. Stiles [Alp88, Jam67]. Stirling [MW20]. Stockton [Hog87]. Stodart [Bra95]. Stoicheff [van19]. Stoker [Wei24]. Stoll [Ruz72]. Stommel [Wun97]. Stoneley [Jef76, MD81]. Stoner [Bat69]. Stopford [CC61]. Storey [Smi69]. Störmer [Cha58]. Stott [Kem75]. Stowell [Bro24]. Stoyle [Ell08]. Strang [Pag01]. Stratten [DS86]. Stratton [Cha61]. Stroud [Gui97]. Struther [Ree15]. Struve [Cow64]. Stuart [Bel66, GS90, HD92, Lov75, MS08]. Stubblefield [Whi01]. Subrahmanyan [Tay96]. Sucksmith [CL82]. Suffolk [HH95]. Sugden [QT86]. Sulston [WF19]. Summers [CEW18]. Susan [RR05]. Sutcliffe [Saw92]. Sutherland [She82, SC99]. Sutton [PSS78, Ske70, Spr95, Whi94a, Wig83]. Svante [Bla85]. Svedberg [CP72]. Svedelius [Sko61]. Sven [Jac00]. Swallow [Cha97]. Swann [Mit91]. Swarup [KC23, Ses62]. Sweden [Whe74]. Swinburne [Fre60]. Sydney [AR09, Cra72, GI00, Mor16, Str56, Whe01, Cow71, DSD20, FSC20, Hut84, Lig90, Par72, WB20]. Sykes [All83]. Sylvia [DM06]. Symons [Atk04, RE08]. Synge [Flo08, Gor96]. Szentágothai [GS13]. Szwarc [JGP06].

Tabor [Fie08]. Tadeus [Rot99]. Tait [DM06, VC18]. Talling [MH19]. Tandy [Phi83]. Tansley [God57]. Taube [FM21]. Taunton [Kib88]. Tayler [Mes98]. Taylor [Bat76, Cul95, Dun68, Kem75, Pay81, Tem56]. Tecwyn [NS82]. Telegdi [Bro09]. Temin [Dul95]. Temple [Kil94, Tho19]. Terence [Kel17]. Thaddeus [Jon23a]. Thatcher [Aga23]. Theodor [Dal83]. Theodore [Dub56, Fei95, Hof07, Tur55b, Whi22, Gol66, QT86]. Theodosius [For77]. Theorell [Dal83]. Third [Ree94]. Thiruvenkata [BR79]. Thoday [MT21, Sti65]. Thode [CCG^{+}00]. Thomas [Ber63, BM24, Bur01, CH19, DS21, Far61, Kin16, Oak66, PR78, Pro68, Pug67, Tay62, WB05, Wel78, Wil94a, You12, Adr66, BC60, BR71, Bon06, Bra92, BW22, Bro77, BT68, Cha85a, CEW18, Dal61, Gui97, Hab71, Hai24, HG70, HM03, HW87, Hol94, Hun03, Law08, Lea91, MH08, Mes91, Mor66, PF99, RH22, Rat60, Rob92, SK66, Yon62]. Thompson [BS04, CS98, Ric85, Tho73b]. Thomson [AGJ^{+}22, Bla60a, Gar70b, Moo77, Tho91]. Thornton [Nut77]. Thorpe [Hin87, Kro95]. Thouless [Leg22]. Tiegs [Pan57]. Tilley [DN74]. Timoshenko [MY73]. Tinbergen [Hin90]. Tiselius [KP74]. Titchmarsh [Car64]. Tizard [Far61]. Tjardus [vC99]. Tobias [McK22]. Todd [And58, Ati96, BK00]. Toit [KS69]. Tolansky [DR74]. Tom [WK03]. Tompkins [KE04]. Townsend [SM24, vE57]. Tracy [Bea82]. Trafford [Gur12]. Travers [Baw63]. Trevan [Gad57]. Trevelyan [Adr63]. Trevor [Wri24, BM22, Den74]. Trueman [Pug58]. Trumpington [BK00]. Tsien [Hua18]. Tukey [McC03]. Tulving [Cra24]. Turing [New55]. Turnbull [Ait62, Bed57]. Turner [Dou64, Ing68]. Turrill [Hub71]. Turton [Wil87]. Tutin [Bra92]. Tutte [You12]. Tuzo [Gar95]. Twyman [Men60]. Tyndall [MP62]. Tyrrell [KTR07].

Ubbelohde [Wei90]. Ugo [BIR12]. Ulf [Bla85]. Ulick [CM81]. Ulrich [CL14]. Ulrik [Tis60]. Underwood [Bla81]. Urey [CRST83]. Urgel [Bun02]. Ursell [AM13]. Ussing [Lar09]. Uvarov [Wig71].

Valentine [Spi91, Bro09]. Valera [Syn76]. Vallance [Ati76]. Vallentine [McK22]. Vane [Mon06]. Vaughan [Eva22, Owe95]. Venetia [Tat94]. Vening [BS67]. Venkata [Bha71]. Verner [CP69]. Verney [DP70]. Vernon [KC22, Mul94, Por68]. Vero [New98]. Vessot [Fos57]. VI [Whe74]. Vickerman [CG21]. Victor [Abr88, Coo99, Den86, Flo20, Man98, PB79, Rat66, Ree94, Tre11, Wis83, Alb87, LBG98]. Vincent [CK14, Cot84, Loc96, Oak62, Fal70, WB03]. Vinogradov [CV85]. Violet [Pep03]. Viscount [Bri58, Hol66, Men70, Tho58]. Vitaly [Lon11b]. Vivian [Kat78, Law01, LMW06]. Vladimir [ADE00, KT18]. Vleck [Ble82]. Vogt [Cut05]. Vol [Ano71a, Lin56, McR76]. Volume [Lon20a, Lon20b, Lon21b, Ano68]. Vulimiri [MT08]. Vynne [Chr72].

W [BH18]. Waddington [Rob77]. Wade [Min22]. Wadia [Stu70]. Waelsch [Pap19]. Wager [Dee67]. Wain [Fow02]. Wakefield [Jon60a]. Walker [Hit06, HU19, Hor14, SS06, Tay62]. Wall [Fit22]. Wallace [SK66, Tod79, Fle55a, Lyn15]. Wallis [PR81]. Walls [KM15]. Walsh [Han00, Pri78]. Walshe [Phi74]. Walter [Abr71, DS21, Gab02, Leg17, MS19, NH13, She18, TH16, Whi62, WZ79, Ada09, Alp88, BS04, Big10, BO58, Bul66, FR98, GW20, HL18, Jam67, KH04, MA21, Mot82, Pic68, Str56, WB05]. Walters [DT24]. Walton [KH04]. Warburg [Kre72]. Warcup [BY16]. Ward [Dom21]. Wareing [Hal99]. Warner [Gar21, Gut21, TJ24, vC99]. Warris [Ric85]. Washington [Zuc83]. Waterhouse [DWS02]. Waterlow [Ano19, Mil18]. Waters [NJ86]. Watson [FP95, MO23, PW74, Rat75, Wat18, Whi66b]. Watson-Watt [Rat75]. Watt [GS90, MM87, Rat75, TM90]. Watts [Nee02]. Waverley [Bri58]. Wayne [Moo15]. Weatherall [Hig24]. Weck [WW86]. Weedon [Pat05]. Weil [Ser99]. Weinberg [DR22]. Weiskrantz [Goo20]. Weisner [SV91]. Weiss [Pro23]. Welford [BS04]. Wells [BC07b, Fog86, Hor95, Tho19]. Welsh [Sto94]. Werner [Ale96, Pan57, Hay65, MP77]. Wesley [Bar11]. West [GT23]. Westcott [May15]. Westheimer [BC18]. Westley [Cha03]. Westmacott [Sou87]. Westoll [PF99]. Westren [Ait62]. Wetherell [All00]. Weyl [New57]. Whatley [Joh23]. Wheeler [CK06, Pig77]. Whelan [Azz24]. Whiddington [Fea71]. Whiffen [McL04]. Whitchurch [Woo95b]. White [MP82, MR96, PM94, Stu85]. Whitehead [New61]. Whitelaw [Lau07]. Whitham [MS15]. Whittaker [EW22, Hay85, Tem56]. Whittard [Bul66]. Whitteridge [GI96]. Whittington [For12]. Whittle [FH98]. Whyte [MB21]. Whytlaw [CH58]. Whytlaw-Gray [CH58]. Widdowson [Ash02]. Wieland [Kar58]. Wigglesworth [Loc96]. Wigner [SVW00]. Wild [FE12]. Wilder [McC03, EF78]. Wilfred [Pay81, Rob76]. Wilfrid [Buc77, LH88, Zuc73]. Wilhelm [Dav70, BB97]. Wilkes [CK14]. Wilkie [Wol01]. Wilkins [AKS06]. Wilkinson [Fox87, GG00]. Willans [Wil60]. William [Arn91a, Baw63, Ble96a, Bod04, Bun94, CEG^{+}19, Coo94, Cow04, DFSV23, Edw20, EP10, Far98, Gad57, Gar06, GR16, Gor24, Gre06, GN19, Han18, Hil72, JS10, Kat82, Kre73, Lyn15, Man03, Mar64, Mat90b, MR96, Nag19, Oak73b, Pea68, PC10, PC11, RP96, CBE09, Sma10, Smi58, Smi98, Stu13a, Tay64, Wes95, Wig83, Yat06, YPS78, Ati76, Azz24, Bak77, Bar86, Baw58, Baw61, Ber63, BS01, BH96, Bul75, Bul69, Cla71, CH79, Con09, Cow57, Cox58, CSB22, Dav71, DAB97, Dou64, DGV11, Dun78, Gar75, Gra04, Har61, Har66, Hil58, Hin87, Hol66, Hub71, Kay71, KR11, LMA80, Mac23, MM87, Mak94, Man60, Man87, MW20, MP07, NJ86, Pag01, Pat79, PM88, Phi79, Poo60, Qua94, Rai08, Rat71, Ray69, RS15, Red56, Ree08, Rei14, Rel61, Rit00]. William [Roc70, Sal61, Sho63, Spi91, Stu74, Sut80, TF67, TH71, Tho73b, Wai97, Web10, Whi66a, Wil75, Woo05, You12]. Williams [BE05, Cru04, Dat88, Die56, HT16, KP78, NS82, New62, SR17, Wig82]. Williamson [Fei95]. Willie [CL82]. Willis [KM70, Tur58, GB71]. Willliam [Haw56]. Wilson [AW88, BBD84, Bha24, Bla60a, Can14, Gar95, Hai24, Jor04, KM70, Son99, TF67, Tho65, Woo97b, Eva66, ML03]. Wilton [Rit00]. Windrush [Hey87]. Wingate [Pie99]. Winge [Wes64]. Wingfield [Vog72]. Winifred [BF91]. Winston [Jon66]. Wittrick [Man87]. Wolcott [Von97]. Wolfe [Coo72]. Wolfendale [EW22]. Wolfgang [CL14, Hay65, Pei60]. Wolfson [Bul94, Ram22]. Wolpert [TS24]. Wood [AB88, BM13, Cla55b, Die56, Lyo22, Wil23b]. Woodham [HP08, HP09]. Woodruff [Mor05]. Woods [Bar98, GF65]. Woodward [TC81]. Woodworth [JB84]. Wooldridge [Tay64]. Woolfson [WD21]. Woolley [McC88]. Work [BG02]. Wormall [MF66]. Worsley [Gib73]. Worthington [MR13]. Wragg [Gar81a]. Wray [Aca08]. Wreyford [DT81]. Wright [Cow57, Hil90b]. Wulf [Adr76]. Wyllie [AS24]. Wyndham [HB22]. Wynne [MW01, New98]. Wynne-Edwards [New98]. Wyrill [SB08b]. Wyssling [Mat90a].

XVI [Ano71a].

Yakov [Gin94]. Yarnton [FGW13]. Yasutomi [YN06]. Yate [Cla86]. Yates [Fin95]. Yonchien [Hua18]. Yonge [Mor92]. Young [Boy98, CR07, DS79, Ran90, SL88]. Yukawa [Kem83].

Zacharias [KS92]. Zachary [Boy98]. Zangwill [Gre01]. Zbigniew [How01]. Zeeman [Ran22]. Zel’Dovich [Gin94]. Zernike [Tol67]. Zewail [Tho20]. Ziegler [Baw75]. Zienkiewicz [Owe09]. Ziman [BN06]. Zuckerman [Kro95].

References

Akhtar:1998:AFJ

[AA98] Muhammad Akhtar and Monika E. Akhtar. Adolf Friedrich Johann Butenandt. 24 March 1903–18 January 1995. Biographical Memoirs of Fellows of the Royal Society, 44:79–92, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0006.

Andrews:2002:AGN

[AA02] Elizabeth B. Andrews and J. A. Allen. Alastair Graham. 6 November 1906–12 December 2000. Biographical Memoirs of Fellows of the Royal Society, 48:129–150, January 2002. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0008.

Altmann:1974:CAC

[AB74] S. L. Altmann and Edmund John Bowen. Charles Alfred Coulson, 1910–1974. Biographical Memoirs of Fellows of the Royal Society, 20:74–134, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1974.
0004.

Aroney:1988:RJW

[AB88] M. J. Aroney and Amyand David Buckingham. Raymond James Wood Le Fèvre, 1 April 1905–26 August 1986. Biographical Memoirs of Fellows of the Royal Society, 34:373–403, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0014.

Anderson:2015:MJR

[AB15] Thomas R. Anderson and Harry L. Bryden FRS. Michael John Robert Fasham. 29 May 1942–7 June 2008. Biographical Memoirs of Fellows of the Royal Society, 61:103–121, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0014.

Abercrombie:1961:RGH

[Abe61] Michael Abercrombie. Ross Granville Harrison, 1870–1959. Biographical Memoirs of Fellows of the Royal Society, 7:110–126, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0009.

Abraham:1971:HWF

[Abr71] Edward Penley Abraham. Howard Walter Florey, Baron Florey of Adelaide and Marston, 1898–1968. Biographical Memoirs of Fellows of the Royal Society, 17:255–302, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1971.0011.

Abraham:1983:EBC

[Abr83] Edward Penley Abraham. Ernst Boris Chain, 19 June 1906– 12 August 1979. Biographical Memoirs of Fellows of the Royal Society, 29:42–91, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1983.
0003.

Abragam:1988:LVP

[Abr88] Anatole Abragam. Louis Victor Pierre Raymond De Broglie, 15 August 1892–19 March 1987. Biographical Memoirs of Fellows of the Royal Society, 34:21–41, December 1988. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/770045;   https:
//royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.
0002.

Allibone:1960:SRM

[AC60] Thomas Edward Allibone and J. R. Clarke. Samuel Roslington Milner, 1875–1958. Biographical Memoirs of Fellows of the Royal Society, 5:128–147, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Amoroso:1972:HME

[AC72] Emmanuel Ciprian Amoroso and George Washington Corner. Herbert McLean Evans, 1882–1971. Biographical Memoirs of Fellows of the Royal Society, 18:82–186, November 1972. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1972.0005.

Acar:2008:GRW

[Aca08] Memş Acar. Gordon Richard Wray. 30 January 1928–17 August 2005. Biographical Memoirs of Fellows of the Royal Society, 54:461– 482, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2008.0014.

Andersen:2020:DCO

[ACW20] Hans C. Andersen, Arup K. Chakraborty, and John D. Weeks. David Chandler. 15 October 1944–18 April 2017. Biographical Memoirs of Fellows of the Royal Society, 68:87–102, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0046.

Ashcroft:2021:SPJ

[AD21] By S. J. H. Ashcroft and R. M. Denton. Sir Philip John Randle. 16 July 1926–26 September 2006. Biographical Memoirs of Fellows of the Royal Society, 70:359–385, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0044.

Adams:1985:MHA

[Ada85] John Frank Adams. Maxwell Herman Alexander Newman, 7 February 1897–22 February 1984. Biographical Memoirs of Fellows of the Royal Society, 31:437–452, November 1985. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1985.0015.

Adams:2009:WES

[Ada09] Alfred Adams FRS. Walter Eric Spear. 20 January 1921– 21 February 2008. Biographical Memoirs of Fellows of the Royal Society, 55:267–289, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.
2003.

Arigoni:2000:VPJ

[ADE00] Duilio Arigoni, Jack D. Dunitz, and Albert Eschenmoser. Vladimir Prelog. 23 July 1906–7 January 1998. Biographical Memoirs of Fellows of the Royal Society, 46:443–464, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0095.

Adrian:1963:GMT

[Adr63] Edgar Douglas Adrian. George Macaulay Trevelyan, 1876–1962. Biographical Memoirs of Fellows of the Royal Society, 9:315–321, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1963.0017.

Adrian:1964:HSG

[Adr64] Edgar Douglas Adrian. Herbert Spencer Gasser, 1888–1963. Biographical Memoirs of Fellows of the Royal Society, 10:75–82, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0005.

Adrian:1966:TGB

[Adr66] Edgar Douglas Adrian. Thomas Graham Brown, 1882–1965. Biographical Memoirs of Fellows of the Royal Society, 12:22–33, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0002.

Adrian:1976:DWB

[Adr76] Richard Hume Adrian. Detlev Wulf Bronk, 13 August 1897– 17 November 1975. Biographical Memoirs of Fellows of the Royal Society, 22:1–9, November 1976. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1976.
0001.

Agar:2023:MHT

[Aga23] Jon Agar. Margaret Hilda Thatcher. 13 October 1925–8 April 2013. Biographical Memoirs of Fellows of the Royal Society, 74:403–418, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2022.0036.

Armstrong:2022:AJT

[AGJ^{+}22] Fraser A. Armstrong, Roger Grinter, Michael K. Johnson, Nick E. Le Brun, and Michael T. Wilson. Andrew James Thomson. 31 July 1940–2 March 2021. Biographical Memoirs of Fellows of the Royal Society, 73:471–497, December 2022. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0006.

Askonas:1997:NKJ

[AH97] Brigitte A. Askonas and James G. Howard. Niels Kaj Jerne. 23 December 1911–7 October 1994. Biographical Memoirs of Fellows of the Royal Society, 43:237–251, January 1997. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0013.

Aitken:1962:HWT

[Ait62] Alexander Craig Aitken. Herbert Westren Turnbull, 1885–1961. Biographical Memoirs of Fellows of the Royal Society, 8:149–158, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1962.0011.

Akhtar:1996:SSM

[Akh96] Muhammad Akhtar. Salimuzzaman Siddiqui, M. B. E., 19 October 1897–14 April 1994. Biographical Memoirs of Fellows of the Royal Society, 42:401–417, November 1996. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0025.

Akhtar:2000:GJP

[Akh00] Muhammad Akhtar. George Joseph Popjàk. 5 May 1914– 30 December 1998. Biographical Memoirs of Fellows of the Royal Society, 46:403–424, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0093.

Arnott:2006:MHF

[AKS06] Struther Arnott, T. W. B. Kibble, and Tim Shallice. Maurice Hugh Frederick Wilkins. 15 December 1916–5 October 2004. Biographical Memoirs of Fellows of the Royal Society, 52:455–478, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0031.

Armitage:2023:JFA

[AL23] Jonathan G. M. Armitage and Malcolm S. Longair CBE FRS. John Frank Allen. 6 May 1908–22 April 2001. Biographical Memoirs of Fellows of the Royal Society, 75:7–27, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0021.

Albery:1987:VGJ

[Alb87] Wyndham John Albery. Victor Gold, 29 June 1922–29 September 1985. Biographical Memoirs of Fellows of the Royal Society, 33:261– 288, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1987.0010.

Alexander:1996:HWL

[Ale96] Robert Mcneill Alexander. Hans Werner Lissmann, 30 April 1909–21 April 1995. Biographical Memoirs of Fellows of the Royal Society, 42:235–245, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0015.

Alexander:2001:OEL

[Ale01] R. McN. Alexander. Otto Egon Lowenstein. 24 October 1906– 31 January 1999. Biographical Memoirs of Fellows of the Royal Society, 47:357–368, January 2001. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0020.

Alexander:2022:RMA

[Ale22] Gordon Alexander. Robert McNeill Alexander. 7 July 1934–21 March 2016. Biographical Memoirs of Fellows of the Royal Society, 72:9–32, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2021.0030.

Allen:1970:HLH

[All70] P. Allen. Herbert Leader Hawkins, 1887–1968. Biographical Memoirs of Fellows of the Royal Society, 16:314–329, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0011.

Allen:1972:LBP

[All72] Norman Percy Allen. Leonard Bessemer Pfeil, 1898–1969. Biographical Memoirs of Fellows of the Royal Society, 18:476–487, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0016.

Allibone:1973:BFJ

[All73] Thomas Edward Allibone. Basil Ferdinand Jamieson Schonland, 1896–1972. Biographical Memoirs of Fellows of the Royal Society, 19:629–653, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1973.
0023.

Allen:1976:FAF

[All76] Peter Allen. Francis Arthur Freeth, 2 January 1884–15 July 1970. Biographical Memoirs of Fellows of the Royal Society, 22:104–118, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1976.0004.

Allibone:1980:DGJ

[All80] Thomas Edward Allibone. Dennis Gabor, 5 June 1900–9 February 1979. Biographical Memoirs of Fellows of the Royal Society, 26:106– 147, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769781;   https://royalsocietypublishing.org/doi/epdf/

10.1098/rsbm.1980.0004. This memoir discusses Leo Szilard’s unsuccessful attempt to convince his close friend Dennis Gabor (Gábor Dénes in Hungarian) to build the electronic microscope that Szilard had invented.

Allibone:1983:CSF

[All83] Thomas Edward Allibone. Charles Sykes, 27 February 1905– 29 January 1982. Biographical Memoirs of Fellows of the Royal Society, 29:552–583, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1983.
0020.

Allibone:1984:CRB

[All84] Thomas Edward Allibone. Cecil Reginald Burch, 12 May 1901–19 July 1983. Biographical Memoirs of Fellows of the Royal Society, 30:3–42, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1984.
0001.

Allibone:1991:JDD

[All91] Thomas Edward Allibone. James Dyson, 10 December 1914– 22 January 1990. Biographical Memoirs of Fellows of the Royal Society, 37:150–174, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0008.

Allibone:1994:JSF

[All94] Thomas Edward Allibone. John Samuel Forrest, 20 August 1907–11 November 1992. Biographical Memoirs of Fellows of the Royal Society, 40:105–126, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.
0031.

Allen:1999:GGC

[All99] Geoffrey Allen. Geoffrey Gee, C.B.E. 6 June 1910–13 December 1996. Biographical Memoirs of Fellows of the Royal Society, 45:183– 194, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0013.

Allaby:2000:AMW

[All00] James V. Allaby. Alan Marmaduke Wetherell. 31 December 1932–13 September 1998. Biographical Memoirs of Fellows of the Royal Society, 46:565–576, January 2000. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0101.

Allan:2013:SWI

[All13] William Allan. Sir William Ian Axford. 2 January 1933–13 March 2010. Biographical Memoirs of Fellows of the Royal Society, 59: 5–31, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2013.0007.

Alpern:1988:WSS

[Alp88] M. Alpern. Walter Stanley Stiles, 15 June 1901–15 December 1985. Biographical Memoirs of Fellows of the Royal Society, 34:815–885, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1988.0026.

Allen:2001:ANA

[AM01] Anthony K. Allen and Helen M. Muir. Albert Neuberger. 15 April 1908–14 August 1996. Biographical Memoirs of Fellows of the Royal Society, 47:369–382, January 2001. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0021.

Abrahams:2013:FJU

[AM13] I. D. Abrahams and P. A. Martin. Fritz Joseph Ursell. 28 April 1923–11 May 2012. Biographical Memoirs of Fellows of the Royal Society, 59:407–421, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2013.
0005.

Anthony:2016:SHD

[AM16] Christopher Anthony and J. Colin Murrell. Sir Howard Dalton. 8 February 1944–12 January 2008. Biographical Memoirs of Fellows of the Royal Society, 62:89–107, January 2016.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0007.

Andrade:1957:HRR

[And57] Edward Neville Da Costa Andrade. Harold Roper Robinson, 1889– 1955. Biographical Memoirs of Fellows of the Royal Society, 3:160– 172, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1957.0011.

Andrewes:1958:CT

[And58] Christopher Howard Andrewes. Charles Todd, 1869–1957. Biographical Memoirs of Fellows of the Royal Society, 4:281–290, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1958.0022.

Andrewes:1967:AI

[And67] Christopher Howard Andrewes. Alick Isaacs, 1921–1967. Biographical Memoirs of Fellows of the Royal Society, 13:204–221, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1967.0010.

Andrewes:1971:FPR

[And71] Christopher Howard Andrewes. Francis Peyton Rous, 1879–1970. Biographical Memoirs of Fellows of the Royal Society, 17:643–662, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0025.

Andrewes:1979:JC

[And79] Christopher Howard Andrewes. James Craigie, 1899–1978. Biographical Memoirs of Fellows of the Royal Society, 25:233–240, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1979.0007.

Andler:2006:JLN

[And06] Martin Andler. Jean Leray. 7 November 1906–10 November 1998. Biographical Memoirs of Fellows of the Royal Society, 52:137–148, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2006.0011.

Anonymous:1960:A

[Ano60] Anonymous. Addendum. Biographical Memoirs of Fellows of the Royal Society, 5:1, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Anonymous:1966:CA

[Ano66] Anonymous. Corrigenda–Addendum. Biographical Memoirs of Fellows of the Royal Society, 12:565, November 1966. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1966.0029.

Anonymous:1968:EV

[Ano68] Anonymous. Erratum to volume 13. Biographical Memoirs of Fellows of the Royal Society, 14:469, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0021.

Anonymous:1971:BRB

[Ano71a] Anonymous. Book review: Biographical Memoirs of Fellows of the Royal Society, Vol. XVI. London: Royal Society. 1970. Pp. 562. The Lancet (London, England), 297(7688):24, January 1971. CODEN LANCAO. ISSN 0140-6736 (print), 1474-547X (electronic).

Anonymous:1971:E

[Ano71b] Anonymous. Errata. Biographical Memoirs of Fellows of the Royal Society, 17:757, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1971.
0031.

Anonymous:1973:C

[Ano73] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 19:695, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1973.

Anonymous:1974:C

[Ano74] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 20:505, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1974.
0022.

Anonymous:1975:C

[Ano75] Anonymous. Corrigendum. Biographical Memoirs of Fellows of the Royal Society, 21:585, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1975.
0020.

Anonymous:1977:E

[Ano77] Anonymous. Erratum. Biographical Memoirs of Fellows of the Royal Society, 23:643, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0024.

Anonymous:1978:C

[Ano78] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 24:605, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1978.
0021.

Anonymous:1979:C

[Ano79] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 25:i, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1979.
0001.

Anonymous:1980:C

[Ano80] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 26:543, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1980.

Anonymous:1981:C

[Ano81] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 27:697, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0026.

Anonymous:1982:C

[Ano82] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 28:718, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0027.

Anonymous:1983:C

[Ano83] Anonymous. Corrigendum. Biographical Memoirs of Fellows of the Royal Society, 29:677, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1983.
0024.

Anonymous:1984:C

[Ano84] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 30:595, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1984.
0020.

Anonymous:1985:C

[Ano85] Anonymous. Corrigendum. Biographical Memoirs of Fellows of the Royal Society, 31:667, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0024.

Anonymous:1986:C

[Ano86] Anonymous. Corrigendum. Biographical Memoirs of Fellows of the Royal Society, 32:677, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.

Anonymous:1987:A

[Ano87] Anonymous. Addenda. Biographical Memoirs of Fellows of the Royal Society, 33:729, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0026.

Anonymous:1988:A

[Ano88] Anonymous. Addendum. Biographical Memoirs of Fellows of the Royal Society, 34:1003, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.
0030.

Anonymous:1990:A

[Ano90a] Anonymous. Addendum. Biographical Memoirs of Fellows of the Royal Society, 36:603, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0047.

Anonymous:1990:C

[Ano90b] Anonymous. Corrigenda. Biographical Memoirs of Fellows of the Royal Society, 35:425–426, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). See [Dac88, Gar88].

Anonymous:1990:E

[Ano90c] Anonymous. Errata. Biographical Memoirs of Fellows of the Royal Society, 36:601–602, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0046.

Anonymous:1991:E

[Ano91] Anonymous. Errata. Biographical Memoirs of Fellows of the Royal Society, 37:491, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0025.

Anonymous:1992:E

[Ano92] Anonymous. Errata. Biographical Memoirs of Fellows of the Royal Society, 38:413, November 1992. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1992.
0021.

Anonymous:1994:E

[Ano94] Anonymous. Errata. Biographical Memoirs of Fellows of the Royal Society, 39:451, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Anonymous:1998:NLF

[Ano98] Anonymous. Norman Leslie Falcon. 29 May 1904–31 May 1996. Biographical Memoirs of Fellows of the Royal Society, 44:161–174, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1998.0011.

Anonymous:2013:CE

[Ano13] Anonymous. Correction for editorial. Biographical Memoirs of Fellows of the Royal Society, 59:437, January 2013. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0043. See [Stu13b].

Anonymous:2019:CJC

[Ano19] Anonymous. Correction for John Conrad Waterlow. 13 JUNE 1916–19 OCTOBER 2010. Biographical Memoirs of Fellows of the Royal Society, 66:479, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.

0037. See [Mil18].

Anthony:2015:JRQ

[Ant15] Christopher Anthony. John Rodney Quayle. 18 November 1926– 26 February 2006. Biographical Memoirs of Fellows of the Royal Society, 61:331–349, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0008.

Apple:2007:NHL

[App07] David J. Apple. Nicholas Harold Lloyd Ridley. 10 July 1906–25 May 2001. Biographical Memoirs of Fellows of the Royal Society, 53:285– 307, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2007.0022.

Ash:2009:EGS

[AR09] Eric A. Ash CBE FRS and E. Peter Raynes FRS. Edward George Sydney Paige. 18 July 1930–20 February 2004. Biographical Memoirs of Fellows of the Royal Society, 55:185–200, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0009.

Arkell:1955:GML

[Ark55] William Joscelyn Arkell. George Martin Lees, 1898–1955. Biographical Memoirs of Fellows of the Royal Society, 1:162–173, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1955.0012.

Arndt:1991:CWB

[Arn91a] Ulrich Wolfgang Arndt. Charles William Bunn, 15 January 1905–13 April 1990. Biographical Memoirs of Fellows of the Royal Society, 37:69–83, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0004.

Arnold:1991:JBH

[Arn91b] M. H. Arnold. Joseph Burtt Hutchinson, 21 March 1902– 16 January 1988. Biographical Memoirs of Fellows of the Royal Society, 37:277–297, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0014.

Arnott:1994:JMR

[Arn94] Struther Arnott. John Monteath Robertson, 24 July 1900–27 December 1989. Biographical Memoirs of Fellows of the Royal Society, 39:349–362, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Ambler:2002:MRP

[AS02] Richard P. Ambler and Kenneth Murray F. R. S. Martin Rivers Pollock. 10 December 1914–21 December 1999. Biographical Memoirs of Fellows of the Royal Society, 48:357–373, January 2002.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0021.

Atkin:2003:FML

[AS03] R. J. Atkin and T. J. Sluckin. Frank Matthews Leslie. 8 March 1935–15 June 2000. Biographical Memoirs of Fellows of the Royal Society, 49:315–333, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2003.
0018.

Aitchison:2016:RHD

[AS16] Ian J. R. Aitchison and Sir Chris Llewellyn Smith FRS. Richard Henry Dalitz. 28 February 1925–13 January 2006. Biographical Memoirs of Fellows of the Royal Society, 62:59–88, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0019.

Ahlberg:2021:JCN

[AS21] P. E. Ahlberg and T. R. Smithson. Jennifer A. Clack. 3 November 1947–26 March 2020. Biographical Memoirs of Fellows of the Royal Society, 71:79–101, December 2021. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0008.

Stanley:2024:ADH

[AS24] Mark J. Arends and Margaret Stanley. Andrew David Hamilton Wyllie. 24 January 1944–26 May 2022. Biographical Memoirs of Fellows of the Royal Society, 76:501–518, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0025.

Ashwell:2002:EMW

[Ash02] Margaret Ashwell O.B.E. Elsie May Widdowson, C.H. 21 October 1906–14 June 2000. Biographical Memoirs of Fellows of the Royal Society, 48:483–506, January 2002. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0028.

Ashmore:2016:PFJ

[Ash16] Jonathan F. Ashmore FRS. Paul Fatt. 13 January 1924– 28 September 2014. Biographical Memoirs of Fellows of the Royal Society, 62:167–186, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2016.
0005.

Ashfold:2022:RND

[Ash22] M. N. R. Ashfold FRS. Richard Newland Dixon. 25 December 1930–25 May 2021. Biographical Memoirs of Fellows of the Royal Society, 73:157–175, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0005.

Askonas:1990:JHH

[Ask90] Brigitte Alice Askonas. John Herbert Humphrey, 16 December 1915–25 December 1987. Biographical Memoirs of Fellows of the Royal Society, 36:273–300, December 1990. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1990.0033.

Arnon:2016:EKK

[ASS16] Ruth Arnon, Michael Sela, and Colin Shindler. Ephraim Katchalski-Katzir. 16 May 1916–30 May 2009. Biographical Memoirs of Fellows of the Royal Society, 62:267–275, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0015.

Astell:2001:MSA

[Ast01] Caroline R. Astell. Michael Smith. 26 April 1932–4 October 2000. Biographical Memoirs of Fellows of the Royal Society, 47:429–441, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2001.0025.

Alderman:1960:DM

[AT60] A. R. Alderman and Cecil Edgar Tilley. Douglas Mawson, 1882– 1958. Biographical Memoirs of Fellows of the Royal Society, 5:119– 127, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Atiyah:1976:WVD

[Ati76] Michael Francis Atiyah. William Vallance Douglas Hodge, 17 June 1903–7 July 1975. Biographical Memoirs of Fellows of the Royal Society, 22:169–192, November 1976. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0007.

Atiyah:1996:JAT

[Ati96] Michael Francis Atiyah. John Arthur Todd, 23 August 1908– 22 December 1994. Biographical Memoirs of Fellows of the Royal Society, 42:483–494, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0029.

Atiyah:2007:RHB

[Ati07] Michael Atiyah. Raoul Harry Bott. 24 September 1923–20 December 2005. Biographical Memoirs of Fellows of the Royal Society, 53: 63–76, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2007.0006.

Atiyah:2014:FEP

[Ati14] Sir Michael Atiyah OM FRS. Friedrich Ernst Peter Hirzebruch. 17 October 1927–27 May 2012. Biographical Memoirs of Fellows of the Royal Society, 60:229–247, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0010.

Atkins:2004:MCR

[Atk04] Peter Atkins. Martyn Christian Raymond Symons. 12 November 1925–28 January 2002. Biographical Memoirs of Fellows of the Royal Society, 50:299–308, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.

Atterling:1991:KMG

[Att91] H. Atterling. Karl Manne Georg Siegbahn, 3 December 1886– 24 September 1978. Biographical Memoirs of Fellows of the Royal Society, 37:427–444, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0022.

Authier:2019:ARL

[Aut19] André Authier. Andrew Richard Lang. 9 September 1924–30 June 2008. Biographical Memoirs of Fellows of the Royal Society, 67:237– 254, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0003.

Avery:2008:GBE

[Ave08] Mary Ellen Avery. Gertrude Belle Elion. 23 January 1918– 21 February 1999. Biographical Memoirs of Fellows of the Royal Society, 54:161–168, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2007.
0051.

Anderson:1988:GSW

[AW88] Ephraim Saul Anderson and Sir Robert Williams. Graham Selby Wilson, 10 September 1895–5 April 1987. Biographical Memoirs of Fellows of the Royal Society, 34:887–919, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0027.

Azzi:2024:WJW

[Azz24] Angelo Azzi. William Joseph Whelan. 14 November 1924–5 June 2021. Biographical Memoirs of Fellows of the Royal Society, 77: 435–457, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Bailey:1988:HMS

[Bai88] H. C. Bailey. Herbert Muggleton Stanley, 20 July 1903–4 July 1987. Biographical Memoirs of Fellows of the Royal Society, 34:787–813, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1988.0025.

Bailey:1994:SMP

[Bai94] A. J. Bailey. Stanley Miles Partridge, 2 August 1913–26 April 1992. Biographical Memoirs of Fellows of the Royal Society, 40:327–346, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1994.0043.

Baker:1976:JSH

[Bak76] John Randal Baker. Julian Sorell Huxley, 22 June 1887– 14 February 1975. Biographical Memoirs of Fellows of the Royal Society, 22:206–238, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1976.
0009.

Baker:1977:WHG

[Bak77] John Fleetwood Baker. William Henry Glanville, 1 February 1900–30 June 1976. Biographical Memoirs of Fellows of the Royal Society, 23:91–113, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0005.

Ballou:2024:OSJ

[Bal24] Byron Ballou. Oliver Smithies. 23 June 1925–10 January 2017. Biographical Memoirs of Fellows of the Royal Society, 76:389–415, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2023.0043.

Bang:1983:NEN

[Ban83] T. Bang. Niels Erik Nørlund, 26 October 1885–4 July 1981. Biographical Memoirs of Fellows of the Royal Society, 29:481–493, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1983.0017.

Bartlett:1960:KSL

[Bar60] Frederic Charles Bartlett. Karl Spencer Lashley, 1890–1958. Biographical Memoirs of Fellows of the Royal Society, 5:107–118, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Barrington:1973:GRB

[Bar73] Ernest James William Barrington. Gavin Rylands de Beer, 1899– 1972. Biographical Memoirs of Fellows of the Royal Society, 19:64– 93, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1973.0003.

Barcroft:1975:IBD

[Bar75a] Henry Barcroft. Ivan de Burgh Daly, 14 April 1893–8 February 1974. Biographical Memoirs of Fellows of the Royal Society, 21:196– 226, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1975.0005.

Barrington:1975:FGW

[Bar75b] Ernest James William Barrington. Francis Gerald William Knowles, 9 March 1915–13 July 1974. Biographical Memoirs of Fellows of the Royal Society, 21:431–446, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1975.0013.

Bartlett:1981:ESP

[Bar81] Maurice Stevenson Bartlett. Egon Sharpe Pearson, 11 August 1895–12 June 1980. Biographical Memoirs of Fellows of the Royal Society, 27:425–443, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0017.

Barlow:1986:WRD

[Bar86] Horace Basil Barlow. William Rushton, 8 December 1901–21 June 1980. Biographical Memoirs of Fellows of the Royal Society, 32:421– 459, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1986.0014.

Barbour:1998:SEW

[Bar98] A. K. Barbour. Stephen Esslemont Woods. 12 January 1912– 18 June 1994. Biographical Memoirs of Fellows of the Royal Society, 44:455–469, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0029.

Barfield:2004:SGR

[Bar04] Norman Barfield. Sir George Robert Edwards OM CBE. 9 July 1908–2 March 2003. Biographical Memoirs of Fellows of the Royal Society, 50:73–92, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0007.

Barber:2011:JWM

[Bar11] D. J. Barber. John Wesley Mitchell. 3 December 1913–12 July 2007. Biographical Memoirs of Fellows of the Royal Society, 57:269–289, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2011.0007.

Bard:2021:EAB

[Bar21] Judith Bard. Edward Arthur Boyse. 11 August 1923–14 July 2007. Biographical Memoirs of Fellows of the Royal Society, 70:41–56, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2020.0035.

Barth:2023:NDS

[Bar23] Peter Thomas Barth. Naomi Datta. 17 September 1922–30 November 2008. Biographical Memoirs of Fellows of the Royal Society, 74: 123–141, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2022.0032.

Bastien:1963:AMG

[Bas63] P. G. Bastien. Albert Marcel Germain René Portevin, 1880–1962. Biographical Memoirs of Fellows of the Royal Society, 9:223–235, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1963.0012.

Bates:1969:ECS

[Bat69] Leslie Fleetwood Bates. Edmund Clifton Stoner, 1899–1968. Biographical Memoirs of Fellows of the Royal Society, 15:201–237, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1969.0011.

Batchelor:1976:GIT

[Bat76] George Keith Batchelor. Geoffrey Ingram Taylor, 7 March 1886–27 June 1975. Biographical Memoirs of Fellows of the Royal Society, 22:565–633, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1976.
0021.

Bateson:2003:GSD

[Bat03] Patrick Bateson FRS. George Salt. 12 December 1903–17 February 2003. Biographical Memoirs of Fellows of the Royal Society, 49:447– 459, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2003.0026.

Bawn:1958:WCM

[Baw58] Cecil Edwin Henry Bawn. William Cudmore McCullagh Lewis, 1885–1956. Biographical Memoirs of Fellows of the Royal Society, 4:192–203, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1958.
0017.

Bawn:1961:WEG

[Baw61] Cecil Edwin Henry Bawn. William Edward Garner, 1889–1960. Biographical Memoirs of Fellows of the Royal Society, 7:85–94, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0007.

Bawn:1963:MWT

[Baw63] Cecil Edwin Henry Bawn. Morris William Travers, 1872–1961. Biographical Memoirs of Fellows of the Royal Society, 9:301–313, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1963.0016.

Bawden:1971:AAP

[Baw71] Frederick Charles Bawden. Alfred Alexander Peter Kleczkowski, 1908–1970. Biographical Memoirs of Fellows of the Royal Society, 17:431–440, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1971.
0016.

Bawn:1975:KZN

[Baw75] Cecil Edwin Henry Bawn. Karl Ziegler, 26 November 1898– 11 August 1973. Biographical Memoirs of Fellows of the Royal Society, 21:569–584, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1975.
0019.

Bolton:1992:EBD

[BB92] T. B. Bolton and Alison F. Brading. Edith Bülbring, 27 December 1903–5 July 1990. Biographical Memoirs of Fellows of the Royal Society, 38:67–95, November 1992. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0004.

Born:1996:HBJ

[BB96] Gustav Victor Rudolf Born and P. Banks. Hugh Blaschko, 4 January 1900–18 April 1993. Biographical Memoirs of Fellows of the Royal Society, 42:40–60, November 1996. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0004.

Bisset:1997:WSF

[BB97] G. W. Bisset and T. V. P. Bliss. Wilhelm Siegmund Feldberg, C. B. E. 19 November 1900–23 October 1993. Biographical Memoirs of Fellows of the Royal Society, 43:145–170, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0009.

Bainbridge:1999:JMH

[BB99a] A. Bainbridge and K. J. Brent. John Malcolm Hirst, D.S.C. 20 April 1921–30 December 1997. Biographical Memoirs of Fellows of the Royal Society, 45:219–238, January 1999. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0015.

Baker:1999:PEM

[BB99b] A. Baker and B. Bollobás. Paul Erdõs. 26 March 1913– 20 September 1996. Biographical Memoirs of Fellows of the Royal Society, 45:147–164, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0011.

Burdekin:2018:JFK

[BB18] Frederick Michael Burdekin OBE FREng FRS and Paul Bowen FREng. John Frederick Knott. 9 December 1938–5 October 2017. Biographical Memoirs of Fellows of the Royal Society, 65:217–234, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2018.0005.

Barford:2022:DLN

[BB22] David Barford and Thomas L. Blundell. Dame Louise Napier Johnson. 26 September 1940–25 September 2012. Biographical Memoirs of Fellows of the Royal Society, 72:221–250, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0038.

Bates:1984:HSW

[BBD84] David Robert Bates, Robert Lewis Fullarton Boyd, and D. G. Davis. Harrie Stewart Wilson Massey, 16 May 1908–27 November 1983. Biographical Memoirs of Fellows of the Royal Society, 30:445– 511, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1984.0016.

Boyden:2013:FJF

[BBM13] Stephen Boyden, Robert Blanden, and Cedric Mims. Frank John Fenner FAA. 21 December 1914–22 November 2010. Biographical Memoirs of Fellows of the Royal Society, 59:125–144, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0009.

Bretscher:1955:EF

[BC55] E. Bretscher and John Douglas Cockcroft. Enrico Fermi, 1901– 1954. Biographical Memoirs of Fellows of the Royal Society, 1:68– 78, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769243;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0006.

Bennet-Clark:1960:TGM

[BC60] Thomas Archibald Bennet-Clark. Thomas Godfrey Mason, 1890– 1959. Biographical Memoirs of Fellows of the Royal Society, 6:183– 189, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1960.0031.

Bennet-Clark:1972:JB

[BC72] Thomas Archibald Bennet-Clark. John Barker, 1901–1970. Biographical Memoirs of Fellows of the Royal Society, 18:35–42, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0003.

Beverton:1992:JAG

[BC92] Raymond John Heaphy Beverton and David Henry Cushing. John Alan Gulland, 16 September 1926–24 June 1990. Biographical Memoirs of Fellows of the Royal Society, 38:163–183, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0009.

Barrett:2007:DGL

[BC07a] Spencer C. H. Barrett and Deborah Charlesworth. David Graham Lloyd. 20 June 1937–30 May 2007. Biographical Memoirs of Fellows of the Royal Society, 53:203–221, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0011.

Burdekin:2007:AAW

[BC07b] Michael Burdekin and Bernard Crossland. Alan Arthur Wells. 1 May 1924–8 November 2005. Biographical Memoirs of Fellows of the Royal Society, 53:365–384, January 2007. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0019.

Benner:2018:FHW

[BC18] Steven Benner and Elias J. Corey ForMemRS. Frank Henry Westheimer. 15 January 1912–14 April 2007. Biographical Memoirs of Fellows of the Royal Society, 65:469–480, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0009.

Briggs:1961:EJM

[BCB61] George Edward Briggs, M. J. P. Canny, and John Barker. Ernest John Maskell, 1895–1958. Biographical Memoirs of Fellows of the Royal Society, 7:161–171, November 1961. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1961.0013.

Briggs:2005:SAW

[BE05] Derek E. G. Briggs and Dianne Edwards. Sir Alwyn Williams. 8 June 1921–4 April 2004. Biographical Memoirs of Fellows of the Royal Society, 51:437–453, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2005.
0028.

Brassell:2024:GEN

[BE24] Simon C. Brassell and Timothy I. Eglinton FRS. Geoffrey Eglinton. 1 November 1927–11 March 2016. Biographical Memoirs of Fellows of the Royal Society, 76:119–143, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0004.

Beale:1982:TMS

[Bea82] Geoffrey Herbert Beale. Tracy Morton Sonneborn, 19 October 1905–26 January 1981. Biographical Memoirs of Fellows of the Royal Society, 28:537–574, November 1982. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0021.

Beale:1995:CAM

[Bea95] Geoffrey Herbert Beale. Charlotte Auerbach, 14 May 1899– 17 March 1994. Biographical Memoirs of Fellows of the Royal Society, 41:20–42, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1995.
0002.

Bedson:1957:HMT

[Bed57] Samuel Phillips Bedson. Hubert Maitland Turnbull, 1875–1955. Biographical Memoirs of Fellows of the Royal Society, 3:289–304, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1957.0020.

Bell:1966:JSF

[Bel66] Robert Edward Bell. John Stuart Foster, 1890–1964. Biographical Memoirs of Fellows of the Royal Society, 12:147–161, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1966.0006.

Bell:1981:EJB

[Bel81] Ronald Percy Bell. Edmund John Bowen, 29 April 1898– 19 November 1980. Biographical Memoirs of Fellows of the Royal Society, 27:83–101, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0004.

Bell:1986:FLE

[Bel86] George Douglas Hutton Bell. Frank Leonard Engledow, 20 August 1890–3 July 1985. Biographical Memoirs of Fellows of the Royal Society, 32:187–219, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0007.

Bell:2001:KGC

[Bel01] J. D. Bell. Keith Gordon Cox. 25 April 1933–27 August 1998. Biographical Memoirs of Fellows of the Royal Society, 47:91–104, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2001.0006.

Bendall:1994:RHA

[Ben94] D. S. Bendall. Robert Hill, 2 April 1899–15 March 1991. Biographical Memoirs of Fellows of the Royal Society, 40:141–170, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1994.0033.

Bennett:2008:CPL

[Ben08] Gary Bennett. Charles Philippe Leblond. 5 February 1910–10 April 2007. Biographical Memoirs of Fellows of the Royal Society, 54:175– 191, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2007.0042.

Bennett:2019:BBJ

[Ben19] Martin A. Bennett FRS FAA. Brice Bosnich. 3 June 1936–13 April 2015. Biographical Memoirs of Fellows of the Royal Society, 67:59– 87, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0019.

Bernal:1963:WTA

[Ber63] John Desmond Bernal. William Thomas Astbury, 1898–1961. Biographical Memoirs of Fellows of the Royal Society, 9:1–35, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1963.0001.

Bergel:1977:AHJ

[Ber77] Franz Bergel. Alexander Haddow, 18 January 1907–21 January 1976. Biographical Memoirs of Fellows of the Royal Society, 23:132– 191, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1977.0007.

Bergersen:2004:EJH

[Ber04] Fraser J. Bergersen. Eric John Hewitt. 27 February 1919–1 January 2002. Biographical Memoirs of Fellows of the Royal Society, 50:147– 155, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2004.0011.

Bergersen:2005:PSN

[Ber05] F. J. Bergersen. Phillip Sadler Nutman. 10 October 1914–4 May 2004. Biographical Memoirs of Fellows of the Royal Society, 51:315– 326, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2005.0020.

Berry:2020:JFN

[Ber20] Michael Berry FRSSir. John Frederick Nye. 26 February 1923–8 January 2019. Biographical Memoirs of Fellows of the Royal Society, 69:425–441, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0002.

Bethe:1968:JRO

[Bet68] Hans Albrecht Bethe. J. Robert Oppenheimer, 1904–1967. Biographical Memoirs of Fellows of the Royal Society, 14:390–416, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769451;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0016.

Beynon:1996:MBN

[Bey96] John Herbert Beynon. Michael Barber, 3 November 1934–8 May 1991. Biographical Memoirs of Fellows of the Royal Society, 42: 2–9, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1996.0001.

Best:1972:DAS

[BF72] Charles Herbert Best and Albert M. Fisher. David Alymer Scott, 1892–1971. Biographical Memoirs of Fellows of the Royal Society, 18:510–524, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1972.
0018.

Boalch:1991:MWP

[BF91] G. T. Boalch and Gordon Elliott Fogg. Mary Winifred Parke, 23 March 1908–17 July 1989. Biographical Memoirs of Fellows of the Royal Society, 37:381–397, November 1991. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1991.0019.

Beech:2024:JDP

[BF24] Peter Beech and Arthur Forer FRSC. Jeremy David Pickett-Heaps. 5 June 1940–11 April 2021. Biographical Memoirs of Fellows of the Royal Society, 77:315–333, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Buhmann:2018:MJD

[BFIT18] M. D. Buhmann, R. Fletcher FRS, A. Iserles, and P. Toint. Michael J. D. Powell. 29 July 1936–19 April 2015. Biographical Memoirs of Fellows of the Royal Society, 64:341–366, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Barker:2019:DJP

[BFO^{+}19] Mary Barker, Caroline H. Fall, Clive Osmond, Cyrus Cooper, Tom P. Fleming, Kent L. Thornburg, and Graham J. Burton. David James Purslove Barker. 29 June 1938–27 August 2013. Biographical Memoirs of Fellows of the Royal Society, 67:29–57, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0021.

Black:1995:SHP

[BG95] Douglas Black and John Archibald Browne Gray. Sir Harold Percival Himsworth, K. C. B., 19 May 1905–1 November 1993. Biographical Memoirs of Fellows of the Royal Society, 41:200–218, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1995.0013.

Bevington:2002:SHW

[BG02] John C. Bevington and Brian G. Gowenlock. Sir Harry Work Melville, K.C.B. 27 April 1908–14 June 2000. Biographical Memoirs of Fellows of the Royal Society, 48:289–308, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0016.

Bodmer:2011:JLM

[BG11] Walter Bodmer FRS and Ann Ganesan. Joshua Lederberg. 23 May 1925–2 February 2008. Biographical Memoirs of Fellows of the Royal Society, 57:229–251, January 2011. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0024.

Brammar:2014:SKM

[BG14] W. J. Brammar and W. B. Gratzer. Sir Kenneth Murray. 30 December 1930–7 April 2013. Biographical Memoirs of Fellows of the Royal Society, 60:331–348, January 2014. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0008.

Batchelor:1988:AEG

[BH88a] George Keith Batchelor and Raymond Hide. Adrian Edmund Gill, 22 February 1937–19 April 1986. Biographical Memoirs of Fellows of the Royal Society, 34:221–258, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0009.

Bethe:1988:PPE

[BH88b] Hans Albrecht Bethe and G. Hildebrandt. Paul Peter Ewald: 23 January 1888–22 August 1985. Biographical Memoirs of Fellows of the Royal Society, 34:133–176, December 1988. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/770049;   https:
//royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.
0006.

Broda:1996:WHJ

[BH96] Paul Broda and Bruce Holloway. William Hayes, 18 January 1913–7 January 1994. Biographical Memoirs of Fellows of the Royal Society, 42:173–189, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0011.

Bondi:1997:RAL

[BH97] Hermann Bondi and Fred Hoyle. Raymond Arthur Lyttleton. 7 May 1911–16 May 1995. Biographical Memoirs of Fellows of the Royal Society, 43:305–319, January 1997. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0017.

Bingham:2008:SHR

[BH08] N. H. Bingham and W. K. Hayman. Sir Harry Raymond Pitt. 3 June 1914–8 October 2005. Biographical Memoirs of Fellows of the Royal Society, 54:257–274, January 2008. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0008.

Beagley:2018:DWJ

[BH18] B. Beagley and J. R. Helliwell. Durward W. J. Cruickshank. 7 March 1924–13 July 2007. Biographical Memoirs of Fellows of the Royal Society, 65:71–87, December 2018. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0018.

Bhagavantam:1971:CVR

[Bha71] S. Bhagavantam. Chandrasekhara Venkata Raman, 1888–1970. Biographical Memoirs of Fellows of the Royal Society, 17:564–592, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0022.

Bhadeshia:2024:GWG

[Bha24] Harshad K. D. H. Bhadeshia FRS. Geoffrey Wilson Greenwood. 3 February 1929–13 December 2022. Biographical Memoirs of Fellows of the Royal Society, 76:221–240, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0022.

Burcham:1996:PBM

[BI96] William Ernest Burcham and G. R. Isaak. Philip Burton Moon, 17 May 1907–9 October 1994. Biographical Memoirs of Fellows of the Royal Society, 42:249–264, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0016.

Bignami:2002:GPS

[Big02] Giovanni Bignami. Giuseppe Paolo Stanislao Occhialini. 5 December 1907–30 December 1993. Biographical Memoirs of Fellows of the Royal Society, 48:331–340, January 2002. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0019.

Biggs:2010:WPJ

[Big10] P. M. Biggs. Walter Plowright. 20 July 1923–20 February 2010. Biographical Memoirs of Fellows of the Royal Society, 56:341–358, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2010.0018.

Bilby:1990:JDE

[Bil90] Bruce Alexander Bilby. John Douglas Eshelby, 21 December 1916–10 December 1981. Biographical Memoirs of Fellows of the Royal Society, 36:125–150, December 1990. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1990.0027.

Birkinshaw:1972:HR

[Bir72] J. H. Birkinshaw. Harold Raistrick, 1890–1971. Biographical Memoirs of Fellows of the Royal Society, 18:488–509, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1972.0017.

Berry:2012:UFJ

[BIR12] R. Stephen Berry, Mitio Inokuti, and A. R. P. Rau. Ugo Fano. 28 July 1912–13 February 2001. Biographical Memoirs of Fellows of the Royal Society, 58:55–66, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0030.

Birgeneau:2017:RAC

[Bir17] Robert J. Birgeneau FRS. Roger Arthur Cowley. 24 February 1939–27 January 2015. Biographical Memoirs of Fellows of the Royal Society, 63:137–157, January 2017. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2017.
0011.

Bishop:1987:ARC

[Bis87] Richard Evelyn Donohue Bishop. Arthur Roderick Collar, 22 February 1908–12 February 1986. Biographical Memoirs of Fellows of the Royal Society, 33:163–185, December 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0006.

Brown:2000:ART

[BK00] Daniel M. Brown and Hans Kornberg. Alexander Robertus Todd, O.M., Baron Todd of Trumpington. 2 October 1907–10 January 1997. Biographical Memoirs of Fellows of the Royal Society, 46:515– 532, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0099.

Burdekin:2017:ESJ

[BK17] Frederick Michael Burdekin OBE FRS FREng and John Frederick Knott OBE FRS FREng. Edwin Smith. 28 July 1931–4 July 2010. Biographical Memoirs of Fellows of the Royal Society, 63:553–565, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2017.0008.

Bliss:2024:PAJ

[BL24] Tim Bliss FRS and Terje Lømo. Per Andersen. 12 January 1930–17 February 2020. Biographical Memoirs of Fellows of the Royal Society, 76:27–47, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0041.

Blackett:1960:CTR

[Bla60a] Patrick Maynard Stuart Blackett. Charles Thomson Rees Wilson, 1869–1959. Biographical Memoirs of Fellows of the Royal Society, 6:269–295, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1960.

Blackett:1960:JFJ

[Bla60b] Patrick Maynard Stuart Blackett. Jean Frédéric Joliot, 1900–1958. Biographical Memoirs of Fellows of the Royal Society, 6:86–105, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1960.0026.

Blackman:1972:GIF

[Bla72] Moses Blackman. George Ingle Finch, 1888–1970. Biographical Memoirs of Fellows of the Royal Society, 18:223–239, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1972.0007.

Blaxter:1977:HDK

[Bla77] Kenneth Lyon Blaxter. Herbert Davenport Kay, 9 September 1893–24 November 1976. Biographical Memoirs of Fellows of the Royal Society, 23:283–310, November 1977. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0011.

Blaxter:1981:EJU

[Bla81] Kenneth Lyon Blaxter. Eric John Underwood, 7 September 1905–19 August 1980. Biographical Memoirs of Fellows of the Royal Society, 27:578–602, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0023.

Blaschko:1985:USE

[Bla85] Hermann Karl Felix Blaschko. Ulf Svante Von Euler, 7 February 1905–10 March 1983. Biographical Memoirs of Fellows of the Royal Society, 31:144–170, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0006.

Black:1994:HOS

[Bla94] James Whyte Black. Heinz Otto Schild, 18 May 1906–15 June 1984. Biographical Memoirs of Fellows of the Royal Society, 39:381–415, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Bleaney:1982:JHV

[Ble82] Brebis Bleaney. John Hasbrouck Van Vleck, 13 March 1899– 27 October 1980. Biographical Memoirs of Fellows of the Royal Society, 28:627–665, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0024.

Bleaney:1996:GWS

[Ble96a] Brebis Bleaney. George William Series, 22 February 1920– 2 January 1995. Biographical Memoirs of Fellows of the Royal Society, 42:387–398, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0024.

Bleaney:1996:HGK

[Ble96b] Brebis Bleaney. Heinrich Gerhard Kuhn, 10 March 1904–26 August 1994. Biographical Memoirs of Fellows of the Royal Society, 42:221– 232, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1996.0014.

Bleaney:1997:ARM

[Ble97] Brebis Bleaney. Allan Roy Mackintosh. 22 January 1936– 20 December 1995. Biographical Memoirs of Fellows of the Royal Society, 43:323–331, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0018.

Bleaney:1999:EMP

[Ble99] Brebis Bleaney. Edward Mills Purcell. 30 August 1912–7 March 1997. Biographical Memoirs of Fellows of the Royal Society, 45:437– 447, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0029.

Bleaney:2001:SMM

[Ble01] Brebis Bleaney. Sir Mark (Marcus Laurence Elwin) Oliphant, A.C., K.B.E. 8 October 1901–14 July 2000. Biographical Memoirs of Fellows of the Royal Society, 47:383–393, January 2001.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0022.

Blow:2004:MFP

[Blo04] D. M. Blow. Max Ferdinand Perutz, OM CH CBE, 19 May 1914–6 February 2002. Biographical Memoirs of Fellows of the Royal Society, 50(3):227–256, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.jstor.

org/stable/4140521;   https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2004.0016.

Brooks:2011:CHM

[BLT11] A. S. Brooks, J. W. G. Lund, and J. F. Talling FRS. Clifford Hiley Mortimer. 27 February 1911–11 May 2010. Biographical Memoirs of Fellows of the Royal Society, 57:291–314, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0006. See correction [BLT12].

Brooks:2012:CCH

[BLT12] A. S. Brooks, J. W. G. Lund FRS, and J. F. Talling FRS. Correction for Clifford Hiley Mortimer. 27 February 1911–11 May 2010. Biographical Memoirs of Fellows of the Royal Society, 58:347, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2012.0035. See [BLT11].

Bishop:1974:MR

[BM74] Ann Bishop and Arnold Ashley Miles. Muriel Robertson, 1883– 1973. Biographical Memoirs of Fellows of the Royal Society, 20:316– 347, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1974.0014.

Barcroft:1981:DHS

[BM81] Henry Barcroft and D. M. Matthews. David Henry Smyth, 9 February 1908–10 September 1979. Biographical Memoirs of Fellows of the Royal Society, 27:524–561, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0021.

Bone:1998:NBM

[BM98] Q. Bone and N. R. Merrett. Norman Bertram Marshall. 5 February 1915–13 February 1996. Biographical Memoirs of Fellows of the Royal Society, 44:281–296, January 1998. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0019.

Bray:2005:GBR

[BM05] Kelvin Bray and Alexander Moulton. Geoffrey Bertram Robert Feilden CBE. 20 February 1917–1 May 2004. Biographical Memoirs of Fellows of the Royal Society, 51:135–148, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0009.

Burland:2013:SAM

[BM13] John B. Burland CBE FRS FREng and Robert J. Mair CBE FRS FREng. Sir Alan Marshall Muir Wood FREng FICE. 8 August 1921–1 February 2009. Biographical Memoirs of Fellows of the Royal Society, 59:261–276, January 2013. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0011.

Bretscher:2017:FHC

[BM17] Mark S. Bretscher FRS and Graeme Mitchison. Francis Harry Compton Crick OM. 8 June 1916–28 July 2004. Biographical Memoirs of Fellows of the Royal Society, 63:159–196, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0010.

Brown:2022:TEA

[BM22] L. M. Brown and I. M. Mills. Trevor Evans. 26 April 1927–10 October 2010. Biographical Memoirs of Fellows of the Royal Society, 72: 161–171, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2021.0033.

Browne:2023:JDH

[BM23] David J. Browne and D. Graham McCartney. John David Hunt. 12 December 1936–8 December 2012. Biographical Memoirs of Fellows of the Royal Society, 75:239–261, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0046.

Brown:2024:JJT

[BM24] Geoff Brown and Bob Michell. John Joseph Thomas Hayes Owen. 7 January 1934–5 January 2023. Biographical Memoirs of Fellows of the Royal Society, 77:301–314, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Brown:1992:EGB

[BMW92] Robert Hanbury Brown, H. C. Minnett, and Frederick William George White. Edward George Bowen, 14 January 1911–12 August 1991. Biographical Memoirs of Fellows of the Royal Society, 38:41–65, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1992.0003.

Bone:1992:JES

[BN92] Quentin Bone and D. Nichols. James Eric Smith, 23 February 1909–3 September 1990. Biographical Memoirs of Fellows of the Royal Society, 38:323–343, November 1992. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0017.

Buckingham:2001:KFO

[BN01] A. D. Buckingham and H. Nakatsuji. Kenichi Fukui. 4 October 1918–9 January 1998. Biographical Memoirs of Fellows of the Royal Society, 47:223–237, January 2001. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0013.

Berry:2006:JMZ

[BN06] Michael Berry and John F. Nye. John Michael Ziman. 16 May 1925–2 January 2005. Biographical Memoirs of Fellows of the Royal Society, 52:479–491, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.

Boyd-Orr:1958:WE

[BO58] John Boyd-Orr. Walter Elliot, 1888–1958. Biographical Memoirs of Fellows of the Royal Society, 4:72–80, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0006.

Bock:2006:EWM

[Boc06] Walter J. Bock. Ernst Walter Mayr. 5 July 1904–3 February 2005. Biographical Memoirs of Fellows of the Royal Society, 52:167–187, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2006.0013.

Boden:2004:SWM

[Bod04] Edward Boden. Sir William Macgregor Henderson. 17 July 1913– 29 November 2000. Biographical Memoirs of Fellows of the Royal Society, 50:133–146, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0010.

Bolt:1997:ILM

[Bol97] Bruce A. Bolt. Inge Lehmann. 13 May 1888–21 February 1993. Biographical Memoirs of Fellows of the Royal Society, 43:287–301, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1997.0016.

Bomford:1967:JGH

[Bom67] G. Bomford. James de Graaff-Hunter, 1881–1967. Biographical Memoirs of Fellows of the Royal Society, 13:78–88, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1967.0004.

Bondi:2006:TGM

[Bon06] Hermann Bondi. Thomas Gold. 22 May 1920–22 June 2004. Biographical Memoirs of Fellows of the Royal Society, 52:117–135, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2006.0009.

Borst:2017:ECS

[Bor17] Piet Borst. Edward Charles Slater. 16 January 1917–26 March 2016. Biographical Memoirs of Fellows of the Royal Society, 63:527– 551, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2016.0024.

Bosnich:2012:AMS

[Bos12] Brice Bosnich FRS. Alan McLeod Sargeson FAA. 13 October 1930–29 December 2008. Biographical Memoirs of Fellows of the Royal Society, 58:265–282, January 2012. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0017.

Boulton:2015:ARK

[Bou15] A. John Boulton. Alan Roy Katritzky. 18 August 1928–10 February 2014. Biographical Memoirs of Fellows of the Royal Society, 61:225– 245, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2015.0001.

Bowen:1958:DLC

[Bow58] Edmund John Bowen. David Leonard Chapman, 1869–1958. Biographical Memoirs of Fellows of the Royal Society, 4:34–44, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1958.0004.

Bowen:1967:DLH

[Bow67] Edmund John Bowen. Dalziel Llewellyn Hammick, 1887–1966. Biographical Memoirs of Fellows of the Royal Society, 13:107–123, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1967.0006.

Bowman:2005:DSF

[Bow05] John C. Bowman. Douglas Scott Falconer. 10 March 1913– 23 February 2004. Biographical Memoirs of Fellows of the Royal Society, 51:119–133, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2005.

Boyd:1956:SFD

[Boy56] John Smith Knox Boyd. Sheldon Francis Dudley, 1884–1956. Biographical Memoirs of Fellows of the Royal Society, 2:85–99, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1956.0006.

Boyd:1963:LR

[Boy63] John Smith Knox Boyd. Leonard Rogers, 1868–1962. Biographical Memoirs of Fellows of the Royal Society, 9:261–285, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1963.0014.

Boyd:1966:NHF

[Boy66] John Smith Knox Boyd. Neil Hamilton Fairley, 1891–1966. Biographical Memoirs of Fellows of the Royal Society, 12:123–145, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0005.

Boycott:1998:JZY

[Boy98] Brian B. Boycott. John Zachary Young. 18 March 1907–4 July 1997. Biographical Memoirs of Fellows of the Royal Society, 44:487– 509, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1998.0031.

Boyd:2023:JHB

[Boy23] Robert Boyd. John Herbert Beynon. 29 December 1923–24 August 2015. Biographical Memoirs of Fellows of the Royal Society, 75:53– 74, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2023.0007.

Burke:1999:JSB

[BP99] Philip G. Burke and Ian C. Percival. John Stewart Bell. 28 July 1928–1 October 1990. Biographical Memoirs of Fellows of the Royal Society, 45:1–17, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.

Binnie:1971:THH

[BR71] Alfred Maurice Binnie and Paul Harry Roberts. Thomas Henry Havelock, 1877–1968. Biographical Memoirs of Fellows of the Royal Society, 17:327–377, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1971.
0013.

Barr:1973:JBC

[BR73] Murray Llewellyn Barr and R. J. Rossiter. James Bertram Collip, 1892–1965. Biographical Memoirs of Fellows of the Royal Society, 19:234–267, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1973.
0009.

Baker:1979:TRS

[BR79] Wilson Baker and S. Rangaswami. Thiruvenkata Rajendra Seshadri, 3 February 1900–27 September 1975. Biographical Memoirs of Fellows of the Royal Society, 25:504–533, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0018.

Bray:2014:JFC

[BR14] K. N. C. Bray FRS and N. Riley. John Frederick Clarke. 1 May 1927–11 June 2013. Biographical Memoirs of Fellows of the Royal Society, 60:87–106, January 2014. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0012.

Berry:2018:RNP

[BR18] Alan Berry and Sheena E. Radford FRS. Richard Nelson Perham. 27 April 1937–14 February 2015. Biographical Memoirs of Fellows of the Royal Society, 65:317–339, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0004.

Bragg:1965:RWJ

[Bra65] William Lawrence Bragg. Reginald William James, 1891–1964. Biographical Memoirs of Fellows of the Royal Society, 11:114–125, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1965.0007.

Brambell:1970:ASG

[Bra70] Francis William Rogers Brambell. Arthur St George Joseph McCarthy Huggett, 1897–1968. Biographical Memoirs of Fellows of the Royal Society, 16:342–364, November 1970. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0013.

Bradshaw:1992:TGT

[Bra92] Anthony David Bradshaw. Thomas Gaskell Tutin, 21 April 1908–7 October 1987. Biographical Memoirs of Fellows of the Royal Society, 38:359–375, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1992.
0019.

Brady:1995:JSK

[Bra95] John Brady. John Stodart Kennedy, 15 May 1912–4 February 1993. Biographical Memoirs of Fellows of the Royal Society, 41:244–260, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0015.

Bradley:2012:GDL

[Bra12] Derek Bradley FRS FREng. Graham Dixon-Lewis. 1 July 1922–5 August 2010. Biographical Memoirs of Fellows of the Royal Society, 58:33–53, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2011.
0024.

Brammar:2015:PHC

[Bra15] W. J. Brammar. Patricia Hannah Clarke. 29 July 1919–28 January 2010. Biographical Memoirs of Fellows of the Royal Society, 61: 39–51, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2015.0012.

Bradley:2022:DDA

[Bra22] Derek Bradley FRS. Duncan Dowson. 31 August 1928–6 January 2020. Biographical Memoirs of Fellows of the Royal Society, 73:177– 196, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2021.0036.

Barton:1968:RPL

[BRE68] Derek Harold Richard Barton, H. N. Rydon, and J. A. Elvidge. Reginald Patrick Linstead, 1902–1966. Biographical Memoirs of Fellows of the Royal Society, 14:308–347, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0014.

Brent:2005:REB

[Bre05] Leslie Baruch Brent. Rupert Everett Billingham. 15 October 1921–16 November 2002. Biographical Memoirs of Fellows of the Royal Society, 51:33–50, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2005.
0003.

Breckenridge:2014:SDJ

[Bre14] Sir Alasdair Breckenridge CBE FRCP FRSE FMedSci. Sir David Jack. 22 February 1924–8 November 2011. Biographical Memoirs of Fellows of the Royal Society, 60:249–260, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0006.

Bridges:1958:JAV

[Bri58] Edward Ettingdean Bridges. John Anderson, Viscount Waverley, 1882–1958. Biographical Memoirs of Fellows of the Royal Society, 4:306–325, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1958.
0024.

Bridges:1968:CRA

[Bri68] Edward Ettingdean Bridges. Clement Richard Attlee, First Earl Attlee, 1883–1967. Biographical Memoirs of Fellows of the Royal Society, 14:15–36, November 1968. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1968.
0002.

Broadbent:1970:FCB

[Bro70] D. E. Broadbent. Frederic Charles Bartlett, 1886–1969. Biographical Memoirs of Fellows of the Royal Society, 16:1–13, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0001.

Brown:1977:TAB

[Bro77] Robert Brown. Thomas Archibald Bennet-Clark, 13 January 1903–24 November 1975. Biographical Memoirs of Fellows of the Royal Society, 23:1–18, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0001.

Broadbent:1980:MBM

[Bro80] Edward Granville Broadbent. Morien Bedford Morgan, 20 December 1912–4 April 1978. Biographical Memoirs of Fellows of the Royal Society, 26:371–410, November 1980. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1980.0012.

Brown:2009:VLT

[Bro09] Laurie M. Brown. Valentine Louis Telegdi. 11 January 1922–8 April 2006. Biographical Memoirs of Fellows of the Royal Society, 55:291– 304, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2009.0012.

Brown:2010:FRN

[Bro10] L. M. Brown. Frank Reginald Nunes Nabarro MBE. 7 March 1916–20 July 2006. Biographical Memoirs of Fellows of the Royal Society, 56:273–283, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.

Brown:2013:SGH

[Bro13] Malcolm W. Brown FRS. Sir Gabriel Horn. 9 December 1927–2 August 2012. Biographical Memoirs of Fellows of the Royal Society, 59:157–170, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2013.
0016.

Brownlee:2015:FSC

[Bro15] George G. Brownlee FRS. Frederick Sanger CBE CH OM. 13 August 1918–19 November 2013. Biographical Memoirs of Fellows of the Royal Society, 61:437–466, January 2015. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0013.

Browning:2016:SBJ

[Bro16] Keith A. Browning FRS. Sir (Basil) John Mason CB. 18 August 1923–6 January 2015. Biographical Memoirs of Fellows of the Royal Society, 62:359–380, January 2016. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0028.

Brown:2019:RCD

[Bro19] John M. Brown FRS. Ronald Charles David Breslow. 14 March 1931–25 October 2017. Biographical Memoirs of Fellows of the Royal Society, 66:53–77, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0039.

Brookfield:2023:BCC

[Bro23] John Brookfield. Bryan Campbell Clarke. 24 June 1932–27 February 2014. Biographical Memoirs of Fellows of the Royal Society, 74: 109–121, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2022.0042.

Brown:2024:MJS

[Bro24] L. M. Brown FRS. Michael James Stowell. 10 July 1935–27 February 2022. Biographical Memoirs of Fellows of the Royal Society, 76: 441–454, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2023.0035.

Bryan:2001:MRH

[Bry01] E. R. Bryan. Michael Rex Horne, O.B.E. 29 December 1921– 6 January 2000. Biographical Memoirs of Fellows of the Royal Society, 47:279–292, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0016.

Bruins:1967:FAV

[BS67] G. J. Bruins and J. G. J. Scholte. Felix Andries Vening Meinesz, 1887–1966. Biographical Memoirs of Fellows of the Royal Society, 13:294–308, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1967.
0015.

Blin-Stoyle:1990:MAG

[BS90] Roger John Blin-Stoyle. Michael Anthony Grace, 13 May 1920–17 May 1988. Biographical Memoirs of Fellows of the Royal Society, 35:127–149, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Blackburn:2001:WDO

[BS01] G. M. Blackburn and I. O. Sutherland. William David Ollis. 22 December 1924–13 June 1999. Biographical Memoirs of Fellows of the Royal Society, 47:395–413, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0023.

Barnett:2004:WTW

[BS04] Michael Barnett and Robin Smith. Walter Thompson Welford. 31 August 1916–18 September 1990. Biographical Memoirs of Fellows of the Royal Society, 50:315–331, January 2004. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2004.0021.

Bradshaw:2005:GLS

[BS05] Anthony D. Bradshaw and Vassiliki B. Smocovitis. George Ledyard Stebbins. 6 January 1906–19 January 2000. Biographical Memoirs of Fellows of the Royal Society, 51:397–408, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0026.

Baer:2008:REM

[BS08] Tomas Baer and Giacinto Scoles. Roger Ervin Miller. 23 July 1952–6 November 2005. Biographical Memoirs of Fellows of the Royal Society, 54:215–228, January 2008. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0007.

Burke:2013:MJS

[BS13] Philip G. Burke CBE FRS and Peter J. Storey. Michael John Seaton. 16 January 1923–29 May 2007. Biographical Memoirs of Fellows of the Royal Society, 59:317–336, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0044.

Bergman:2014:PMC

[BS14] George Bergman and Trevor Stuart FRS. Paul Moritz Cohn. 8 January 1924–20 April 2006. Biographical Memoirs of Fellows of the Royal Society, 60:127–150, January 2014. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0016.

Buckingham:2016:ACC

[BS16] A. D. Buckingham CBE FRS and R. J. Saykally. Alan Carrington CBE. 6 January 1934–31 August 2013. Biographical Memoirs of Fellows of the Royal Society, 62:7–17, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0024.

Seckl:2024:KDB

[BS24] Richard Begent and Michael Seckl. Kenneth Dawson Bagshawe. 17 August 1925–27 December 2022. Biographical Memoirs of Fellows of the Royal Society, 76:49–67, June 2024. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0050.

Brockwell:2013:FJB

[BSD13] John Brockwell, Janet I. Sprent, and David A. Day. Fraser John Bergersen AM FAA. 26 May 1929–3 October 2011. Biographical Memoirs of Fellows of the Royal Society, 59:33–58, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0004.

Bateson:2018:RAH

[BSHC18] Sir Patrick Bateson FRS, Joan Stevenson-Hinde, and Tim Clutton-Brock FRS. Robert Aubrey Hinde CBE. 26 October 1923– 23 December 2016. Biographical Memoirs of Fellows of the Royal Society, 65:151–177, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0011.

Bullen:1968:TMC

[BT68] Keith Edward Bullen and George Frederick James Temple. Thomas MacFarland Cherry, 1898–1966. Biographical Memoirs of Fellows of the Royal Society, 14:117–138, November 1968. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0006.

Birch:2005:AFM

[BT05] Bryan J. Birch and Martin J. Taylor. Albrecht Fröhlich. 22 May 1916–8 November 2001. Biographical Memoirs of Fellows of the Royal Society, 51:149–168, January 2005. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0010.

Burton:2020:LEF

[BT20] G. R. Burton and J. F. Toland FRS. Ludwig Edward Fraenkel. 28 May 1927–27 April 2019. Biographical Memoirs of Fellows of the Royal Society, 69:175–201, December 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0014.

Buckingham:1977:JWL

[Buc77] Amyand David Buckingham. John Wilfrid Linnett, 3 August 1913–7 November 1975. Biographical Memoirs of Fellows of the Royal Society, 23:311–343, November 1977. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0012.

Buckingham:2006:SJA

[Buc06] A. D. Buckingham. Sir John Anthony Pople. 31 October 1925–15 March 2004. Biographical Memoirs of Fellows of the Royal Society, 52:299–314, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0021.

Buchanan:2010:SJB

[Buc10] J. Grant Buchanan. Sir James Baddiley. 15 May 1918–19 November 2008. Biographical Memoirs of Fellows of the Royal Society, 56:3– 23, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2010.0010.

Budden:1988:JAR

[Bud88] Kenneth George Budden. John Ashworth Ratcliffe, 12 December 1902–25 October 1987. Biographical Memoirs of Fellows of the Royal Society, 34:669–711, December 1988. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0022.

Bullard:1957:GPL

[Bul57] Edward Crisp Bullard. Gerald Ponsonby Lenox-Conyngham, 1866– 1956. Biographical Memoirs of Fellows of the Royal Society, 3:129– 140, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1957.0009.

Bulman:1966:WFW

[Bul66] Oliver Meredith Boone Bulman. Walter Frederick Whittard, 1902– 1966. Biographical Memoirs of Fellows of the Royal Society, 12:531– 542, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1966.0027.

Bullard:1967:MNH

[Bul67] Edward Crisp Bullard. Maurice Neville Hill, 1919–1966. Biographical Memoirs of Fellows of the Royal Society, 13:192–203, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1967.0009.

Bulman:1969:WGF

[Bul69] Oliver Meredith Boone Bulman. William George Fearnsides, 1879– 1968. Biographical Memoirs of Fellows of the Royal Society, 15:83– 98, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1969.0004.

Bullard:1975:WME

[Bul75] Edward Crisp Bullard. William Maurice Ewing, 12 May 1906–4 May 1974. Biographical Memoirs of Fellows of the Royal Society, 21:268–311, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1975.
0007.

Bullock:1994:IWB

[Bul94] Alan Bullock. Isaac Wolfson, BT., 17 September 1897–20 June 1991. Biographical Memoirs of Fellows of the Royal Society, 40:421– 427, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1994.0048.

Bunting:1994:GWC

[Bun94] A. H. Bunting. George William Cooke, 6 January 1916–10 February 1992. Biographical Memoirs of Fellows of the Royal Society, 39:91– 109, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Bundle:2002:RUL

[Bun02] David R. Bundle. Raymond Urgel Lemieux. 16 June 1920–22 July 2000. Biographical Memoirs of Fellows of the Royal Society, 48:251– 273, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2002.0014.

Burkill:1968:AEI

[Bur68] John Charles Burkill. Albert Edward Ingham, 1900–1967. Biographical Memoirs of Fellows of the Royal Society, 14:271–286, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1968.0012.

Burkill:1971:ASB

[Bur71] John Charles Burkill. Abram Samoilovitch Besicovitch, 1891–1970. Biographical Memoirs of Fellows of the Royal Society, 17:1–16, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0001.

Burkill:1978:JEL

[Bur78] John Charles Burkill. John Edensor Littlewood, 9 June 1885– 6 September 1977. Biographical Memoirs of Fellows of the Royal Society, 24:323–367, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1978.
0010.

Burgen:1994:FCM

[Bur94] Arnold Stanley Vincent Burgen. Frank Campbell MacIntosh, 24 December 1909–11 September 1992. Biographical Memoirs of Fellows of the Royal Society, 40:253–264, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1994.0038.

Burdekin:2001:STA

[Bur01] F. M. Burdekin. Sir (Thomas) Angus Lyall Paton, C.M.G. 10 May 1905–7 April 1999. Biographical Memoirs of Fellows of the Royal Society, 47:415–427, January 2001. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0024.

Burbidge:2003:SFH

[Bur03] Geoffrey Burbidge FRS. Sir Fred Hoyle. 24 June 1915–20 August 2001 Elected FRS 1957. Biographical Memoirs of Fellows of the Royal Society, 49:213–247, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0013.

Burdekin:2019:SBC

[Bur19] Michael Burdekin OBE, FREng, FRS. Sir Bernard Crossland CBE. 20 October 1923–17 January 2011. Biographical Memoirs of Fellows of the Royal Society, 66:123–142, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0032.

Burns:2021:DJG

[Bur21] Ian G. Burns. Duncan Joseph Greenwood. 16 October 1932– 13 February 2010. Biographical Memoirs of Fellows of the Royal Society, 70:189–209, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0037.

Busemann:1960:LP

[Bus60] A. Busemann. Ludwig Prandtl, 1875–1953. Biographical Memoirs of Fellows of the Royal Society, 5:193–205, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Butterworth:2001:SCC

[But01] Ian Butterworth. Sir Clifford Charles Butler. 20 May 1922–30 June 1999. Biographical Memoirs of Fellows of the Royal Society, 47:39– 54, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2001.0003.

Bulbring:1984:JHB

[BW84] Edith Bülbring and J. M. Walker. Joshua Harold Burn, 6 March 1892–13 July 1981. Biographical Memoirs of Fellows of the Royal Society, 30:45–89, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1984.
0002.

Born:1990:RGM

[BW90] Gustav Victor Rudolf Born and David John Weatherall. Robert Gwyn MacFarlane, 26 June 1907–26 March 1987. Biographical Memoirs of Fellows of the Royal Society, 35:209–245, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Briscoe:2022:TMJ

[BW22] James Briscoe and John N. Wood. Thomas Michael Jessell. 2 August 1951–28 April 2019. Biographical Memoirs of Fellows of the Royal Society, 72:197–219, June 2022. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0037.

Battersby:2016:SJW

[BY16] Sir Alan R. Battersby FRS and Douglas W. Young FRSE. Sir John Warcup Cornforth AC CBE. 7 September 1917–8 December 2013. Biographical Memoirs of Fellows of the Royal Society, 62: 19–57, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2015.0016.

Byalko:2022:IMK

[Bya22] Alexei V. Byalko. Isaak Markovich Khalatnikov. 17 October 1919–9 January 2021. Biographical Memoirs of Fellows of the Royal Society, 72:251–273, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0047.

Butler:1967:JH

[BZ67] John Alfred Valentine Butler and P. Zuman. Jaroslav Heyrovský, 1890–1967. Biographical Memoirs of Fellows of the Royal Society, 13:167–191, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1967.
0008.

Crookshank:1956:LLF

[CA56] H. Crookshank and J. B. Auden. Lewis Leigh Fermor, 1880–1954. Biographical Memoirs of Fellows of the Royal Society, 2:100–116, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1956.0007.

Curtis:2001:SJC

[CA01] David R. Curtis and Per Andersen. Sir John Carew Eccles, A.C. 27 January 1903–2 May 1997. Biographical Memoirs of Fellows of the Royal Society, 47:159–187, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0010.

Calladine:2021:EHM

[Cal21] C. R. Calladine. Eric Harold Mansfield. 24 May 1923–20 October 2016. Biographical Memoirs of Fellows of the Royal Society, 71:335– 359, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2020.0049.

Cameron:1960:RM

[Cam60] Gordon Roy Cameron. Robert Muir, 1864–1959. Biographical Memoirs of Fellows of the Royal Society, 5:148–173, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Campbell:1994:EHL

[Cam94] Fergus William Campbell. Edwin Herbert Land, 7 May 1909– 1 March 1991. Biographical Memoirs of Fellows of the Royal Society, 40:197–219, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.
0035.

Cann:2014:ACW

[Can14] Rebecca L. Cann. Allan Charles Wilson. 18 October 1934–21 July 1991. Biographical Memoirs of Fellows of the Royal Society, 60:455– 473, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2013.0006.

Caplin:1999:BRC

[Cap99] David Caplin. Bryan Randell Coles. 9 June 1926–24 February 1997. Biographical Memoirs of Fellows of the Royal Society, 45:51–66, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1999.0005.

Cartwright:1964:ECT

[Car64] Mary Lucy Cartwright. Edward Charles Titchmarsh, 1899–1963. Biographical Memoirs of Fellows of the Royal Society, 10:305–324, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0018.

Cartwright:1965:JH

[Car65] Mary Lucy Cartwright. Jacques Hadamard, 1865–1963. Biographical Memoirs of Fellows of the Royal Society, 11:75–99, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1965.0005.

Cartwright:1999:HCC

[Car99] D. E. Cartwright. Henry Charnock, C.B.E. 25 December 1920– 27 November 1997. Biographical Memoirs of Fellows of the Royal Society, 45:35–50, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0004.

Carter:2014:GBM

[Car14] Philip B. Carter. George Bellamy Mackaness. 20 August 1922–4 March 2007. Biographical Memoirs of Fellows of the Royal Society, 60:277–298, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2014.
0017.

Cassels:1973:LJM

[Cas73] John William Scott Cassels. Louis Joel Mordell, 1888–1972. Biographical Memoirs of Fellows of the Royal Society, 19:493–520, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0018.

Cawley:2024:DJE

[Caw24] Peter Cawley FRS. David J. Ewins. 25 March 1942–25 January 2023. Biographical Memoirs of Fellows of the Royal Society, 77: 163–183, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Collar:1981:GHR

[CB81] Arthur Roderick Collar and A. R. W. Broadway. Gordon Hindle Rawcliffe, 2 June 1910–3 September 1979. Biographical Memoirs of Fellows of the Royal Society, 27:479–503, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0019.

Chorin:2022:GIB

[CB22] Alexandre J. Chorin and Zdeněk P. Bažant ForMemRS. Grigory Isaakovich Barenblatt. 10 July 1927–22 June 2018. Biographical Memoirs of Fellows of the Royal Society, 72:33–53, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0024.

Smallman:2009:SRW

[CBE09] R. E. Smallman CBE FRS. Sir Robert William Kerr Honeycombe KBE. 2 May 1921–14 September 2007. Biographical Memoirs of Fellows of the Royal Society, 55:91–106, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0020. See correction [Sma10].

Croxall:2015:RML

[CBPC15] John Croxall CBE FRS, Ian Boyd FRSE, Ian Parker, and Geoffrey Cook. Richard Maitland Laws CBE. 23 April 1926–7 October 2014. Biographical Memoirs of Fellows of the Royal Society, 61:265–304, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2015.0006.

Clark:1961:JSB

[CC61] Wilfrid Edward Le Gros Clark and W. Mansfield Cooper. John Sebastian Bach Stopford, Baron Stopford of Fallowfield, 1888–1961. Biographical Memoirs of Fellows of the Royal Society, 7:271–279, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0021.

Chory:2010:CJL

[CC10] Joanne Chory and Rod Casey. Christopher John Lamb CBE. 19 March 1950–21 August 2009. Biographical Memoirs of Fellows of the Royal Society, 56:189–213, January 2010. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0014.

Clarke:2000:HGT

[CCG^{+}00] W. B. Clarke, J. H. Crocket, R. J. Gillespie, H. R. Krouse, D. M. Shaw, and H. P. Schwarcz. Henry George Thode, M.B.E. 10 September 1910–22 March 1997. Biographical Memoirs of Fellows of the Royal Society, 46:499–514, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0098.

Cochran:1981:NFN

[CD81] William Cochran and Samuel Devons. Norman Feather, 16 November 1904–14 August 1978. Biographical Memoirs of Fellows of the Royal Society, 27:254–282, November 1981. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0011.

Cadogan:1988:DHH

[CD88] John Ivan George Cadogan and D. I. Davies. Donald Holroyde Hey, 12 September 1904–21 January 1987. Biographical Memoirs of Fellows of the Royal Society, 34:293–320, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0011.

Court:2012:JRH

[CDS12] Geoffrey Court, John Dainton FRS, and Terry Sloan. John Riley Holt. 15 February 1918–6 January 2009. Biographical Memoirs of Fellows of the Royal Society, 58:113–128, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0029.

Cushing:1996:RJH

[CE96] David Henry Cushing and Ronald Walter Edwards. Raymond John Heaphy Beverton, C. B. E., 29 August 1922–23 July 1995. Biographical Memoirs of Fellows of the Royal Society, 42:24–38, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1996.0003.

Carr:2019:SWH

[CEG^{+}19] Bernard J. Carr, George F. R. Ellis FRS, Gary W. Gibbons FRS, James B. Hartle, Thomas Hertog, Roger Penrose FRS, Malcolm J. Perry, and Kip S. Thorne. Stephen William Hawking CH CBE. 8 January 1942–14 March 2018. Biographical Memoirs of Fellows of the Royal Society, 66:267–308, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0001.

Chadwick:2015:AJM

[CEP15] P. Chadwick FRS, A. H. England, and D. F. Parker. Anthony James Merrill Spencer. 23 August 1929–26 January 2008. Biographical Memoirs of Fellows of the Royal Society, 61:505–529, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0003.

Cervero:2019:AIA

[Cer19] Fernando Cervero DSc. Ainsley Iggo. 2 August 1924–25 March 2012. Biographical Memoirs of Fellows of the Royal Society, 67:217– 236, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0023.

Cresser:2018:TSW

[CEW18] Malcolm S. Cresser, Anthony C. Edwards, and Thomas Lawson West. Thomas Summers West. 18 November 1927–9 January 2010. Biographical Memoirs of Fellows of the Royal Society, 65:449–468, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2018.0016.

Cassels:1976:HAH

[CF76] John William Scott Cassels and Albrecht Fröhlich. Hans Arnold Heilbronn, 8 October 1908–28 April 1975. Biographical Memoirs of Fellows of the Royal Society, 22:119–135, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0005.

Corrigan:2020:DIO

[CG20] Edward Corrigan FRS and Peter Goddard CBE FRS. David Ian Olive. 16 April 1937–7 November 2012. Biographical Memoirs of Fellows of the Royal Society, 69:467–498, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0024.

Cox:2021:KVM

[CG21] Francis Cox and Keith Gull CBE FRS. Keith Vickerman. 21 March 1933–28 June 2016. Biographical Memoirs of Fellows of the Royal Society, 71:491–511, December 2021. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0019.

Catalan:2022:JFS

[CGZ22] Gustau Catalan, J. Marty Gregg, and Pavlo Zubko. James Floyd Scott. 4 May 1942–6 April 2020. Biographical Memoirs of Fellows of the Royal Society, 73:395–420, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0048.

Cox:1958:RWG

[CH58] Ernest Gordon Cox and J. Hume. Robert Whytlaw-Gray, 1877– 1958. Biographical Memoirs of Fellows of the Royal Society, 4:327– 339, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1958.0025.

Cartwright:1971:EFC

[CH71] Mary Lucy Cartwright and Walter Kurt Hayman. Edward Foyle Collingwood, 1900–1970. Biographical Memoirs of Fellows of the Royal Society, 17:139–158, November 1971. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1971.0005.

Clapham:1979:WOJ

[CH79] Arthur Roy Clapham and John Laker Harley. William Owen James, 1900–1978. Biographical Memoirs of Fellows of the Royal Society, 25:336–364, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1979.
0012.

Charlesworth:2005:JMS

[CH05] Brian Charlesworth and Paul Harvey. John Maynard Smith. 6 January 1920–19 April 2004. Biographical Memoirs of Fellows of the Royal Society, 51:253–265, January 2005. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0016.

Crowther:2019:JTF

[CH19] R. A. Crowther FRS FMEDSCI and K. C. Holmes FRS. John Thomas Finch. 28 February 1930–5 December 2017. Biographical Memoirs of Fellows of the Royal Society, 66:183–199, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0028.

Chaundy:1955:ALD

[Cha55] T. W. Chaundy. Arthur Lee Dixon, 1867–1955. Biographical Memoirs of Fellows of the Royal Society, 1:33–36, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0004.

Chapman:1958:FCM

[Cha58] Sydney Chapman. Fredrik Carl Mülertz Störmer, 1874–1957. Biographical Memoirs of Fellows of the Royal Society, 4:257–279, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1958.0021.

Chadwick:1961:FJM

[Cha61] James Chadwick. Frederick John Marrian Stratton, 1881–1960. Biographical Memoirs of Fellows of the Royal Society, 7:280–293, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0022.

Chaloner:1985:TMH

[Cha85a] William Gilbert Chaloner. Thomas Maxwell Harris, 8 January 1903–1 May 1983. Biographical Memoirs of Fellows of the Royal Society, 31:228–260, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0009.

Charnock:1985:GER

[Cha85b] Henry Charnock. George Edward Raven Deacon, 21 March 1906–16 November 1984. Biographical Memoirs of Fellows of the Royal Society, 31:112–142, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0005.

Charig:1990:FRP

[Cha90] A. J. Charig. Francis Rex Parrington, 20 February 1905–17 April 1981. Biographical Memoirs of Fellows of the Royal Society, 36:359– 378, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1990.0036.

Charnock:1997:JCS

[Cha97] Henry Charnock. John Crossley Swallow. 11 October 1923– 3 December 1994. Biographical Memoirs of Fellows of the Royal Society, 43:505–519, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0028.

Chadwick:2001:AEG

[Cha01] P. Chadwick. Albert Edward Green. 11 November 1912–12 August 1999. Biographical Memoirs of Fellows of the Royal Society, 47:255– 278, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2001.0015.

Chadwick:2002:INS

[Cha02] P. Chadwick F.R.S. Ian Naismith Sneddon, O.B.E. 8 December 1919–4 November 2000. Biographical Memoirs of Fellows of the Royal Society, 48:417–437, January 2002. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0025.

Chandler:2003:SAW

[Cha03] Richard J. Chandler. Sir Alec Westley Skempton. 4 June 1914–9 August 2001. Biographical Memoirs of Fellows of the Royal Society, 49:509–519, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2003.
0030.

Cherry:1962:AGM

[Che62] Thomas Macfarland Cherry. Anthony George Maldon Michell, 1870–1959. Biographical Memoirs of Fellows of the Royal Society, 8:90–103, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1962.
0007.

Chick:1956:CJM

[Chi56] Harriette Chick. Charles James Martin, 1866–1955. Biographical Memoirs of Fellows of the Royal Society, 2:172–208, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1956.0013.

Chibnall:1964:KB

[Chi64] Albert Charles Chibnall. Kenneth Bailey, 1909–1963. Biographical Memoirs of Fellows of the Royal Society, 10:1–13, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1964.0001.

ChilverofCranfield:1999:SAG

[Chi99] Lord Chilver of Cranfield. Sir Alfred Grenville Pugsley, O.B.E., F.R.Eng. 13 May 1903–7 March 1998. Biographical Memoirs of Fellows of the Royal Society, 45:417–435, January 1999.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0028.

Christophers:1956:JAS

[Chr56] Samuel Rickard Christophers. John Alexander Sinton, 1884–1956. Biographical Memoirs of Fellows of the Royal Society, 2:268–290, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1956.0019.

Christopherson:1972:RVS

[Chr72] Derman Guy Christopherson. Richard Vynne Southwell, 1888– 1970. Biographical Memoirs of Fellows of the Royal Society, 18:549– 565, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1972.0020.

Carrington:1974:NKA

[CHW74] Alan Carrington, G. J. Hills, and K. R. Webb. Neil Kensington Adam, 1891–1973. Biographical Memoirs of Fellows of the Royal Society, 20:1–26, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1974.
0001.

Ciarlet:2002:JLL

[Cia02] Philippe G. Ciarlet. Jacques-Louis Lions. 2 May 1928–17 May 2001. Biographical Memoirs of Fellows of the Royal Society, 48:275–287, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2002.0015.

Campbell:1999:DHC

[CJ99] K. S. W. Campbell and J. S. Jell. Dorothy Hill, A.C., C.B.E. 10 September 1907–23 April 1997. Biographical Memoirs of Fellows of the Royal Society, 45:195–217, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0014.

Cox:2001:RPB

[CJ01] B. G. Cox and J. H. Jones. Ronald Percy Bell. 24 November 1907–9 January 1996. Biographical Memoirs of Fellows of the Royal Society, 47:19–38, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0002.

Casselton:2012:DLD

[CJ12] Lorna A. Casselton FRS and David A. Jones. Dan Lewis. 30 December 1910–30 September 2009. Biographical Memoirs of Fellows of the Royal Society, 58:163–178, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0023.

Clarke:2023:ALD

[CJ23] Ann Clarke and Martin H. Johnson FRS. Anne Laura Dorinthea McLaren DBE. 26 April 1927–7 July 2007. Biographical Memoirs of Fellows of the Royal Society, 74:309–337, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0053.

Campbell:1980:JPK

[CK80] N. Campbell and Charles Kemball. James Pickering Kendall, 30 July 1889–14 June 1978. Biographical Memoirs of Fellows of the Royal Society, 26:255–273, November 1980. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1980.0007.

Campbell-Kelly:2006:DJW

[CK06] Martin Campbell-Kelly. David John Wheeler. 9 February 1927– 13 December 2004. Biographical Memoirs of Fellows of the Royal Society, 52:437–453, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0030.

Campbell-Kelly:2014:SMV

[CK14] Martin Campbell-Kelly. Sir Maurice Vincent Wilkes. 26 June 1913–29 November 2010. Biographical Memoirs of Fellows of the Royal Society, 60:433–454, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2013.

Clary:2015:NCH

[CKT15] David C. Clary FRS, Peter J. Knowles, and David J. Tozer. Nicholas Charles Handy. 17 June 1941–2 October 2012. Biographical Memoirs of Fellows of the Royal Society, 61:145–160, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0002.

Crangle:1982:WSS

[CL82] J. Crangle and Henry Solomon Lipson. Willie Sucksmith, 21 September 1896–16 September 1981. Biographical Memoirs of Fellows of the Royal Society, 28:575–587, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0022.

Chisholm:2008:FAC

[CL08] Malcolm H. Chisholm and Lord Lewis of Newnham. Frank Albert Cotton. 9 April 1930–20 February 2007. Biographical Memoirs of Fellows of the Royal Society, 54:95–115, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0003.

Crowther:2014:UWA

[CL14] R. A. Crowther FRS FMedSci and A. G. W. Leslie FRS. Ulrich Wolfgang Arndt. 23 April 1924–24 March 2006. Biographical Memoirs of Fellows of the Royal Society, 60:39–55, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0003.

Clark:1955:AK

[Cla55a] Wilfrid Edward Le Gros Clark. Arthur Keith, 1866–1955. Biographical Memoirs of Fellows of the Royal Society, 1:145–161, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1955.0011.

Clark:1955:FWJ

[Cla55b] Wilfrid Edward Le Gros Clark. Frederic Wood Jones, 1879–1954. Biographical Memoirs of Fellows of the Royal Society, 1:119–134, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1955.0009.

Clark:1961:CEK

[Cla61] Walter Clark. Charles Edward Kenneth Mees, 1882–1960. Biographical Memoirs of Fellows of the Royal Society, 7:172–197, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0014.

Clapham:1970:GRS

[Cla70] Arthur Roy Clapham. George Robert Sabine Snow, 1897–1969. Biographical Memoirs of Fellows of the Royal Society, 16:498–522, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1970.0021.

Clapham:1971:WHP

[Cla71] Arthur Roy Clapham. William Harold Pearsall, 1891–1964. Biographical Memoirs of Fellows of the Royal Society, 17:511–540, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0020.

Clarke:1977:PMS

[Cla77] Cyril Clarke. Philip Macdonald Sheppard, 27 July 1921–17 October 1976. Biographical Memoirs of Fellows of the Royal Society, 23:464– 500, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1977.0018.

Clapham:1980:EJS

[Cla80] Arthur Roy Clapham. Edward James Salisbury, 16 April 1886– 10 November 1978. Biographical Memoirs of Fellows of the Royal Society, 26:502–541, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1980.
0014.

Clarke:1985:RRR

[Cla85] Cyril Clarke. Robert Russel Race, 28 November 1907–15 April 1984. Biographical Memoirs of Fellows of the Royal Society, 31:454– 492, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1985.0016.

Clarke:1986:RYS

[Cla86] Patricia Hannah Clarke. Roger Yate Stanier, 22 October 1916– 29 January 1982. Biographical Memoirs of Fellows of the Royal Society, 32:541–568, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0018.

Clarke:1995:EBF

[Cla95] Bryan Campbell Clarke. Edmund Brisco Ford, 23 April 1901– 21 January 1988. Biographical Memoirs of Fellows of the Royal Society, 41:146–168, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1995.
0010.

Clarke:2000:MDL

[Cla00] Patricia H. Clarke. Malcolm Douglas Lilly. 9 August 1936–18 May 1998. Biographical Memoirs of Fellows of the Royal Society, 46:317– 332, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0087.

Clarke:2008:AJC

[Cla08] Bryan C. Clarke. Arthur James Cain. 25 July 1921–20 August 1999. Biographical Memoirs of Fellows of the Royal Society, 54: 47–57, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2007.0034.

Clarke:2009:SEJ

[Cla09] Malcolm R. Clarke FRS. Sir Eric James Denton CBE. 30 September 1923–2 January 2007. Biographical Memoirs of Fellows of the Royal Society, 55:59–76, January 2009. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0001.

Clark:2021:CJJ

[Cla21] Norman G. Clark. Calestous Juma. 9 June 1953–15 December 2017. Biographical Memoirs of Fellows of the Royal Society, 71:293–315, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2021.0009.

Clement:2020:AML

[Cle20] Charles Clement. Anthony Milner Lane. 27 July 1928–9 February 2011. Biographical Memoirs of Fellows of the Royal Society, 68: 297–314, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0044.

Cottrell:1981:URE

[CM81] Alan Howard Cottrell and J. E. O. Mayne. Ulick Richardson Evans, 31 March 1889–3 April 1980. Biographical Memoirs of Fellows of the Royal Society, 27:234–253, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0010.

Caro:1987:LHM

[CMO87] D. E. Caro, R. L. Martin, and Marcus Laurence Elwin Oliphant. Leslie Harold Martin, 21 December 1900–1 February 1983. Biographical Memoirs of Fellows of the Royal Society, 33:387–409, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1987.0015.

Cronin:2023:MFL

[CNO23] Thomas Cronin, Dan-Eric Nilsson, and Daniel Osorio. Michael Francis Land. 12 April 1942–14 December 2020. Biographical Memoirs of Fellows of the Royal Society, 75:279–297, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0023.

Clary:2022:ADB

[CO22] David C. Clary FRS and Brian J. Orr. Amyand David Buckingham. 28 January 1930–4 February 2021. Biographical Memoirs of Fellows of the Royal Society, 72:77–99, June 2022.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0026.

Cockcroft:1963:NHD

[Coc63] John Douglas Cockcroft. Niels Henrik David Bohr, 1885–1962. Biographical Memoirs of Fellows of the Royal Society, 9:37–53, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1963.0002.

Cockcroft:1967:GH

[Coc67] John Douglas Cockcroft. George De Hevesy, 1885–1966. Biographical Memoirs of Fellows of the Royal Society, 13:125–166, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1967.0007.

Cocking:2003:RBJ

[Coc03] Edward C. Cocking FRS. Robert Brown. 29 July 1908–13 July 1999. Biographical Memoirs of Fellows of the Royal Society, 49: 69–81, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2003.0004.

Colebrook:1956:AF

[Col56] Leonard Colebrook. Alexander Fleming, 1881–1955. Biographical Memoirs of Fellows of the Royal Society, 2:117–127, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1956.0008.

Colebrook:1964:GD

[Col64] Leonard Colebrook. Gerhard Domagk, 1895–1964. Biographical Memoirs of Fellows of the Royal Society, 10:39–50, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1964.0003.

Collar:1971:EFR

[Col71] Arthur Roderick Collar. Ernest Frederick Relf, 1888–1970. Biographical Memoirs of Fellows of the Royal Society, 17:593–616, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0023.

Collar:1978:AF

[Col78] Arthur Roderick Collar. Arthur Fage, 1890–1977. Biographical Memoirs of Fellows of the Royal Society, 24:32–53, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0003.

Collinson:2002:SKR

[Col02] D. W. Collinson. Stanley Keith Runcorn. 19 November 1922– 5 December 1995. Biographical Memoirs of Fellows of the Royal Society, 48:391–403, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2002.
0023.

Collins:2018:SRK

[Col18] Peter Collins. Sir Ralph Kohn. 9 December 1927–11 November 2016. Biographical Memoirs of Fellows of the Royal Society, 65:235– 247, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2018.0023.

Connerade:2009:WRS

[Con09] Jean-Patrick Connerade. William Reginald Stephen Garton. 7 March 1912–28 August 2002. Biographical Memoirs of Fellows of the Royal Society, 55:77–90, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.
0023.

Conder:2022:GHJ

[Con22] Marston D. E. Conder FRSNZ. Graham Higman. 19 January 1917–8 April 2008. Biographical Memoirs of Fellows of the Royal Society, 73:277–290, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.

Cook:1958:ELK

[Coo58] James Wilfred Cook. Ernest Laurence Kennaway, 1881–1958. Biographical Memoirs of Fellows of the Royal Society, 4:138–154, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1958.0013.

Cook:1960:IMH

[Coo60] Alan Hugh Cook. Ian Morris Heilbron, 1886–1959. Biographical Memoirs of Fellows of the Royal Society, 6:65–85, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1960.0025.

Cooper:1972:HWH

[Coo72] Leslie Hugh Norman Cooper. Hildebrand Wolfe Harvey, 1887–1970. Biographical Memoirs of Fellows of the Royal Society, 18:330–347, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0011.

Cook:1990:SHJ

[Coo90] Alan Hugh Cook. Sir Harold Jeffreys, 2 April 1891–18 March 1989. Biographical Memoirs of Fellows of the Royal Society, 36:301–333, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1990.0034.

Coope:1994:FWS

[Coo94] G. R. Coope. Frederick William Shotton, 8 October 1906–21 July 1990. Biographical Memoirs of Fellows of the Royal Society, 39:417– 432, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Cook:1998:LEB

[Coo98] Alan Cook. Louis Essen, O. B. E. 6 September 1908–24 August 1997. Biographical Memoirs of Fellows of the Royal Society, 44:143– 158, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1998.0010.

Cook:1999:RVJ

[Coo99] Alan Cook. Reginald Victor Jones, C.H., C.B., C.B.E. 29 September 1911–17 December 1997. Biographical Memoirs of Fellows of the Royal Society, 45:239–254, January 1999. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0016.

Cook:2001:HCV

[Coo01] Alan Cook. Hendrik Christoffel Van De Hulst Ridder In De Orde Van Nederlandse Leeuw. 19 November 1918–31 July 2000. Biographical Memoirs of Fellows of the Royal Society, 47:465–479, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2001.0028.

Corner:1960:JDR

[Cor60] George Washington Corner. John Davison Rockefeller, Jr, 1874– 1960. Biographical Memoirs of Fellows of the Royal Society, 6:247– 257, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1960.0035.

Corner:1990:HME

[Cor90] Edred John Henry Corner. His Majesty Emperor Hirihito of Japan, K. G., 29 April 1901–7 January 1989. Biographical Memoirs of Fellows of the Royal Society, 36:241–272, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1990.0032.

Cottrell:1972:ENC

[Cot72] Alan Howard Cottrell. Edward Neville da Costa Andrade, 1887– 1971. Biographical Memoirs of Fellows of the Royal Society, 18: 1–20, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1972.0001.

Cottrell:1984:GVR

[Cot84] Alan Howard Cottrell. Geoffrey Vincent Raynor, 2 October 1913–20 October 1983. Biographical Memoirs of Fellows of the Royal Society, 30:546–563, November 1984. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1984.
0018.

Coulson:1973:SFB

[Cou73] Charles Alfred (Charles Alfred) Coulson. Samuel Francis Boys, 1911–1972. Biographical Memoirs of Fellows of the Royal Society, 19:94–115, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1973.
0004.

Cowan:1957:WWS

[Cow57] J. M. Cowan. William Wright Smith, 1875–1956. Biographical Memoirs of Fellows of the Royal Society, 3:192–202, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1957.0013.

Cowling:1964:OS

[Cow64] Thomas George Cowling. Otto Struve, 1897–1963. Biographical Memoirs of Fellows of the Royal Society, 10:283–304, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1964.0017.

Cowling:1971:SC

[Cow71] T. G. Cowling. Sydney Chapman, 1888–1970. Biographical Memoirs of Fellows of the Royal Society, 17:53–89, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1971.0003.

Cowley:2004:SEW

[Cow04] Roger A. Cowley. Sir Edgar William John Mitchell CBE. 25 September 1925–30 October 2002. Biographical Memoirs of Fellows of the Royal Society, 50:171–181, January 2004. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2004.0013.

Cowley:2005:BNB

[Cow05] Roger Cowley. Bertram Neville Brockhouse. 15 July 1918–13 October 2003. Biographical Memoirs of Fellows of the Royal Society, 51:

51–65, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2005.0004.

Cowsik:2021:MGK

[Cow21] Ramanath Cowsik. Mambillikalathil Govind Kumar Menon. 28 August 1928–22 November 2016. Biographical Memoirs of Fellows of the Royal Society, 71:399–422, December 2021. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0013.

Cox:1957:LFS

[Cox57] Leslie Reginald Cox. Leonard Frank Spath, 1882–1957. Biographical Memoirs of Fellows of the Royal Society, 3:217–226, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1957.0015.

Cox:1958:WJA

[Cox58] Leslie Reginald Cox. William Joscelyn Arkell, 1904–1958. Biographical Memoirs of Fellows of the Royal Society, 4:1–14, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1958.0001.

Cox:1995:IGG

[Cox95] Keith Gordon Cox. Ian Graham Gass, 20 March 1926–8 October 1992. Biographical Memoirs of Fellows of the Royal Society, 41:170– 182, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1995.0011.

Cox:2003:HED

[Cox03] David Cox FRS. Henry Ellis Daniels. 2 October 1912–16 April 2000. Biographical Memoirs of Fellows of the Royal Society, 49:133–146, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2003.0008.

Chick:1969:EVM

[CP69] Harriette Chick and Rudolph Albert Peters. Elmer Verner Mc-Collum, 1879–1967. Biographical Memoirs of Fellows of the Royal Society, 15:159–171, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1969.
0008.

Claesson:1972:S

[CP72] Stig Claesson and Kai O. Pedersen. The Svedberg, 1884–1971. Biographical Memoirs of Fellows of the Royal Society, 18:594–627, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0022.

Cuthbertson:1994:ELS

[CP94] W. F. J. Cuthbertson and J. E. Page. Ernest Lester Smith, 7 August 1904–6 November 1992. Biographical Memoirs of Fellows of the Royal Society, 40:347–365, November 1994. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1994.0044.

Chatt:1983:ANN

[CR83] Joseph Chatt and M. I. Rybinskaya. Aleksandr Nikolaevich Nesmeyanov, 9 September 1899–17 January 1980. Biographical Memoirs of Fellows of the Royal Society, 29:399–480, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1983.0016.

Clarke:2007:ADY

[CR07] J. F. Clarke and N. Riley. Alec David Young. 15 August 1913–27 January 2005. Biographical Memoirs of Fellows of the Royal Society, 53:401–418, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2007.
0004.

Campbell:2016:BSR

[CR16a] Charles T. Campbell and Ruth A. Rabinovitch. Benton Seymour Rabinovitch. 19 February 1919–2 August 2014. Biographical Memoirs of Fellows of the Royal Society, 62:505–524, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0021.

Clark:2016:JLB

[CR16b] Robin J. H. Clark CNZM FRS and Paul R. Raithby. Jack Lewis, Baron Lewis of Newnham HonFRSC. 13 February 1928–17 July 2014. Biographical Memoirs of Fellows of the Royal Society, 62:299– 322, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2015.0022.

Craigie:1957:AF

[Cra57] James Craigie. Arthur Felix, 1887–1956. Biographical Memoirs of Fellows of the Royal Society, 3:53–79, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1957.0005.

Craig:1972:RSN

[Cra72] David Parker Craig. Ronald Sydney Nyholm, 1917–1971. Biographical Memoirs of Fellows of the Royal Society, 18:445–475, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1972.0015.

Craft:2020:PMC

[Cra20] Sir Alan W. Craft. Patrick Mollison CBE. 17 March 1914– 26 November 2011. Biographical Memoirs of Fellows of the Royal Society, 68:315–331, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0022.

Crabtree:2021:MHG

[Cra21] Robert H. Crabtree FRS. Malcolm L. H. Green. 16 April 1936–24 July 2020. Biographical Memoirs of Fellows of the Royal Society, 70:175–188, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0038.

Crabtree:2023:PMM

[Cra23] Robert H. Crabtree FRS. Peter M. Maitlis. 15 January 1933–18 May 2022. Biographical Memoirs of Fellows of the Royal Society, 75:327–343, December 2023. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0048.

Craik:2024:ETM

[Cra24] Fergus I. M. Craik FRS. Endel Tulving. 26 May 1927–11 September 2023. Biographical Memoirs of Fellows of the Royal Society, 77:375– 393, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Crew:1967:RCP

[Cre67] Francis Albert Eley Crew. Reginald Crundall Punnett, 1875–1967. Biographical Memoirs of Fellows of the Royal Society, 13:309–326, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1967.0016.

Crease:2024:FJM

[Cre24] Robert P. Crease. Francis James Macdonald Farley. 13 October 1920–16 July 2018. Biographical Memoirs of Fellows of the Royal Society, 77:185–199, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Crow:1997:MKN

[Cro97] James F. Crow. Motoo Kimura. 13 November 1924–13 November 1994. Biographical Memoirs of Fellows of the Royal Society, 43:255– 265, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1997.0014.

Crombie:2000:RAR

[Cro00] Leslie Crombie. Ralph Alexander Raphael, C.B.E. 1 January 1921–27 April 1998. Biographical Memoirs of Fellows of the Royal Society, 46:465–481, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0096.

Crossland:2010:JEH

[Cro10] Ian Crossland. John Edwin Harris. 2 June 1932–3 February 2009. Biographical Memoirs of Fellows of the Royal Society, 56:149–169, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2010.0004.

Crompton:2019:DHN

[Cro19] D. W. T. Crompton. Donald Henry Northcote. 27 December 1921–7 January 2004. Biographical Memoirs of Fellows of the Royal Society, 67:357–370, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0020.

Crowther:2020:SAK

[Cro20] R. A. Crowther FRS FMedSci. Sir Aaron Klug OM. 11 August 1926–20 November 2018. Biographical Memoirs of Fellows of the Royal Society, 68:273–296, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0034.

Cohen:1983:HCU

[CRST83] K. P. Cohen, Stanley Keith Runcorn, H. E. Suess, and Henry George Thode. Harold Clayton Urey, 29 April 1893– 5 January 1981. Biographical Memoirs of Fellows of the Royal Society, 29:622–659, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1983.
0022.

Cruft:1976:ESJ

[Cru76] H. J. Cruft. Edgar Stedman,12 July 1890–8 May 1975. Biographical Memoirs of Fellows of the Royal Society, 22:528–553, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1976.0019.

Cruickshank:2000:SEG

[Cru00] D. W. J. Cruickshank. Sir Ernest Gordon Cox, K.B.E. 24 April 1906–23 June 1996. Biographical Memoirs of Fellows of the Royal Society, 46:65–84, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0074.

Crumpton:2004:AFW

[Cru04] Michael J. Crumpton. Alan Frederick Williams. 25 May 1945–9 April 1992. Biographical Memoirs of Fellows of the Royal Society, 50:351–366, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0023.

Campbell:1998:RHS

[CS98] P. N. Campbell and L. A. Stocken. Robert Henry Stewart Thompson, C. B. E.. 2 February 1912–16 January 1998. Biographical Memoirs of Fellows of the Royal Society, 44:419–431, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0027.

Crane:2022:WGC

[CSB22] Peter R. Crane FRS, Andrew C. Scott, and David J. Beerling FRS. William Gilbert Chaloner. 22 November 1928–13 October 2016. Biographical Memoirs of Fellows of the Royal Society, 73:107–126, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2022.0022.

Clary:2017:JNM

[CST17] Sir David C. Clary FRS, Anthony J. Stace FRS, and Jonathan Tennyson FRS. John Norman Murrell. 2 March 1932–25 January 2016. Biographical Memoirs of Fellows of the Royal Society, 63:467– 486, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2016.0026.

Crowdy:2014:PGS

[CT14] Darren Crowdy and Saleh Tanveer. Philip Geoffrey Saffman. 19 March 1931–17 August 2008. Biographical Memoirs of Fellows of the Royal Society, 60:375–395, January 2014. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0021.

Cartwright:1976:JPD

[CU76] David Edgar Cartwright and Fritz Joseph Ursell. Joseph Proudman, 30 December 1888–26 June 1975. Biographical Memoirs of Fellows of the Royal Society, 22:319–333, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0014.

Cullen:1990:HEM

[Cul90] Alexander Lamb Cullen. Harold Everard Monteagle Barlow, 15 November 1899–20 April 1989. Biographical Memoirs of Fellows of the Royal Society, 36:17–42, December 1990. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1990.0022.

Cullen:1995:SGT

[Cul95] J. Cullen. Sir George Taylor, 15 February 1904–12 November 1993. Biographical Memoirs of Fellows of the Royal Society, 41:458–469, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0027.

Cunningham:1964:CGD

[Cun64] D. J. C. Cunningham. Claude Gordon Douglas, 1882–1963. Biographical Memoirs of Fellows of the Royal Society, 10:51–74, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0004.

Curran:1984:PID

[Cur84] Samuel Crowe Curran. Philip Ivor Dee, 8 April 1904–17 April 1983. Biographical Memoirs of Fellows of the Royal Society, 30:139–166, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1984.0005.

Curtis:2022:JHC

[Cur22] Robert Turner Curtis. John Horton Conway. 26 December 1937–11 April 2020. Biographical Memoirs of Fellows of the Royal Society, 72:117–138, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0034.

Cushing:2005:RDP

[Cus05] David Cushing. Reginald Dawson Preston. 21 July 1908–3 May 2000. Biographical Memoirs of Fellows of the Royal Society, 51:347– 353, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2005.0022.

Cuthbert:2005:MLV

[Cut05] Alan W. Cuthbert. Marthe Louise Vogt. 8 September 1903– 9 September 2003. Biographical Memoirs of Fellows of the Royal Society, 51:409–423, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2005.
0027.

Cassels:1985:IMV

[CV85] John William Scott Cassels and Robert Charles Vaughan. Ivan Matveevich Vinogradov, 14 September 1891–20 March 1983. Biographical Memoirs of Fellows of the Royal Society, 31:611–631, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1985.0021.

Coates:1994:KMJ

[CV94] John Henry Coates and A. J. Van der Poorten. Kurt Mahler, 26 July 1903–26 February 1988. Biographical Memoirs of Fellows of the Royal Society, 39:263–279, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Chen:2017:KFR

[CV17] William Chen and Robert Vaughan FRS. Klaus Friedrich Roth. 29 October 1925–10 November 2015. Biographical Memoirs of Fellows of the Royal Society, 63:487–525, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0014.

Creeth:2002:RAK

[CVW02] J. Michael Creeth, Leon Vallet, and Winifred M. Watkins F.R.S. Ralph Ambrose Kekwick. 11 November 1908–17 January 2000. Biographical Memoirs of Fellows of the Royal Society, 48:233–249, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2002.0013.

Connor:2011:DRM

[CW11] John Connor FRS and Colin Windsor FRS. Derek Robinson. 27 May 1941–2 December 2002. Biographical Memoirs of Fellows of the Royal Society, 57:395–422, January 2011.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0012.

Dixon:1997:WCP

[DAB97] R. N. Dixon, D. M. Agar, and R. E. Burge. William Charles Price. 1 April 1909–10 March 1993. Biographical Memoirs of Fellows of the Royal Society, 43:431–442, January 1997. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0023.

Dacie:1988:HLJ

[Dac88] John Vivian Dacie. Hermann Lehmann, 8 July 1910–13 July 1985. Biographical Memoirs of Fellows of the Royal Society, 34:405–449, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1988.0015. See corrigendum [Ano90b].

Dainton:1985:GBK

[Dai85] Frederick Sydney Dainton. George Bogdan Kistiakowsky, 18 November 1900–7 December 1982. Biographical Memoirs of Fellows of the Royal Society, 31:376–408, November 1985. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1985.0013.

Dainton:1990:NNS

[Dai90] Frederick Sydney Dainton. Nikolai Nikolaevich Semenov, 16 April 1896–25 September 1986. Biographical Memoirs of Fellows of the Royal Society, 36:526–546, December 1990. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1990.0042.

Dale:1955:EM

[Dal55] Henry Hallett Dale. Edward Mellanby, 1884–1955. Biographical Memoirs of Fellows of the Royal Society, 1:192–222, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0015.

Dale:1957:PH

[Dal57] Henry Hallett Dale. Percival Hartley, 1881–1957. Biographical Memoirs of Fellows of the Royal Society, 3:80–100, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1957.0006.

Dale:1958:AJE

[Dal58] Henry Hallett Dale. Arthur James Ewins, 1882–1958. Biographical Memoirs of Fellows of the Royal Society, 4:81–91, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0007.

Dale:1961:TRE

[Dal61] Henry Hallett Dale. Thomas Renton Elliott, 1877–1961. Biographical Memoirs of Fellows of the Royal Society, 7:52–74, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1961.0005.

Dale:1962:OL

[Dal62] Henry Hallett Dale. Otto Loewi, 1873–1961. Biographical Memoirs of Fellows of the Royal Society, 8:67–89, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1962.0006.

Dalziel:1983:AHT

[Dal83] Keith Dalziel. Axel Hugo Theodor Theorell, 6 July 1903–15 August 1982. Biographical Memoirs of Fellows of the Royal Society, 29:584– 621, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1983.0021.

Dalgarno:1997:SDR

[Dal97] Alexander Dalgarno. Sir David Robert Bates. 18 November 1916–5 January 1994. Biographical Memoirs of Fellows of the Royal Society, 43:49–71, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.

Darwin:1958:DRH

[Dar58] Charles Galton Darwin. Douglas Rayner Hartree, 1897–1958. Biographical Memoirs of Fellows of the Royal Society, 4:102–116, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1958.0010.

Darlington:1961:OR

[Dar61] Cyril Dean Darlington. Otto Renner, 1883–1960. Biographical Memoirs of Fellows of the Royal Society, 7:206–220, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1961.0016.

Darwin:1962:EG

[Dar62] Charles Galton Darwin. Ezer Griffiths, 1888–1962. Biographical Memoirs of Fellows of the Royal Society, 8:41–48, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1962.0004.

Darlington:1972:HNB

[Dar72] Cyril Dean Darlington. Horace Newton Barber, 1914–1971. Biographical Memoirs of Fellows of the Royal Society, 18:21–33, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0002.

Datta:1988:HWS

[Dat88] Naomi Datta. Herbert Williams Smith, 3 May 1919–16 June 1987. Biographical Memoirs of Fellows of the Royal Society, 34:754–786, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1988.0024.

Davies:1970:PJW

[Dav70] Mansel Davies. Peter Joseph Wilhelm Debye, 1884–1966. Biographical Memoirs of Fellows of the Royal Society, 16:175–232, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1970.0007.

Davidson:1971:WOK

[Dav71] James Norman Davidson. William Ogilvy Kermack, 1898–1970. Biographical Memoirs of Fellows of the Royal Society, 17:399–429, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0015.

Davies:2018:ALC

[Dav18] J. Brian Davies. Alexander Lamb Cullen OBE. 30 April 1920– 27 December 2013. Biographical Memoirs of Fellows of the Royal Society, 64:131–148, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Davies:2023:RJG

[Dav23] Alwyn G. Davies FRS. Ronald James Gillespie. 21 August 1924–26 February 2021. Biographical Memoirs of Fellows of the Royal Society, 75:129–147, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0001.

Duke-Elder:1958:JHP

[DE58] William Stewart Duke-Elder. John Herbert Parsons, 1868–1957. Biographical Memoirs of Fellows of the Royal Society, 4:204–214, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1958.0018.

Denholm:1995:RMS

[DE95] I. Denholm and Michael Elliott. Roman Mieczyslaw Sawicki, 20 April 1930–22 July 1990. Biographical Memoirs of Fellows of the Royal Society, 41:397–418, November 1995. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0024.

Donkin:2019:JAG

[DE19] Stephen Donkin and Karin Erdmann. James Alexander Green. 26 February 1926–7 April 2014. Biographical Memoirs of Fellows of the Royal Society, 67:173–190, December 2019. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0012.

Deacon:1963:JAE

[Dea63] George Edward Raven Deacon. John Augustine Edgell, 1880– 1962. Biographical Memoirs of Fellows of the Royal Society, 9:87– 90, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1963.0005.

Deacon:1966:HP

[Dea66] George Edward Raven Deacon. Hans Pettersson, 1888–1966. Biographical Memoirs of Fellows of the Royal Society, 12:405–421, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0019.

Deer:1967:LRW

[Dee67] William Alexander Deer. Laurence Rickard Wager, 1904–1965. Biographical Memoirs of Fellows of the Royal Society, 13:358–385, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1967.0019.

Denton:1974:TIS

[Den74] Eric James Denton. Trevor Ian Shaw, 1928–1972. Biographical Memoirs of Fellows of the Royal Society, 20:359–380, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1974.0016.

Denton:1981:PCC

[Den81] Eric James Denton. Peter Christopher Caldwell, 25 January 1927–7 June 1979. Biographical Memoirs of Fellows of the Royal Society, 27:153–172, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0007.

Denbigh:1986:PVD

[Den86] Kenneth George Denbigh. Peter Victor Danckwerts, 14 October 1916–25 October 1984. Biographical Memoirs of Fellows of the Royal Society, 32:97–114, December 1986. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1986.0004.

Dietmann:2023:JWS

[DFSV23] Rainer Dietmann, Tom Fisher, Chris Smyth, and Sanju Velani. John William Scott (‘Ian’) Cassels. 11 July 1922–27 July 2015. Biographical Memoirs of Fellows of the Royal Society, 74:87–108, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2022.0035.

Daly:1970:CAL

[DG70] Ivan De Burgh Daly and R. A. Gregory. Charles Arthur Lovatt Evans, 1884–1968. Biographical Memoirs of Fellows of the Royal Society, 16:233–252, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1970.
0008.

Denton:2016:SJH

[DG16] J. D. Denton FRS FREng and J. P. Gostelow. Sir John Harold Horlock FREng. 19 April 1928–22 May 2015. Biographical Memoirs of Fellows of the Royal Society, 62:213–232, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0009.

Dewey:2023:JGR

[DG23] J. F. Dewey FRS and R. Graham. John Graham Ramsay. 17 June 1931–12 January 2021. Biographical Memoirs of Fellows of the Royal Society, 74:363–385, April 2023. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0031.

Davies:2016:SAC

[DGL16] Rodney D. Davies FRS, Sir Francis Graham-Smith FRS, and Andrew G. Lyne FRS. Sir Alfred Charles Bernard Lovell OBE. 31 August 1913–6 August 2012. Biographical Memoirs of Fellows of the Royal Society, 62:323–344, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0026.

Diggle:2018:JEB

[DGS18] Peter J. Diggle, Peter J. Green FRS, and Bernard W. Silverman FRS. Julian Ernst Besag. 26 March 1945–6 August 2010. Biographical Memoirs of Fellows of the Royal Society, 64:27–50, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Dowell:2011:WEB

[DGV11] J. D. Dowell FRS, J. A. R. Griffith, and W. F. Vinen FRS. William Ernest Burcham CBE. 1 August 1913–5 November 2008. Biographical Memoirs of Fellows of the Royal Society, 57:63–78, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0020.

Dowson:2004:SDG

[DH04] Duncan Dowson and Gordon Higginson. Sir Derman Guy Christopherson. 6 September 1915–7 November 2000. Biographical Memoirs of Fellows of the Royal Society, 50:47–59, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2004.0005.

Dickens:1975:ECD

[Dic75] Frank Dickens. Edward Charles Dodds, 13 October 1899– 16 December 1973. Biographical Memoirs of Fellows of the Royal Society, 21:227–267, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1975.
0006.

Dieke:1956:RWW

[Die56] G. H. Dieke. Robert Williams Wood, 1868–1955. Biographical Memoirs of Fellows of the Royal Society, 2:326–345, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1956.0022.

Dunitz:2013:LEO

[DJ13] Jack D. Dunitz FRS and Gerald F. Joyce. Leslie Eleazer Orgel. 12 January 1927–27 October 2007. Biographical Memoirs of Fellows of the Royal Society, 59:277–289, January 2013.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0002.

Davis:2003:RHB

[DL03] John Davis and Bernard Lovell FRS. Robert Hanbury Brown. 31 August 1916–16 January 2002 Elected FRS 1960. Biographical Memoirs of Fellows of the Royal Society, 49:83–106, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0005.

Dewey:2004:RMS

[DL04] John Frederick Dewey and Bernard Elgey Leake. Robert Millner Shackleton. 30 December 1909–3 May 2001. Biographical Memoirs of Fellows of the Royal Society, 50:285–297, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2004.0018.

Daly:1968:KJF

[DM68] Ivan De Burgh Daly and R. G. Macbeth. Kenneth James Franklin, 1897–1966. Biographical Memoirs of Fellows of the Royal Society, 14:223–242, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1968.
0010.

Denton:2006:SAS

[DM06] Derek A. Denton and Iain MacIntyre. Sylvia Agnes Sophia Tait. 8 January 1917–28 February 2003. Biographical Memoirs of Fellows of the Royal Society, 52:379–399, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0026.

Davidson:2009:DAR

[DM09] Walter F. Davidson FRSC and A. R. W. McKellar FRSC. Donald Allan Ramsay CM. 11 July 1922–25 October 2007. Biographical Memoirs of Fellows of the Royal Society, 55:221–239, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0007.

Davies:2020:RJH

[DM20] Alwyn G. Davies FRS and Paul F. McMillan. Robin Jon Hawes Clark. 16 February 1935–6 December 2018. Biographical Memoirs of Fellows of the Royal Society, 68:103–129, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0037.

Deer:1974:CET

[DN74] William Alexander Deer and Stephen Robert Nockolds. Cecil Edgar Tilley, 1894–1973. Biographical Memoirs of Fellows of the Royal Society, 20:381–400, December 1974. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1974.0017.

Dockray:1998:RAG

[Doc98] G. J. Dockray. Roderic Alfred Gregory, C. B. E.. 29 December 1913-5 September 1990. Biographical Memoirs of Fellows of the Royal Society, 44:207–216, January 1998. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0014.

Dodds:1957:AEA

[Dod57] Edward Charles Dodds. Alexander, Earl of Athlone, 1874–1957. Biographical Memoirs of Fellows of the Royal Society, 3:1–5, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1957.0001.

Dodson:2002:DMC

[Dod02] Guy Dodson F.R.S. Dorothy Mary Crowfoot Hodgkin, O.M. 12 May 1910–29 July 1994. Biographical Memoirs of Fellows of the Royal Society, 48:179–219, January 2002. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0011.

Dodson:2020:GDJ

[Dod20] Eleanor Dodson FRS. Guy Dodson. 13 January 1937–24 December 2012. Biographical Memoirs of Fellows of the Royal Society, 68:

151–174, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0042.

Doll:1994:ABH

[Dol94] William Richard Shaboe Doll. Austin Bradford Hill, 8 July 1897–18 April 1991. Biographical Memoirs of Fellows of the Royal Society, 40:127–140, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.
0032.

Dollery:1995:SJM

[Dol95] Colin Dollery. Sir John McMichael, 25 July 1904–3 March 1993. Biographical Memoirs of Fellows of the Royal Society, 41:282–296, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0017.

Domb:1998:GSR

[Dom98] Cyril Domb. George Stanley Rushbrooke. 19 January 1915-14 December 1995. Biographical Memoirs of Fellows of the Royal Society, 44:365–384, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0024.

Dombey:2021:JCW

[Dom21] Norman Dombey. John Clive Ward. 1 August 1924–6 May 2000. Biographical Memoirs of Fellows of the Royal Society, 70:419–440, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2020.0023.

Dorey:1966:HJG

[Dor66] Stanley Fabes Dorey. Herbert John Gough, 1890–1965. Biographical Memoirs of Fellows of the Royal Society, 12:181–194, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1966.0008.

Dornan:2018:IBC

[Dor18] Peter Dornan FRS. Ian Butterworth CBE. 3 December 1930–29 November 2013. Biographical Memoirs of Fellows of the Royal Society, 64:69–87, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Douglas:1964:WES

[Dou64] R. W. Douglas. William Ernest Stephen Turner, 1881–1963. Biographical Memoirs of Fellows of the Royal Society, 10:325–355, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0019.

Dougherty:2007:KGB

[Dou07] J. P. Dougherty. Kenneth George Budden. 23 June 1915– 4 September 2005. Biographical Memoirs of Fellows of the Royal Society, 53:109–126, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2007.
0009.

Douglas:2019:SDC

[Dou19] Angela E. Douglas. Sir David Cecil Smith. 21 May 1930–29 June 2018. Biographical Memoirs of Fellows of the Royal Society, 67:401– 419, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0011.

Downie:1970:SPB

[Dow70] Allan Watt Downie. Samuel Phillips Bedson, 1886–1969. Biographical Memoirs of Fellows of the Royal Society, 16:14–35, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0002.

Downs:1994:AAC

[Dow94] Diarmuid Downs. Alexander Arnold Constantine Issigonis, 18 November 1906–2 October 1988. Biographical Memoirs of Fellows of the Royal Society, 39:199–211, February 1994. CODEN BM-FRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Daly:1970:EBV

[DP70] Ivan De Burgh Daly and Lillian Mary Pickford. Ernest Basil Verney, 1894–1967. Biographical Memoirs of Fellows of the Royal Society, 16:523–542, November 1970. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1970.
0022.

Dalitz:1986:PAM

[DP86] Richard Henry Dalitz, F.R.S. and Sir Rudolf Ernst Peierls, F.R.S. Paul Adrien Maurice Dirac. 8 August 1902–20 October 1984. elected F.R.S. 1930. Biographical Memoirs of Fellows of the Royal Society, 32:137–185, December 1986. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/770111;   https:
//royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0006.

Ditchburn:1974:ST

[DR74] Robert William Ditchburn and George Dixon Rochester. Samuel Tolansky, 1907–1973. Biographical Memoirs of Fellows of the Royal Society, 20:429–455, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1974.
0019.

Dunn-Rankin:2022:FJW

[DR22] D. Dunn-Rankin. Felix Jiri Weinberg. 2 April 1928–5 December 2012. Biographical Memoirs of Fellows of the Royal Society, 72: 389–406, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2021.0012.

Day:1979:CAJ

[DS79] R. L. Day and Alfred Spinks. Christopher Alwyne Jack Young, 7 March 1912–20 January 1978. Biographical Memoirs of Fellows of the Royal Society, 25:554–573, November 1979. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0020.

Denton:1986:FSR

[DS86] Eric James Denton and A. J. Southward. Frederick Stratten Russell, 3 November 1897–5 June 1984. Biographical Memoirs of Fellows of the Royal Society, 32:461–493, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1986.0015.

Duff:2021:STW

[DS21] M. J. Duff FRS and K. S. Stelle. Sir Thomas Walter Bannerman Kibble. 23 December 1932–2 June 2016. Biographical Memoirs of Fellows of the Royal Society, 70:225–244, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0040.

Davidson:2020:SPS

[DSD20] P. J. Davidson, M. M. Sharma, and J. F. Davidson FRS. Sydney Percy Smith Andrew. 16 May 1926–2 November 2011. Biographical Memoirs of Fellows of the Royal Society, 68:7–22, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0005.

Downie:1985:DGE

[DST85] Allan Watt Downie FRS, C. E. Gordon Smith, and J. O’H. Tobin. David Gwynne Evans, 6 September 1909–13 June 1984. Biographical Memoirs of Fellows of the Royal Society, 31:172–196, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1985.0007.

Dainton:1981:RGW

[DT81] Frederick Sydney Dainton and Brian Arthur Thrush. Ronald George Wreyford Norrish, 9 November 1897–7 June 1978. Biographical Memoirs of Fellows of the Royal Society, 27:379–424, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1981.0016.

Dewey:2020:SMM

[DT20] J. F. Dewey FRS and P. N. Taylor. Stephen Moorbath. 9 May 1929–16 October 2016. Biographical Memoirs of Fellows of the Royal Society, 69:365–391, December 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0016.

Davies:2024:KWS

[DT24] A. Russell Davies and Roger I. Tanner. Kenneth Walters. 14 September 1934–28 March 2022. Biographical Memoirs of Fellows of the Royal Society, 77:395–415, December 2024. CODEN BM-FRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Dubos:1956:OTA

[Dub56] R. J. Dubos. Oswald Theodore Avery, 1877–1955. Biographical Memoirs of Fellows of the Royal Society, 2:35–48, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1956.0003.

Duckworth:1991:IMM

[Duc91] W. E. Duckworth. Ieuan Maddock, 29 March 1917–29 December 1988. Biographical Memoirs of Fellows of the Royal Society, 37:321– 340, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1991.0016.

Dulbecco:1995:HMT

[Dul95] Renato Dulbecco. Howard M. Temin, 10 December 1934– 9 February 1994. Biographical Memoirs of Fellows of the Royal Society, 41:471–480, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1995.
0028.

Dunham:1966:AH

[Dun66] Kingsley Charles Dunham. Arthur Holmes, 1890–1965. Biographical Memoirs of Fellows of the Royal Society, 12:290–310, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1966.0013.

Dunham:1968:JHT

[Dun68] Kingsley Charles Dunham. James Haward Taylor, 1909–1968. Biographical Memoirs of Fellows of the Royal Society, 14:443–448, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1968.0019.

Dunham:1978:WBS

[Dun78] Kingsley Charles Dunham. William Bullerwell, 27 September 1916–25 November 1977. Biographical Memoirs of Fellows of the Royal Society, 24:1–13, November 1978. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0001.

Dunham:1982:LHA

[Dun82] Kingsley Charles Dunham. Leonard Hawkes, 6 August 1891– 29 October 1981. Biographical Memoirs of Fellows of the Royal Society, 28:125–139, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0006.

Dunham:1983:FDA

[Dun83a] Kingsley Charles Dunham. Frank Dixey, 7 April 1892–1 November 1982. Biographical Memoirs of Fellows of the Royal Society, 29:158– 176, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1983.0006.

Dunham:1983:SHH

[Dun83b] Kingsley Charles Dunham. Sidney Henry Haughton, 7 May 1888–24 May 1982. Biographical Memoirs of Fellows of the Royal Society, 29:245–267, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1983.
0011.

Dunitz:1996:LCP

[Dun96] Jack David Dunitz. Linus Carl Pauling, 28 February 1901– 19 August 1994. Biographical Memoirs of Fellows of the Royal Society, 42:317–338, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0020.

Day:2002:DFW

[DWS02] Maxwell F. C. Day, Maxwell J. Whitten, and Don P. A. Sands. Douglas Frew Waterhouse, C.M.G. 3 June 1916–1 December 2000. Biographical Memoirs of Fellows of the Royal Society, 48:459–481, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2002.0027.

Dyson:1999:FDK

[Dys99] J. E. Dyson. Franz Daniel Kahn. 13 May 1926–8 February 1998. Biographical Memoirs of Fellows of the Royal Society, 45:255–267, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1999.0017.

Dyson:2011:NKD

[Dys11] Freeman Dyson FRS. Nicholas Kemmer. 7 December 1911–21 October 1998. Biographical Memoirs of Fellows of the Royal Society, 57(1):189–204, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.jstor.

org/stable/41412878;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2011.0008.

Dyson:2022:FJD

[Dys22] George Dyson. Freeman John Dyson. 15 December 1923– 28 February 2020. Biographical Memoirs of Fellows of the Royal Society, 73:197–226, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0050.

Eaborn:1990:HGM

[Eab90] C. Eaborn. Henry Gilman, 9 May 1893–7 November 1986. Biographical Memoirs of Fellows of the Royal Society, 36:151–172, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1990.0028.

Elderfield:2008:JME

[EB08] Henry Elderfield and Edward A. Boyle. John Marmion Edmond. 27 April 1943–10 April 2001. Biographical Memoirs of Fellows of the Royal Society, 54:137–159, January 2008. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0011.

Edmunds:2020:BEJ

[Edm20] M. G. Edmunds. Bernard Ephraim Julius Pagel. 4 January 1930–14 July 2007. Biographical Memoirs of Fellows of the Royal Society, 69:499–516, December 2020. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0004.

Edwards:1994:BLM

[Edw94] A. P. J. Edwards. Ben Lockspeiser, 9 March 1891–18 October 1990. Biographical Memoirs of Fellows of the Royal Society, 39:245–261, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Edwards:1996:PCS

[Edw96] Robert Geoffrey Edwards. Patrick Christopher Steptoe, C. B. E., 9 June 1913–22 March 1988. Biographical Memoirs of Fellows of the Royal Society, 42:433–452, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0027.

Edwards:2018:JPC

[Edw18] Kenneth Edwards. John Philip Cooper CBE. 16 December 1923– 18 November 2011. Biographical Memoirs of Fellows of the Royal Society, 64:89–106, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Edwardson:2020:AWC

[Edw20] J. Michael Edwardson. Alan William Cuthbert. 7 May 1932–27 August 2016. Biographical Memoirs of Fellows of the Royal Society, 68: 131–149, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0036.

Edwards:2021:LLC

[Edw21] A. W. F. Edwards FRS. Luigi Luca Cavalli-Sforza. 25 January 1922–31 August 2018. Biographical Memoirs of Fellows of the Royal Society, 70:79–105, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0015.

Epstein:1995:DPB

[EE95] Michael Anthony Epstein and M. A. Eastwood. Denis Parsons Burkitt, 28 February 1911–23 March 1993. Biographical Memoirs of Fellows of the Royal Society, 41:88–102, November 1995.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0006.

Eccles:1978:WGP

[EF78] John Carew Eccles and William Feindel. Wilder Graves Penfield, 26 January 1891–5 April 1976. Biographical Memoirs of Fellows of the Royal Society, 24:473–513, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0015.

Edmunds:2017:FGF

[EFP17] D. E. Edmunds, L. E. Fraenkel FRS, and M. Pemberton. Frederick Gerard Friedlander. 25 December 1917–20 May 2001. Biographical Memoirs of Fellows of the Royal Society, 63:273–307, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0027.

Eaborn:1996:JCN

[EL96] C. Eaborn and G. J. Leigh. Joseph Chatt, 6 November 1914–19 May 1994. Biographical Memoirs of Fellows of the Royal Society, 42:95– 110, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1996.0007.

Eley:1976:EKR

[Ele76] Daniel Douglas Eley. Eric Keightley Rideal, 11 April 1890– 25 September 1974. Biographical Memoirs of Fellows of the Royal Society, 22:381–413, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1976.
0017.

Elliott:2008:RJB

[Ell08] J. P. Elliott. Roger John Blin-Stoyle. 24 December 1924–31 January 2007. Biographical Memoirs of Fellows of the Royal Society, 54:31– 45, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2007.0041.

Elliott:2009:BBJ

[Ell09] Roger Elliott FRS. Brebis Bleaney. 6 June 1915–4 November 2006. Biographical Memoirs of Fellows of the Royal Society, 55:3–11, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2008.0018.

Ellis:2022:JDB

[Ell22] George F. R. Ellis FRS. John David Barrow. 29 November 1952–26 September 2020. Biographical Memoirs of Fellows of the Royal Society, 73:41–63, December 2022. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0007.

Elsden:1982:RMJ

[Els82] S. R. Elsden. Roy Markham, 29 January 1916–16 November 1979. Biographical Memoirs of Fellows of the Royal Society, 28:319–345, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1982.0013.

Elvidge:1991:AHC

[Elv91] J. A. Elvidge. Arthur Herbert Cook, 10 July 1911–26 July 1988. Biographical Memoirs of Fellows of the Royal Society, 37:85–102, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1991.0005.

Emeleus:1960:FAP

[Eme60] Harry Julius Emeleus. Friedrich Adolf Paneth, 1887–1958. Biographical Memoirs of Fellows of the Royal Society, 6:227–246, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1960.0034.

Ellis:2010:DWS

[EP10] George F. R. Ellis and Sir Roger Penrose. Dennis William Sciama. 18 November 1926–19 December 1999. Biographical Memoirs of Fellows of the Royal Society, 56:401–422, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0023.

Elliott:2005:MHL

[ES05] Roger Elliott and J. H. Sanders. Maurice Henry Lecorney Pryce. 24 January 1913–24 July 2003. Biographical Memoirs of Fellows of the Royal Society, 51:355–366, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0023.

Evans:1966:WS

[Eva66] David Gwynne Evans. Wilson Smith, 1897–1965. Biographical Memoirs of Fellows of the Royal Society, 12:478–487, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1966.0023.

Evans:1976:CLO

[Eva76] David Gwynne Evans. Cyril Leslie Oakley, 20 June 1907–27 March 1975. Biographical Memoirs of Fellows of the Royal Society, 22:295– 305, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1976.0012.

Evans:1999:SOH

[Eva99] Lloyd T. Evans. Sir Otto Herzberg Frankel. 4 November 1900– 21 November 1998. Biographical Memoirs of Fellows of the Royal Society, 45:165–181, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0012.

Evans:2005:HMC

[Eva05] E. Raymond Evans. Henton Morrogh CBE FRS FREng. 29 September 1917–20 September 2003. Biographical Memoirs of Fellows of the Royal Society, 51:303–314, January 2005. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0019.

Evans:2020:DLB

[Eva20] Neil Wyn Evans. Donald Lynden-Bell. 5 April 1935–6 February 2018. Biographical Memoirs of Fellows of the Royal Society, 69:333– 363, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2020.0008.

Evans:2022:SVJ

[Eva22] David E. Evans FLSW. Sir Vaughan Jones. 31 December 1952– 6 September 2020. Biographical Memoirs of Fellows of the Royal Society, 73:333–356, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0051.

Everett:1984:EGA

[Eve84] Douglas Hugh Everett. Eugen Glueckauf, 9 April 1906– 12 September 1981. Biographical Memoirs of Fellows of the Royal Society, 30:192–224, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1984.
0007.

Ellis:2022:SAW

[EW22] Richard S. Ellis and Alan A. Watson. Sir Arnold Whittaker Wolfendale. 25 June 1927–21 December 2020. Biographical Memoirs of Fellows of the Royal Society, 72:407–430, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0028.

Ewald:1960:ML

[Ewa60] Paul Peter Ewald. Max von Laue, 1879–1960. Biographical Memoirs of Fellows of the Royal Society, 6:134–156, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1960.0028.

Eyre:2013:LEJ

[Eyr13] Brian Eyre CBE FRS FREng. Lewis Edward John Roberts CBE. 31 January 1922–10 April 2012. Biographical Memoirs of Fellows of the Royal Society, 59:291–315, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0001.

Friedel:2003:LEF

[FA03] Jacques Friedel ForMemRS and Pierre Averbuch. Louis Eugène Félix Néel. 22 November 1904–17 November 2000. Biographical Memoirs of Fellows of the Royal Society, 49:367–384, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0021.

Falcon:1970:VCI

[Fal70] Norman Leslie Falcon. Vincent Charles Illing, 1890–1969. Biographical Memoirs of Fellows of the Royal Society, 16:365–384, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1970.0014.

Fowden:2021:SLF

[FAL21] Abigail L. Fowden, John W. Anderson, and Peter J. Lea. Sir Leslie Fowden. 13 October 1925–16 December 2008. Biographical Memoirs of Fellows of the Royal Society, 71:171–196, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2021.0002.

Farren:1961:HTT

[Far61] William Scott Farren. Henry Thomas Tizard, 1885–1959. Biographical Memoirs of Fellows of the Royal Society, 7:313–348, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1961.0024.

Farara:1998:JWF

[Far98] C. J. Farara. John William Fozard, O. B. E.. 16 January 1928–17 July 1996. Biographical Memoirs of Fellows of the Royal Society, 44:193–204, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0013.

Fawcett:2018:GRJ

[Faw18] James Fawcett. Geoffrey Raisman. 28 June 1939–27 January 2017. Biographical Memoirs of Fellows of the Royal Society, 65:341–355, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2018.0001.

Falcon:1987:PEP

[FD87] Norman Leslie Falcon and Kingsley Charles Dunham. Percy Edward (Peter) Kent, 18 March 1913–9 July 1986. Biographical Memoirs of Fellows of the Royal Society, 33:343–373, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0013.

Frater:2012:JPW

[FE12] R. H. Frater AO FAA FTSE and R. D. Ekers FRS FAA. John Paul Wild AC CBE FAA FTSE. 17 May 1923–10 May 2008. Biographical Memoirs of Fellows of the Royal Society, 58:327–346, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0034.

Feather:1971:RW

[Fea71] Norman Feather. Richard Whiddington, 1885–1970. Biographical Memoirs of Fellows of the Royal Society, 17:741–756, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1971.0030.

Feast:2004:CHB

[Fea04] W. J. Feast. Clement Henry Bamford. 10 October 1912–7 November 1999. Biographical Memoirs of Fellows of the Royal Society, 50: 1–13, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2004.0001.

Feast:2018:ALC

[Fea18] W. J. Feast. Anthony Ledwith CBE. 14 August 1933–5 January 2015. Biographical Memoirs of Fellows of the Royal Society, 64: 285–298, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Fedoroff:1994:BMJ

[Fed94] Nina V. Fedoroff. Barbara McClintock, 16 June 1902–2 September 1992. Biographical Memoirs of Fellows of the Royal Society, 40:265– 280, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1994.0039.

Feilden:1995:DTN

[Fei95] Geoffrey Bertram Robert Feilden. David Theodore Nelson Williamson, 15 February 1923–10 May 1992. Biographical Memoirs of Fellows of the Royal Society, 41:516–532, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0031.

Feilden:2005:LHA

[Fei05] G. B. R. Feilden. Lionel Haworth. 4 August 1912–12 April 2000. Biographical Memoirs of Fellows of the Royal Society, 51:195–220, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2005.0013.

Feldberg:1960:BM

[Fel60] Wilhelm Siegmund Feldberg. Bruno Mendel, 1897–1959. Biographical Memoirs of Fellows of the Royal Society, 6:191–199, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1960.0032.

Feldberg:1967:JHG

[Fel67] Wilhelm Siegmund Feldberg. John Henry Gaddum, 1900–1965. Biographical Memoirs of Fellows of the Royal Society, 13:56–77, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1967.0003.

Feldberg:1970:HHD

[Fel70] Wilhelm Siegmund Feldberg. Henry Hallett Dale, 1875–1968. Biographical Memoirs of Fellows of the Royal Society, 16:77–174, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1970.0006.

Fenner:1987:FMB

[Fen87] Frank John Fenner. Frank Macfarlane Burnet, 3 September 1899–31 August 1985. Biographical Memoirs of Fellows of the Royal Society, 33:99–162, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0005.

Ferguson-Smith:2017:JHE

[Fer17] Malcolm A. Ferguson-Smith FRS. John Hilton Edwards. 26 March 1928–11 October 2007. Biographical Memoirs of Fellows of the Royal Society, 63:215–242, January 2017. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0005.

Falconer:2015:CAR

[FGOS15] Kenneth Falconer, Peter M. Gruber, Adam Ostaszewski, and Trevor Stuart. Claude Ambrose Rogers. 1 November 1920– 5 December 2005. Biographical Memoirs of Fellows of the Royal Society, 61:403–435, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0007.

Frater:2013:BYM

[FGW13] R. H. Frater, W. M. Goss, and H. W. Wendt. Bernard Yarnton Mills AC FAA. 8 August 1920–25 April 2011. Biographical Memoirs of Fellows of the Royal Society, 59:215–239, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0015.

Feilden:1998:SFW

[FH98] G. B. R. Feilden and William Hawthorne. Sir Frank Whittle, O. M., K. B. E.. 1 June 1907–9 August 1996. Biographical Memoirs of Fellows of the Royal Society, 44:435–452, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0028.

Field:2008:DTO

[Fie08] John Field. David Tabor. 23 October 1913–26 November 2005. Biographical Memoirs of Fellows of the Royal Society, 54:425–459, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2007.0031.

Fildes:1956:RFJ

[Fil56] Paul Gordon Fildes. Richard Friedrich Johannes Pfeiffer, 1858– 1945. Biographical Memoirs of Fellows of the Royal Society, 2:237– 247, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1956.0016.

Finney:1995:FYM

[Fin95] David John Finney. Frank Yates, 12 May 1902–17 June 1994. Biographical Memoirs of Fellows of the Royal Society, 41:554–573, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0033.

Fitter:2010:ADB

[Fit10] A. H. Fitter. Anthony David Bradshaw. 17 January 1926–21 August 2008. Biographical Memoirs of Fellows of the Royal Society, 56: 25–39, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2010.0015.

Fitzgerald:2022:PDW

[Fit22] Maria Fitzgerald FMedSci FRS. Patrick David Wall. 5 April 1925–8 August 2001. Biographical Memoirs of Fellows of the Royal Society, 72:371–388, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0014.

Fincham:1995:DGC

[FJ95] John Robert Stanley Fincham and B. John. David Guthrie Catcheside, 31 May 1907–1 June 1994. Biographical Memoirs of Fellows of the Royal Society, 41:118–134, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0008.

Flavell:2003:SRR

[Fla03] Richard B. Flavell FRS. Sir Ralph Riley. 23 October 1924– 27 August 1999. Biographical Memoirs of Fellows of the Royal Society, 49:385–396, January 2003. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2003.
0022.

Fleck:1955:WAA

[Fle55a] Alexander Fleck. Wallace Alan Akers, 1888–1954. Biographical Memoirs of Fellows of the Royal Society, 1:1–4, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0001.

Fleure:1955:EMP

[Fle55b] Herbert John Fleure. Emmanuel Marie Pierre Martin Jacquin de Margerie, 1862–1953. Biographical Memoirs of Fellows of the Royal Society, 1:185–191, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1955.
0014.

Fleck:1957:FSB

[Fle57] Alexander Fleck. Frederick Soddy. Born Eastbourne 2 September 1877. Died Brighton 26 September 1956. Biographical Memoirs of Fellows of the Royal Society, 3:203–216, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL http://rsbm.royalsocietypublishing.org/
content/roybiogmem/3/203;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1957.0014.

Fleming:1971:EM

[Fle71] Charles Alexander Fleming. Ernest Marsden, 1889–1970. Biographical Memoirs of Fellows of the Royal Society, 17:462–496, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1971.0018.

Fletcher:1999:SSC

[Fle99] W. Fletcher. Sir Samuel Crowe Curran. 23 May 1912–25 February 1998. Biographical Memoirs of Fellows of the Royal Society, 45:95– 109, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0008.

Flood:1963:OM

[Flo63] E. A. Flood. Otto Maass, 1890–1961. Biographical Memoirs of Fellows of the Royal Society, 9:183–204, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1963.0010.

Florides:2008:JLS

[Flo08] Petros S. Florides. John Lighton Synge. 23 March 1897–30 March 1995. Biographical Memoirs of Fellows of the Royal Society, 54:401– 424, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2007.0040.

Flower:2020:GVR

[Flo20] Rod Flower FRS. Gustav Victor Rudolph Born. 29 July 1921–16 April 2018. Biographical Memoirs of Fellows of the Royal Society, 68:23–47, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0026.

Fantes:2019:JMM

[FM19] Peter Fantes and Sally Mitchison. J. Murdoch Mitchison. 11 June 1922–17 March 2011. Biographical Memoirs of Fellows of the Royal Society, 67:279–306, December 2019. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0006.

Ford:2021:HTN

[FM21] Peter C. Ford and Thomas J. Meyer. Henry Taube. 30 November 1915–16 November 2005. Biographical Memoirs of Fellows of the Royal Society, 70:409–418, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0042.

Fogg:1986:GPW

[Fog86] Gordon Elliott Fogg. George Philip Wells, 17 July 1901– 27 September 1985. Biographical Memoirs of Fellows of the Royal Society, 32:649–676, December 1986. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0022.

Ford:1977:TGD

[For77] Edmund Brisco Ford. Theodosius Grigorievich Dobzhansky, 25 January 1900–18 December 1975. Biographical Memoirs of Fellows of the Royal Society, 23:58–89, November 1977. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0004.

Ford:1995:SOS

[For95] Hugh Ford. Sir Owen Saunders, 24 September 1904–10 October 1993. Biographical Memoirs of Fellows of the Royal Society, 41:378– 394, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1995.0023.

Fortey:1999:CPO

[For99] R. A. Fortey. Colin Patterson. 13 October 1933–9 March 1998. Biographical Memoirs of Fellows of the Royal Society, 45:365–377, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1999.0025.

Ford:2000:FBM

[For00] Hugh Ford. Frank Brian Mercer, O.B.E. 22 December 1927– 22 November 1998. Biographical Memoirs of Fellows of the Royal Society, 46:345–363, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0089.

Fortey:2012:HBW

[For12] Richard A. Fortey FRS. Harry Blackmore Whittington. 24 March 1916–20 June 2010. Biographical Memoirs of Fellows of the Royal Society, 58:299–325, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2012.

Foster:1957:LVK

[Fos57] John Stuart Foster. Louis Vessot King, 1886–1956. Biographical Memoirs of Fellows of the Royal Society, 3:101–108, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1957.0007.

Fowden:2002:RLW

[Fow02] Leslie Fowden F.R.S. Ralph Louis Wain, C.B.E. 29 May 1911– 14 December 2000. Biographical Memoirs of Fellows of the Royal Society, 48:439–458, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2002.
0026.

Fox:1987:JHW

[Fox87] L. Fox. James Hardy Wilkinson, 27 September 1919–5 October 1986. Biographical Memoirs of Fellows of the Royal Society, 33:669– 708, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769967;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0024.

Frank:1971:CFP

[FP71] Frederick Charles Frank and Donald Hill Perkins. Cecil Frank Powell, 1903–1969. Biographical Memoirs of Fellows of the Royal Society, 17:541–563, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1971.
0021.

Fettes:1995:JWS

[FP95] D. J. Fettes and J. A. Plant. Janet Watson, 1 September 1923–29 March 1985. Biographical Memoirs of Fellows of the Royal Society, 41:500–514, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1995.
0030.

Fleming:2004:GPK

[FP04] Graham R. Fleming and David Phillips. George Porter KT OM, Lord Porter of Luddenham. 6 December 1920–31 August 2002. Biographical Memoirs of Fellows of the Royal Society, 50:257–283, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2004.0017.

Fishlock:1998:WCM

[FR98] D. Fishlock and L. E. J. Roberts. Walter Charles Marshall, C. B. E., Lord Marshall of Goring. 5 March 1932–20 February 1996. Biographical Memoirs of Fellows of the Royal Society, 44:299–312, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1998.0020.

Fisher:2024:MFL

[FR24] Elizabeth Fisher and Sohaila Rastan. Mary Frances Lyon. 15 May 1925–25 December 2014. Biographical Memoirs of Fellows of the Royal Society, 77:241–259, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Franklin:1960:FJC

[Fra60] Kenneth James Franklin. Francis Joseph Cole, 1872–1959. Biographical Memoirs of Fellows of the Royal Society, 5:37–47, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Freeth:1957:FGD

[Fre57] Francis Arthur Freeth. Frederick George Donnan, 1870–1956. Biographical Memoirs of Fellows of the Royal Society, 3:23–39, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1957.0003.

Freeth:1960:JS

[Fre60] Francis Arthur Freeth. James Swinburne, 1858–1958. Biographical Memoirs of Fellows of the Royal Society, 5:253–268, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Frisch:1970:LM

[Fri70] Otto Robert Frisch. Lise Meitner, 1878–1968. Biographical Memoirs of Fellows of the Royal Society, 16:405–420, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769597;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0016.

Fryer:1980:SMM

[Fry80] Geoffrey Fryer. Sidnie Milana Manton, 4 May 1902–2 January 1979. Biographical Memoirs of Fellows of the Royal Society, 26:327–356, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1980.0010.

Fogg:1992:DJC

[FS92] Gordon Elliott Fogg and A. J. Southward. Dennis John Crisp, 29 April 1916–18 January 1990. Biographical Memoirs of Fellows of the Royal Society, 38:111–129, November 1992. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0006.

Flavell:2020:MDG

[FS20] Richard B. Flavell DSc CBE FRS and John W. Snape PhD. Michael Denis Gale. 25 August 1943–18 July 2009. Biographical Memoirs of Fellows of the Royal Society, 69:203–223, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0011.

Field:2020:SLA

[FSC20] D. Field, J. P. Simons FRS, and Charles Cockell. Sydney Leach. 11 April 1924–24 December 2019. Biographical Memoirs of Fellows of the Royal Society, 69:313–332, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0018.

Fraser:1983:EHK

[FW83] D. R. Fraser and Elsie May Widdowson. Egon Hynek Kodicek, 3 August 1908–27 July 1982. Biographical Memoirs of Fellows of the Royal Society, 29:297–331, November 1983. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1983.0013.

Wong:1994:CSH

[fWT94] J. T. f. Wong and W. Thompson. Charles Samuel Hanes, 21 May 1903–6 July 1990. Biographical Memoirs of Fellows of the Royal Society, 39:147–155, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Gabathuler:2002:SAA

[Gab02] E. Gabathuler O.B.E., F.R.S. Sir Alexander [Alec] Walter Merrison, D.L. 20 March 1924–19 February 1989. Biographical Memoirs of Fellows of the Royal Society, 48:309–322, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0017.

Gaddum:1956:GA

[Gad56] John Henry Gaddum. Gleb Anrep, 1891–1955. Biographical Memoirs of Fellows of the Royal Society, 2:19–34, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1956.0002.

Gaddum:1957:JWT

[Gad57] John Henry Gaddum. John William Trevan, 1887–1956. Biographical Memoirs of Fellows of the Royal Society, 3:273–288, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1957.0019.

Gadian:2021:SRE

[Gad21] David G. Gadian. Sir Rex Edward Richards. 28 October 1922–15 July 2019. Biographical Memoirs of Fellows of the Royal Society, 70:387–407, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0039.

Gait:2019:DMB

[Gai19] Michael J. Gait. Daniel McGillivray Brown. 3 February 1923–24 April 2012. Biographical Memoirs of Fellows of the Royal Society, 66:79–100, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2018.0008.

Gall:2003:HGC

[Gal03] Joseph G. Gall. Harold Garnet Callan. 5 March 1917–3 November 1993 Elected FRS 1963. Biographical Memoirs of Fellows of the Royal Society, 49:107–118, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0006.

Garnett:1970:HJF

[Gar70a] Alice Garnett. Herbert John Fleure, 1877–1969. Biographical Memoirs of Fellows of the Royal Society, 16:253–278, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0009.

Garnham:1970:RTL

[Gar70b] Percy Cyril Claude Garnham. Robert Thomson Leiper, 1881–1969. Biographical Memoirs of Fellows of the Royal Society, 16:385–404, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1970.0015.

Garnham:1974:EH

[Gar74] Percy Cyril Claude Garnham. Edward Hindle, 1886–1973. Biographical Memoirs of Fellows of the Royal Society, 20:217–234, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1974.0009.

Garrett:1975:WBF

[Gar75] Stephen Denis Garrett. William Brown, 17 February 1888– 18 January 1975. Biographical Memoirs of Fellows of the Royal Society, 21:155–174, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1975.
0003.

Garnham:1980:AJH

[Gar80] Percy Cyril Claude Garnham. Alexander John Haddow, 27 December 1912–26 December 1978. Biographical Memoirs of Fellows of the Royal Society, 26:225–254, November 1980. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1980.0006.

Garrett:1981:PWB

[Gar81a] Stephen Denis Garrett. Percy Wragg Brian, 5 September 1910–17 August 1979. Biographical Memoirs of Fellows of the Royal Society, 27:102–130, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0005.

Garton:1981:AGS

[Gar81b] William Reginald Stephen Garton. Allen Goodrich Shenstone, 27 July 1893–16 February 1980. Biographical Memoirs of Fellows of the Royal Society, 27:504–523, November 1981. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0020.

Garnham:1988:HES

[Gar88] Percy Cyril Claude Garnham. Henry Edward Shortt, 15 April 1887–9 November 1987. Biographical Memoirs of Fellows of the Royal Society, 34:713–751, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.

0023. See corrigendum [Ano90b].

Garland:1995:JTW

[Gar95] G. D. Garland. John Tuzo Wilson, 24 October 1908–15 April 1993. Biographical Memoirs of Fellows of the Royal Society, 41:534–552, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0032.

Garrett:2006:RWS

[Gar06] Chris Garrett. Robert William Stewart. 21 August 1923– 19 January 2005. Biographical Memoirs of Fellows of the Royal Society, 52:349–363, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0024.

Gardner:2015:SRG

[Gar15] Sir Richard Gardner FRS. Sir Robert Geoffrey Edwards CBE. 27 September 1925–10 April 2013. Biographical Memoirs of Fellows of the Royal Society, 61:81–102, January 2015. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0020.

Garside:2021:SFE

[Gar21] John Garside. Sir Frederick Edward Warner. 31 March 1910–3 July 2010. Biographical Memoirs of Fellows of the Royal Society, 71:513– 528, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2020.0048.

Gabor:1971:WJB

[GB71] Dennis Gabor and J. Brown. Willis Jackson, Baron Jackson of Burnley, 1904–1970. Biographical Memoirs of Fellows of the Royal Society, 17:379–398, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1971.
0014.

Gluckman:2013:SGC

[GB13] Sir Peter Gluckman KNZM FRS FMedSci FRSNZ and Tatjana Buklijas. Sir Graham Collingwood (Mont) Liggins. 24 June 1926–24 August 2010. Biographical Memoirs of Fellows of the Royal Society, 59:193–214, January 2013. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0039.

Gann:2014:NEM

[GB14] Alexander Gann and Jean Beggs CBE FRS. Noreen Elizabeth Murray CBE. 26 February 1935–12 May 2011. Biographical Memoirs of Fellows of the Royal Society, 60:349–374, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0009.

Goodwin:1985:CAH

[GBC85] Leonard George Goodwin and L. J. Bruce-Chwatt. Cecil Arthur Hoare, 6 March 1892–23 August 1984. Biographical Memoirs of Fellows of the Royal Society, 31:294–323, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1985.0011.

Gibbon:2018:MDK

[GCJM18] J. D. Gibbon, S. C. Cowley FRS, N. Joshi, and M. A. H. Mac-Callum. Martin David Kruskal. 28 September 1925–26 December 2006. Biographical Memoirs of Fellows of the Royal Society, 64: 261–284, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Green:2017:ADM

[GCK17] Malcolm L. H. Green FRS, Christopher C. Cummins, and James F. Kronauge. Alan Davison. 24 March 1936–14 November 2015. Biographical Memoirs of Fellows of the Royal Society, 63:197–213, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2017.0004.

Gale:1965:DDW

[GF65] Ernest Frederick Gale and Paul Gordon Fildes. Donald Devereux Woods, 1912–1964. Biographical Memoirs of Fellows of the Royal Society, 11:202–219, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1965.
0014.

Green:2000:SGW

[GG00] M. L. H. Green and W. P. Griffith. Sir Geoffrey Wilkinson. 14 July 1921–26 September 1996. Biographical Memoirs of Fellows of the Royal Society, 46:593–606, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0103.

Gordon:1996:DWJ

[GI96] G. Gordon and Ainsley Iggo. David Whitteridge, 22 June 1912–15 June 1994. Biographical Memoirs of Fellows of the Royal Society, 42:525–538, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0031.

Gray:2000:FSD

[GI00] P. Gray and K. J. Ivin. Frederick Sydney Dainton. 11 November 1914–5 December 1997. Biographical Memoirs of Fellows of the Royal Society, 46:85–124, January 2000. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0075.

Giannelli:2016:PEP

[Gia16] F. Giannelli FRCP. Paul Emanuel Polani. 1 January 1914– 18 February 2006. Biographical Memoirs of Fellows of the Royal Society, 62:461–481, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2016.
0003.

Gibson:1973:FJW

[Gib73] Quentin Howieson Gibson. Francis John Worsley Roughton, 1899– 1972. Biographical Memoirs of Fellows of the Royal Society, 19:562– 582, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1973.0020.

Gillespie:2020:FFB

[Gil20] T. A. Gillespie FRSE. Frank Featherstone Bonsall. 31 March 1920–22 February 2011. Biographical Memoirs of Fellows of the Royal Society, 69:63–77, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0007.

Ginzburg:1994:YBZ

[Gin94] Vitaly Lazarevich Ginzburg. Yakov Borissovich Zel’Dovich, 8 March 1914–2 December 1987. Biographical Memoirs of Fellows of the Royal Society, 40:429–441, November 1994. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1994.0049.

Gladstone:1973:PGF

[GKW73] G. P. Gladstone, B. C. J. G. Knight, and Graham Selby Wilson. Paul Gordon Fildes, 1882–1971. Biographical Memoirs of Fellows of the Royal Society, 19:317–347, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1973.0013.

Graham-Smith:2018:RDD

[GLD18] Francis Graham-Smith FRS, Andrew G. Lyne FRS, and Clive Dickinson. Rodney Deane Davies CBE. 8 January 1930–8 November 2015. Biographical Memoirs of Fellows of the Royal Society, 64: 149–162, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Glover:2020:HHR

[Glo20] Keith Glover FRS FREng. Howard Harry Rosenbrock. 16 December 1920–21 October 2010. Biographical Memoirs of Fellows of the Royal Society, 68:353–369, June 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0039.

Gray:2007:SCA

[GLR07] G. W. Gray, G. R. Luckhurst, and E. P. Raynes. Sivaramakrishna Chandrasekhar. 6 August 1930–8 March 2004. Biographical Memoirs of Fellows of the Royal Society, 53:127–141, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0015.

Glueckauf:1977:RSO

[Glu77] Eugen Glueckauf. Robert Spence, 7 October 1905–10 March 1976. Biographical Memoirs of Fellows of the Royal Society, 23:501–528, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1977.0019.

Glynn:2011:RDK

[Gly11] Ian Glynn FRS. Richard Darwin Keynes CBE. 14 August 1919–12 June 2010. Biographical Memoirs of Fellows of the Royal Society, 57:205–227, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2011.
0001.

Grunberg-Manago:1997:SOS

[GM97] Marianne Grunberg-Manago. Severo Ochoa. 24 September 1905– 1 November 1993. Biographical Memoirs of Fellows of the Royal Society, 43:351–365, January 1997. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0020.

Gregory:2006:HCL

[GM06] Richard L. Gregory and John N. Murrell. Hugh Christopher Longuet-Higgins. 11 April 1923–27 March 2004. Biographical Memoirs of Fellows of the Royal Society, 52:149–166, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0012.

Goswami:2021:DLF

[GM21] J. N. Goswami and J. D. Macdougall. Devendra Lal. 14 February 1929–1 December 2012. Biographical Memoirs of Fellows of the Royal Society, 70:263–281, June 2021. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0047.

Glover:2024:SAG

[GMP24] Keith Glover FRS, Jan Maciejowski, and Ian Postlethwaite. Sir Alistair G. J. MacFarlane. 9 May 1931–2 November 2021. Biographical Memoirs of Fellows of the Royal Society, 76:241–258, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2023.0034.

Greitzer:2019:SWR

[GN19] E. M. Greitzer FREng NAE and D. E. Newland FREng. Sir William Rede Hawthorne. 22 May 1913–16 September 2011. Biographical Memoirs of Fellows of the Royal Society, 66:309–328, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2017.0036.

Napper:2024:JRP

[GN24] Dougal Goodman and Simon Napper. John Roundell Palmer, fourth Earl of Selborne. 24 March 1940–12 February 2021. Biographical Memoirs of Fellows of the Royal Society, 76:277–289, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2023.0040.

Godwin:1957:AGT

[God57] Harry Godwin. Arthur George Tansley, 1871–1955. Biographical Memoirs of Fellows of the Royal Society, 3:227–246, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1957.0016.

Gold:1965:GCS

[Gol65] Ernest Gold. George Clarke Simpson, 1878–1965. Biographical Memoirs of Fellows of the Royal Society, 11:156–175, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1965.0011.

Goldstein:1966:TK

[Gol66] Sydney Goldstein. Theodore Von Kármán, 1881–1963. Biographical Memoirs of Fellows of the Royal Society, 12:334–365, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1966.0016.

Goldman:2017:AAD

[Gol17] Maurice Goldman. Anatole Abragam. 15 December 1914–8 June 2011. Biographical Memoirs of Fellows of the Royal Society, 63: 7–21, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2017.0026.

Goodeve:1972:FES

[Goo72] Charles Frederick Goodeve. Frank Edward Smith, 1876–1970. Biographical Memoirs of Fellows of the Royal Society, 18:525–548, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0019.

Goodwin:1982:JSK

[Goo82] Leonard George Goodwin. John Smith Knox Boyd, 18 September 1891–10 June 1981. Biographical Memoirs of Fellows of the Royal Society, 28:27–57, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.

Goodhew:2016:RES

[Goo16] P. J. Goodhew FREng. Raymond Edward Smallman CBE FREng. 4 August 1929–25 February 2015. Biographical Memoirs of Fellows of the Royal Society, 62:525–539, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0030.

Goodale:2020:LWM

[Goo20] Melvyn Alan Goodale FRS. Lawrence Weiskrantz. 28 March 1926–27 January 2018. Biographical Memoirs of Fellows of the Royal Society, 69:539–559, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0045.

Gordon:1996:RLM

[Gor96] Hugh Gordon. Richard Laurence Millington Synge, 28 October 1914–18 August 1994. Biographical Memoirs of Fellows of the Royal Society, 42:455–479, November 1996. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0028.

Gordon:2024:SWS

[Gor24] David Gordon. Sir William Stanley Peart. 31 March 1922–14 March 2019. Biographical Memoirs of Fellows of the Royal Society, 76: 291–314, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2023.0028.

Gowing:1990:LHB

[Gow90] Margaret Mary Gowing. Lord Hinton Bankside, 12 May 1901–22 June 1983. Biographical Memoirs of Fellows of the Royal Society, 36:217–239, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0031.

Groenewege:1983:JMB

[GP83] M. P. Groenewege and A. F. Peerdeman. Johannes Martin Bijvoet, 23 January 1892–4 March 1980. Biographical Memoirs of Fellows of the Royal Society, 29:26–41, November 1983.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1983.0002.

Glover:1978:RAM

[GPPG78] J. Glover, J. F. Pennock, G. A. J. Pitt, and Trevor Walworth Goodwin. Richard Alan Morton, 1899–1977. Biographical Memoirs of Fellows of the Royal Society, 24:408–442, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0013.

Granit:1985:HKH

[GR85] Ragnar Arthur Granit and F. Ratliff. Haldan Keffer Hartline, 22 December 1903–18 March 1983. Biographical Memoirs of Fellows of the Royal Society, 31:262–292, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1985.0010.

Greenfield:2000:HBO

[GR00a] A. D. M. Greenfield and I. C. Roddie. Henry Barcroft. 18 October 1904–11 January 1998. Biographical Memoirs of Fellows of the Royal Society, 46:1–17, January 2000. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0069.

Gurdon:2000:JNC

[GR00b] J. B. Gurdon and Barbara Rodbard. Joseph Needham, C.H. 9 December 1900–24 March 1995. Biographical Memoirs of Fellows of the Royal Society, 46:365–376, January 2000. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0091.

Goodby:2016:GWG

[GR16] John W. Goodby FRS and Peter Raynes FRS. George William Gray CBE MRIA FRSE. 4 September 1926–12 May 2013. Biographical Memoirs of Fellows of the Royal Society, 62:187–211, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2016.0001.

Gray:1969:JEH

[Gra69] James Gray. John Edward Harris, 1910–1968. Biographical Memoirs of Fellows of the Royal Society, 15:99–108, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1969.0005.

Grant:1982:GFM

[Gra82] J. K. Grant. Guy Frederic Marrian, 3 March 1904–24 July 1981. Biographical Memoirs of Fellows of the Royal Society, 28:347–378, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1982.0014.

Gray:1990:BHC

[Gra90] John Archibald Browne Gray. Bryan Harold Cabot Matthews, 14 June 1906–23 July 1986. Biographical Memoirs of Fellows of the Royal Society, 35:263–279, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Grafen:2004:WDH

[Gra04] Alan Grafen. William Donald Hamilton. 1 August 1936–7 March 2000. Biographical Memoirs of Fellows of the Royal Society, 50:109– 132, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2004.0009.

Gratzer:2010:SJR

[Gra10] Walter Gratzer. Sir John Royden Maddox. 27 November 1925–12 April 2009. Biographical Memoirs of Fellows of the Royal Society, 56:237–255, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.
0024.

Grace:2020:PGJ

[Gra20] John Grace FRSE. Paul Gordon Jarvis. 23 May 1935–5 February 2013. Biographical Memoirs of Fellows of the Royal Society, 68: 251–272, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0027.

Greenwood:1996:HJE

[Gre96] Norman Neill Greenwood. Harry Julius Emeléus, C. B. E., 22 June 1903–2 December 1993. Biographical Memoirs of Fellows of the Royal Society, 42:125–150, November 1996. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0009.

Green:1998:AER

[Gre98] D. H. Green. Alfred Edward Ringwood. 19 April 1930–12 November 1993. Biographical Memoirs of Fellows of the Royal Society, 44:351– 362, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1998.0023.

Green:2000:DFE

[Gre00a] M. L. H. Green. Dennis Frederick Evans. 27 March 1928– 6 November 1990. Biographical Memoirs of Fellows of the Royal Society, 46:165–175, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0078.

Greenwood:2000:JHC

[Gre00b] G. W. Greenwood. John Hugh Chesters, O.B.E. 6 October 1906–14 December 1994. Biographical Memoirs of Fellows of the Royal Society, 46:37–48, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0072.

Gregory:2001:OLZ

[Gre01] Richard L. Gregory. Oliver Louis Zangwill. 29 October 1913– 12 October 1987. Biographical Memoirs of Fellows of the Royal Society, 47:515–524, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0031.

Green:2006:SJW

[Gre06] John T. Green. Sir James William Longman Beament. 17 November 1921–10 March 2005. Biographical Memoirs of Fellows of the Royal Society, 52:15–27, January 2006. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0003.

Greenspan:2012:SBO

[Gre12] Ralph J. Greenspan. Seymour Benzer. 15 October 1921–30 November 2007. Biographical Memoirs of Fellows of the Royal Society, 58: 23–32, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2012.0031. See correction [Gre13].

Greenspan:2013:CSB

[Gre13] Ralph J. Greenspan. Correction for Seymour Benzer. 15 October 1921–30 November 2007. Biographical Memoirs of Fellows of the Royal Society, 59:437, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2012.

0045. See [Gre12].

Greenwood:2014:BAB

[Gre14] G. W. Greenwood FRS FREng. Bruce Alexander Bilby. 3 September 1922–20 November 2013. Biographical Memoirs of Fellows of the Royal Society, 60:57–73, January 2014. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0015.

Gregory:2019:DJG

[Gre19] Peter J. Gregory. Dennis James Greenland. 13 June 1930– 23 December 2012. Biographical Memoirs of Fellows of the Royal Society, 66:225–241, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0030.

Grillner:1995:RGO

[Gri95] S. Grillner. Ragnar Granit, 30 October 1900–11 March 1991. Biographical Memoirs of Fellows of the Royal Society, 41:184–197, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0012.

Grinter:2011:SFM

[Gri11] Roger Grinter. Stephen Finney Mason. 6 July 1923–11 December 2007. Biographical Memoirs of Fellows of the Royal Society, 57:253– 268, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2011.0010.

Grierson:2018:HSS

[Gri18] Donald Grierson OBE FRSB FRS. Harry Smith. 19 September 1935–9 February 2015. Biographical Memoirs of Fellows of the Royal Society, 64:387–399, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Green:1984:PHA

[GRT84] James Alexander Green, J. E. Roseblade, and John Griggs Thompson. Philip Hall, 11 April 1904–30 December 1982. Biographical Memoirs of Fellows of the Royal Society, 30:250–279, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1984.0009.

Graham-Smith:1986:MRS

[GS86] Francis Graham-Smith. Martin Ryle, 27 September 1918– 14 October 1984. Biographical Memoirs of Fellows of the Royal Society, 32:495–524, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0016.

Greig-Smith:1990:ASW

[GS90] P. Greig-Smith. Alexander Stuart Watt, 21 June 1892–2 March 1985. Biographical Memoirs of Fellows of the Royal Society, 35: 403–423, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Gulyas:2013:JSO

[GS13] Balázs Gulyás and Peter Somogyi FRS. János Szentágothai. 31 October 1912–8 September 1994. Biographical Memoirs of Fellows of the Royal Society, 59:383–406, January 2013. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0038.

Gardner:2018:SHH

[GS18] Sir Richard Gardner FRS and Eric Sidebottom. Sir Henry Harris. 28 January 1925–31 October 2014. Biographical Memoirs of Fellows of the Royal Society, 65:133–149, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0014.

Groner:2019:LSO

[GSL19] Yoram Groner, Pnina Sachs, and Joseph Lotem. Leo Sachs. 14 October 1924–12 December 2013. Biographical Memoirs of Fellows of the Royal Society, 66:355–375, June 2019. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0027.

Gibbard:2023:RGW

[GT23] P. L. Gibbard and C. Turner. Richard Gilbert West. 31 May 1926–30 December 2020. Biographical Memoirs of Fellows of the Royal Society, 74:457–480, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0017.

Greenland:2011:PHN

[GTK11] D. J. Greenland FRS, P. B. Tinker, and G. J. D. Kirk. Peter Hague Nye. 16 September 1921–13 February 2009. Biographical Memoirs of Fellows of the Royal Society, 57:315–326, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0017.

Guillery:1997:TPS

[Gui97] R. W. Guillery. Thomas Philip Stroud Powell. 13 July 1923– 8 February 1996. Biographical Memoirs of Fellows of the Royal Society, 43:413–427, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0022.

Guillery:2002:EGG

[Gui02] R. W. Guillery F.R.S. Edward George Gray. 11 January 1924–14 August 1999. Biographical Memoirs of Fellows of the Royal Society, 48:151–165, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2002.
0009.

Gunning:2000:JHH

[Gun00] Brian E. S. Gunning. John Heslop-Harrison. 10 February 1920–7 May 1998. Biographical Memoirs of Fellows of the Royal Society, 46:197–217, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0080.

Gurr:2012:ATJ

[Gur12] Michael Gurr. Anthony Trafford James. 6 March 1922–7 December 2006. Biographical Memoirs of Fellows of the Royal Society, 58:129– 150, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2011.0018.

Gutfreund:1996:KDA

[Gut96] Herbert Gutfreund. Keith Dalziel, 24 August 1921–7 January 1994. Biographical Memoirs of Fellows of the Royal Society, 42:112–122, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1996.0008.

Guthrie:2021:AEW

[Gut21] Sarah Guthrie. Anne Elizabeth Warner. 25 August 1940–16 May 2012. Biographical Memoirs of Fellows of the Royal Society, 70: 441–462, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2020.0046.

Goodwin:1992:ABD

[GV92] Leonard George Goodwin and Keith Vickerman. Ann Bishop, 19 December 1899–7 May 1990. Biographical Memoirs of Fellows of the Royal Society, 38:27–39, November 1992. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0002.

Golov:2018:HEH

[GV18] A. I. Golov and W. F. Vinen FRS. Henry Edgar Hall. 28 September 1928–4 December 2015. Biographical Memoirs of Fellows of the Royal Society, 65:115–132, December 2018. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0020.

Gibbons:2019:IRG

[GVS19] Wendy E. Gibbons, Ronald D. Vale, and Winfield S. Sale. Ian Read Gibbons. 30 October 1931–30 January 2018. Biographical Memoirs of Fellows of the Royal Society, 66:201–223, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0034.

Griffin:1976:ROR

[GW76] R. F. Griffin and Richard Van Der Riet Woolley. Roderick Oliver Redman, 17 July 1905–6 March 1975. Biographical Memoirs of Fellows of the Royal Society, 22:334–357, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0015.

Grimmett:2007:JMH

[GW07] Geoffrey Grimmett and Dominic Welsh. John Michael Hammersley. 21 March 1920–2 May 2004. Biographical Memoirs of Fellows of the Royal Society, 53:163–183, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0001.

Garrett:2020:WHM

[GW20] Chris Garrett FRS and Carl Wunsch ForMemRS. Walter Heinrich Munk. 19 October 1917–8 February 2019. Biographical Memoirs of Fellows of the Royal Society, 69:393–424, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0003.

Garratt:2011:CHB

[GWM11] J. R. Garratt, E. K. Webb, and S. McCarthy. Charles Henry Brian Priestley. 8 July 1915–18 May 1998. Biographical Memoirs of Fellows of the Royal Society, 57:349–378, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0015.

Gates:1962:HBS

[GY62] Sidney Barrington Gates and Alec David Young. Herbert Brian Squire, 1909–1961. Biographical Memoirs of Fellows of the Royal Society, 8:118–135, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1962.
0009.

Habell:1971:TS

[Hab71] K. J. Habell. Thomas Smith, 1883–1969. Biographical Memoirs of Fellows of the Royal Society, 17:681–687, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1971.0027.

Haines:2024:TWM

[Hai24] Sir Andy Haines. Thomas Wilson Meade. 21 January 1936– 24 October 2022. Biographical Memoirs of Fellows of the Royal Society, 76:259–276, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0036.

Hall:1999:PFW

[Hal99] Michael A. Hall. Philip Frank Wareing. 27 April 1914–29 March 1996. Biographical Memoirs of Fellows of the Royal Society, 45:507– 518, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0033.

Hannaford:2000:SAW

[Han00] Peter Hannaford. Sir Alan Walsh. 19 December 1916–3 August 1998. Biographical Memoirs of Fellows of the Royal Society, 46:533– 564, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0100.

Hancock:2018:IWM

[Han18] Gus Hancock. Ian William Murison Smith. 15 June 1937–8 November 2016. Biographical Memoirs of Fellows of the Royal Society, 64: 401–419, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Harington:1956:HK

[Har56] Charles Robert Harington. Harold King, 1887–1956. Biographical Memoirs of Fellows of the Royal Society, 2:157–171, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1956.0012.

Harland:1961:WLB

[Har61] Sydney Cross Harland. William Lawrence Balls, 1882–1960. Biographical Memoirs of Fellows of the Royal Society, 7:1–16, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1961.0001.

Harris:1963:HHT

[Har63] Thomas Maxwell Harris. Hugh Hamshaw Thomas, 1885–1962. Biographical Memoirs of Fellows of the Royal Society, 9:287–299, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1963.0015.

Hartley:1966:WAS

[Har66] Harold Brewer Hartley. William Arthur Stanier, 1876–1965. Biographical Memoirs of Fellows of the Royal Society, 12:488–502, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0024.

Harris:1973:LSP

[Har73] Harry Harris. Lionel Sharples Penrose, 1898–1972. Biographical Memoirs of Fellows of the Royal Society, 19:521–561, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1973.0019.

Harley:1981:GEB

[Har81] John Laker Harley. Geoffrey Emett Blackman, 17 April 1903– 8 February 1980. Biographical Memoirs of Fellows of the Royal Society, 27:45–82, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0003.

Harris:1982:JHM

[Har82] Henry Harris. Joachim Hämmerling, 9 March 1901–5 August 1980. Biographical Memoirs of Fellows of the Royal Society, 28:111–124, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1982.0005.

Harrison:1987:LHM

[Har87] Richard John Harrison. Leonard Harrison Matthews, 12 June 1901–27 November 1986. Biographical Memoirs of Fellows of the Royal Society, 33:411–442, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0016.

Harley:1991:SDG

[Har91] John Laker Harley. Stephen Denis Garrett, 1 November 1906– 26 December 1989. Biographical Memoirs of Fellows of the Royal Society, 37:177–195, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0009.

Harrison:1999:REF

[Har99] Bryan D. Harrison. Richard Ellis Ford Matthews, O.N.Z. 20 November 1921–19 February 1995. Biographical Memoirs of Fellows of the Royal Society, 45:295–313, January 1999. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0020.

Hargreaves:2004:HBG

[Har04] C. M. Hargreaves. Hendrik Brugt Gerhard Casimir Knight of the Order of the Nederlandse Leeuw Commander in the Order of Orange Nassau. 15 July 1909–4 May 2000. Biographical Memoirs of Fellows of the Royal Society, 50:39–45, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2004.0004.

Hardingham:2018:IHM

[Har18] Tim Hardingham. Isabella Helen Mary Muir CBE. 20 August 1920–28 November 2005. Biographical Memoirs of Fellows of the Royal Society, 65:283–298, December 2018. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0042.

Harris:2020:FSS

[Har20] Neil R. P. Harris. Frank Sherwood ‘Sherry’ Rowland. 28 June 1927–10 March 2012. Biographical Memoirs of Fellows of the Royal Society, 68:371–383, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0032.

Hastings:2016:JBM

[Has16] Stuart Hastings. John Bryce McLeod. 23 December 1929– 20 August 2014. Biographical Memoirs of Fellows of the Royal Society, 62:381–407, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0031.

Haughton:1956:ALH

[Hau56] Sidney Henry Haughton. Arthur Lewis Hall, 1872–1955. Biographical Memoirs of Fellows of the Royal Society, 2:139–148, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1956.0010.

Hawkins:1956:WKS

[Haw56] Herbert Leader Hawkins. Willliam Kingdon Spencer, 1878–1955. Biographical Memoirs of Fellows of the Royal Society, 2:291–298, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1956.0020.

Haworth:1960:GMB

[Haw60] Robert Downs Haworth. George Macdonald Bennett, 1892–1959. Biographical Memoirs of Fellows of the Royal Society, 5:23–36, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Hawthorne:1976:HRR

[Haw76] William Rede Hawthorne. Harry Ralph Ricardo, 26 January 1885–18 May 1974. Biographical Memoirs of Fellows of the Royal Society, 22:358–380, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1976.
0016.

Hayman:1965:WWR

[Hay65] Walter Kurt Hayman. Werner Wolfgang Rogosinski, 1894–1964. Biographical Memoirs of Fellows of the Royal Society, 11:135–145, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1965.0009.

Hayes:1982:MLH

[Hay82] William Hayes. Max Ludwig Henning Delbrück, 4 September 1906–10 March 1981. Biographical Memoirs of Fellows of the Royal Society, 28:59–90, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0003.

Hayman:1985:JMW

[Hay85] Walter Kurt Hayman. John Macnaghten Whittaker, 7 March 1905–29 January 1984. Biographical Memoirs of Fellows of the Royal Society, 31:653–666, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0023.

Hayman:2000:DML

[Hay00] W. K. Hayman. Dame Mary (Lucy) Cartwright, D.B.E. 17 December 1900–3 April 1998. Biographical Memoirs of Fellows of the Royal Society, 46:19–35, January 2000. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0070.

Hayman:2002:ACO

[Hay02] W. K. Hayman F.R.S. Albert Cyril Offord. 9 June 1906–4 June 2000. Biographical Memoirs of Fellows of the Royal Society, 48:341– 355, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2002.0020.

Halberstam:1971:HD

[HB71] H. Halberstam and D. A. Burgess. Harold Davenport, 1907–1969. Biographical Memoirs of Fellows of the Royal Society, 17:159–192, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0006.

Heath-Brown:2020:CHA

[HB20] D. R. Heath-Brown. Christopher Hooley. 7 August 1928– 13 December 2018. Biographical Memoirs of Fellows of the Royal Society, 69:225–246, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0027.

Hitchman:2022:WJA

[HB22] Michael L. Hitchman FRSE and Philip N. Bartlett FRS. Wyndham John Albery. 5 April 1936–3 December 2013. Biographical Memoirs of Fellows of the Royal Society, 73:13–39, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0018.

Horton:2018:JMA

[HCB18] Peter Horton FRS, Wah Soon Chow, and Christopher Barrett. Joan Mary Anderson. 12 May 1932–28 August 2015. Biographical Memoirs of Fellows of the Royal Society, 65:7–29, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0006.

Hawthorne:1973:HC

[HCH73] William Rede Hawthorne, H. Cohen, and A. R. Howell. Hayne Constant, 1904–1968. Biographical Memoirs of Fellows of the Royal Society, 19:268–279, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1973.
0010.

Hyde:1992:JSA

[HD92] B. G. Hyde and Peter Day. John Stuart Anderson, 9 January 1908–25 December 1990. Biographical Memoirs of Fellows of the Royal Society, 38:1–26, November 1992. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0001.

Heap:1998:SBA

[HD98] R. B. Heap and R. G. Dyer. Sir Barry Albert Cross, C. B. E. 17 March 1925–27 April 1994. Biographical Memoirs of Fellows of the Royal Society, 44:95–108, January 1998. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0007.

Heitler:1961:ES

[Hei61] Walter Heinrich Heitler. Erwin Schrödinger, 1887–1961. Biographical Memoirs of Fellows of the Royal Society, 7:221–228, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769408;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1961.0017.

Henderson:1990:HAP

[Hen90] William Macgregor Henderson. Hugh Algernon Percy, 10th Duke of Northumberland, 6 April 1914–11 October 1988. Biographical Memoirs of Fellows of the Royal Society, 35:355–359, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Herzberg:1955:HK

[Her55] Gerhard Herzberg. Heinrich Kayser, 1853–1940. Biographical Memoirs of Fellows of the Royal Society, 1:135–143, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0010.

Herzberg:1981:CSB

[Her81] Gerhard Herzberg. Caryle Smith Beals, 29 June 1899–2 July 1979. Biographical Memoirs of Fellows of the Royal Society, 27:28–44, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1981.0002.

Herzberg:1982:AED

[Her82] Gerhard Herzberg. Alexander Edgar Douglas, 12 April 1916–26 July 1981. Biographical Memoirs of Fellows of the Royal Society, 28:91–109, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0004.

Herring:2014:MRC

[Her14] Peter J. Herring. Malcolm Roy Clarke. 24 October 1930–10 May 2013. Biographical Memoirs of Fellows of the Royal Society, 60:107– 125, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2014.0011.

Hewitt:1999:JGC

[Hew99] G. F. Hewitt. John Gordon Collier, F.R.Eng. 22 January 1935– 18 November 1995. Biographical Memoirs of Fellows of the Royal Society, 45:67–75, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0006.

Hewish:2002:JSH

[Hew02] Antony Hewish F.R.S. James Stanley Hey, M.B.E. 3 May 1909–27 February 2000. Biographical Memoirs of Fellows of the Royal Society, 48:167–178, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2002.

Heyman:1987:JFB

[Hey87] J. Heyman. John Fleetwood Baker, Baron Baker of Windrush, 19 March–9 September 1985. Biographical Memoirs of Fellows of the Royal Society, 33:1–20, December 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0001.

Heyde:1991:PAP

[Hey91] C. C. Heyde. Patrick Alfred Pierce Moran, 14 July 1917– 19 September 1988. Biographical Memoirs of Fellows of the Royal Society, 37:365–379, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0018.

Holliday:2006:JRS

[HF06] Robin Holliday and Richard B. Flavell. John Robert Stanley Fincham. 11 August 1926–9 February 2005. Biographical Memoirs of Fellows of the Royal Society, 52:83–95, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0007. See correction [HF08].

Hinde:2007:JJR

[HF07] R. A. Hinde and J. L. Finney. Joseph (Józef) Rotblat. 4 November 1908–31 August 2005. Biographical Memoirs of Fellows of the Royal Society, 53:309–326, January 2007. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0023.

Holliday:2008:CHF

[HF08] Robin Holliday and Richard B. Flavell. Correction for Holliday and Flavell, John Robert Stanley Fincham. 11 August 1926–9 February 2005. Biographical Memoirs of Fellows of the Royal Society, 54: 483, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2008.3000. See [HF06].

Henderson:2009:DMB

[HF09] R. Henderson FRS and N. P. Franks. David Mervyn Blow. 27 June 1931–8 June 2004. Biographical Memoirs of Fellows of the Royal Society, 55:13–35, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.
0022.

Hartley:1970:TRM

[HG70] Harold Brewer Hartley and Dennis Gabor. Thomas Ralph Merton, 1888–1969. Biographical Memoirs of Fellows of the Royal Society, 16:421–440, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1970.
0017.

Heap:2011:ADB

[HG11] Sir Brian Heap CBE FRS and Gregory Gregoriadis. Alec Douglas Bangham. 10 November 1921–9 March 2010. Biographical Memoirs of Fellows of the Royal Society, 57:25–43, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0004.

Hutchison:1981:CDE

[HGM81] William Kenneth Hutchison, J. A. Gray, and Harrie Stewart Wilson Massey. Charles Drummond Ellis, 11 August 1895–10 January 1980. Biographical Memoirs of Fellows of the Royal Society, 27:199– 233, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1981.0009.

Hutchison:1974:FJD

[HH74] William Kenneth Hutchison and D. Hebden. Frederick James Dent, 1905–1973. Biographical Memoirs of Fellows of the Royal Society, 20:154–180, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1974.
0006.

Hirst:1988:BKO

[HH88] John Malcolm Hirst and Bryan Desmond Harrison. Basil Kassanis, 16 October 1911–23 March 1985. Biographical Memoirs of Fellows of the Royal Society, 34:345–371, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0013.

Heslop-Harrison:1995:EAB

[HH95] John Heslop-Harrison. Eric Ashby, Baron Ashby, of Brandon, Suffolk, Kt, 24 August 1904–22 October 1992. Biographical Memoirs of Fellows of the Royal Society, 41:2–18, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0001.

Hey:2022:DDE

[HHP22] Michael J. Hey, Jonathan D. Hirst, and Martyn Poliakoff FRS FREng. Daniel Douglas Eley. 1 October 1914–3 September 2015. Biographical Memoirs of Fellows of the Royal Society, 73:227–249, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2022.0004.

Hibbert:2021:PGB

[Hib21] Alan Hibbert. Philip George Burke. 18 October 1932–4 June 2019. Biographical Memoirs of Fellows of the Royal Society, 70:57–77, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2020.0043.

Higgs:2024:SDJ

[Hig24] Douglas R. Higgs FRS. Sir David John Weatherall. 9 March 1933–8 December 2018. Biographical Memoirs of Fellows of the Royal Society, 76:455–476, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0031.

Hills:1958:WNB

[Hil58] Edwin Sherbon Hills. William Noel Benson, 1885–1957. Biographical Memoirs of Fellows of the Royal Society, 4:26–33, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0003.

Hill:1972:LWM

[Hil72] Robert Hill. Leslie William Mapson, 1907–1970. Biographical Memoirs of Fellows of the Royal Society, 18:427–444, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1972.0014.

Hill:1987:ESH

[Hil87] Dorothy Hill. Edwin Sherbon Hills, 31 August 1906–2 May 1986. Biographical Memoirs of Fellows of the Royal Society, 33:289–323, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1987.0011.

Hill:1988:NLF

[Hil88] John Mcgregor Hill. Norman Laurence Franklin, 1 September 1924–7 November 1986. Biographical Memoirs of Fellows of the Royal Society, 34:205–219, December 1988. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0008.

Hill:1990:ARF

[Hil90a] William George Hill. Alan Robertson, 21 February 1920–25 April 1989. Biographical Memoirs of Fellows of the Royal Society, 36:463– 488, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1990.0040.

Hill:1990:SWD

[Hil90b] William George Hill. Sewall Wright, 21 December 1889–3 March 1988. Biographical Memoirs of Fellows of the Royal Society, 36:567– 579, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1990.0044.

Hiley:1997:DJB

[Hil97] B. J. Hiley. David Joseph Bohm. 20 December 1917–27 October 1992. Biographical Memoirs of Fellows of the Royal Society, 43:107– 131, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1997.0007.

Hilsum:2011:SGG

[Hil11] Cyril Hilsum FRS. Sir Gareth Gwyn Roberts. 16 May 1940– 6 February 2007. Biographical Memoirs of Fellows of the Royal Society, 57:379–393, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2011.
0016.

Hindle:1958:JGK

[Hin58] Edward Hindle. John Graham Kerr, 1869–1957. Biographical Memoirs of Fellows of the Royal Society, 4:155–166, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0014.

Hinton:1960:CDG

[Hin60] Christopher Hinton. Claude Dixon Gibb, 1898–1959. Biographical Memoirs of Fellows of the Royal Society, 5:87–94, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Hinde:1987:WHT

[Hin87] Robert Aubrey Hinde. William Homan Thorpe, 1 April 1902–7 April 1986. Biographical Memoirs of Fellows of the Royal Society, 33:619–639, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0022.

Hinde:1990:NTA

[Hin90] Robert Aubrey Hinde. Nikolaas Tinbergen, 15 April 1907– 21 December 1988. Biographical Memoirs of Fellows of the Royal Society, 36:547–565, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0043.

Hirst:1963:JR

[Hir63] Edmund Langley Hirst. John Read, 1884–1963. Biographical Memoirs of Fellows of the Royal Society, 9:237–260, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1963.0013.

Hirst:1990:PHG

[Hir90] John Malcolm Hirst. Philip Herries Gregory, 24 July 1907–9 February 1986. Biographical Memoirs of Fellows of the Royal Society, 35:

151–177, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Hirsch:2015:DJH

[Hir15] Sir Peter Hirsch FRS. David John Hugh Cockayne. 19 March 1942–22 December 2010. Biographical Memoirs of Fellows of the Royal Society, 61:53–79, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2014.
0025.

Hitchin:2006:AGW

[Hit06] Nigel J. Hitchin. Arthur Geoffrey Walker. 17 July 1909–31 March 2001. Biographical Memoirs of Fellows of the Royal Society, 52:413– 421, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2006.0028.

Hitchin:2014:SSC

[Hit14] Nigel J. Hitchin FRS. Shiing-Shen Chern. 26 October 1911– 3 December 2004. Biographical Memoirs of Fellows of the Royal Society, 60:75–85, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2014.
0018.

Hitchin:2020:SMA

[Hit20] Nigel Hitchin. Sir Michael Atiyah OM. 22 April 1929–11 January 2019. Biographical Memoirs of Fellows of the Royal Society, 69: 9–35, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2020.0001.

Houston:2011:PNR

[HJ11] Peter A. Houston and Roy Jackson FRS. Peter Neville Robson OBE FREng. 23 November 1930–10 January 2010. Biographical Memoirs of Fellows of the Royal Society, 57:423–434, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0013.

Holmes:2023:AGM

[HKB23] Andrew B. Holmes FRS, Michael L. Klein FRS, and Ray H. Baughman. Alan Graham MacDiarmid. 14 April 1927–7 February 2007.

Biographical Memoirs of Fellows of the Royal Society, 74:259–282, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2022.0045.

Haworth:1963:JIO

[HL63] Robert Downs Haworth and A. H. Lamberton. James Irvine Orme Masson, 1887–1962. Biographical Memoirs of Fellows of the Royal Society, 9:205–221, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1963.
0011.

Hohenberg:2018:WKM

[HL18] Pierre C. Hohenberg and James S. Langer. Walter Kohn. 9 March 1923–19 April 2016. Biographical Memoirs of Fellows of the Royal Society, 64:249–259, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Hammock:2023:JEC

[HL23] Bruce D. Hammock and Qing X. Li. John Edward Casida. 22 December 1929–30 June 2018. Biographical Memoirs of Fellows of the Royal Society, 74:63–86, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0016.

Hartmann:1982:EHA

[HLH82] H. Hartmann and Hugh Christopher Longuet-Higgins. Erich Hückel, 9 August 1896–16 February 1980. Biographical Memoirs of Fellows of the Royal Society, 28:153–162, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0008.

Hall:1977:BMJ

[HM77] Arnold Alexander Hall and Morien Bedford Morgan. Bennett Melvill Jones, 28 January 1887–31 October 1975. Biographical Memoirs of Fellows of the Royal Society, 23:252–282, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0010.

Haslam:2003:TSS

[HM03] Edwin Haslam and David G. Morris. Thomas Stevens Stevens. 8 October 1900–13 November 2000. Biographical Memoirs of Fellows of the Royal Society, 49:521–535, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0031.

Haworth:1994:RRJ

[HNW94] Lionel Haworth, M. A. Nedham, and G. L. Wilde. Robin Ralph Jamison, 12 July 1912–18 March 1991. Biographical Memoirs of Fellows of the Royal Society, 40:171–194, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1994.0034.

Hodge:1956:HFB

[Hod56] William Vallance Douglas Hodge. Henry Frederick Baker, 1866– 1956. Biographical Memoirs of Fellows of the Royal Society, 2:49– 68, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1956.0004.

Hodge:1973:SL

[Hod73] William Vallance Douglas Hodge. Solomon Lefschetz, 1884–1972. Biographical Memoirs of Fellows of the Royal Society, 19:433–453, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0016.

Hodgkin:1975:KLJ

[Hod75] Dorothy Mary Crowfoot Hodgkin. Kathleen Lonsdale, 28 January 1903–1 April 1971. Biographical Memoirs of Fellows of the Royal Society, 21:447–484, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1975.
0014.

Hodgkin:1979:EDA

[Hod79] Alan Hodgkin. Edgar Douglas Adrian, Baron Adrian of Cambridge, 30 November 1889–4 August 1977. Biographical Memoirs of Fellows of the Royal Society, 25:1–73, November 1979.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0002.

Hodgkin:1980:JDB

[Hod80] Dorothy Mary Crowfoot Hodgkin. John Desmond Bernal, 10 May 1901–15 September 1971. Biographical Memoirs of Fellows of the Royal Society, 26:16–84, November 1980. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1980.0002.

Hoffman:2007:GTS

[Hof07] Darleane C. Hoffman. Glenn Theodore Seaborg. 19 April 1912– 25 February 1999. Biographical Memoirs of Fellows of the Royal Society, 53:327–338, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2007.
0021.

Hogben:1974:FAE

[Hog74] Lancelot Thomas Hogben. Francis Albert Eley Crew, 1886–1973. Biographical Memoirs of Fellows of the Royal Society, 20:135–153, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1974.0005.

Hogg:1987:MHM

[Hog87] Quintin Mcgarel Hogg. Maurice Harold Macmillan, First Earl of Stockton, 10 February 1894–29 December 1986. Biographical Memoirs of Fellows of the Royal Society, 33:375–385, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0014.

Holder:1966:WRM

[Hol66] Douglas William Holder. William Richard Morris, First Viscount Nuffield, 1877–1963. Biographical Memoirs of Fellows of the Royal Society, 12:387–404, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1966.

Holroyd:1971:AFB

[Hol71] Ronald Holroyd. Alexander Fleck, Baron Fleck of Saltcoats, 1889– 1968. Biographical Memoirs of Fellows of the Royal Society, 17:242– 254, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1971.0010.

Holt:1994:TGP

[Hol94] John Riley Holt. Thomas Gerald Pickavance, 19 October 1915– 12 November 1991. Biographical Memoirs of Fellows of the Royal Society, 39:303–323, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Holt:1996:JMC

[Hol96] John Riley Holt. James Macdonald Cassels, 9 September 1924– 19 October 1994. Biographical Memoirs of Fellows of the Royal Society, 42:62–78, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0005.

Holmes:2001:SJC

[Hol01] K. C. Holmes. Sir John Cowdery Kendrew. 24 March 1917– 23 August 1997. Biographical Memoirs of Fellows of the Royal Society, 47:311–332, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0018.

Honeycombe:1995:NJP

[Hon95] Robert William Kerr Honeycombe. Norman James Petch, 13 February 1917–9 December 1992. Biographical Memoirs of Fellows of the Royal Society, 41:342–357, November 1995. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0021.

Hopkinson:1996:HHS

[Hop96] D. A. Hopkinson. Harry Harris, 30 September 1919–17 July 1994. Biographical Memoirs of Fellows of the Royal Society, 42:153–170, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1996.0010.

Hope:2023:RMG

[Hop23] Ronald Anthony Hope. Raymond Michael Gaze. 22 June 1927– 11 September 2012. Biographical Memoirs of Fellows of the Royal Society, 74:185–209, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0026.

Horne:1986:OAK

[Hor86] Michael Rex Horne. Oleg Alexander Kerensky, 16 April 1905–25 June 1984. Biographical Memoirs of Fellows of the Royal Society, 32:321–353, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0010.

Horowitz:1995:GWB

[Hor95] N. H. Horowitz. George Wells Beadle, 23 October 1903–9 June 1989. Biographical Memoirs of Fellows of the Royal Society, 41:44– 54, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1995.0003.

Horton:2014:DAW

[Hor14] Peter Horton FRS. David Alan Walker. 18 August 1928– 13 February 2012. Biographical Memoirs of Fellows of the Royal Society, 60:413–432, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2014.
0007.

Howie:2001:ZSB

[How01] Archie Howie. Zbigniew Stanislaw Basinski, O.C. 28 April 1928–12 August 1999. Biographical Memoirs of Fellows of the Royal Society, 47:1–17, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0001.

Howie:2015:AKC

[How15] A. Howie CBE FRS. Anthony Kelly CBE DL. 25 January 1929–3 June 2014. Biographical Memoirs of Fellows of the Royal Society, 61:247–264, January 2015. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0009.

Harrison:1960:AG

[HP60] G. P. Harrison and Alfred John Sutton Pippard. Alexander Gibb, 1872–1958. Biographical Memoirs of Fellows of the Royal Society, 5: 75–86, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Huppert:2001:DGC

[HP01] Herbert E. Huppert and Nigel Peake. David George Crighton. 15 November 1942–12 April 2000. Biographical Memoirs of Fellows of the Royal Society, 47:105–123, January 2001. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0007.

Hirsch:2008:SJW

[HP08] Peter Hirsch and D. W. Pashley. Sir James Woodham Menter. 22 August 1921–18 July 2006. Biographical Memoirs of Fellows of the Royal Society, 54:193–214, January 2008. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0015. See correction [HP09].

Hirsch:2009:CSJ

[HP09] Peter Hirsch FRS and D. W. Pashley FRS. Correction for Sir James Woodham Menter. 22 August 1921–18 July 2006. Biographical Memoirs of Fellows of the Royal Society, 55:361, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2008.0060. See [HP08].

Himsworth:1972:CRH

[HPR72] Harold Percival Himsworth and Rosalind Venetia Pitt-Rivers. Charles Robert Harington, 1897–1972. Biographical Memoirs of Fellows of the Royal Society, 18:266–308, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1972.0009.

Hush:2018:DPC

[HR18] Noel S. Hush AO FRS FAA and Leo Radom FAA. David Parker Craig AO FAA. 23 December 1919–1 July 2015. Biographical Memoirs of Fellows of the Royal Society, 64:107–129, June 2018. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Hattersley-Smith:2007:DJM

[HS07] Geoffrey Hattersley-Smith. Digby Johns McLaren. 11 December 1919–8 December 2004. Biographical Memoirs of Fellows of the Royal Society, 53:237–246, January 2007. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0007.

Hirst:1970:SP

[HT70] Edmund Langley Hirst and J. R. Turvey. Stanley Peat, 1902–1969. Biographical Memoirs of Fellows of the Royal Society, 16:441–462, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1970.0018.

Hughes-Jones:2003:RAS

[HT03] Nevin Hughes-Jones FRS and Patricia Tippett. Ruth Ann Sanger. 6 June 1918–4 June 2001. Biographical Memoirs of Fellows of the Royal Society, 49:461–473, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0027.

Hartl:2014:JFC

[HT14] Daniel L. Hartl and Rayla Greenberg Temin. James Franklin Crow. 18 January 1916–4 January 2012. Biographical Memoirs of Fellows of the Royal Society, 60:151–167, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0004.

Hill:2016:RJP

[HT16] H. A. O. Hill FRS and A. J. Thomson OBE FRS. Robert Joseph Paton Williams MBE. 25 February 1926–21 March 2015. Biographical Memoirs of Fellows of the Royal Society, 62:541–570, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0020.

Hlusko:2019:ACW

[HU19] Leslea J. Hlusko and Peter S. Ungar. Alan Cyril Walker. 23 August 1938–20 November 2017. Biographical Memoirs of Fellows of the Royal Society, 67:449–467, December 2019. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0017.

Huang:2018:RYT

[Hua18] Christopher L.-H. Huang DM (Oxford), ScD (Cambridge), FESC. Roger Yonchien Tsien. 1 February 1952–24 August 2016. Biographical Memoirs of Fellows of the Royal Society, 65:405–428, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2018.0013.

Hubbard:1971:WBT

[Hub71] C. E. Hubbard. William Bertram Turrill, 1890–1961. Biographical Memoirs of Fellows of the Royal Society, 17:689–712, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1971.0028.

Hubbard:1975:JHA

[Hub75] C. E. Hubbard. John Hutchinson, 7 April 1884–2 September 1972. Biographical Memoirs of Fellows of the Royal Society, 21:345–365, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1975.0009.

Hulst:1994:JHO

[Hul94] Hendrik Christoffel Van De Hulst. Jan Hendrik Oort, 28 April 1900–5 November 1992. Biographical Memoirs of Fellows of the Royal Society, 40:319–326, November 1994. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1994.0042.

Humble:2011:AKH

[Hum11] Peter Humble. Alan Kenneth Head AO FAA. 10 August 1925–9 January 2010. Biographical Memoirs of Fellows of the Royal Society, 57:147–166, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2010.
0013.

Hunt:2003:TBB

[Hun03] J. C. R. Hunt FRS. Thomas Brooke Benjamin. 15 April 1929–16 August 1995 Elected FRS 1966. Biographical Memoirs of Fellows of the Royal Society, 49:39–67, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0003.

Hunter:2008:EJC

[Hun08] R. H. F. Hunter. Ernest John Christopher Polge. 16 August 1926–17 August 2006. Biographical Memoirs of Fellows of the Royal Society, 54:275–296, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.
0006.

Hunter:2017:LJC

[Hun17] Michael Hunter FBA. Lisa Jardine CBE. 12 April 1944–25 October 2015. Biographical Memoirs of Fellows of the Royal Society, 63:363– 375, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2017.0015.

Hutchinson:1984:SCH

[Hut84] Joseph Burtt Hutchinson. Sydney Cross Harland, 19 July 1891– 8 November 1982. Biographical Memoirs of Fellows of the Royal Society, 30:297–316, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1984.
0011.

Hutchinson:2023:AGE

[Hut23] John W. Hutchinson ForMemRS. Anthony G. Evans. 4 December 1942–9 September 2009. Biographical Memoirs of Fellows of the Royal Society, 74:143–159, April 2023. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0027.

Huxley:1992:KSC

[Hux92] Andrew Fielding Huxley. Kenneth Stewart Cole, 10 July 1900–18 April 1984. Biographical Memoirs of Fellows of the Royal Society, 38:97–110, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1992.
0005.

Huxley:1997:RHA

[Hux97] Andrew Huxley. Richard Hume Adrian, D. L. 2nd Baron Adrian of Cambridge. 16 October 1927–4 April 1995. Biographical Memoirs of Fellows of the Royal Society, 43:15–30, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0002.

Huxley:2000:SAL

[Hux00] Sir Andrew Huxley. Sir Alan Lloyd Hodgkin, O.M., K.B.E. 5 February 1914–20 December 1998. Biographical Memoirs of Fellows of the Royal Society, 46:219–241, January 2000. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0081.

Huxley:2003:DKH

[Hux03] Andrew Huxley FRS. David Keynes Hill. 23 July 1915–18 August 2002 Elected FRS 1972. Biographical Memoirs of Fellows of the Royal Society, 49:197–212, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0012.

Hartquist:2020:ADJ

[HV20] Thomas W. Hartquist and Ewine F. Van Dishoeck ForMemRS. Alexander Dalgarno. 5 January 1928–9 April 2015. Biographical Memoirs of Fellows of the Royal Society, 69:145–174, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0009.

Haworth:1971:AR

[HW71] Robert Downs Haworth and A. E. Whalley. Alexander Robertson, 1896–1970. Biographical Memoirs of Fellows of the Royal Society, 17:617–642, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1971.
0024.

Houghton:1977:GMB

[HW77] John Theodore Houghton and C. D. Walshaw. Gordon Miller Bourne Dobson, 25 February 1889–11 March 1976. Biographical Memoirs of Fellows of the Royal Society, 23:41–57, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0003.

Hirschfeld:1987:TGR

[HW87] J. W. P. Hirschfeld and G. E. Wall. Thomas Gerald Room, 10 November 1902–2 April 1986. Biographical Memoirs of Fellows of the Royal Society, 33:573–601, December 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0020.

Hoare:2004:RMN

[HW04] Tony Hoare and Maurice V. Wilkes. Roger Michael Needham CBE FREng. 9 February 1935–1 March 2003. Biographical Memoirs of Fellows of the Royal Society, 50:183–199, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2004.0014.

Hansen:2010:JRK

[HW10] David E. Hansen and Robert J. P. Williams. Jeremy Randall Knowles CBE. 28 April 1935–3 April 2008. Biographical Memoirs of Fellows of the Royal Society, 56:171–187, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0022.

Holmes:2017:HEH

[HW17] Kenneth C. Holmes FRS and Alan Weeds. Hugh Esmor Huxley MBE. 25 February 1924–25 July 2013. Biographical Memoirs of Fellows of the Royal Society, 63:309–344, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0011.

Hyndman:2015:EIF

[Hyn15] Roy Hyndman FRSC. Edward Irving FRSC CM. 27 May 1927– 25 February 2014. Biographical Memoirs of Fellows of the Royal Society, 61:183–201, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0004.

Isacker:2021:JPE

[IG21] Piet Van Isacker and William Gelletly. James Philip Elliott. 27 July 1929–21 October 2008. Biographical Memoirs of Fellows of the Royal Society, 71:123–143, December 2021. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0006.

Iggo:2006:ASP

[Igg06] Ainsley Iggo. Autar Singh Paintal. 24 September 1925– 21 December 2004. Biographical Memoirs of Fellows of the Royal Society, 52:251–262, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0018.

Imura:1995:SNF

[Imu95] Hiroo Imura. Shosaku Numa, 7 February 1929–15 February 1992. Biographical Memoirs of Fellows of the Royal Society, 41:310–314, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0019.

Ingold:1964:EDH

[Ing64] Christopher Kelk Ingold. Edward David Hughes, 1906–1963. Biographical Memoirs of Fellows of the Royal Society, 10:147–182, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0010.

Ingold:1968:EET

[Ing68] Christopher Kelk Ingold. Eustace Ebenezer Turner, 1893–1966. Biographical Memoirs of Fellows of the Royal Society, 14:449–467, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1968.0020.

Irvine:2023:RMC

[Irv23] Robin Irvine FRS. Rex Malcolm Chaplin Dawson. 3 June 1924–29 March 2021. Biographical Memoirs of Fellows of the Royal Society, 75:75–92, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0040.

Isler:1978:PKA

[Isl78] Otto Isler. Paul Karrer, 21 April 1889–18 June 1971. Biographical Memoirs of Fellows of the Royal Society, 24:245–321, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1978.0009.

Iversen:2006:JAM

[Ive06] Leslie Iversen. Julius Axelrod. 30 May 1912–29 December 2004. Biographical Memoirs of Fellows of the Royal Society, 52:1–13, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2006.0002.

Jacobson:2000:SOH

[Jac00] C.-O. Jacobson. Sven Otto Hörstadius. 18 February 1898–16 June 1996. Biographical Memoirs of Fellows of the Royal Society, 46:243– 256, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0082.

James:1967:WS

[Jam67] William Owen James. Walter Stiles, 1886–1966. Biographical Memoirs of Fellows of the Royal Society, 13:343–357, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1967.0018.

James:1990:JFA

[Jam90] Ioan Mackenzie James. John Frank Adams, 5 November 1930–7 January 1989. Biographical Memoirs of Fellows of the Royal Society, 36:1–16, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0021.

James:2009:CES

[Jam09] Ioan James FRS. Claude Elwood Shannon: 30 April 1916— 24 February 2001. Biographical Memoirs of Fellows of the Royal Society, 55:257–265, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.
0015.

Jones:1984:AWJ

[JB84] Ewart Ray Herbert Jones and Raymond Bonnett. Alan Woodworth Johnson, 29 September 1917–5 December 1982. Biographical Memoirs of Fellows of the Royal Society, 30:319–348, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1984.0012.

Joanny:2019:PGG

[JC19] Jean-François Joanny and Michael Cates. Pierre-Gilles de Gennes. 24 October 1932–18 May 2007. Biographical Memoirs of Fellows of the Royal Society, 66:143–158, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0033.

Jeffreys:1976:RSM

[Jef76] Harold Jeffreys. Robert Stoneley, 14 May 1894–2 February 1976. Biographical Memoirs of Fellows of the Royal Society, 22:554–564, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1976.0020.

Jeffreys:1977:KEB

[Jef77] Harold Jeffreys. Keith Edward Bullen, 29 June 1906–23 September 1976. Biographical Memoirs of Fellows of the Royal Society, 23:19– 39, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1977.0002.

Jeffes:1985:FDR

[Jef85] J. H. E. Jeffes. Frederick Denys Richardson, 17 September 1913–10 September 1983. Biographical Memoirs of Fellows of the Royal Society, 31:494–521, November 1985. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1985.0017.

Jellicoe:1999:LEA

[Jel99] The Earl Jellicoe. Lord Edward Arthur Alexander Shackleton. 15 July 1911–22 September 1994. Biographical Memoirs of Fellows of the Royal Society, 45:485–505, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0032.

Jones:1982:FSM

[JG82] Ewart Ray Herbert Jones and Peter Garratt. Franz Sondheimer, 17 May 1926–11 February 1981. Biographical Memoirs of Fellows of the Royal Society, 28:505–536, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0020.

Jacob:1998:AML

[JG98] F. Jacob and M. Girard. Andre Michel Lwoff. 8 May 1902– 30 September 1994. Biographical Memoirs of Fellows of the Royal Society, 44:255–263, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0017.

Jones:2013:PHA

[JG13] Dorothy Jones and William D. Grant. Peter Henry Andrews Sneath. 17 November 1923–9 September 2011. Biographical Memoirs of Fellows of the Royal Society, 59:337–357, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0008.

Jagur-Grodzinski:2006:MSJ

[JGP06] Joseph Jagur-Grodzinski and Stanislaw Penczek. Michael Szwarc. 19 June 1909–4 August 2000. Biographical Memoirs of Fellows of the Royal Society, 52:365–377, January 2006. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0025.

Jack:2020:KHJ

[JH20] David Jack and Stuart Hampshire. Kenneth Henderson Jack. 12 October 1918–28 January 2013. Biographical Memoirs of Fellows of the Royal Society, 69:247–272, December 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0026.

Jones:2014:SEP

[JJ14] David S. Jones and John H. Jones. Sir Edward Penley Abraham CBE. 10 June 1913–9 May 1999. Biographical Memoirs of Fellows of the Royal Society, 60:5–22, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0002.

Jones:2022:AHC

[JKL^{+}22] Richard A. Jones, Richard A. Kemp, Jonathan G. Lasch, Michael Lattman, and Nicholas C. Norman. Alan Herbert Cowley. 29 January 1934–2 August 2020. Biographical Memoirs of Fellows of the Royal Society, 72:139–160, June 2022. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0027.

Johnson:1977:JDR

[Joh77] Alan Woodworth Johnson. John Donald Rose, 2 January 1911–14 October 1976. Biographical Memoirs of Fellows of the Royal Society, 23:449–463, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.

Johnson:2000:DCP

[Joh00] L. N. Johnson. David Chilton Phillips, Lord Phillips of Ellesmere, K.B.E. 7 March 1924–23 February 1999. Biographical Memoirs of Fellows of the Royal Society, 46:377–401, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0092.

Johnson:2003:SKC

[Joh03] G. A. L. Johnson. Sir Kingsley Charles Dunham. 2 January 1910–5 April 2001 Elected FRS 1955. Biographical Memoirs of Fellows of the Royal Society, 49:147–162, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0009.

John:2023:FRW

[Joh23] Philip John. Frederick Robert Whatley. 26 January 1924– 14 November 2020. Biographical Memoirs of Fellows of the Royal Society, 74:481–499, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0021.

Jones:1960:HWB

[Jon60a] H. Jones. Herbert Wakefield Banks Skinner, 1900–1960. Biographical Memoirs of Fellows of the Royal Society, 6:259–268, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1960.0036.

Jones:1960:JJ

[Jon60b] Harold Spencer Jones. John Jackson, 1887–1958. Biographical Memoirs of Fellows of the Royal Society, 5:95–106, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Jones:1966:WLS

[Jon66] Reginald Victor Jones. Winston Leonard Spencer Churchill, 1874– 1965. Biographical Memoirs of Fellows of the Royal Society, 12:35– 105, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1966.0003.

Jones:1974:ACN

[Jon74] John Kenyon Netherton Jones. Arthur Charles Neish, 1916–1973. Biographical Memoirs of Fellows of the Royal Society, 20:294–315, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1974.0013.

Jones:1979:AEO

[Jon79] Douglas Samuel Jones. Arthur Erdélyi, 2 October 1908– 12 December 1977. Biographical Memoirs of Fellows of the Royal Society, 25:266–286, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1979.
0009.

Jones:1983:EEM

[Jon83] Francis Edgar Jones. Eric Eastwood, 12 March 1910–6 October 1981. Biographical Memoirs of Fellows of the Royal Society, 29:177– 195, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1983.0007.

Jones:1991:RDH

[Jon91] Ewart Ray Herbert Jones. Robert Downs Haworth, 15 March 1898–19 May 1990. Biographical Memoirs of Fellows of the Royal Society, 37:263–276, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0013.

Jones:2003:SER

[Jon03] J. H. Jones. Sir Ewart Ray Herbert Jones 16 March 1911– 7 May 2002 Elected FRS 1950. Biographical Memoirs of Fellows of the Royal Society, 49:263–282, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0015.

Jones:2010:HRD

[Jon10] R. Neil Jones. Hubert Rees DFC. 2 October 1923–13 September 2009. Biographical Memoirs of Fellows of the Royal Society, 56:359– 376, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2010.0003.

Jones:2021:BLA

[Jon21] J. C. Jones. Basil Lythgoe. 18 August 1913–18 April 2009. Biographical Memoirs of Fellows of the Royal Society, 70:283–295, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2019.0043.

Jones:2023:TRR

[Jon23a] Roy Jones FRSB. Thaddeus Robert Rudolph Mann. 4 December 1908–27 November 1993. Biographical Memoirs of Fellows of the Royal Society, 74:283–294, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0051.

Jonsen:2023:KJP

[Jon23b] Paul Jonsen FRS. Kenneth John Packer. 18 May 1938– 18 September 2021. Biographical Memoirs of Fellows of the Royal Society, 75:361–385, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0050.

Jordan:2004:SRW

[Jor04] C. Jordan. Sir Robert Wilson CBE. 16 April 1927–2 September 2002. Biographical Memoirs of Fellows of the Royal Society, 50:367– 386, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2004.0024.

Joyce:1992:RCR

[Joy92] R. J. V. Joyce. Reginald Charles Rainey, 18 June 1913–18 January 1990. Biographical Memoirs of Fellows of the Royal Society, 38:287– 303, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1992.0015.

Joyce:2016:RCN

[Joy16] Bruce A. Joyce FRS. Ronald Charles Newman FInstP. 10 December 1931–30 July 2014. Biographical Memoirs of Fellows of the Royal Society, 62:447–459, January 2016. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0004.

Johnson:1981:GSA

[JP81] P. Johnson and Max Ferdinand Perutz. Gilbert Smithson Adair, 21 September 1896–22 June 1979. Biographical Memoirs of Fellows of the Royal Society, 27:1–27, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0001.

Johnson:1984:ASF

[JRS84] Alan Woodworth Johnson, Francis Leslie Rose, and Charles Walter Suckling. Alfred Spinks, 25 February 1917–11 February 1982. Biographical Memoirs of Fellows of the Royal Society, 30:566–594, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1984.0019.

Joyce:2010:DWP

[JS10] Bruce A. Joyce and Michael J. Stowell. Donald William Pashley. 21 April 1927–16 May 2009. Biographical Memoirs of Fellows of the Royal Society, 56:317–340, January 2010. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0009.

Jack:1997:BAH

[JW97] David Jack and Tom Walker. Benjamin Arthur Hems. 29 June 1912–2 July 1995. Biographical Memoirs of Fellows of the Royal Society, 43:217–233, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0012.

Jencks:2000:FAL

[JW00] William P. Jencks and Richard V. Wolfenden. Fritz Albert Lipmann. 12 June 1899–24 July 1986. Biographical Memoirs of Fellows of the Royal Society, 46:333–344, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0088.

Johnson:2020:NNG

[JWW^{+}20] Brian F. G. Johnson FRS, Alan J. Welch, J. Derek Woollins, Charles Johnson, and Catherine E. Housecroft. Norman Neill Greenwood. 19 January 1925–14 November 2012. Biographical Memoirs of Fellows of the Royal Society, 68:195–218, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0015.

Karrer:1958:HW

[Kar58] Paul Karrer. Heinrich Wieland, 1877–1957. Biographical Memoirs of Fellows of the Royal Society, 4:341–352, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0026.

Kassanis:1982:KMS

[Kas82] Basil Kassanis. Kenneth Manley Smith, 13 November 1892–11 June 1981. Biographical Memoirs of Fellows of the Royal Society, 28:451– 477, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1982.0018.

Katz:1978:AVH

[Kat78] Bernard Katz. Archibald Vivian Hill, 26 September, 1886–3 June 1977. Biographical Memoirs of Fellows of the Royal Society, 24:71– 149, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1978.0005.

Katz:1982:SWK

[Kat82] Bernard Katz. Stephen William Kuffler, 24 August 1913– 11 October 1980. Biographical Memoirs of Fellows of the Royal Society, 28:225–259, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0011.

Kay:1968:REC

[Kay68] Herbert Davenport Kay. Rupert Edward Cecil Lee Guinness, Second Earl of Iveagh, 1874–1967. Biographical Memoirs of Fellows of the Royal Society, 14:287–307, November 1968.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0013.

Kay:1971:WKS

[Kay71] Herbert Davenport Kay. William Kershaw Slater, 1893–1970. Biographical Memoirs of Fellows of the Royal Society, 17:663–680, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0026.

Kay:1972:JBO

[Kay72] Herbert Davenport Kay. John Boyd Orr, Baron Boyd Orr of Brechin Mearns, 1880–1971. Biographical Memoirs of Fellows of the Royal Society, 18:43–81, November 1972. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1972.0004.

Kendall:1982:JNA

[KBP82] David George Kendall, Maurice Stevenson Bartlett, and T. L. Page. Jerzy Neyman, 16 April 1894–5 August 1981. Biographical Memoirs of Fellows of the Royal Society, 28:379–412, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1982.0015.

Kilmartin:2022:VMI

[KC22] J. V. Kilmartin FRS and R. A. Crowther FRS FMedSci. Vernon Martin Ingram. 19 May 1924–17 August 2006. Biographical Memoirs of Fellows of the Royal Society, 73:315–331, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0023.

Kembhavi:2023:GSM

[KC23] Ajit K. Kembhavi and Jayaram N. Chengalur. Govind Swarup. 23 March 1929–7 September 2020. Biographical Memoirs of Fellows of the Royal Society, 75:455–478, December 2023. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0049.

Krebs:1982:FLA

[KD82] Hans Adolf Krebs and K. Decker. Feodor Lynen, 6 April 1911–6 August 1979. Biographical Memoirs of Fellows of the Royal Society, 28:261–317, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0012.

King:2004:FCT

[KE04] David King and John Enderby. Frederick Clifford Tompkins. 29 August 1910–5 November 1995. Biographical Memoirs of Fellows of the Royal Society, 50:309–313, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2004.0020.

Keilin:1960:MJG

[Kei60] David Keilin. Maurice Jules Gaston Corneille Caullery, 1868–1958. Biographical Memoirs of Fellows of the Royal Society, 6:13–31, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1960.0022.

Kekwick:1981:AND

[Kek81] Ralph Ambrose Kekwick. Alan Nigel Drury, 3 November 1889–2 August 1980. Biographical Memoirs of Fellows of the Royal Society, 27:173–198, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0008.

Kelly:2017:SPT

[Kel17] Michael J. Kelly FRS FREng. Sir Paul Terence Callaghan FRS PCNZM. 19 August 1947–24 March 2012. Biographical Memoirs of Fellows of the Royal Society, 63:79–98, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0006.

Kemball:1975:HST

[Kem75] Charles Kemball. Hugh Stott Taylor, 6 February 1890–17 April 1974. Biographical Memoirs of Fellows of the Royal Society, 21:517– 547, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1975.0017.

Kemmer:1983:HYJ

[Kem83] Nicholas Kemmer. Hideki Yukawa, 23 January 1907–8 September 1981. Biographical Memoirs of Fellows of the Royal Society, 29:660– 676, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1983.0023.

Kent:1982:MMP

[Ken82] Percy Edward Kent. Montague Mattinson Pennell, 20 March 1916–30 December 1981. Biographical Memoirs of Fellows of the Royal Society, 28:413–429, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0016.

Kendall:1991:ANK

[Ken91a] David George Kendall. Andrei Nikolaevich Kolmogorov, 25 April 1903–20 October 1987. Biographical Memoirs of Fellows of the Royal Society, 37:299–319, November 1991. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1991.0015.

Kenn:1991:RAB

[Ken91b] Maurice J. Kenn. Ralph Alger Bagnold, 3 April 1896–28 May 1990. Biographical Memoirs of Fellows of the Royal Society, 37:55–68, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1991.0003.

Kerensky:1979:GRF

[Ker79] Oleg Alexander Kerensky. Gilbert Roberts, 18 February 1899– 1 January 1978. Biographical Memoirs of Fellows of the Royal Society, 25:477–503, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1979.
0017.

Keverne:2005:SRJ

[Kev05] E. B. Keverne. Sir Richard John Harrison. 8 October 1920– 17 October 1999. Biographical Memoirs of Fellows of the Royal Society, 51:185–193, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2005.
0012.

Keynes:1990:DAH

[Key90] Richard Darwin Keynes. Denis Arthur Haydon, 21 February 1930–29 November 1988. Biographical Memoirs of Fellows of the Royal Society, 36:199–216, December 1990. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1990.0030.

Kuhn:1983:DAJ

[KH83] Heinrich Gerhard Kuhn and Christopher Hartley. Derek Ainslie Jackson, 23 June 1906–20 February 1982. Biographical Memoirs of Fellows of the Royal Society, 29:268–296, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1983.0012.

Knight:1990:PFB

[KH90] D. E. Knight and Alan Lloyd Hodgkin. Peter Frederick Baker, 11 March 1939–10 March 1987. Biographical Memoirs of Fellows of the Royal Society, 35:1–35, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Kelly:1997:JDB

[KH97] Anthony Kelly and A. J. Howard. James Derek Birchall, O. B. E. 7 October 1930–7 December 1995. Biographical Memoirs of Fellows of the Royal Society, 43:89–104, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0006.

Kelly:2004:WLM

[KH04] John S. Kelly and John H. Horlock. Walter Laing Macdonald Perry KT OBE, Baron Perry of Walton. 21 June 1921–17 July 2003. Biographical Memoirs of Fellows of the Royal Society, 50:201–225, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2004.0015.

Hassell:2021:LRM

[KHG21] Lord (John) Krebs, Michael Hassell, and Sir Charles Godfray. Lord Robert May of Oxford OM. 8 January 1936–28 April 2020. Biographical Memoirs of Fellows of the Royal Society, 71:375–398, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2021.0007.

Kibble:1988:PTM

[Kib88] Thomas Walter Bannerman Kibble. Paul Taunton Matthews, 19 November 1919–26 February 1987. Biographical Memoirs of Fellows of the Royal Society, 34:553–580, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0018.

Kibble:1998:MAS

[Kib98] T. W. B. Kibble. Muhammad Abdus Salam, K. B. E.. 29 January 1926–21 November 1996. Biographical Memoirs of Fellows of the Royal Society, 44:387–401, January 1998. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0025.

Kilmister:1994:GFJ

[Kil94] C. W. Kilmister. George Frederick James Temple, 2 September 1901–30 January 1992. Biographical Memoirs of Fellows of the Royal Society, 40:383–400, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.
0046.

King:1956:SOO

[Kin56] Harold King. Sigmungd Otto Orsenhelm, 1871–1955. Biographical Memoirs of Fellows of the Royal Society, 2:256–267, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1956.0018.

Kinloch:2000:NAB

[Kin00] Anthony J. Kinloch. Norman Adrian De Bruyne. 8 November 1904–7 March 1997. Biographical Memoirs of Fellows of the Royal Society, 46:125–143, January 2000. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0076.

Kingman:2009:DGK

[Kin09] John Kingman FRS. David George Kendall. 15 January 1918– 23 October 2007. Biographical Memoirs of Fellows of the Royal Society, 55:121–138, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.
0017.

King:2016:LTE

[Kin16] Roderick W. King. Lloyd Thomas Evans AO FAA. 6 August 1927–23 March 2015. Biographical Memoirs of Fellows of the Royal Society, 62:125–146, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2016.
0008.

King:2022:ACP

[Kin22] Roger Antony King. Andrew Clennel Palmer. 26 May 1938– 21 December 2019. Biographical Memoirs of Fellows of the Royal Society, 73:379–393, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0015.

Kapitza:1969:LDL

[KL69] Piotr Leonidovich Kapitza and Evgenii Mikhailovich Lifshitz. Lev Davydovitch Landau, 1908–1968. Biographical Memoirs of Fellows of the Royal Society, 15:140–158, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1969.0007.

Keeler:2019:AAN

[KL19] Richard K. Keeler and Michel Lefebvre. Alan Astbury. 27 November 1934–21 July 2014. Biographical Memoirs of Fellows of the Royal Society, 66:7–23, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2016.

Kennedy:1969:JER

[KM69] William Quarrier Kennedy and A. G. MacGregor. James Ernest Richey, 1886–1968. Biographical Memoirs of Fellows of the Royal Society, 15:185–200, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1969.
0010.

Kent:1970:DWW

[KM70] L. H. Kent and Walter Thomas James Morgan. David Willis Wilson Henderson, 1903–1968. Biographical Memoirs of Fellows of the Royal Society, 16:330–341, November 1970. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0012.

Knight:2015:DFW

[KM15] Sir Peter Knight FRS and Gerard J. Milburn. Daniel Frank Walls FRSNZ. 13 September 1942–12 May 1999. Biographical Memoirs of Fellows of the Royal Society, 61:531–540, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0019.

Keeler:2024:RFJ

[KM24] James Keeler and Gareth A. Morris FRS. Ray Freeman. 6 January 1932–1 May 2022. Biographical Memoirs of Fellows of the Royal Society, 76:201–220, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0026.

Kreisel:1969:LEJ

[KN69] Georg Kreisel and Maxwell Herman Alexander Newman. Luitzen Egbertus Jan Brouwer, 1881–1966. Biographical Memoirs of Fellows of the Royal Society, 15:39–68, November 1969. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1969.0002.

King:1999:PJM

[KO99] James F. King and Karl Overton. Paul Jose De Mayo. 8 August 1924–26 July 1994. Biographical Memoirs of Fellows of the Royal Society, 45:111–128, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0009.

Kothari:1960:MS

[Kot60] D. S. Kothari. Meghnad Saha, 1893–1956. Biographical Memoirs of Fellows of the Royal Society, 5:216–236, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Kekwick:1974:AT

[KP74] Ralph Ambrose Kekwick and Kai O. Pedersen. Arne Tiselius, 1902– 1971. Biographical Memoirs of Fellows of the Royal Society, 20:401– 428, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1974.0018.

Kilburn:1978:FCW

[KP78] Tom Kilburn and L. S. Piggott. Frederic Calland Williams, 26 June 1911- 11 August 1977. Biographical Memoirs of Fellows of the Royal Society, 24:583–604, November 1978. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0020.

Kirsch:2011:WPJ

[KR11] Jack F. Kirsch and John P. Richard. William Platt Jencks. 15 August 1927–3 January 2007. Biographical Memoirs of Fellows of the Royal Society, 57:179–188, January 2011. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0019.

Krebs:1972:OHW

[Kre72] Hans Adolf Krebs. Otto Heinrich Warburg, 1883–1970. Biographical Memoirs of Fellows of the Royal Society, 18:628–699, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1972.0023.

Kreisel:1973:BAW

[Kre73] Georg Kreisel. Bertrand Arthur William Russell, Earl Russell, 1872–1970. Biographical Memoirs of Fellows of the Royal Society, 19:583–620, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1973.
0021.

Kreisel:1980:KGA

[Kre80] Georg Kreisel. Kurt Gödel, 28 April 1906–14 January 1978. Biographical Memoirs of Fellows of the Royal Society, 26:148–224, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1980.0005.

Krikorian:1995:FCS

[Kri95] Abraham D. Krikorian. Frederick Campion Steward, 16 June 1904–13 September 1993. Biographical Memoirs of Fellows of the Royal Society, 41:420–437, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1995.
0025.

Krohn:1995:SZB

[Kro95] Peter Leslie Krohn. Solly Zuckerman, Baron Zuckerman, of Burnham Thorpe, O. M., K. C. B., 30 May 1904–1 April 1993. Biographical Memoirs of Fellows of the Royal Society, 41:576–598, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1995.0034.

Kingwill:1969:PJD

[KS69] D. G. Kingwill and Basil Ferdinand Jamieson Schonland. Petrus Johann du Toit, 1888–1967. Biographical Memoirs of Fellows of the Royal Society, 15:247–266, November 1969. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1969.0013.

Kemmer:1971:MB

[KS71] Nicholas Kemmer and R. Schlapp. Max Born, 1882–1970. Biographical Memoirs of Fellows of the Royal Society, 17:17–52, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/769700;   http:/
/www.jstor.org/stable/pdfplus/769700.pdf;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1971.0002.

Krebs:1992:KZL

[KS92] John Richard Krebs and S. Sjölander. Konrad Zacharias Lorenz, 7 November 1903–27 February 1989. Biographical Memoirs of Fellows of the Royal Society, 38:209–228, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0011.

Khesin:2018:VIA

[KT18] Boris Khesin and Sergei Tabachnikov. Vladimir Igorevich Arnold. 12 June 1937–3 June 2010. Biographical Memoirs of Fellows of the Royal Society, 64:7–26, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Kahneman:2020:AMT

[KT20] Daniel Kahneman and Deborah Treisman. Anne Marie Treisman. 27 February 1935–9 February 2018. Biographical Memoirs of Fellows of the Royal Society, 68:407–430, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0035.

Kerr:2007:DAJ

[KTR07] J. R. Kerr and D. Taylor-Robinson. David Arthur John Tyrrell. 19 June 1925–2 May 2005. Biographical Memoirs of Fellows of the Royal Society, 53:349–363, January 2007. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0014.

Kuhn:1965:JF

[Kuh65] Heinrich Gerhard Kuhn. James Franck, 1882–1964. Biographical Memoirs of Fellows of the Royal Society, 11:53–74, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1965.0004.

Kurti:1958:FES

[Kur58] Nicholas Kurti. Franz Eugen Simon, 1893–1956. Biographical Memoirs of Fellows of the Royal Society, 4:224–256, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0020.

Kurti:1983:LFB

[Kur83] Nicholas Kurti. Leslie Fleetwood Bates, 7 March 1897–20 January 1978. Biographical Memoirs of Fellows of the Royal Society, 29:1– 25, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1983.0001.

Kushmerick:2001:RED

[Kus01] Martin J. Kushmerick. Robert Ernest DAVIES. 17 August 1919–7 March 1993. Biographical Memoirs of Fellows of the Royal Society, 47:141–157, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0009.

Kistiakowsky:1979:JBC

[KW79] G. B. Kistiakowsky and Frank Henry Westheimer. James Bryant Conant, 26 March 1893–11 February 1978. Biographical Memoirs of Fellows of the Royal Society, 25:208–232, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0006.

Kornberg:1984:HAK

[KW84] Hans Leo Kornberg and D. H. Williamson. Hans Adolf Krebs, 25 August 1900–22 November 1981. Biographical Memoirs of Fellows of the Royal Society, 30:351–385, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1984.0013.

Laing:2020:CRS

[Lai20] William Laing. Charles Roger Slack. 22 April 1937–24 October 2016. Biographical Memoirs of Fellows of the Royal Society, 69:517– 537, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0007.

Langlands:1985:HCO

[Lan85] Robert Phelan Langlands. Harish-Chandra, 11 October 1923– 16 October 1983. Biographical Memoirs of Fellows of the Royal Society, 31:198–225, November 1985. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0008.

Larsen:2009:HHU

[Lar09] Erik Hviid Larsen. Hans Henriksen Ussing. 30 December 1911– 22 December 2000. Biographical Memoirs of Fellows of the Royal Society, 55:305–335, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.
0002.

Launder:2007:JHW

[Lau07] Brian E. Launder. James Hunter Whitelaw. 28 January 1936– 16 August 2006. Biographical Memoirs of Fellows of the Royal Society, 53:385–399, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2007.
0024.

Launder:2023:PDR

[Lau23] Brian Launder FRS. Peter Damian Richardson. 22 August 1935–21 April 2020. Biographical Memoirs of Fellows of the Royal Society, 75:411–432, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0013.

Laws:2001:SVE

[Law01] R. M. Laws. Sir Vivian Ernest Fuchs. 11 February 1908– 11 November 1999. Biographical Memoirs of Fellows of the Royal Society, 47:203–222, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0012.

Lawrenson:2008:TEA

[Law08] Peter J. Lawrenson. Thomas Edward Allibone. 11 November 1903–9 September 2003. Biographical Memoirs of Fellows of the Royal Society, 54:3–30, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.

Lonsdale:1967:KSK

[LB67] Kathleen Yardley Lonsdale and Homi Jehangir Bhabha. Kariamanikkam Srinivasa Krishnan, 1898–1961. Biographical Memoirs of Fellows of the Royal Society, 13:244–255, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1967.0012.

Lee:2007:HAB

[LB07] Sabine Lee and Gerry E. Brown. Hans Albrecht Bethe. 2 July 1906–6 March 2005. Biographical Memoirs of Fellows of the Royal Society, 53:1–20, January 2007. CODEN BM-FRA3. ISBN 0-85403-648-2. ISSN 0080-4606 (print), 1748-8494

(electronic). URL http://www.jstor.org/stable/20461364;
https://royalsocietypublishing.org/doi/epdf/10.1098/rsbm.
2007.0018.

Lynden-Bell:1998:VAA

[LBG98] D. Lynden-Bell and V. Gurzadyan. Victor Amazaspovich Ambartsumian. 18 September 1908–12 August 1996. Biographical Memoirs of Fellows of the Royal Society, 44:23–34, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0002.

Longair:2017:SDJ

[LC17] Malcolm Longair CBE, FRS, FRSE and Michael Cates FRSE, FRS. Sir David John Cameron MacKay FRS. 22 April 1967–14 April 2016. Biographical Memoirs of Fellows of the Royal Society, 63:443– 465, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2017.0013.

Lemieux:1980:LEM

[LE80] Raymond Urgel Lemieux and O. E. Edwards. Léo Edmond Marion, 22 March 1899–16 July 1979. Biographical Memoirs of Fellows of the Royal Society, 26:357–370, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1980.0011.

Leake:1991:TNG

[Lea91] Bernard E. Leake. Thomas Neville George, 13 May 1904–18 June 1980. Biographical Memoirs of Fellows of the Royal Society, 37:198– 217, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1991.0010.

Lee:2007:REP

[Lee07] Sabine Lee. Rudolf Ernst Peierls. 5 June 1907–19 September 1995. Biographical Memoirs of Fellows of the Royal Society, 53:265–284, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL http://rsbm.royalsocietypublishing.
org/content/53/265;   http://rsbm.royalsocietypublishing.
org/content/suppl/2009/05/01/53.0.265.DC1;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2007.0003.

Legget:1985:CJM

[Leg85] R. F. Legget. Chalmers Jack Mackenzie, 10 July 1888–26 February 1984. Biographical Memoirs of Fellows of the Royal Society, 31:410– 434, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1985.0014.

Legon:2017:SHW

[Leg17] Anthony C. Legon FRS. Sir Harold Walter Kroto. 7 October 1939–30 April 2016. Biographical Memoirs of Fellows of the Royal Society, 63:413–442, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2017.
0003.

Leggett:2022:DJT

[Leg22] Anthony J. Leggett. David James Thouless. 21 September 1934–6 April 2019. Biographical Memoirs of Fellows of the Royal Society, 72:337–358, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0049.

Lehman:2012:AKM

[Leh12] I. Robert Lehman. Arthur Kornberg. 3 March 1918–26 October 2007. Biographical Memoirs of Fellows of the Royal Society, 58:151– 161, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2012.0032.

Lewis:1967:JAG

[Lew67] Wilfrid Bennett Lewis. Joseph Alexander Gray, 1884–1966. Biographical Memoirs of Fellows of the Royal Society, 13:89–106, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1967.0005.

Lewis:1983:CDD

[Lew83] Dan Lewis. Cyril Dean Darlington, 19 December 1903–26 March 1981. Biographical Memoirs of Fellows of the Royal Society, 29:113– 157, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1983.0005.

Lewis:1986:LFC

[Lew86] Dan Lewis. Leonard Francis La Cour, 28 July 1907–3 November 1984. Biographical Memoirs of Fellows of the Royal Society, 32:355– 375, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1986.0011.

Lewis:1992:KMJ

[Lew92] Dan Lewis. Kenneth Mather, 22 June 1911–20 March 1990. Biographical Memoirs of Fellows of the Royal Society, 38:247–266, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1992.0013.

Lewis:2002:QMH

[Lew02] Geoffrey Lewis. Quintin Mcgarel Hogg, Lord Hailsham of St Marylebone. 9 October 1907–12 October 2001. Biographical Memoirs of Fellows of the Royal Society, 48:221–231, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0012.

Louis:2020:FCK

[LG20] Christos Louis and Marian R. Goldsmith. Fotis C. Kafatos. 16 April 1940–18 November 2017. Biographical Memoirs of Fellows of the Royal Society, 69:273–289, December 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0010.

Lewis:1984:HGM

[LH84] Dan Lewis and D. M. Hunt. Hans Grüneberg, 26 May 1907– 23 October 1982. Biographical Memoirs of Fellows of the Royal Society, 30:226–247, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1984.
0008.

Lovell:1988:WBL

[LH88] Alfred Charles Bernard Lovell and D. G. Hurst. Wilfrid Bennett Lewis, 24 June 1908–10 January 1987. Biographical Memoirs of Fellows of the Royal Society, 34:451–509, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0016.

Longuet-Higgins:1990:RSM

[LH90] Hugh Christopher Longuet-Higgins. Robert Sanderson Mulliken, 7 June 1896–31 October 1986. Biographical Memoirs of Fellows of the Royal Society, 35:327–354, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Land:2018:RLG

[LH18] Michael F. Land FRS and Priscilla Heard. Richard Langton Gregory. 24 July 1923–17 May 2010. Biographical Memoirs of Fellows of the Royal Society, 64:163–182, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Longuet-Higgins:1978:LON

[LHF78] Hugh Christopher Longuet-Higgins and Michael Ellis Fisher. Lars Onsager, 27 November–5 October 1976. Biographical Memoirs of Fellows of the Royal Society, 24:443–471, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0014.

Lighthill:1985:KSS

[Lig85] Michael James Lighthill. Keith Stewartson, 20 September 1925–7 May 1983. Biographical Memoirs of Fellows of the Royal Society, 31:545–569, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0019.

Lighthill:1990:SGD

[Lig90] Michael James Lighthill. Sydney Goldstein, 3 December 1903– 22 January 1989. Biographical Memoirs of Fellows of the Royal Society, 36:173–197, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0029.

Liggins:1998:GSD

[Lig98] Graham Liggins. Geoffrey Sharman Dawes, C. B. E. 21 January 1918–6 May 1996. Biographical Memoirs of Fellows of the Royal Society, 44:111–125, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0008.

Lindsay:1956:RBM

[Lin56] R. B. Lindsay. Review: Biographical Memoirs of Fellows of the Royal Society: Vol. 1. Physics Today, 9(10):36, October 1956. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Lindsay:2001:JAH

[Lin01] Colonel O. J. M. Lindsay. John Anthony Hardinge Giffard, 3rd Earl of Halsbury. 4 June 1908–14 January 2000. Biographical Memoirs of Fellows of the Royal Society, 47:239–253, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0014.

Lipson:1973:AJB

[Lip73] Henry Solomon Lipson. Albert James Bradley, 1899–1972. Biographical Memoirs of Fellows of the Royal Society, 19:116–128, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0005.

Lissmann:1978:JGO

[Lis78] Hans Werner Lissmann. James Gray, 14 October 1891– 14 December 1975. Biographical Memoirs of Fellows of the Royal Society, 24:55–70, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1978.
0004.

LewisofNewnham:1997:CCA

[LJ97] Lord Lewis of Newnham and B. F. G. Johnson. Cyril Clifford Addison. 28 November 1913–1 April 1994. Biographical Memoirs of Fellows of the Royal Society, 43:3–12, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0001.

Lew:2023:IMG

[LK23] Virgilio L. Lew and Steven J. D. Karlish. Ian Michael Glynn. 3 June 1928–7 July 2022. Biographical Memoirs of Fellows of the Royal Society, 75:149–174, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0011.

Lainson:1997:PCC

[LKK97] R. Lainson and R. Killick-Kendrick. Percy Cyril Claude Garnham, C. M. G. 15 January 1901–25 December 1994. Biographical Memoirs of Fellows of the Royal Society, 43:173–192, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0010.

Luthert:2005:NHA

[LL05] Philip J. Luthert and Cynthia Medford Langley. Norman Henry Ashton CBE. 11 September 1913–4 January 2000. Biographical Memoirs of Fellows of the Royal Society, 51:1–13, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0001.

Lawson:2023:SDA

[LL23] Chris Lawson FRSC and Peter Lillford CBE FRSC. Sir David Allan (Dai) Rees. 28 April 1936–10 June 2021. Biographical Memoirs of Fellows of the Royal Society, 74:387–402, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0043.

Lyon:1994:JFL

[LM94] Mary Frances Lyon and Patrick Loudon Mollison. John Freeman Loutit, 19 February 1910–11 June 1992. Biographical Memoirs of Fellows of the Royal Society, 40:237–252, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1994.0037.

Ley:2002:SDH

[LM02] Steven V. Ley C.B.E. and F. R. S. and Rebecca M. Myers. Sir Derek Harold Richard Barton. 8 September 1918–16 March 1998. Biographical Memoirs of Fellows of the Royal Society, 48:1–23, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2002.0001.

Lyle:1980:WSD

[LMA80] T. K. Lyle, Stephen Miller, and Norman Henry Ashton. William Stewart Duke-Elder, 22 April 1898–27 March 1978. Biographical Memoirs of Fellows of the Royal Society, 26:85–105, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1980.0003.

Lewis:2006:SJV

[LMW06] Mitchell Lewis, Patrick Mollison, and David Weatherall. Sir John Vivian Dacie. 20 July 1912–12 February 2005. Biographical Memoirs of Fellows of the Royal Society, 52:67–82, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0006.

Leigh:2016:MFL

[LN16] G. Jeffery Leigh OBE and John F. Nixon FRS. Michael Franz Lappert. 31 December 1928–28 March 2014. Biographical Memoirs of Fellows of the Royal Society, 62:277–298, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0014.

Locke:1996:SVB

[Loc96] Michael Locke. Sir Vincent Brian Wigglesworth, C. B. E., 17 April 1899–12 February 1994. Biographical Memoirs of Fellows of the Royal Society, 42:541–553, November 1996. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0032.

Longair:2011:JEB

[Lon11a] Malcolm S. Longair CBE FRS FRSE. John Evan Baldwin. 6 December 1931–7 December 2010. Biographical Memoirs of Fellows of the Royal Society, 57:3–23, January 2011. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0011.

Longair:2011:VLG

[Lon11b] Malcolm S. Longair CBE FRS FRSE. Vitaly Lazarevich Ginzburg. 4 October 1916–8 November 2009. Biographical Memoirs of Fellows of the Royal Society, 57:129–146, January 2011. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0002.

Longair:2016:E

[Lon16] Malcolm Longair CBE FRS FRSE. Editorial. Biographical Memoirs of Fellows of the Royal Society, 62:1–6, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0023.

Longair:2017:E

[Lon17] Malcolm Longair CBE FRS FRSE. Editorial. Biographical Memoirs of Fellows of the Royal Society, 63:1–6, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0030.

Longair:2018:Ea

[Lon18a] Malcolm Longair CBE FRS FRSE. Editorial. Biographical Memoirs of Fellows of the Royal Society, 64:1–6, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Longair:2018:Eb

[Lon18b] Malcolm Longair CBE FRS FRSE. Editorial. Biographical Memoirs of Fellows of the Royal Society, 65:1–6, December 2018.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0029.

Longair:2019:Ea

[Lon19a] Malcolm Longair CBE. Editorial. Biographical Memoirs of Fellows of the Royal Society, 66:1–6, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0004.

Longair:2019:Eb

[Lon19b] Malcolm Longair CBE FRS FRSE. Editorial. Biographical Memoirs of Fellows of the Royal Society, 67:1–8, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0030.

Longair:2020:EBMa

[Lon20a] Malcolm Longair. Editorial: Biographical Memoirs, Volume 68. Biographical Memoirs of Fellows of the Royal Society, 68:1–6, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0005.

Longair:2020:EBMb

[Lon20b] Malcolm Longair. Editorial: Biographical Memoirs, Volume 69. Biographical Memoirs of Fellows of the Royal Society, 69:1–7, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2020.0033.

Longair:2021:E

[Lon21a] Malcolm Longair. Editorial. Biographical Memoirs of Fellows of the Royal Society, 71:1–8, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0031.

Longair:2021:EBM

[Lon21b] Malcolm Longair. Editorial: Biographical Memoirs, Volume 70. Biographical Memoirs of Fellows of the Royal Society, 70:1–8, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0001.

Longair:2021:HRH

[Lon21c] Malcolm Longair. His Royal Highness Prince Philip, the Duke of Edinburgh. 10 June 1921–9 April 2021. Biographical Memoirs of Fellows of the Royal Society, 71:9–10, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0032.

Longair:2022:HMQ

[Lon22a] Malcolm Longair. Her Majesty Queen Elizabeth II. 21 April 1926–8 September 2022. Biographical Memoirs of Fellows of the Royal Society, 73:1–2, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0034.

Longair:2022:AHM

[Lon22b] Malcolm S. Longair. Antony Hewish. 11 May 1924–13 September 2021. Biographical Memoirs of Fellows of the Royal Society, 72: 173–196, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2021.0045.

Longair:2022:Ea

[Lon22c] Malcolm Longair CBE FRS FRSE. Editorial. Biographical Memoirs of Fellows of the Royal Society, 72:1–8, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0001.

Longair:2022:Eb

[Lon22d] Malcolm Longair CBE FRS FRSE. Editorial. Biographical Memoirs of Fellows of the Royal Society, 73:3–12, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0029.

Longair:2023:Ea

[Lon23a] Malcolm S. Longair CBE FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 74:1–8, April 2023. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0006.

Longair:2023:REH

[Lon23b] Malcolm S. Longair CBE FRS. Richard Edwin Hills. 30 September 1945–5 June 2022. Biographical Memoirs of Fellows of the Royal Society, 74:237–257, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0044.

Longair:2023:Eb

[Lon23c] Malcolm S. Longair FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 75:1, December 2023. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0027.

Longair:2024:E

[Lon24a] Malcolm S. Longair CBE FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 77:1–8, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Longair:2024:Ea

[Lon24b] Malcolm S. Longair FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 76:1–8, June 2024. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2024.0001.

Selborne:1998:RMM

[Lor98] Lord Selborne. Roger Mellor Makins, G.C.B., G.C.M.G., D.L., the 1st Baron Sherfield. 3 February 1904–9 November 1996. Biographical Memoirs of Fellows of the Royal Society, 44:267–278, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0018.

Lovell:1964:JLP

[Lov64] Alfred Charles Bernard Lovell. Joseph Lade Pawsey, 1908–1962. Biographical Memoirs of Fellows of the Royal Society, 10:229–243, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0014.

Lovell:1975:PMS

[Lov75] Sir Bernard Lovell, FRS. Patrick Maynard Stuart Blackett, Baron Blackett, of Chelsea, 18 November 1897–13 July 1974. Biographical Memoirs of Fellows of the Royal Society, 21:1–115, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769678;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1975.0001.

Lovelock:2004:AJP

[Lov04] James Lovelock. Archer John Porter Martin CBE. 1 March 1910–28 July 2002. Biographical Memoirs of Fellows of the Royal Society, 50:157–170, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0012.

Lovejoy:2011:GEH

[Lov11] Thomas E. Lovejoy. George Evelyn Hutchinson. 13 January 1903–17 May 1991. Biographical Memoirs of Fellows of the Royal Society, 57:167–177, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2010.
0016.

Loudon:1991:AFG

[LP91] Rodney Loudon and Edward George Sydney Paige. Alan Frank Gibson, 30 May 1923–27 March 1988. Biographical Memoirs of Fellows of the Royal Society, 37:219–244, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1991.0011.

Launder:2019:DBS

[LPP19] Brian E. Launder FRS, FREng, Suhas V. Patankar, and Andrew Pollard FCAE. Dudley Brian Spalding. 9 January 1923–27 November 2016. Biographical Memoirs of Fellows of the Royal Society, 66: 377–402, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2018.0024.

Longair:2017:GRB

[LR17] Malcolm Longair CBE FRS FRSE and Martin Rees (Baron Rees of Ludlow) OM FRS FREng FMedSci. Geoffrey Ronald Burbidge. 24 September 1925–26 January 2010. Biographical Memoirs of Fellows of the Royal Society, 63:55–78, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0002.

Laland:2019:SPB

[LR19] Kevin N. Laland and Steven Rose. Sir Patrick Bateson FRS. 31 March 1938–1 August 2017. Biographical Memoirs of Fellows of the Royal Society, 66:25–51, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0040.

Loutit:1966:LHG

[LS66] John Freeman Loutit and O. C. A. Scott. Louis Harold Gray, 1905– 1965. Biographical Memoirs of Fellows of the Royal Society, 12:195– 217, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1966.0009.

Lythgoe:1985:GRC

[LS85] Basil Lythgoe and G. A. Swan. George Roger Clemo, 2 August 1889–2 March 1983. Biographical Memoirs of Fellows of the Royal Society, 31:66–86, November 1985. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1985.0003.

Lynden-Bell:2012:ARS

[LS12] Donald Lynden-Bell CBE FRS and François Schweizer. Allan Rex Sandage. 18 June 1926–13 November 2010. Biographical Memoirs of Fellows of the Royal Society, 58:245–264, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0021.

Lachmann:2009:RRA

[LW09a] Peter Lachmann FRS and Herman Waldmann FRS. Robert Royston Amos (Robin) Coombs. 9 January 1921–25 January 2006. Biographical Memoirs of Fellows of the Royal Society, 55:45–58, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2008.0021.

Longair:2009:SAB

[LW09b] Malcolm S. Longair CBE FRS FRSE and John R. Waldram FInstP. Sir Alfred Brian Pippard. 7 September 1920–21 September 2008. Biographical Memoirs of Fellows of the Royal Society, 55:201–220, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2009.0014.

Lewis:2020:POJ

[LW20] David J. Lewis and Richard E. P. Winpenny. Paul O’Brien. 22 January 1954–16 October 2018. Biographical Memoirs of Fellows of the Royal Society, 69:443–466, December 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0041.

Lwoff:1977:JLM

[Lwo77] Andre Michel Lwoff. Jacques Lucien Monod, 9 February 1910–31 May 1976. Biographical Memoirs of Fellows of the Royal Society, 23:384–412, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0015.

Lamb:2024:DAB

[LY24] Trevor D. Lamb FRS and King-Wai Yau. Denis Aristide Baylor. 30 January 1940–16 March 2022. Biographical Memoirs of Fellows of the Royal Society, 77:25–48, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Lynden-Bell:2015:WLW

[Lyn15] Donald Lynden-Bell CBE FRS. Wallace Leslie William Sargent. 15 February 1935–29 October 2012. Biographical Memoirs of Fellows of the Royal Society, 61:467–483, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0018.

Lyon:2001:CEF

[Lyo01] Mary F. Lyon. Charles Edmund Ford. 24 October 1912–7 January 1999. Biographical Memoirs of Fellows of the Royal Society, 47:189– 201, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2001.0011.

Lyon:2022:GWF

[Lyo22] Stuart Lyon. Graham Wood. 6 February 1934–4 November 2016. Biographical Memoirs of Fellows of the Royal Society, 72:447–470, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2021.0025.

Morata:2021:WJG

[MA21] Ginés Morata ForMemRS and Markus Affolter. Walter Jakob Gehring. 20 March 1939–29 May 2014. Biographical Memoirs of Fellows of the Royal Society, 71:197–212, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0011.

Mabberley:1999:EJH

[Mab99] David J. Mabberley. Edred John Henry Corner, C.B.E. 12 January 1906–14 September 1996. Biographical Memoirs of Fellows of the Royal Society, 45:77–93, January 1999. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0007.

Macfarlane:1977:MMS

[Mac77] Robert Gwyn Macfarlane. Montague Maizels, 30 September 1899–10 February 1976. Biographical Memoirs of Fellows of the Royal Society, 23:344–366, November 1977. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0013.

MacIntyre:2000:DHC

[Mac00] Iain MacIntyre. Douglas Harold Copp. 16 January 1915–17 March 1998. Biographical Memoirs of Fellows of the Royal Society, 46: 49–64, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0073.

Macfarlane:2003:JFC

[Mac03] Alistair Macfarlane CBE FRS FREng. John Flavell Coales CBE 14 September 1907–5 June 1999 Elected FRS 1970. Biographical Memoirs of Fellows of the Royal Society, 49:119–131, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0007.

MacLeod:2012:MMG

[Mac12] Roy MacLeod. Margaret Mary Gowing CBE FBA. 26 April 1921–7 November 1998. Biographical Memoirs of Fellows of the Royal Society, 58:67–111, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2012.
0027.

Mackie:2022:QBA

[Mac22] George Mackie. Quentin Bone. 17 August 1931–6 July 2021. Biographical Memoirs of Fellows of the Royal Society, 72:55–76, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0042.

Mackay:2023:WGH

[Mac23] Trudy F. C. Mackay FRS. William George Hill. 7 August 1940– 17 December 2021. Biographical Memoirs of Fellows of the Royal Society, 75:219–237, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0003.

Maddrell:1990:JAR

[Mad90] Simon Hugh Piper Maddrell. James Arthur Ramsay, 6 September 1909–5 February 1988. Biographical Memoirs of Fellows of the Royal Society, 36:419–433, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.

Maddock:2006:JAC

[Mad06] Linda Maddock. Joseph Arthur Colin Nicol. 5 December 1915– 20 December 2004. Biographical Memoirs of Fellows of the Royal Society, 52:203–218, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0015.

Mahajan:2024:APM

[Mah24] Kushal Kumar Mahajan. Ashesh Prosad Mitra. 21 February 1927– 3 September 2007. Biographical Memoirs of Fellows of the Royal Society, 77:285–300, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Maizels:1969:EJC

[Mai69] Montague Maizels. Edward Joseph Conway, 1894–1968. Biographical Memoirs of Fellows of the Royal Society, 15:69–82, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1969.0003.

Makins:1994:WGP

[Mak94] Roger Mellor Makins. William George Penney, O. M., K. B. E., Baron Penney of East Hendred, 24 June 1909–3 March 1991. Biographical Memoirs of Fellows of the Royal Society, 39:281–302, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Mak:2018:EAM

[Mak18] Tak Wah Mak. Ernest Armstrong McCulloch. 21 April 1926–20 January 2011. Biographical Memoirs of Fellows of the Royal Society, 64:317–339, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Mann:1960:WHM

[Man60] Frederick George Mann. William Hobson Mills, 1873–1959. Biographical Memoirs of Fellows of the Royal Society, 6:201–225, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1960.0033.

Mann:1964:DK

[Man64] Thaddeus Robert Rudolph Mann. David Keilin, 1887–1963. Biographical Memoirs of Fellows of the Royal Society, 10:183–205, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0011.

Mansfield:1987:WHW

[Man87] Eric Harold Mansfield. William Henry Wittrick, 29 October 1922–2 July 1986. Biographical Memoirs of Fellows of the Royal Society, 33:709–728, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0025.

Mansfield:1998:OVS

[Man98] Terry A. Mansfield. Oscar Victor Sayer Heath. 26 July 1903–16 June 1997. Biographical Memoirs of Fellows of the Royal Society, 44:219–235, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0015.

Mander:2003:CWS

[Man03] Lewis N. Mander FRS. Charles William Shoppee. 4 February 1904–20 October 1994. Biographical Memoirs of Fellows of the Royal Society, 49:495–507, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2003.
0029.

Marion:1964:EWR

[Mar64] Leo Edmond Marion. Edgar William Richard Steacie, 1900–1962. Biographical Memoirs of Fellows of the Royal Society, 10:257–281, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0016.

Marston:1966:AGD

[Mar66] Hedley Ralph Marston. Albert Gherbury David Rivett, 1885–1961. Biographical Memoirs of Fellows of the Royal Society, 12:437–455, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0021.

Marshall:1979:FCF

[Mar79] Norman Bertram Marshall. Francis Charles Fraser, 16 June 1903–21 October 1978. Biographical Memoirs of Fellows of the Royal Society, 25:287–317, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1979.
0010.

Marshall:1986:ACH

[Mar86] Norman Bertram Marshall. Alister Clavering Hardy, 10 February 1896–22 May 1985. Biographical Memoirs of Fellows of the Royal Society, 32:223–273, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0008.

Marrian:1988:JSM

[Mar88] D. H. Marrian. Joseph Stanley Mitchell, 22 July 1909–22 February 1987. Biographical Memoirs of Fellows of the Royal Society, 34:581– 607, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1988.0019.

Martin:2012:IMA

[Mar12] T. John Martin FRS. Iain MacIntyre. 30 August 1924– 18 September 2008. Biographical Memoirs of Fellows of the Royal Society, 58:179–201, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2011.
0025.

Massey:1971:DFM

[Mas71] Harrie Stewart Wilson Massey. David Forbes Martyn, 1906–1970. Biographical Memoirs of Fellows of the Royal Society, 17:497–510, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0019.

Mason:2002:JSS

[Mas02] John Mason C.B., F.R.S. John Stanley Sawyer. 19 June 1916– 19 September 2000. Biographical Memoirs of Fellows of the Royal Society, 48:405–416, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2002.
0024.

Masser:2023:ABA

[Mas23] David Masser FRS. Alan Baker. 19 August 1939–4 February 2018. Biographical Memoirs of Fellows of the Royal Society, 74:9–36, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2022.0033.

Mather:1988:JLJ

[Mat88a] Kenneth Mather. John Leonard Jinks, 21 October 1929–6 June 1987. Biographical Memoirs of Fellows of the Royal Society, 34:321– 344, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1988.0012.

Matthews:1988:CAF

[Mat88b] Richard Ellis Ford Matthews. Charles Alexander Fleming, 9 September 1916–11 September 1987. Biographical Memoirs of Fellows of the Royal Society, 34:178–203, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0007.

Matile:1990:AFW

[Mat90a] Philippe Matile. Albert Frey-Wyssling, 8 November 1900–30 August 1988. Biographical Memoirs of Fellows of the Royal Society, 35: 113–126, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Matty:1990:EJW

[Mat90b] A. J. Matty. Ernest James William Barrington, 17 February 1909– 15 December 1985. Biographical Memoirs of Fellows of the Royal Society, 35:37–54, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Mayneord:1979:JAV

[May79] William Valentine Mayneord. John Alfred Valentine Butler, 14 February 1899–16 July 1977. Biographical Memoirs of Fellows of the Royal Society, 25:144–178, November 1979. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0004.

Mayne:2015:JHW

[May15] David Q. Mayne FRS FREng. John Hugh Westcott. 3 November 1920–10 October 2014. Biographical Memoirs of Fellows of the Royal Society, 61:541–554, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0017.

McGrath:2021:SJW

[MB21] John Christie McGrath and Richard A. Bond. Sir James Whyte Black OM. 14 June 1924–22 March 2010. Biographical Memoirs of Fellows of the Royal Society, 70:23–40, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0047.

Mitchison:2024:NAM

[MB24] Timothy J. Mitchison FRS and Peter C. L. Beverley FMedSci. Nicholas Avrion Mitchison. 5 May 1928–28 December 2022. Biographical Memoirs of Fellows of the Royal Society, 77:261–283, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Misson:1999:AEM

[MBW99] Gary P. Misson, A. Clive Bishop, and Winifred M. Watkins. Arthur Ernest Mourant. 11 April 1904–29 August 1994. Biographical Memoirs of Fellows of the Royal Society, 45:329–348, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0023.

McCance:1960:CHD

[McC60] Andrew McCance. Cecil Henry Desch, 1874–1958. Biographical Memoirs of Fellows of the Royal Society, 5:48–68, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

McCance:1965:JMM

[McC65] Andrew McCance. James McFadyen McNeill, 1892–1964. Biographical Memoirs of Fellows of the Royal Society, 11:126–134, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1965.0008.

McCrea:1981:HHP

[McC81] William Hunter McCrea. Harry Hemley Plaskett, 5 July 1893– 26 January 1980. Biographical Memoirs of Fellows of the Royal Society, 27:444–478, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1981.
0018.

McCrea:1988:RVR

[McC88] William Hunter McCrea. Richard van der Riet Woolley, 24 April 1906–24 December 1986. Biographical Memoirs of Fellows of the Royal Society, 34:921–982, December 1988. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0028.

McCullagh:2003:JWT

[McC03] Peter McCullagh FRS. John Wilder Tukey. 16 June 1915–26 July 2000. Biographical Memoirs of Fellows of the Royal Society, 49:537– 555, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
3650242;   https://royalsocietypublishing.org/doi/epdf/

10.1098/rsbm.2003.0032. Supplement contains a 16-page biography of 431 publications by J. Wilder Tukey.

McCulloch:2008:SHC

[McC08] J. S. G. McCulloch. Sir Herbert Charles Pereira. 12 May 1913– 19 December 2004. Biographical Memoirs of Fellows of the Royal Society, 54:245–256, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.
0001.

McCubbin:2022:GHS

[McC22] N. A. McCubbin. Godfrey Harry Stafford. 15 April 1920–29 July 2013. Biographical Memoirs of Fellows of the Royal Society, 73:445– 469, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2022.0008.

McKenzie:1987:ECB

[McK87] Dan Peter McKenzie. Edward Crisp Bullard, 21 September 1907–3 April 1980. Biographical Memoirs of Fellows of the Royal Society, 33:65–98, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0004.

McKee:2022:PVT

[McK22] Jeffrey K. McKee. Phillip Vallentine Tobias. 14 October 1925–7 June 2012. Biographical Memoirs of Fellows of the Royal Society, 72:359–370, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0041.

McLauchlan:2000:HMP

[McL00] K. A. McLauchlan. Herbert Marcus Powell. 7 August 1906– 10 March 1991. Biographical Memoirs of Fellows of the Royal Society, 46:425–442, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0094.

McLauchlan:2004:DHW

[McL04] K. A. McLauchlan. David Hardy Whiffen. 15 August 1922– 2 December 2002. Biographical Memoirs of Fellows of the Royal Society, 50:333–349, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0022.

McRobert:1976:BRB

[McR76] George R. McRobert. Book review: Biographical Memoirs of Fellows of the Royal Society, Vol 21 (pp 584) The Royal Society, 1975. British Medical Journal, 1(6016):1021, April 1976. CODEN BMJOAE. ISSN 0007-1447.

Massey:1981:EHS

[MD81] Harrie Stewart Wilson Massey and D. H. Davis. Eric Henry Stoneley Burhop, 31 January 1911–22 January 1980. Biographical Memoirs of Fellows of the Royal Society, 27:131–152, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1981.0006.

Hatch:2003:SRN

[MDHW03] Barry (C. B.) Osmond FRS Marshall D. Hatch FRS and Joseph T. Wiskich. Sir Rutherford Ness Robertson. 29 September 1913– 5 March 2001 Elected FRS 1961. Biographical Memoirs of Fellows of the Royal Society, 49:397–413, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0023.

McCave:2011:SNJ

[ME11] Ian Nicholas McCave and Henry Elderfield FRS. Sir Nicholas John Shackleton. 23 June 1937–24 January 2006. Biographical Memoirs of Fellows of the Royal Society, 57:435–462, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0005.

Meade:2008:E

[Mea08] T. W. Meade. Editorial. Biographical Memoirs of Fellows of the Royal Society, 54:1–2, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.
0019.

Meade:2009:E

[Mea09] T. W. Meade. Editorial. Biographical Memoirs of Fellows of the Royal Society, 55:1–2, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.
0017.

Meade:2010:E

[Mea10] Tom Meade. Editorial. Biographical Memoirs of Fellows of the Royal Society, 56:1–2, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2010.
0021.

Meade:2011:E

[Mea11] Tom Meade FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 57:1–2, January 2011. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2011.
0020.

Medawar:1961:PAG

[Med61] Peter Brian Medawar. Peter Alfred Gorer, 1907–1961. Biographical Memoirs of Fellows of the Royal Society, 7:95–109, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1961.0008.

Medawar:1980:MAA

[Med80] Peter Brian Medawar. Michael Abercrombie, 14 August 1912– 28 May 1979. Biographical Memoirs of Fellows of the Royal Society, 26:1–15, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1980.
0001.

Mehra:1975:SB

[Meh75] Jagdish Mehra. Satyendra Nath Bose, 1 January 1894–4 February 1974. Biographical Memoirs of Fellows of the Royal Society, 21:116– 154, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1975.0002.

Mehra:2002:RPF

[Meh02] Jagdish Mehra. Richard Phillips Feynman. 11 May 1918–15 February 1988. Biographical Memoirs of Fellows of the Royal Society, 48: 97–128, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
3650252;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0007.

Menzies:1960:FT

[Men60] A. C. Menzies. Frank Twyman, 1876–1959. Biographical Memoirs of Fellows of the Royal Society, 5:269–279, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Mendelssohn:1964:DKC

[Men64] Kurt Alfred Georg Mendelssohn. David Keith Chalmers Mac-Donald, 1920–1963. Biographical Memoirs of Fellows of the Royal Society, 10:207–220, November 1964. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1964.
0012.

Menzies:1970:SMB

[Men70] Robert Gordon Menzies. Stanley Melbourne Bruce — Viscount Bruce of Melbourne, 1883–1967. Biographical Memoirs of Fellows of the Royal Society, 16:57–62, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0004.

Mestel:1991:TGC

[Mes91] Leon Mestel. Thomas George Cowling, 17 June 1906–16 June 1990. Biographical Memoirs of Fellows of the Royal Society, 37:103–125, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1991.0006.

Mestel:1998:RJT

[Mes98] L. Mestel. Roger John Tayler, O. B. E.. 25 October 1929– 23 January 1997. Biographical Memoirs of Fellows of the Royal Society, 44:405–416, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0026.

Mestel:1999:MSM

[Mes99] L. Mestel. Martin Schwarzschild. 31 May 1912–10 April 1997. Biographical Memoirs of Fellows of the Royal Society, 45:469–484, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1999.0031.

Morgan:1966:AW

[MF66] Walter Thomas James Morgan and G. E. Francis. Arthur Wormall, 1900–1964. Biographical Memoirs of Fellows of the Royal Society, 12:543–564, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1966.

Massey:1976:JCO

[MF76] Harrie Stewart Wilson Massey and Norman Feather. James Chadwick, 20 October 1891–24 July 1974. Biographical Memoirs of Fellows of the Royal Society, 22:10–70, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0002.

Michell:2018:MGL

[MF18] Bob Michell FRS FMedSci and Mike Ferguson CBE FRS FRSE FMedSci. Martin Geoffrey Low. 27 July 1950–6 August 2013. Biographical Memoirs of Fellows of the Royal Society, 65:267–282, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2017.0041.

MacFarlane:1990:FEJ

[MH90] George G. MacFarlane and Cyril Hilsum. Francis Edgar Jones, 16 January 1914–10 April 1988. Biographical Memoirs of Fellows of the Royal Society, 35:179–199, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

May:2008:TRE

[MH08] Robert M. May and Michael P. Hassell. Thomas Richard Edmund Southwood. 20 June 1931–26 October 2005. Biographical Memoirs of Fellows of the Royal Society, 54:333–374, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0005.

Maberly:2019:JFT

[MH19] S. C. Maberly and S. I. Heaney. John Francis Talling. 23 March 1929–20 June 2017. Biographical Memoirs of Fellows of the Royal Society, 66:447–461, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0031.

Hodell:2024:HEA

[MH24] Ian Nicholas McCave and David Arnold Hodell. Henry Elderfield. 25 April 1943–19 April 2016. Biographical Memoirs of Fellows of the Royal Society, 76:145–168, June 2024.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0046.

Midwinter:2021:SCK

[Mid21] John Midwinter. Sir Charles Kuen Kao. 4 November 1933– 23 September 2018. Biographical Memoirs of Fellows of the Royal Society, 70:211–224, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0006.

Martin:2024:JHL

[MIH24] Paul Martin and David Ish-Horowicz. Julian Hart Lewis. 12 August 1946–30 April 2014. Biographical Memoirs of Fellows of the Royal Society, 77:221–239, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Millar:1984:FGM

[Mil84] I. T. Millar. Frederick George Mann, 29 June 1897–29 March 1982. Biographical Memoirs of Fellows of the Royal Society, 30:407–441, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1984.0015.

Miller:1997:SKR

[Mil97] David Miller. Sir Karl Raimund Popper, C. H., F. B. A. 28 July 1902–17 September 1994. Biographical Memoirs of Fellows of the Royal Society, 43:369–409, January 1997. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0021.

Millward:2018:JCW

[Mil18] D. Joe Millward. John Conrad Waterlow. 13 June 1916–19 October 2010. Biographical Memoirs of Fellows of the Royal Society, 65:429– 448, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2018.0010. See correction [Ano19].

Miller:2020:DCG

[Mil20] Christopher Miller. David Christopher Gadsby. 26 March 1947–9 March 2019. Biographical Memoirs of Fellows of the Royal Society, 68:175–193, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0048.

Mingos:2022:KWO

[Min22] D. Michael P. Mingos FRS. Kenneth Wade. 13 October 1932–16 March 2014. Biographical Memoirs of Fellows of the Royal Society, 73:499–520, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0009.

Mitchell:1961:PGH

[Mit61] George Hoole Mitchell. Percy George Hamnall Boswell, 1886– 1960. Biographical Memoirs of Fellows of the Royal Society, 7:17– 30, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1961.0002.

Mitchison:1990:PBM

[Mit90] Nicholas Avrion Mitchison. Peter Brian Medawar, 28 February 1915–2 October 1987. Biographical Memoirs of Fellows of the Royal Society, 35:281–301, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Mitchison:1991:MMS

[Mit91] John Murdoch Mitchison. Michael Meredith Swann, Baron Swann of Coln St Denys, 1 March 1920–22 September 1990. Biographical Memoirs of Fellows of the Royal Society, 37:445–460, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1991.0023.

Mitchison:2000:JGH

[Mit00] N. A. Mitchison. James Griffiths Howard. 25 September 1927– 6 October 1998. Biographical Memoirs of Fellows of the Royal Society, 46:257–267, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.

Machesky:2024:AHM

[MJR24] Laura M. Machesky, Aron B. Jaffe, and Anne J. Ridley FRS. Alan Hall. 19 May 1952–3 May 2015. Biographical Memoirs of Fellows of the Royal Society, 77:201–219, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Mather:1985:MBC

[ML85] Kenneth Mather and William John Cooper Lawrence. Morley Benjamin Crane, 17 March 1890–17 September 1983. Biographical Memoirs of Fellows of the Royal Society, 31:88–110, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1985.0004.

McOmie:2003:WBJ

[ML03] J. F. W. McOmie and D. M. G. Lloyd. Wilson Baker. 24 January 1900–3 June 2002. Biographical Memoirs of Fellows of the Royal Society, 49:15–38, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2003.
0002.

May:2019:SPA

[ML19] Alex May and Malcolm Longair. Sir Patrick Alfred Caldwell-Moore. 4 March 1923–9 December 2012. Biographical Memoirs of Fellows of the Royal Society, 67:143–152, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0029.

Mair:1987:WWA

[MM87] W. N. Mair and Eric Harold Mansfield. William Watt, 14 April 1912–11 August 1985. Biographical Memoirs of Fellows of the Royal Society, 33:641–667, December 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0023.

Milner:1996:DOH

[MM96] Peter M. Milner and Brenda Atkinson Milner. Donald Olding Hebb, 22 July 1904–20 August 1985. Biographical Memoirs of Fellows of the Royal Society, 42:193–204, November 1996.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0012.

Meyer:2021:HSJ

[MM21] Axel Meyer and Jürgen Mittelstraß. Hubert Simon (Jim) Markl. 17 August 1938–8 January 2015. Biographical Memoirs of Fellows of the Royal Society, 71:361–373, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0010.

McMichael:2018:BAA

[MO18] Andrew J. McMichael FRS and Bridget M. Ogilvie FRS. Brigitte Alice Askonas. 1 April 1923–9 January 2013. Biographical Memoirs of Fellows of the Royal Society, 65:31–45, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0007.

McKellar:2023:JKG

[MO23] A. R. W. McKellar FRSC and Takeshi Oka FRS. James Kay Graham Watson. 20 April 1936–18 December 2020. Biographical Memoirs of Fellows of the Royal Society, 74:419–438, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0038.

Moffatt:1985:JAS

[Mof85] Henry Keith Moffatt. John Arthur Shercliff, 10 September 1927–6 December 1983. Biographical Memoirs of Fellows of the Royal Society, 31:524–541, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0018.

Moffatt:2002:GKB

[Mof02] H. K. Moffatt F.R.S. George Keith Batchelor. 8 March 1920–30 March 2000. Biographical Memoirs of Fellows of the Royal Society, 48:25–41, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2002.

Monteith:1986:HLP

[Mon86] John Lennox Monteith. Howard Latimer Penman, 10 April 1909–13 October 1984. Biographical Memoirs of Fellows of the Royal Society, 32:377–404, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0012.

Moncada:2006:SJR

[Mon06] Salvador Moncada. Sir John Robert Vane. 29 March 1927– 19 November 2004. Biographical Memoirs of Fellows of the Royal Society, 52:401–411, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0027.

Moon:1977:GPT

[Moo77] Philip Burton Moon. George Paget Thomson, 3 May 1892– 10 September 1975. Biographical Memoirs of Fellows of the Royal Society, 23:529–556, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0020.

Moody:2015:CWR

[Moo15] Christopher J. Moody. Charles Wayne Rees CBE. 15 October 1927–21 September 2006. Biographical Memoirs of Fellows of the Royal Society, 61:351–378, January 2015. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0023.

Morton:1966:TPH

[Mor66] Richard Alan Morton. Thomas Percy Hilditch, 1886–1965. Biographical Memoirs of Fellows of the Royal Society, 12:259–289, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0012.

Morgan:1988:JDM

[Mor88] B. L. Morgan. James Dwyer McGee, 17 December 1903– 28 February 1987. Biographical Memoirs of Fellows of the Royal Society, 34:511–551, December 1988. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.
0017.

Morton:1992:CMY

[Mor92] B. Morton. Charles Maurice Yonge, 9 December 1899–17 March 1986. Biographical Memoirs of Fellows of the Royal Society, 38:377– 412, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1992.0020.

Morris:1997:FGR

[Mor97] J. Gareth Morris. Florence Gwendolen Rees. 3 July 1906–4 October 1994. Biographical Memoirs of Fellows of the Royal Society, 43:445– 459, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1997.0024.

Morris:2005:SMF

[Mor05] Peter Morris. Sir Michael Francis Addison Woodruff. 3 April 1911–10 March 2001. Biographical Memoirs of Fellows of the Royal Society, 51:455–471, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2005.
0030.

Morris:2016:LSD

[Mor16] A. J. Morris FREng FRAeS. Leslie Sydney Dennis Morley FREng FRAeS. 23 May 1924–16 June 2011. Biographical Memoirs of Fellows of the Royal Society, 62:433–445, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0029.

Morange:2017:FJJ

[Mor17] Michel Morange. François Jacob. 17 June 1920–19 April 2013. Biographical Memoirs of Fellows of the Royal Society, 63:345–361, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2016.0021.

Morrison:2019:SDD

[Mor19] David Morrison BTech CEng FIMechE. Sir Diarmuid Downs CBE. 23 April 1922–12 February 2014. Biographical Memoirs of Fellows of the Royal Society, 66:159–181, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0036.

Morris:2021:SPM

[Mor21] P. G. Morris. Sir Peter Mansfield. 9 October 1933–8 February 2017. Biographical Memoirs of Fellows of the Royal Society, 70:313–334, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2020.0031.

Mott:1955:JEL

[Mot55] Nevill Francis Mott. John Edward Lennard-Jones, 1894–1954. Biographical Memoirs of Fellows of the Royal Society, 1:174–184, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1955.0013.

Mott:1982:WHH

[Mot82] Nevill Francis Mott. Walter Heinrich Heitler, 2 January 1904– 15 November 1981. Biographical Memoirs of Fellows of the Royal Society, 28:141–151, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0007.

Mott:1987:HJA

[Mot87] Nevill Francis Mott. Harry Jones, 12 April 1905–15 December 1986. Biographical Memoirs of Fellows of the Royal Society, 33:325–342, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1987.0012.

Mott:1992:HFD

[Mot92] Nevill Francis Mott. Herbert Fröhlich, 9 December 1905– 23 January 1991. Biographical Memoirs of Fellows of the Royal Society, 38:145–162, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1992.

Mott:1962:AMT

[MP62] Nevill Francis Mott and Cecil Frank Powell. Arthur Mannering Tyndall, 1881–1961. Biographical Memoirs of Fellows of the Royal Society, 8:159–165, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1962.
0012.

MacIntosh:1974:GLB

[MP74] Frank Campbell MacIntosh and William Drummond Macdonald Paton. George Lindor Brown, 1903–1971. Biographical Memoirs of Fellows of the Royal Society, 20:41–73, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1974.0003.

Montgomery:1975:FKO

[MP75] H. B. S. Montgomery and Adrian Frank Posnette. Franklin Kidd, 12 October 1890–7 May 1974. Biographical Memoirs of Fellows of the Royal Society, 21:406–430, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1975.0012.

Mott:1977:WHD

[MP77] Sir Nevill Mott and Sir Rudolf Peierls. Werner Heisenberg, 5 December 1901–1 February 1976. Biographical Memoirs of Fellows of the Royal Society, 23:212–251, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769614;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0009.

McMichael:1982:GWP

[MP82] John McMichael and William Stanley Peart. George White Pickering, 26 June 1904–3 September 1980. Biographical Memoirs of Fellows of the Royal Society, 28:431–449, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0017.

Mestel:2007:WHM

[MP07] Leon Mestel and Bernard E. J. Pagel. William Hunter Mc-Crea. 13 December 1904–25 April 1999. Biographical Memoirs of Fellows of the Royal Society, 53:223–236, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0005.

Maughan:2023:SML

[MP23] Barbara Maughan and Andrew Pickles. Sir Michael Llewellyn Rutter. 15 August 1933–23 October 2021. Biographical Memoirs of Fellows of the Royal Society, 75:433–454, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0018.

Parker:2024:SCB

[MP24] Zoltán Molnár and Andrew Parker. Sir Colin Brian Blakemore. 1 June 1944–27 June 2022. Biographical Memoirs of Fellows of the Royal Society, 76:95–118, June 2024. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0029.

Minnett:1996:SFW

[MR96] H. C. Minnett and Rutherford Ness Robertson. Sir Frederick William George White, C. B. E., 26 May 1905–17 August 1994. Biographical Memoirs of Fellows of the Royal Society, 42:497–521, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1996.0030.

Mansfield:2013:BSW

[MR13] Sir Peter Mansfield FRS and Eric J. Roebuck. Brian Stewart Worthington. 9 June 1938–9 December 2007. Biographical Memoirs of Fellows of the Royal Society, 59:423–435, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0017.

Morrison:2019:LMP

[MR19] John F. B. Morrison and John A. Russell. Lillian Mary Pickford. 14 August 1902–14 August 2002. Biographical Memoirs of Fellows of the Royal Society, 67:371–400, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0008.

Moore:2022:SPM

[MR22] Chris Moore and Eileen C. Rees. Sir Peter Markham Scott CH. 14 September 1909–29 August 1989. Biographical Memoirs of Fellows of the Royal Society, 73:421–443, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0019.

MacIntosh:1990:JHQ

[MS90] Frank Campbell MacIntosh and T. L. Sourkes. Juda Hirsch Quastel, 2 October 1899–15 October 1987. Biographical Memoirs of Fellows of the Royal Society, 36:379–418, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1990.0037.

Mason:1996:FPS

[MS96] Basil John Mason and F. B. Smith. Frank Pasquill, 8 September 1914–15 October 1994. Biographical Memoirs of Fellows of the Royal Society, 42:277–288, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0018.

McCombie:1998:HHH

[MS98] Charles W. McCombie and Joseph C. Smith. Harold Horace Hopkins. 6 December 1918-22 October 1994. Biographical Memoirs of Fellows of the Royal Society, 44:239–252, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0016.

Moodie:2006:JMC

[MS06] Alexander Moodie and John C. H. Spence. John Maxwell Cowley. 18 February 1923–18 May 2004. Biographical Memoirs of Fellows of the Royal Society, 52:29–43, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0005.

Mitchell:2008:FSS

[MS08] Vivien Mitchell and John Spring. Frank Stuart Spring. 5 September 1907–2 March 1997. Biographical Memoirs of Fellows of the Royal Society, 54:375–382, January 2008. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0032.

Minzoni:2015:GBW

[MS15] A. A. Minzoni and N. F. Smyth. Gerald Beresford Whitham. 13 December 1927–26 January 2014. Biographical Memoirs of Fellows of the Royal Society, 61:555–577, January 2015. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0026.

McCann:2019:CWS

[MS19] Michael J. McCann and Colin J. Suckling. Charles Walter Suckling. 24 July 1920–30 October 2013. Biographical Memoirs of Fellows of the Royal Society, 66:423–446, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0025.

Massey:1978:DCM

[MT78] Harrie Stewart Wilson Massey and Harold Warris Thompson. David Christie Martin, 1914–1976. Biographical Memoirs of Fellows of the Royal Society, 24:390–407, November 1978. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0012.

Menon:2008:VRA

[MT08] M. G. K. Menon and P. N. Tandon. Vulimiri Ramalingaswami. 8 August 1921–28 May 2001. Biographical Memoirs of Fellows of the Royal Society, 54:297–317, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0033.

Munn:2019:RNH

[MT19] R. W. Munn and D. R. Taylor. Robert Neville Haszeldine. 3 May 1925–13 October 2016. Biographical Memoirs of Fellows of the Royal Society, 66:243–265, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0035.

MacDonald:2021:JMT

[MT21] D. W. MacDonald and J. N. Thompson. John Marion Thoday. 30 August 1916–25 August 2008. Biographical Memoirs of Fellows of the Royal Society, 71:479–490, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0004.

Muir:1994:SRN

[Mui94] I. D. Muir. Stephen Robert Nockolds, 10 May 1909–7 February 1990. Biographical Memoirs of Fellows of the Royal Society, 40:307– 317, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1994.0041.

Muir:2023:RRO

[Mui23] Tom W. Muir FRS. Robert Ramage. 4 October 1935–16 October 2019. Biographical Memoirs of Fellows of the Royal Society, 75:387– 409, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2023.0010.

Mulvey:1994:VEC

[Mul94] T. Mulvey. Vernon Ellis Cosslett, 16 June 1908–21 November 1990. Biographical Memoirs of Fellows of the Royal Society, 40:61–84, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1994.0029.

Muller:2022:PJC

[Mül22] Rolf Müller. Paul Jozef Crutzen. 3 December 1933–28 January 2021. Biographical Memoirs of Fellows of the Royal Society, 73:127– 156, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2022.0011.

Murrell:1998:MJS

[Mur98] John. N. Murrell. Michael James Steuart Dewar. 24 September 1918–11 October 1997. Biographical Memoirs of Fellows of the Royal Society, 44:129–140, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0009.

Maxwell:2001:CGW

[MW01] Jonathan Maxwell and Prudence M. J. H. Wormell. Charles Gorrie Wynne. 18 May 1911–1 October 1999. Biographical Memoirs of Fellows of the Royal Society, 47:497–514, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0030.

Martin:2020:WJS

[MW20] A. D. Martin FRS and B. R. Webber FRS. William James Stirling CBE. 4 February 1953–9 November 2018. Biographical Memoirs of Fellows of the Royal Society, 68:385–406, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0031.

Mansfield:1973:SPT

[MY73] Eric Harold Mansfield and D. H. Young. Stephen Prokofievitch Timoshenko, 1878–1972. Biographical Memoirs of Fellows of the Royal Society, 19:679–694, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1973.
0025.

Nabarro:1995:EOA

[NA95] Frank Reginald Nunes Nabarro and A. S. Argon. Egon Orowan, 2 August 1902–3 August 1989. Biographical Memoirs of Fellows of the Royal Society, 41:316–340, November 1995. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0020.

Neuberger:2005:CMC

[NA05] Michael S. Neuberger and Brigitte A. Askonas. César Milstein CH. 8 October 1927–24 March 2002. Biographical Memoirs of Fellows of the Royal Society, 51:267–289, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0017.

Nagley:2019:AWL

[Nag19] Phillip Nagley AM. Anthony William Linnane. 17 July 1930– 11 November 2017. Biographical Memoirs of Fellows of the Royal Society, 67:255–278, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0025.

Naismith:2021:IDC

[Nai21] James H. Naismith. Iain Donald Campbell. 24 April 1941–5 March 2014. Biographical Memoirs of Fellows of the Royal Society, 71:59– 77, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2021.0003.

North:2021:GBM

[NC21] R. Alan North and Marcello Costa. Geoffrey Burnstock. 10 May 1929–3 June 2020. Biographical Memoirs of Fellows of the Royal Society, 71:37–58, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0016.

Needham:2002:DWD

[Nee02] Roger M. Needham F.R.ENG., F.R.S. Donald Watts Davies, C.B.E. 7 June 1924–28 May 2000. Biographical Memoirs of Fellows of the Royal Society, 48:87–96, January 2002. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0006.

Neuberger:1973:JND

[Neu73] Albert Neuberger. James Norman Davidson, 1911–1972. Biographical Memoirs of Fellows of the Royal Society, 19:280–303, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1973.0011.

Neuberger:1978:CED

[Neu78] Albert Neuberger. Charles Enrique Dent, 25 August 1911– 19 September 1976. Biographical Memoirs of Fellows of the Royal Society, 24:15–31, November 1978. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1978.
0002.

Neuberger:1984:RKC

[Neu84]  Albert Neuberger.  Robert Kenneth Callow, 15 February

1901–12 April 1983. Biographical Memoirs of Fellows of the Royal Society, 30:91–116, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1984.
0003.

Neuberger:1990:AAM

[Neu90] Albert Neuberger. Arnold Ashley Miles, 20 March 1904–11 February 1988. Biographical Memoirs of Fellows of the Royal Society, 35: 303–326, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Newman:1955:AMT

[New55] Maxwell Herman Alexander Newman. Alan Mathison Turing, 1912–1954. Biographical Memoirs of Fellows of the Royal Society, 1:253–263, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1955.
0019.

Newman:1957:HW

[New57] Maxwell Herman Alexander Newman. Hermann Weyl, 1885–1955. Biographical Memoirs of Fellows of the Royal Society, 3:305–328, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1957.0021.

Newitt:1960:ACG

[New60] Dudley Maurice Newitt. Alfred Charles Glyn Egerton, 1886–1959. Biographical Memoirs of Fellows of the Royal Society, 6:39–64, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1960.0024.

Newman:1961:JHC

[New61] Maxwell Herman Alexander Newman. John Henry Constantine Whitehead, 1904–1960. Biographical Memoirs of Fellows of the Royal Society, 7:349–363, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1961.
0025.

Newitt:1962:PWB

[New62] Dudley Maurice Newitt. Percy Williams Bridgman, 1882–1961. Biographical Memoirs of Fellows of the Royal Society, 8:26–40, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1962.0003.

Newton:1998:VCW

[New98] Ian Newton. Vero Copner Wynne-Edwards, C. B. E.. 4 July 1906–5 January 1997. Biographical Memoirs of Fellows of the Royal Society, 44:473–484, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0030.

Neuberger:1996:CRN

[NG96] Albert Neuberger and Abraham Goldberg. Claude Rimington, 17 November 1902–8 August 1993. Biographical Memoirs of Fellows of the Royal Society, 42:365–378, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0022.

North:2013:HWK

[NH13] R. Alan North FRS and John Hughes FRS. Hans Walter Kosterlitz. 27 April 1903–26 October 1996. Biographical Memoirs of Fellows of the Royal Society, 59:171–192, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0037.

Nicola:2016:DMA

[Nic16] Nicos A. Nicola. Donald Metcalf AC. 26 February 1929– 15 December 2014. Biographical Memoirs of Fellows of the Royal Society, 62:409–431, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2016.

Norman:1986:WAW

[NJ86] Richard Oswald Chandler Norman and J. H. Jones. William Alexander Waters, 8 May 1903–23 January 1985. Biographical Memoirs of Fellows of the Royal Society, 32:597–627, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1986.0020.

Newman:2022:KNJ

[NL22] Andy Newman and Ben Luisi. Kiyoshi Nagai. 25 June 1949– 27 September 2019. Biographical Memoirs of Fellows of the Royal Society, 72:275–291, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0043.

Nabarro:2000:SFC

[NN00] F. R. N. Nabarro and J. F. Nye. Sir (Frederick) Charles Frank, O.B.E. 6 March 1911–5 April 1998. Biographical Memoirs of Fellows of the Royal Society, 46:177–196, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0079.

Northcote:1991:HKP

[Nor91] D. H. Northcote. Helen Kemp Porter, 10 November 1899– 7 December 1987. Biographical Memoirs of Fellows of the Royal Society, 37:399–409, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0020.

Neuberger:1982:RTW

[NS82] Albert Neuberger and R. L. Smith. Richard Tecwyn Williams, 20 February 1909–29 December 1979. Biographical Memoirs of Fellows of the Royal Society, 28:685–717, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0026.

Nair:2024:PMB

[NS24] Venugopal Nair and Martin W. Shirley. Peter Martin Biggs. 13 August 1926–27 December 2021. Biographical Memoirs of Fellows of the Royal Society, 77:49–67, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Nutman:1977:HGT

[Nut77] Phillip Sadler Nutman. Henry Gerard Thornton, 22 January 1892–6 February 1977. Biographical Memoirs of Fellows of the Royal Society, 23:557–574, November 1977. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0021.

Nutman:1990:GBF

[Nut90] Phillip Sadler Nutman. George Bond, 21 February 1906–5 January 1988. Biographical Memoirs of Fellows of the Royal Society, 36:59– 82, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1990.0024.

Nisbet:1984:JCA

[NW84] N. W. Nisbet and Michael Francis Addison Woodruff. John Charnley, 29 August 1911–5 August 1982. Biographical Memoirs of Fellows of the Royal Society, 30:117–137, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1984.0004.

Oakley:1962:JJB

[Oak62] Cyril Leslie Oakley. Jules Jean Baptiste Vincent Bordet, 1870– 1961. Biographical Memoirs of Fellows of the Royal Society, 8:18– 25, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1962.0002.

Oakley:1966:ATG

[Oak66] Cyril Leslie Oakley. Alexander Thomas Glenny, 1882–1965. Biographical Memoirs of Fellows of the Royal Society, 12:162–180, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0007.

Oakley:1968:GRC

[Oak68] Cyril Leslie Oakley. Gordon Roy Cameron, 1899–1966. Biographical Memoirs of Fellows of the Royal Society, 14:83–116, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1968.0005.

Oakley:1971:LC

[Oak71] Cyril Leslie Oakley. Leonard Colebrook, 1883–1967. Biographical Memoirs of Fellows of the Royal Society, 17:90–138, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1971.0004.

Oakley:1973:CHB

[Oak73a] Cyril Leslie Oakley. Carl Hamilton Browning, 1881–1972. Biographical Memoirs of Fellows of the Royal Society, 19:172–215, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1973.0007.

Oakley:1973:FWR

[Oak73b] Cyril Leslie Oakley. Francis William Rogers Brambell, 1901–1970. Biographical Memoirs of Fellows of the Royal Society, 19:129–171, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0006.

Ochoa:1990:LFL

[Och90] Severo Ochoa. Luis Federico Leloir, 6 September 1906–3 December 1987. Biographical Memoirs of Fellows of the Royal Society, 35: 201–208, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Olson:2014:QHG

[OG14] John S. Olson and H. Gutfreund FRS. Quentin Howieson Gibson. 9 December 1918–16 March 2011. Biographical Memoirs of Fellows of the Royal Society, 60:169–210, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0018.

Ogden:2020:PCM

[Ogd20] R. W. Ogden FRS. Peter Chadwick. 23 March 1931–12 August 2018. Biographical Memoirs of Fellows of the Royal Society, 69:109– 131, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2020.0012.

Ogston:1973:HBH

[Ogs73] Alexander George Ogston. Harold Brewer Hartley, 1878–1972. Biographical Memoirs of Fellows of the Royal Society, 19:349–373, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0014.

OHara:2006:SFH

[O’H06] M. J. O’Hara. Sir Frederick Henry Stewart. 16 January 1916– 9 December 2001. Biographical Memoirs of Fellows of the Royal Society, 52:331–348, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0023.

Oliva:2021:GHD

[OKC21] Rafael Oliva, Stephen A. Krawetz, and GHD Consortium. Gordon Henry Dixon. 25 March 1930–24 July 2016. Biographical Memoirs of Fellows of the Royal Society, 70:131–149, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0019.

Oliverio:1994:DBM

[Oli94] A. Oliverio. Daniel Bovet, 23 March 1907–8 April 1992. Biographical Memoirs of Fellows of the Royal Society, 39:59–70, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Owen:1980:DKS

[OM80] Paul Robert Owen and E. C. Maskell. Dietrich Kücherman, 11 September 1911–23 February 1976. Biographical Memoirs of Fellows of the Royal Society, 26:304–326, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1980.0009.

Oliphant:1968:JDC

[OP68] Marcus Laurence Elwin Oliphant and William George Penney. John Douglas Cockcroft, 1897–1967. Biographical Memoirs of Fellows of the Royal Society, 14:139–188, November 1968.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0007.

Ostriker:2007:LSJ

[Ost07] Jeremiah P. Ostriker. Lyman Spitzer. 26 June 1914–31 March 1997. Biographical Memoirs of Fellows of the Royal Society, 53:339–348, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2007.0020.

Ottewill:2004:DHE

[Ott04] Ronald Ottewill. Douglas Hugh Everett MBE. 26 December 1916–25 June 2002. Biographical Memoirs of Fellows of the Royal Society, 50:93–107, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0008.

Overend:1997:MSC

[Ove97] W. G. Overend. Maurice Stacey, C. B. E. 8 April 1907–9 October 1994. Biographical Memoirs of Fellows of the Royal Society, 43:473– 489, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1997.0026.

Owen:1995:DJM

[Owe95] Maureen E. Owen. Dame Janet Maria Vaughan, D. B. E., 18 October 1899–9 January 1993. Biographical Memoirs of Fellows of the Royal Society, 41:482–498, November 1995. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0029.

Owen:2009:OOC

[Owe09] D. Roger J. Owen FRS. Olgierd (Olek) Cecil Zienkiewicz CBE. 18 May 1921–2 January 2009. Biographical Memoirs of Fellows of the Royal Society, 55:337–359, January 2009. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0008.

Oxley:2023:LAB

[Oxl23] B. W. Oxley. Lionel Alexander Bethune (Alastair) Pilkington. 7 January 1920–5 May 1995. Biographical Memoirs of Fellows of the Royal Society, 74:339–361, April 2023. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0025.

Page:2001:WJS

[Pag01] Frederick Page. William John Strang, C.B.E., F.R.Eng., F.R.Ae.S. 29 June 1921–14 September 1999. Biographical Memoirs of Fellows of the Royal Society, 47:443–451, January 2001. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0026.

Palmer:2009:ENL

[Pal09] T. N. Palmer. Edward Norton Lorenz. 23 May 1917–16 April 2008. Biographical Memoirs of Fellows of the Royal Society, 55:139–155, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2009.0004.

Pantin:1957:OWT

[Pan57] Carl Frederick Abel Pantin. Oscar Werner Tiegs, 1897–1956. Biographical Memoirs of Fellows of the Royal Society, 3:247–255, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1957.0017.

Pantin:1965:RBS

[Pan65] Carl Frederick Abel Pantin. Robert Beresford Seymour Sewell, 1880–1964. Biographical Memoirs of Fellows of the Royal Society, 11:146–155, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1965.
0010.

Papaioannou:2019:SGW

[Pap19] Virginia E. Papaioannou PhD. Salome Gluecksohn-Waelsch. 6 October 1907–7 November 2007. Biographical Memoirs of Fellows of the Royal Society, 67:153–171, December 2019. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0024.

Parkes:1972:SJF

[Par72] Alan Sterling Parkes. Sydney John Folley, 1906–1970. Biographical Memoirs of Fellows of the Royal Society, 18:240–265, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1972.0008.

Pashley:1987:MBD

[Pas87] Donald William Pashley. Moses Blackman, 6 December 1908– 3 June 1983. Biographical Memoirs of Fellows of the Royal Society, 33:47–64, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0003.

Passingham:2019:ACA

[Pas19] Dick Passingham FRS. Alan Cowey. 28 April 1935–19 December 2012. Biographical Memoirs of Fellows of the Royal Society, 66: 101–122, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2018.0017.

Paton:1979:WHA

[Pat79] Thomas Angus Lyall Paton. William Hudson, 27 April 1896– 12 September 1978. Biographical Memoirs of Fellows of the Royal Society, 25:318–335, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1979.
0011.

Paterson:1982:JCJ

[Pat82] M. S. Paterson. John Conrad Jaeger, 30 July 1907–15 May 1979. Biographical Memoirs of Fellows of the Royal Society, 28:163–203, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1982.0009.

Patterson:1990:EHO

[Pat90] Colin Patterson. Erik Helge Osvald Stensiö, 2 October 1891–11 January 1984. Biographical Memoirs of Fellows of the Royal Society, 35:361–380, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Patterson:1997:PHG

[Pat97] C. Patterson. Peter Humphry Greenwood. 21 April 1927–3 March 1995. Biographical Memoirs of Fellows of the Royal Society, 43:195– 213, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1997.0011.

Pattenden:2001:LCJ

[Pat01] Gerald Pattenden. Leslie Crombie. 10 June 1923–3 August 1999. Biographical Memoirs of Fellows of the Royal Society, 47:125–140, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2001.0008.

Pattenden:2005:BCL

[Pat05] Gerald Pattenden. Basil Charles Leicester Weedon. 18 July 1923–10 October 2003. Biographical Memoirs of Fellows of the Royal Society, 51:425–436, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2005.
0029.

Paulson:2018:MJC

[Pau18] Lawrence C. Paulson FRS. Michael John Caldwell Gordon. 28 February 1948–22 August 2017. Biographical Memoirs of Fellows of the Royal Society, 65:89–113, December 2018. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0019.

Payne:1981:EWT

[Pay81] B. O. Payne. Edward Wilfred Taylor, 29 April 1891–1 November 1980. Biographical Memoirs of Fellows of the Royal Society, 27:562– 577, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1981.0022.

Partridge:1979:PVE

[PB79] Stanley Miles Partridge and Birger Blombäck. Pehr Victor Edman, 14 April 1916–19 March 1977. Biographical Memoirs of Fellows of the Royal Society, 25:241–265, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0008.

Petch:1984:AMM

[PB84] Norman James Petch and L. Barnard. Andrew McCance, 30 March 1889–11 June 1983. Biographical Memoirs of Fellows of the Royal Society, 30:388–405, November 1984. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1984.0014.

Peto:2010:SRD

[PB10] Richard Peto and Valerie Beral. Sir Richard Doll CH OBE. 28 October 1912–24 July 2005. Biographical Memoirs of Fellows of the Royal Society, 56:63–83, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2010.
0019.

Powlson:2017:DSJ

[PB17] David Powlson and Phil Brookes. David Stewart Jenkinson. 25 February 1928–16 February 2011. Biographical Memoirs of Fellows of the Royal Society, 63:377–411, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0007.

Pines:2020:ELH

[PB20] Alexander Pines and Dmitry Budker. Erwin Louis Hahn. 9 June 1921–20 September 2016. Biographical Memoirs of Fellows of the Royal Society, 68:219–230, June 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0038.

Pittard:2010:FWE

[PC10] A. J. Pittard and G. B. Cox. Frank William Ernest Gibson. 22 July 1923–11 July 2008. Biographical Memoirs of Fellows of the Royal Society, 56:85–103, January 2010. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0020. See correction [PC11].

Pittard:2011:CFW

[PC11] A. J. Pittard AM FAA and G. B. Cox FAA. Correction for Frank William Ernest Gibson. 22 July 1923–11 July 2008. Biographical Memoirs of Fellows of the Royal Society, 57:463, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0022. See [PC10].

Pettigrew:2018:POB

[PD18] Jack D. Pettigrew FRS and Bogdan Dreher. Peter Orlebar Bishop. 14 June 1917–3 June 2012. Biographical Memoirs of Fellows of the Royal Society, 64:51–68, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Peacock:1968:JWH

[Pea68] A. D. Peacock. John William Heslop Harrison, 1881–1967. Biographical Memoirs of Fellows of the Royal Society, 14:243–270, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1968.0011.

Peart:1995:MMM

[Pea95] William Stanley Peart. Malcolm Milne, 22 May 1915–3 April 1991. Biographical Memoirs of Fellows of the Royal Society, 41:298–307, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0018.

Pease:1998:HOG

[Pea98] R. S. Pease. Hannes Olof Gosta Alfven. 30 May 1908–2 April 1995. Biographical Memoirs of Fellows of the Royal Society, 44:3–19, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1998.0001.

Pearce:2018:NJS

[Pea18] John M. Pearce. Nicholas John Seymour Mackintosh. 9 July 1935– 8 February 2015. Biographical Memoirs of Fellows of the Royal Society, 64:299–316, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Pedley:2001:SMJ

[Ped01] T. J. Pedley. Sir (Michael) James Lighthill. 23 January 1924–17 July 1998. Biographical Memoirs of Fellows of the Royal Society, 47:333–356, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0019.

Peierls:1960:WEP

[Pei60] Rudolf Ernst Peierls. Wolfgang Ernst Pauli, 1900–1958. Biographical Memoirs of Fellows of the Royal Society, 5:174–192, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/769285.

Peierls:1981:ORF

[Pei81] Sir Rudolf Peierls, F.R.S. Otto Robert Frisch, 1 October 1904–22 September 1979. Biographical Memoirs of Fellows of the Royal Society, 27:283–306, November 1981. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/10.2307/769874;
 http://www.jstor.org/stable/769874;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1981.0012.

Penney:1967:HJB

[Pen67] William George Penney. Homi Jehangir Bhabha, 1909–1966. Biographical Memoirs of Fellows of the Royal Society, 13:35–55, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1967.0002.

Pepys:2003:DSP

[Pep03] Mark Pepys FRS. Dame Sheila Patricia Violet Sherlock. 31 March 1918–30 December 2001 Elected FRS 2001. Biographical Memoirs of Fellows of the Royal Society, 49:475–493, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0028.

Pereira:1982:BAK

[Per82] Herbert Charles Pereira. Bernard Augustus Keen, 5 September 1890–5 August 1981. Biographical Memoirs of Fellows of the Royal Society, 28:205–223, November 1982. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1982.
0010.

Perry:1987:RRP

[Per87] Samuel Victor Perry. Rodney Robert Porter, 8 October 1917– 6 September 1985. Biographical Memoirs of Fellows of the Royal Society, 33:443–489, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0017.

Perham:1988:MDA

[Per88] Richard Nelson Perham. Malcolm Dixon, 18 April 1899– 7 December 1985. Biographical Memoirs of Fellows of the Royal Society, 34:97–131, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.
0005.

Peters:1958:JBL

[Pet58] Rudolph Albert Peters. John Beresford Leathes, 1864–1956. Biographical Memoirs of Fellows of the Royal Society, 4:185–191, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1958.0016.

Petch:1992:SHM

[Pet92] Norman James Petch. Sir Harold Montague Finniston, 15 August 1912–2 February 1991. Biographical Memoirs of Fellows of the Royal Society, 38:131–144, November 1992. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0007.

Petersen:2011:KBJ

[Pet11] George B. Petersen ONZM FRSNZ. Kenneth Burton. 26 June 1926–22 November 2010. Biographical Memoirs of Fellows of the Royal Society, 57:79–96, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2011.

Petersen:2024:SMJ

[Pet24] Ole H. Petersen FRS. Sir Michael John Berridge. 22 October 1938–13 February 2020. Biographical Memoirs of Fellows of the Royal Society, 76:69–93, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0047.

Patterson:1999:TSW

[PF99] Colin Patterson and R. A. Fortey. Thomas Stanley Westoll. 3 July 1912–19 September 1995. Biographical Memoirs of Fellows of the Royal Society, 45:531–546, January 1999. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0035.

Pierce:2021:JPG

[PF21] Simon Pierce and Jason D. Fridley. John Philip Grime. 30 April 1935–19 April 2021. Biographical Memoirs of Fellows of the Royal Society, 71:249–270, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0021.

Pfeil:1960:CAE

[Pfe60] Leonard Bessemer Pfeil. Charles Alfred Edwards, 1882–1960. Biographical Memoirs of Fellows of the Royal Society, 6:32–38, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1960.0023.

Phillips:1992:HGJ

[PG92] Charles Garrett Phillips and Rainer Walter Guillery. Henricus Gerardus Jacobus Maria Kuypers, 9 September 1925–26 September 1989. Biographical Memoirs of Fellows of the Royal Society, 38:185– 207, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1992.0010.

Phillips:1974:FMR

[Phi74] Charles Garrett Phillips. Francis Martin Rouse Walshe, 1885–1973. Biographical Memoirs of Fellows of the Royal Society, 20:456–481, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1974.0020.

Phillips:1979:WLB

[Phi79] David Chilton Phillips. William Lawrence Bragg, 31 March 1890– 1 July 1971. Biographical Memoirs of Fellows of the Royal Society, 25:74–143, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://doi.org/10.

1098/rsbm.1979.0003;   https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1979.0003.

Phillips:1983:EGT

[Phi83] Charles Garrett Phillips. Edward George Tandy Liddell, 25 March 1895–17 August 1981. Biographical Memoirs of Fellows of the Royal Society, 29:332–359, November 1983. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1983.0014.

Phillips:1995:DHD

[Phi95] C. S. G. Phillips. Denis Henry Desty, 21 October 1923–18 January 1994. Biographical Memoirs of Fellows of the Royal Society, 41:136– 143, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1995.0009.

Phillips:2013:NEB

[Phi13] Ronald L. Phillips. Norman Ernest Borlaug. 25 March 1914– 12 September 2009. Biographical Memoirs of Fellows of the Royal Society, 59:59–72, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2013.
0012.

Pickering:1968:WRB

[Pic68] George White Pickering. Walter Russell Brain, First Baron Brain of Eynsham, 1895–1966. Biographical Memoirs of Fellows of the Royal Society, 14:61–82, November 1968. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0004.

Pickering:1974:MLR

[Pic74] George White Pickering. Max Leonard Rosenheim, Baron Rosenheim of Camden, 1908–1972. Biographical Memoirs of Fellows of the Royal Society, 20:348–358, December 1974. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1974.0015.

Pickett:2016:MEC

[Pic16] John Pickett FRS. Michael Elliott CBE. 30 September 1924– 17 October 2007. Biographical Memoirs of Fellows of the Royal Society, 62:109–123, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2016.
0018.

Pierpoint:1999:NWP

[Pie99] W. S. Pierpoint. Norman Wingate Pirie. 1 July 1907–29 March 1997. Biographical Memoirs of Fellows of the Royal Society, 45:397– 415, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0027.

Piggott:1977:REM

[Pig77] S. Piggott. Robert Eric Mortimer Wheeler, 10 September 1890–22 July 1976. Biographical Memoirs of Fellows of the Royal Society, 23:623–642, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0023.

Pippard:1994:JBM

[Pip94] Alfred Brian Pippard. John Bardeen, 23 May 1908–30 January 1991. Biographical Memoirs of Fellows of the Royal Society, 39:19– 34, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Pippard:1998:SNF

[Pip98] Brian Pippard. Sir Nevill Francis Mott, C. H. 30 September 1905–8 August 1996. Biographical Memoirs of Fellows of the Royal Society, 44:315–328, January 1998. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0021.

Pippard:2005:DSJ

[Pip05] Sir Brian Pippard. David Shoenberg. 4 January 1911–10 March 2004. Biographical Memoirs of Fellows of the Royal Society, 51:379– 395, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2005.0025.

Pirie:1966:JBS

[Pir66] Norman Wingate Pirie. John Burdon Sanderson Haldane, 1892– 1964. Biographical Memoirs of Fellows of the Royal Society, 12:218– 249, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1966.0010.

Pirie:1973:FCB

[Pir73] Norman Wingate Pirie. Frederick Charles Bawden, 1908–1972. Biographical Memoirs of Fellows of the Royal Society, 19:19–63, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0002.

Pirie:1990:JBM

[Pir90] Norman Wingate Pirie. Jean Brachet, 19 March 1909–10 August 1988. Biographical Memoirs of Fellows of the Royal Society, 36:83– 99, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1990.0025.

Pitt:1994:JCB

[Pit94] Harry Raymond Pitt. John Charles Burkill, 1 February 1900–6 April 1993. Biographical Memoirs of Fellows of the Royal Society, 40:43–59, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.

0028. Reprinted in [Pit98].

Pitt:1998:JCB

[Pit98] H. R. Pitt. John Charles Burkill. Bulletin of the London Mathematical Society, 30(1):85–98, January 1998. CODEN LMSBBT. ISSN 0024-6093 (print), 1469-2120 (electronic). URL https://doi.org/

10.1112/s0024609397003767. Reprinted from [Pit94].

Prelog:1980:LRS

[PJ80] Vladimir Prelog and Oskar Jeger. Leopold Ruzicka, 13 September 1887–26 September 1976. Biographical Memoirs of Fellows of the Royal Society, 26:411–501, November 1980. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1980.0013.

Pepper:2024:SDH

[PK24] Michael Pepper FRS and Michael Kelly FRS. Sir Derek Harry Roberts. 28 March 1932–17 February 2021. Biographical Memoirs of Fellows of the Royal Society, 76:341–363, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0039.

Penney:1988:WRJ

[PM88] William George Penney and V. H. B. Macklen. William Richard Joseph Cook, 10 April 1905–16 September 1987. Biographical Memoirs of Fellows of the Royal Society, 34:43–61, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0003.

Peacock:1994:MJD

[PM94] William James Peacock and D. McCann. Michael James Denham White, 20 August 1910–16 December 1983. Biographical Memoirs of Fellows of the Royal Society, 40:401–419, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1994.0047.

Peaker:2021:LGG

[PO21] Malcolm Peaker FRS and Dame Bridget M. Ogilvie FRS. Leonard George Goodwin. 11 July 1915–25 November 2008. Biographical Memoirs of Fellows of the Royal Society, 71:213–227, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0050.

Polani:1992:JAF

[Pol92] Paul Emanuel Polani. John Alexander Fraser Roberts, 8 September 1899–15 January 1987. Biographical Memoirs of Fellows of the Royal Society, 38:305–322, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1992.
0016.

Polge:2000:LEA

[Pol00] Christopher Polge. Lionel Edward Aston Rowson, O.B.E. 28 May 1914–26 July 1989. Biographical Memoirs of Fellows of the Royal Society, 46:483–497, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0097.

Polge:2006:SAS

[Pol06] Christopher Polge. Sir Alan Sterling Parkes. 10 September 1900–17 July 1990. Biographical Memoirs of Fellows of the Royal Society, 52:263–283, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0019.

Pontecorvo:1968:HJM

[Pon68] Guido Pontecorvo. Hermann Joseph Muller, 1890–1967. Biographical Memoirs of Fellows of the Royal Society, 14:348–389, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1968.0015.

Poole:1960:WRG

[Poo60] Horace H. Poole. William Ringrose Gelston Atkins, 1884–1959. Biographical Memoirs of Fellows of the Royal Society, 5:1–22, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Porter:1966:FJR

[Por66] Helen Kemp Porter. Francis John Richards, 1901–1965. Biographical Memoirs of Fellows of the Royal Society, 12:422–436, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1966.0020.

Porter:1968:VHB

[Por68] Helen Kemp Porter. Vernon Herbert Blackman, 1872–1967. Biographical Memoirs of Fellows of the Royal Society, 14:37–60, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1968.0003.

Porter:1996:CGP

[Por96] Robert Porter. Charles Garrett Phillips, 13 October 1916– 9 September 1994. Biographical Memoirs of Fellows of the Royal Society, 42:341–362, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0021.

Powell:1987:EML

[Pow87] Michael James David Powell. Evelyn Martin Lansdowne Beale, 8 September 1928–23 December 1985. Biographical Memoirs of Fellows of the Royal Society, 33:21–45, December 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0002.

Porter:1963:FGG

[PR63] Helen Kemp Porter and Francis John Richards. Frederick Gugenheim Gregory, 1893–1961. Biographical Memoirs of Fellows of the Royal Society, 9:131–153, November 1963. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1963.0007.

Porter:1978:MTD

[PR78] Helen Kemp Porter and S. L. Ranson. Meirion Thomas, 28 December 1894–5 April 1977. Biographical Memoirs of Fellows of the Royal Society, 24:547–568, November 1978. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0018.

Pugsley:1981:BNW

[PR81] Alfred Grenville Pugsley and N. E. Rowe. Barnes Neville Wallis, 26 September 1887–30 October 1979. Biographical Memoirs of Fellows of the Royal Society, 27:603–627, November 1981.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0024.

Peters:2022:BMC

[PR22] Jo Peters and Sohaila Rastan. Bruce Macintosh Cattanach. 5 November 1932–8 April 2020. Biographical Memoirs of Fellows of the Royal Society, 73:85–106, December 2022. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0013.

Praeger:2010:BHN

[Pra10] Cheryl E. Praeger. Bernhard Hermann Neumann AC. 15 October 1909–21 October 2002. Biographical Memoirs of Fellows of the Royal Society, 56:285–316, January 2010. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0002.

Preston:1990:IMA

[Pre90] Reginald Dawson Preston. Irene Manton, 17 April 1904–13 May 1988. Biographical Memoirs of Fellows of the Royal Society, 35: 247–261, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Prebble:2010:JBH

[Pre10] John N. Prebble. Jeffrey Barry Harborne. 1 September 1928– 21 July 2002. Biographical Memoirs of Fellows of the Royal Society, 56:131–147, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2010.
0012.

Pringle:1968:RJP

[Pri68] John William Sutton Pringle. Richard Julius Pumphrey, 1906– 1967. Biographical Memoirs of Fellows of the Royal Society, 14:435– 442, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1968.0018.

Price:1978:ADW

[Pri78] William Charles Price. Arthur Donald Walsh, 8 August 1916–23 April 1977. Biographical Memoirs of Fellows of the Royal Society, 24:569–582, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1978.
0019.

Price:1994:RED

[Pri94] William Geraint Price. Richard Evelyn Donohue Bishop, 1 January 1925–12 September 1989. Biographical Memoirs of Fellows of the Royal Society, 40:1–29, November 1994. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1994.0026.

Proudman:1964:GRG

[Pro64] Joseph Proudman. George Ridsdale Goldsbrough, 1881–1963. Biographical Memoirs of Fellows of the Royal Society, 10:107–116, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0007.

Proudman:1968:ATD

[Pro68] Joseph Proudman. Arthur Thomas Doodson, 1890–1968. Biographical Memoirs of Fellows of the Royal Society, 14:189–205, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1968.0008.

Proctor:2023:NOW

[Pro23] Michael Proctor FRS. Nigel Oscar Weiss. 16 December 1936– 24 June 2020. Biographical Memoirs of Fellows of the Royal Society, 74:439–456, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0039.

Potter:1997:MLB

[PS97] Paul Potter and Hubert Soltan. Murray Llewellyn Barr, O. C. 20 June 1908–4 May 1995. Biographical Memoirs of Fellows of the Royal Society, 43:33–46, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.

Parker:2021:RMS

[PSCV21] Ian Parker FRS, Clarke Slater, Stuart Cull-Candy FRS, and Angela Vincent FRS. Ricardo Miledi. 15 September 1927–18 December 2017. Biographical Memoirs of Fellows of the Royal Society, 71:423– 450, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2021.0020.

Pasquill:1978:OGS

[PSS78] Frank Pasquill, Percival Albert Sheppard, and Reginald Cockcroft Sutcliffe. Oliver Graham Sutton, 4 February 1903–26 May 1977. Biographical Memoirs of Fellows of the Royal Society, 24:529–546, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1978.0017.

Pooley:2016:BEC

[PSW16] Derek Pooley CBE, George Smith FRS, and Colin Windsor FRS. Brian Eyre CBE FREng. 29 November 1933–28 July 2014. Biographical Memoirs of Fellows of the Royal Society, 62:147–166, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2016.0016.

Preer:2011:GHB

[PT11] John R. Preer and Andrew Tait FRSE. Geoffrey Herbert Beale MBE. 11 June 1913–16 October 2009. Biographical Memoirs of Fellows of the Royal Society, 57:45–62, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0025.

Puddephatt:2017:MHC

[Pud17] Richard J. Puddephatt OC FRS. Malcolm Harold Chisholm. 15 October 1945–20 November 2015. Biographical Memoirs of Fellows of the Royal Society, 63:117–136, January 2017. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0025.

Puddephatt:2021:CEH

[Pud21] Richard J. Puddephatt FRS. Cecil Edwin Henry Bawn. 6 November 1908–19 September 2003. Biographical Memoirs of Fellows of the Royal Society, 70:9–21, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0034.

Puddephatt:2024:AGD

[Pud24] Richard J. Puddephatt. Alwyn George Davies. 13 May 1926–1 September 2023. Biographical Memoirs of Fellows of the Royal Society, 77:107–127, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Pugh:1958:AET

[Pug58] William John Pugh. Arthur Elijah Trueman, 1894–1956. Biographical Memoirs of Fellows of the Royal Society, 4:291–305, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1958.0023.

Pugsley:1961:RAF

[Pug61] Alfred Grenville Pugsley. Robert Alexander Frazer, 1891–1959. Biographical Memoirs of Fellows of the Royal Society, 7:75–84, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0006.

Pugh:1967:OTJ

[Pug67] William John Pugh. Owen Thomas Jones, 1878–1967. Biographical Memoirs of Fellows of the Royal Society, 13:222–243, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1967.0011.

Pugsley:1978:ARJ

[Pug78] Alfred Grenville Pugsley. Andrew Robertson, 30 January 1883– 22 October 1977. Biographical Memoirs of Fellows of the Royal Society, 24:514–528, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1978.
0016.

Parrington:1974:DMS

[PW74] Francis Rex Parrington and Thomas Stanley Westoll. David Meredith Seares Watson, 1886–1973. Biographical Memoirs of Fellows of the Royal Society, 20:482–504, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1974.0021.

Quayle:2003:LRC

[QG03] J. Rodney Quayle FRS and Geoffrey W. Greenwood FRENG FRS. Leonard Rotherham CBE. 31 August 1913–23 March 2001. Biographical Memoirs of Fellows of the Royal Society, 49:431–446, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2003.0025.

Quinn:2012:CDM

[QRJ12] Niall Quinn, John Rothwell, and Peter Jenner. Charles David Marsden. 15 April 1938–29 September 1998. Biographical Memoirs of Fellows of the Royal Society, 58:203–228, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0026.

Quinn:1986:TMS

[QT86] C. P. Quinn and Brian Arthur Thrush. Theodore Morris Sugden, 31 December 1919–3 January 1984. Biographical Memoirs of Fellows of the Royal Society, 32:569–596, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1986.0019.

Quayle:1994:WCE

[Qua94] John Rodney Quayle. William Charles Evans, 1 October 1911–24 July 1988. Biographical Memoirs of Fellows of the Royal Society, 40:85–103, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.
0030.

Quinn:2001:DCM

[Qui01] Peter J. Quinn. Dennis Chapman. 6 May 1927–28 October 1999. Biographical Memoirs of Fellows of the Royal Society, 47:55–66, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2001.0004.

Quinn:2005:SAH

[Qui05] Terry Quinn. Sir Alan Hugh Cook. 2 December 1922–23 July 2004. Biographical Memoirs of Fellows of the Royal Society, 51:87–100, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2005.0006.

Rae:2012:ALJ

[Rae12] Ian D. Rae. Athelstan Laurence Johnson Beckwith OA FAA. 20 February 1930–15 May 2010. Biographical Memoirs of Fellows of the Royal Society, 58:3–21, January 2012. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2011.0022.

Raisman:2008:WMC

[Rai08] Geoffrey Raisman. William Maxwell Cowan. 27 September 1931–30 June 2002. Biographical Memoirs of Fellows of the Royal Society, 54:117–136, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.
0002.

Ramsbottom:1964:GHC

[Ram64] J. Ramsbottom. Gordon Herriot Cunningham, 1892–1962. Biographical Memoirs of Fellows of the Royal Society, 10:15–37, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0002.

Ramsbottom:2022:LWM

[Ram22] Paul Ramsbottom. Lord Wolfson of Marylebone. 11 November 1927–20 May 2010. Biographical Memoirs of Fellows of the Royal Society, 72:431–445, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0035.

Randall:1975:EJH

[Ran75] John Turton Randall. Emmeline Jean Hanson, 14 November 1919–10 August 1973. Biographical Memoirs of Fellows of the Royal Society, 21:313–344, November 1975. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1975.
0008.

Randle:1986:CFC

[Ran86] Philip John Randle. Carl Ferdinand Cori, 5 December 1896– 20 October 1984. Biographical Memoirs of Fellows of the Royal Society, 32:65–95, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0003.

Randle:1990:FGY

[Ran90] Philip John Randle. Frank George Young, 25 March 1908– 20 September 1988. Biographical Memoirs of Fellows of the Royal Society, 36:581–599, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0045.

Rang:2009:JMR

[Ran09] H. P. Rang FRS. Joseph Murdoch Ritchie. 10 June 1925–9 July 2008. Biographical Memoirs of Fellows of the Royal Society, 55:241– 255, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2009.0011.

Rand:2022:SEC

[Ran22] David A. Rand. Sir Erik Christopher Zeeman. 4 February 1925–13 February 2016. Biographical Memoirs of Fellows of the Royal Society, 73:521–547, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0012.

Rao:1973:PCM

[Rao73] Calyampudi Radhakrishna Rao. Prasantha Chandra Mahalanobis, 1893–1972. Biographical Memoirs of Fellows of the Royal Society, 19:455–492, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1973.

Ratcliffe:1960:TLE

[Rat60] John Ashworth Ratcliffe. Thomas Lydwell Eckersley, 1886–1959. Biographical Memoirs of Fellows of the Royal Society, 5:69–74, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Ratcliffe:1964:SKM

[Rat64] John Ashworth Ratcliffe. Sisir Kumar Mitra, 1890–1963. Biographical Memoirs of Fellows of the Royal Society, 10:221–227, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1964.0013.

Ratcliffe:1966:EVA

[Rat66] John Ashworth Ratcliffe. Edward Victor Appleton, 1892–1965. Biographical Memoirs of Fellows of the Royal Society, 12:1–21, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0001.

Ratcliffe:1971:WHE

[Rat71] John Ashworth Ratcliffe. William Henry Eccles, 1875–1966. Biographical Memoirs of Fellows of the Royal Society, 17:195–214, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0008.

Ratcliffe:1975:RAW

[Rat75] John Ashworth Ratcliffe. Robert Alexander Watson-Watt, 13 April 1892–5 December 1973. Biographical Memoirs of Fellows of the Royal Society, 21:548–568, November 1975. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1975.0018.

Raynor:1969:WHR

[Ray69] Geoffrey Vincent Raynor. William Hume-Rothery, 1899–1968. Biographical Memoirs of Fellows of the Royal Society, 15:109–139, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1969.0006.

Raynor:1976:ARP

[Ray76] Geoffrey Vincent Raynor. Alan Richard Powell, 6 March 1894– 11 October 1975. Biographical Memoirs of Fellows of the Royal Society, 22:306–318, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1976.
0013.

Richards:2020:BDB

[RB20] Christopher D. Richards and Timothy V. P. Bliss FRS. Benedict Delisle Burns. 22 February 1915–6 September 2001. Biographical Memoirs of Fellows of the Royal Society, 68:49–70, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0049.

Rickards:2007:AJB

[RC07] Rodney W. Rickards and John Cornforth. Arthur John Birch. 3 August 1915–8 December 1995. Biographical Memoirs of Fellows of the Royal Society, 53:21–44, January 2007. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0017.

Ramakrishnan:2024:SFJ

[RC24] Lalita Ramakrishnan FRS and Brendan P. Cormack. Stanley Falkow. 24 January 1934–5 May 2018. Biographical Memoirs of Fellows of the Royal Society, 76:169–199, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0045.

Riley:2004:GDH

[RE04] Ralph Riley and John Enderby. George Douglas Hutton Bell. 18 October 1905–27 June 1993. Biographical Memoirs of Fellows of the Royal Society, 50:35–38, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0003.

Rogers:2008:RHS

[RE08] George E. Rogers and William H. Elliott. Robert Henry Symons. 20 March 1934–4 October 2006. Biographical Memoirs of Fellows of the Royal Society, 54:383–400, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0012.

Read:2019:RHM

[Rea19] P. L. Read. Raymond Hide. 17 May 1929–6 September 2016. Biographical Memoirs of Fellows of the Royal Society, 67:191–215, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2019.0016.

Redman:1956:WMH

[Red56] Roderick Oliver Redman. William Michael Herbert Greaves, 1897– 1955. Biographical Memoirs of Fellows of the Royal Society, 2:128– 138, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1956.0009.

Reeve:1994:NMV

[Ree94] Suzanne Reeve. Nathaniel Mayer Victor Rothschild, G. B. E., G. M., Third Baron Rothschild, 31 October 1910–20 March 1990. Biographical Memoirs of Fellows of the Royal Society, 39:363–380, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Rees:1998:RMB

[Ree98] Lovat V. C. Rees. Richard Maling Barrer. 16 June 1910–12 September 1996. Biographical Memoirs of Fellows of the Royal Society, 44: 37–49, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1998.0003.

Rees:2008:WPJ

[Ree08] S. M. Rees. William Parry. 3 July 1934–20 August 2006. Biographical Memoirs of Fellows of the Royal Society, 54:229–243, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0035.

Rees:2015:SAS

[Ree15] Sir Dai Rees FRS. Struther Arnott. 25 September 1934–20 April 2013. Biographical Memoirs of Fellows of the Royal Society, 61:

5–22, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2015.0011.

Reid:2014:WJF

[Rei14] S. R. Reid FREng FRSE. William Johnson FREng. 20 April 1922–13 June 2010. Biographical Memoirs of Fellows of the Royal Society, 60:261–275, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2014.
0001.

Reithmeier:2021:DHM

[Rei21] Reinhart Reithmeier. David Herman MacLennan. 3 July 1937–24 June 2020. Biographical Memoirs of Fellows of the Royal Society, 71:317–333, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2021.
0005.

Reid:2024:SDR

[Rei24] Nancy Reid. Sir David R. Cox. 15 July 1924–18 January 2022. Biographical Memoirs of Fellows of the Royal Society, 77:69–91, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Relf:1961:WJD

[Rel61] Ernest Frederick Relf. William Jolly Duncan, 1894–1960. Biographical Memoirs of Fellows of the Royal Society, 7:37–51, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1961.0004.

Reynolds:2007:EFG

[Rey07] P. E. Reynolds. Ernest Frederick Gale. 15 July 1914–7 March 2005. Biographical Memoirs of Fellows of the Royal Society, 53:143–161, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2007.0012.

Rimington:1976:MRL

[RG76] Claude Rimington and C. H. Gray. Max Rudolf Lemberg, 19 October 1896–10 April 1975. Biographical Memoirs of Fellows of the Royal Society, 22:256–294, November 1976. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0011.

Rothwell:2006:PAM

[RG06] John Rothwell and Ian Glynn. Patrick Anthony Merton. 8 October 1920–13 June 2000. Biographical Memoirs of Fellows of the Royal Society, 52:189–201, January 2006. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0014.

Richards:2018:DCS

[RH18] Randal W. Richards,OBE, DSc, PhD, FRSC and Philip Hodge, PhD, FRSC. David Colin Sherrington. 5 March 1945–4 October 2014. Biographical Memoirs of Fellows of the Royal Society, 64: 367–385, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Ramakrishnan:2022:TAS

[RH22] V. Ramakrishnan FRS and Richard Henderson FRS. Thomas Arthur Steitz. 23 August 1940–9 October 2018. Biographical Memoirs of Fellows of the Royal Society, 72:311–336, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0029.

Richardson:1981:CFG

[Ric81] Frederick Denys Richardson. Charles Frederick Goodeve, 21 February 1904–7 April 1980. Biographical Memoirs of Fellows of the Royal Society, 27:307–353, November 1981. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0013.

Richards:1985:HWT

[Ric85] Rex Edward Richards. Harold Warris Thompson, 15 February 1908–31 December 1983. Biographical Memoirs of Fellows of the Royal Society, 31:573–610, November 1985. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1985.0020.

Rickinson:2014:HSC

[Ric14] Alan Rickinson. Harry Smith CBE. 7 August 1921–10 December 2011. Biographical Memoirs of Fellows of the Royal Society, 60:397– 411, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2014.0014.

Rickard:2015:MJO

[Ric15] David Rickard. Michael John O’Hara. 22 February 1933– 24 November 2014. Biographical Memoirs of Fellows of the Royal Society, 61:305–330, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0019.

Riley:1960:HEK

[Ril60] N. D. Riley. Heinrich Ernst Karl Jordan, 1861–1959. Biographical Memoirs of Fellows of the Royal Society, 6:106–133, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1960.0027.

Ritchie:2000:WWD

[Rit00] J. M. Ritchie. William Wilton Douglas. 15 August 1922–2 July 1998. Biographical Memoirs of Fellows of the Royal Society, 46:145– 164, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0077.

Rivers:2020:EOR

[Riv20] Rodney P. A. Rivers. Edward Osmund Royle Reynolds CBE. 3 February 1933–24 April 2017. Biographical Memoirs of Fellows of the Royal Society, 68:333–351, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0018.

Rutter:2010:SMR

[RK10] Michael Rutter and Aaron Klug. Sir Martin Roth. 6 November 1917–26 September 2006. Biographical Memoirs of Fellows of the Royal Society, 56:377–389, January 2010. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0018.

Robinson:1960:JLS

[Rob60] Robert Robinson. John Lionel Simonsen, 1884–1957. Biographical Memoirs of Fellows of the Royal Society, 5:237–252, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Roberts:1964:RRG

[Rob64] John Alexander Fraser Roberts. Reginald Ruggles Gates, 1882– 1962. Biographical Memoirs of Fellows of the Royal Society, 10:83– 106, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1964.0006.

Robertson:1976:JWC

[Rob76] John Monteath Robertson. James Wilfred Cook, 10 December– 21 October 1975. Biographical Memoirs of Fellows of the Royal Society, 22:71–103, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1976.
0003.

Robertson:1977:CHW

[Rob77] Alan Robertson. Conrad Hal Waddington, 8 November 1905– 26 September 1975. Biographical Memoirs of Fellows of the Royal Society, 23:575–622, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0022.

Robertson:1986:GEB

[Rob86] Rutherford Ness Robertson. George Edward Briggs, 25 June 1893–7 February 1985. Biographical Memoirs of Fellows of the Royal Society, 32:35–64, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0002.

Roberts:1992:TNM

[Rob92] Lewis Edward John Roberts. Thomas Nelson Marsham, 10 November 1923–12 October 1989. Biographical Memoirs of Fellows of the Royal Society, 38:229–246, November 1992. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0012.

Robson:1995:REB

[Rob95] J. M. Robson. Robert Edward Bell, 29 November 1918–1 April 1992. Biographical Memoirs of Fellows of the Royal Society, 41:56– 69, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1995.0004.

Roberts:2000:CKC

[Rob00] M. Wyn Roberts. Charles Kemball, C.B.E. 27 March 1923– 4 September 1998. Biographical Memoirs of Fellows of the Royal Society, 46:285–298, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0085.

Robbins:2023:LLI

[Rob23] Trevor W. Robbins FRS. Leslie Lars Iversen. 31 October 1937–30 July 2020. Biographical Memoirs of Fellows of the Royal Society, 75:263–278, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0008.

Rochester:1970:WEC

[Roc70] George Dixon Rochester. William Edward Curtis, 1889–1969. Biographical Memoirs of Fellows of the Royal Society, 16:63–76, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1970.0005.

Rogers:1991:RRA

[Rog91] Claude Ambrose Rogers. Richard Rado, 28 April 1906–23 December 1989. Biographical Memoirs of Fellows of the Royal Society, 37:411–426, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.

Rooke:1996:SKH

[Roo96] Denis Eric Rooke. Sir Kenneth Hutchison, C. B. E., F. Eng., 30 October 1903–28 November 1989. Biographical Memoirs of Fellows of the Royal Society, 42:207–218, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0013.

Rose:1974:RH

[Ros74] John Donald Rose. Ronald Holroyd, 1904–1973. Biographical Memoirs of Fellows of the Royal Society, 20:235–245, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1974.0010.

Rotherham:1972:HK

[Rot72] Leonard Rotherham. Hans Kronberger, 1920–1970. Biographical Memoirs of Fellows of the Royal Society, 18:413–426, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1972.0013.

Rothschild:1999:TRJ

[Rot99] Miriam Rothschild. Tadeus Reichstein. 20 July 1897–1 August 1996. Biographical Memoirs of Fellows of the Royal Society, 45:449– 467, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0030.

Rowlinson:1996:JAB

[Row96] John Shipley Rowlinson. John Adair Barker, 24 March 1925– 27 October 1995. Biographical Memoirs of Fellows of the Royal Society, 42:12–22, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0002.

Rowlinson:2004:KGD

[Row04] John Rowlinson. Kenneth George Denbigh. 30 May 1911– 23 January 2004. Biographical Memoirs of Fellows of the Royal Society, 50:61–71, January 2004. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0006.

Roxburgh:2007:SHB

[Rox07] I. W. Roxburgh. Sir Hermann Bondi KCB. 1 November 1919– 10 September 2005: Elected FRS 1959. Biographical Memoirs of Fellows of the Royal Society, 53:45–61, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0008.

Rang:1996:SWD

[RP96] Humphrey Peter Rang and Walter Laing Macdonald Perry. Sir William Drummond MacDonald Paton, C. B. E., 5 May 1917–17 October 1993. Biographical Memoirs of Fellows of the Royal Society, 42:291–314, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1996.
0019.

Redman:2020:DRC

[RP20] S. J. Redman and R. Porter. David Roderick Curtis. 3 June 1927–11 December 2017. Biographical Memoirs of Fellows of the Royal Society, 69:133–143, December 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0025.

Rastan:2005:RSP

[RR05] Sohaila Rastan and Elizabeth Robertson. Rosa Susan Penelope Beddington. 23 March 1956–18 May 2001. Biographical Memoirs of Fellows of the Royal Society, 51:15–32, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0002.

Robinson:2024:CMD

[RR24] Carol V. Robinson FRS and Colin G. Robinson. Christopher Martin Dobson. 8 October 1949–8 September 2019. Biographical Memoirs of Fellows of the Royal Society, 77:129–143, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Rawson:2020:RCC

[RRS20] Peter F. Rawson, Adrian W. A. Rushton, and Martin I. Simpson. Raymond Charles Casey. 10 October 1917–26 April 2016. Biographical Memoirs of Fellows of the Royal Society, 68:71–86, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0050.

Rees:2007:DGN

[RS07a] D. Rees and R. Y. Sharp. Douglas Geoffrey Northcott. 31 December 1916–8 April 2005. Biographical Memoirs of Fellows of the Royal Society, 53:247–263, January 2007. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0010.

Rowlands:2007:FBJ

[RS07b] David J. Rowlands and John J. Skehel. Fred Brown. 31 January 1925–20 February 2004. Biographical Memoirs of Fellows of the Royal Society, 53:93–108, January 2007. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2007.0016.

Reddoch:2015:WGS

[RS15] Allan H. Reddoch and Willem Siebrand FRSC. William George Schneider OC FRSC. 1 June 1915–18 February 2013. Biographical Memoirs of Fellows of the Royal Society, 61:485–504, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0023.

Robson:2016:JRP

[RSD16] Rob Robson, Barry Smith, and Ray Dixon FRS. John Raymond Postgate FIBiol. 24 June 1922–22 October 2014. Biographical Memoirs of Fellows of the Royal Society, 62:483–504, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0006.

Ramachandran:2021:CGN

[RSG21] Prema Ramachandran, Malini Seshadri, and Sarath Gopalan. Coluthur Gopalan. 28 November 1918–3 October 2019. Biographical Memoirs of Fellows of the Royal Society, 71:229–248, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2021.0022.

Rubbra:1958:AJR

[Rub58] A. A. Rubbra. Arthur John Rowledge, 1876–1957. Biographical Memoirs of Fellows of the Royal Society, 4:215–223, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1958.0019.

Rubbra:1964:AAG

[Rub64] A. A. Rubbra. Alan Arnold Griffith, 1893–1963. Biographical Memoirs of Fellows of the Royal Society, 10:117–136, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1964.0008.

Rubbia:1991:EAS

[Rub91] Carlo Rubbia. Edoardo Amaldi, 5 September 1908–5 December 1989. Biographical Memoirs of Fellows of the Royal Society, 37:1– 31, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1991.0001.

Ruffles:2002:SCM

[Ruf02] Philip Charles Ruffles C.B.E., F.R.S. Stewart Crichton Miller, C.B.E. 2 July 1934–7 August 1999. Biographical Memoirs of Fellows of the Royal Society, 48:323–329, January 2002. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0018.

Russell:1961:RGS

[Rus61] Edward John Russell. Reginald George Stapledon, 1882–1960. Biographical Memoirs of Fellows of the Royal Society, 7:249–270, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0020.

Russell:1968:CFA

[Rus68] Frederick Stratten Russell. Carl Frederick Abel Pantin, 1899–1967. Biographical Memoirs of Fellows of the Royal Society, 14:417–434, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1968.0017.

Rushton:1977:HHM

[Rus77] William Albert Hugh Rushton. Hamilton Hartridge, 7 May 1886–13 January 1976. Biographical Memoirs of Fellows of the Royal Society, 23:192–211, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0008.

Russell:1978:SMM

[Rus78] Frederick Stratten Russell. Sheina Macalister Marshall, 20 April 1896–7 April 1977. Biographical Memoirs of Fellows of the Royal Society, 24:369–389, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1978.
0011.

Russell:2024:SGA

[Rus24] John Russell. Sir Geoffrey Allen. 29 October 1928–15 March 2023. Biographical Memoirs of Fellows of the Royal Society, 77:9–23, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Ruzicka:1972:AS

[Ruz72] Leopold Ruzicka. Arthur Stoll, 1887–1971. Biographical Memoirs of Fellows of the Royal Society, 18:566–593, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1972.0021.

Roberts:2006:HSM

[RW06] Siobhan Roberts and Asia Ivić Weiss. Harold Scott MacDonald Coxeter. 9 February 1907–31 March 2003. Biographical Memoirs of Fellows of the Royal Society, 52:45–66, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0004.

Robinson:2018:PGH

[RW18] John Robinson and Alan H. Welsh FAA. Peter Gavin Hall. 20 November 1951–9 January 2016. Biographical Memoirs of Fellows of the Royal Society, 64:207–229, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Rothstein:1966:LRC

[RWN66] Andrew Rothstein, Errol Ivor White, and C. P. Nuttall. Leslie Reginald Cox, 1897–1965. Biographical Memoirs of Fellows of the Royal Society, 12:107–121, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1966.
0004.

Sleeman:2015:DSJ

[SA15] B. D. Sleeman and I. D. Abrahams. Douglas Samuel Jones MBE. 10 January 1922–26 November 2013. Biographical Memoirs of Fellows of the Royal Society, 61:203–224, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0005.

Sackston:1994:JHC

[Sac94] W. E. Sackston. John Hubert Craigie, 8 December 1887–26 February 1989. Biographical Memoirs of Fellows of the Royal Society, 39:129–145, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Sadler:2023:HAO

[Sad23] Peter J. Sadler FRS. Hugh Allen Oliver Hill. 23 May 1937–30 July 2021. Biographical Memoirs of Fellows of the Royal Society, 75:199– 217, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2023.0002.

Sakmann:2007:BKM

[Sak07] Bert Sakmann. Bernard Katz. 26 March 1911–20 April 2003. Biographical Memoirs of Fellows of the Royal Society, 53:185–202, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2007.0013.

Salisbury:1957:HNR

[Sal57] Edward James Salisbury. Henry Nicholas Ridley, 1855–1956. Biographical Memoirs of Fellows of the Royal Society, 3:141–159, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1957.0010.

Salisbury:1961:WHL

[Sal61] Edward James Salisbury. William Henry Lang, 1874–1960. Biographical Memoirs of Fellows of the Royal Society, 7:146–160, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0012.

Salisbury:1964:CJF

[Sal64] Edward James Salisbury. Carl Johan Fredrik Skottsberg, 1880– 1963. Biographical Memoirs of Fellows of the Royal Society, 10:245– 256, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1964.0015.

Salisbury:1966:RGH

[Sal66] Edward James Salisbury. Ronald George Hatton,1886–1965. Biographical Memoirs of Fellows of the Royal Society, 12:250–258, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0011.

Salt:1978:HEH

[Sal78] George Salt. Howard Everest Hinton, 24 August 1912–2 August 1977. Biographical Memoirs of Fellows of the Royal Society, 24:151– 182, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1978.0006.

Sanders:2000:NKC

[San00] J. H. Sanders. Nicholas Kurti, C.B.E. 14 May 1908–24 November 1998. Biographical Memoirs of Fellows of the Royal Society, 46:299– 315, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0086.

Sandford:2021:RDC

[San21] Graham Sandford. Richard Dickinson Chambers. 16 March 1935–18 April 2019. Biographical Memoirs of Fellows of the Royal Society, 70:107–130, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0036.

Saunders:1961:DRP

[Sau61] O. A. Saunders. David Randall Pye, 1886–1960. Biographical Memoirs of Fellows of the Royal Society, 7:199–205, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1961.0015.

Saunders:1987:DMS

[Sau87] O. A. Saunders. David MacLeish Smith, 9 June 1900–3 August 1986. Biographical Memoirs of Fellows of the Royal Society, 33:603– 617, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1987.0021.

Saunders:1994:ADL

[Sau94] D. S. Saunders. Anthony David Lees, 27 February 1917–3 October 1992. Biographical Memoirs of Fellows of the Royal Society, 40:221– 236, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1994.0036.

Savage:1963:MAC

[Sav63] R. J. G. Savage. Martin Alister Campbell Hinton, 1883–1961. Biographical Memoirs of Fellows of the Royal Society, 9:155–170, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1963.0008.

Sawyer:1992:RCS

[Saw92] John Stanley Sawyer. Reginald Cockcroft Sutcliffe, 16 November 1904–28 May 1991. Biographical Memoirs of Fellows of the Royal Society, 38:345–358, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1992.

Sutcliffe:1977:EGJ

[SB77] Reginald Cockcroft Sutcliffe and A. C. Best. Ernest Gold, 24 July 1881–30 January 1976. Biographical Memoirs of Fellows of the Royal Society, 23:115–131, November 1977. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1977.0006.

Seaborg:2008:MCA

[SB08a] Glenn T. Seaborg and Andrew A. Benson. Melvin Calvin. 8 April 1911–8 January 1997. Biographical Memoirs of Fellows of the Royal Society, 54:59–70, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2007.
0050.

Smith:2008:JWC

[SB08b] George D. W. Smith and Harshad K. D. H. Bhadeshia. John Wyrill Christian. 9 April 1926–27 February 2001. Biographical Memoirs of Fellows of the Royal Society, 54:71–94, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0016.

Sukumar:2022:DMB

[SB22] C. V. Sukumar and A. Bonaccorso. David Maurice Brink. 20 July 1930–8 March 2021. Biographical Memoirs of Fellows of the Royal Society, 73:65–83, December 2022. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0020.

Stoneham:2010:GDJ

[SBG10] Marshall Stoneham and Ian Buckley-Golder. Geoffrey Dearnaley. 22 June 1930–5 May 2009. Biographical Memoirs of Fellows of the Royal Society, 56:41–61, January 2010. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0001. See correction [SBG11].

Stoneham:2011:CGD

[SBG11] Marshall Stoneham FRS and Ian Buckley-Golder. Correction for Geoffrey Dearnaley. 22 June 1930–5 May 2009. Biographical Memoirs of Fellows of the Royal Society, 57:463, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0023. See [SBG10].

Southwood:1999:CSE

[SC99] Richard Southwood and J. R. Clarke. Charles Sutherland Elton. 29 March 1900–1 May 1991. Biographical Memoirs of Fellows of the Royal Society, 45:129–146, January 1999. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0010.

Samson:2023:HJF

[SC23] Leona D. Samson and Will Cairns. Hugh John Forster Cairns. 21 November 1922–12 November 2018. Biographical Memoirs of Fellows of the Royal Society, 74:37–62, April 2023. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0047.

Scazzocchio:2018:JAJ

[Sca18] Claudio Scazzocchio. John Arthur Joseph Pateman. 18 May 1926–18 May 2011. Biographical Memoirs of Fellows of the Royal Society, 65:299–315, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0003.

Schuster:1955:RSC

[Sch55] George Schuster. Richard Stafford Cripps, 1889–1952. Biographical Memoirs of Fellows of the Royal Society, 1:11–32, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0003.

Schonland:1964:MPA

[Sch64] Basil Ferdinand Jamieson Schonland. Maurice Pascal Alers Hankey, First Baron Hankey of the Chart, 1877–1963. Biographical Memoirs of Fellows of the Royal Society, 10:137–146, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1964.0009.

Schmidt:1967:ANR

[Sch67] C. F. Schmidt. Alfred Newton Richards, 1876–1966. Biographical Memoirs of Fellows of the Royal Society, 13:327–342, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1967.0017.

Scott:1956:RCL

[Sco56] Hugh Scott. Robert Cyril Layton Perkins, 1866–1955. Biographical Memoirs of Fellows of the Royal Society, 2:215–236, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1956.0015.

Scott:2014:PGA

[Sco14] S. K. Scott. Peter Gray. 25 August 1926–7 June 2012. Biographical Memoirs of Fellows of the Royal Society, 60:211–227, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2014.0013.

Scriver:1999:NJB

[Scr99] Charles R. Scriver. Norman John Berrill. 28 April 1903–15 October 1996. Biographical Memoirs of Fellows of the Royal Society, 45:19– 34, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0002.

Sutton:2015:JFF

[SD15] Adrian P. Sutton FRS and Olivier Hardouin Duparc FRS. Jacques Friedel. 11 February 1921–27 August 2014. Biographical Memoirs of Fellows of the Royal Society, 61:123–143, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0020.

Sutton:2019:DGP

[SDV19] Adrian P. Sutton FRS, Ralf Drautz, and Vaclav Vitek. David Godfrey Pettifor. 9 March 1945–16 October 2017. Biographical Memoirs of Fellows of the Royal Society, 66:329–353, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0038.

Slater:1965:JH

[SE65] William Kershaw Slater and John Hilton Edwards. John Hammond, 1889–1964. Biographical Memoirs of Fellows of the Royal Society, 11:100–113, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1965.
0006.

Searle:2020:SAS

[Sea20] Roger C. Searle. Sir Anthony Seymour Laughton. 29 April 1927– 27 September 2019. Biographical Memoirs of Fellows of the Royal Society, 69:291–311, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0021.

Sellers:2007:JBB

[Sel07] R. F. Sellers. John Burns Brooksby. 25 December 1914–17 December 1998. Biographical Memoirs of Fellows of the Royal Society, 53: 77–92, January 2007. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2007.0002.

Senn:2019:JAN

[Sen19] Stephen Senn. John Ashworth Nelder. 8 October 1924–7 August 2010. Biographical Memoirs of Fellows of the Royal Society, 67:307– 326, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0013.

Series:1988:RWD

[Ser88] George William Series. Robert William Ditchburn, 14 January 1903–8 April 1987. Biographical Memoirs of Fellows of the Royal Society, 34:63–95, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.
0004.

Serre:1999:AWM

[Ser99] Jean-Pierre Serre. André Weil. 6 May 1906–6 August 1998. Biographical Memoirs of Fellows of the Royal Society, 45:519–529, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1999.0034.

Serre:2009:HPC

[Ser09] Jean-Pierre Serre ForMemRS. Henri Paul Cartan. 8 July 1904–13 August 2008. Biographical Memoirs of Fellows of the Royal Society, 55:37–44, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.
0005.

Seshadri:1962:SSB

[Ses62] Thiruvenkata Rajendra Seshadri. Shanti Swarup Bhatnagar, 1894– 1955. Biographical Memoirs of Fellows of the Royal Society, 8:1– 17, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1962.0001.

Sewell:2015:RHJ

[Sew15] Michael J. Sewell. Rodney Hill. 11 June 1921–2 February 2011. Biographical Memoirs of Fellows of the Royal Society, 61:161–181, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2014.0024.

Sumpter:1991:JMD

[SF91] J. P. Sumpter and Brian Keith Follett. James Munro Dodd, 26 May 1915–15 December 1986. Biographical Memoirs of Fellows of the Royal Society, 37:127–147, November 1991. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1991.0007.

Shortt:1979:SRC

[SG79] Henry Edward Shortt and Percy Cyril Claude Garnham. Samuel Rickard Christophers, 27 November 1873–19 February 1978. Biographical Memoirs of Fellows of the Royal Society, 25:179–207, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1979.0005.

Sajjadi:2018:MSL

[SH18] Shahrdad G. Sajjadi and Julian C. R. Hunt FRS. Michael Selwyn Longuet-Higgins. 8 December 1925–26 February 2016. Biographical Memoirs of Fellows of the Royal Society, 65:249–265, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2017.0031.

Snowden:2021:PJB

[SH21a] Christopher Snowden and Yang Hao. Peter John Bell Clarricoats. 6 April 1932–17 January 2020. Biographical Memoirs of Fellows of the Royal Society, 71:103–122, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0045.

Spedding:2021:CJP

[SH21b] Geoffrey Spedding and Anders Hedenström. Colin James Pennycuick. 11 June 1933–9 December 2019. Biographical Memoirs of Fellows of the Royal Society, 71:451–478, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0023.

Shepherd:2024:DHC

[SH24] John Shepherd FRS and Joe Horwood. David Henry Cushing. 14 March 1920–14 March 2008. Biographical Memoirs of Fellows of the Royal Society, 77:93–106, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Sharp:2015:DRM

[Sha15] R. Y. Sharp. David Rees. 29 May 1918–16 August 2013. Biographical Memoirs of Fellows of the Royal Society, 61:379–401, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0010.

Shaw:2021:RLF

[Sha21] Jeffrey Jon Shaw. Ralph Lainson. 21 February 1927–5 May 2015. Biographical Memoirs of Fellows of the Royal Society, 70:245–262, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2020.0032.

Sheppard:1982:GBB

[She82] Norman Sheppard. Gordon Brims Black McIvor Sutherland, 8 April 1907–27 June 1980. Biographical Memoirs of Fellows of the Royal Society, 28:589–626, November 1982. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0023.

Sheppard:2003:ERA

[She03] Norman Sheppard FRS. Edward Raymond Andrew. 27 June 1921–27 May 2001. Biographical Memoirs of Fellows of the Royal Society, 49:1–14, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2003.
0001.

Sherman:2018:RWG

[She18] S. Murray Sherman. Rainer Walter Guillery. 28 August 1929–7 April 2017. Biographical Memoirs of Fellows of the Royal Society, 64:183–206, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Shotton:1963:WBR

[Sho63] Frederick William Shotton. William Bernard Robinson King, 1889– 1963. Biographical Memoirs of Fellows of the Royal Society, 9:171– 182, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1963.0009.

Shortt:1967:SA

[Sho67] Henry Edward Shortt. Saul Adler, 1895–1966. Biographical Memoirs of Fellows of the Royal Society, 13:1–34, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1967.0001.

Shoenberg:1971:HL

[Sho71] David Shoenberg. Heinz London, 1907–1970. Biographical Memoirs of Fellows of the Royal Society, 17:441–461, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1971.0017.

Shoppee:1972:CKI

[Sho72] Charles William Shoppee. Christopher Kelk Ingold, 1893–1970. Biographical Memoirs of Fellows of the Royal Society, 18:348–411, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0012.

Shoenberg:1983:KAG

[Sho83] David Shoenberg. Kurt Alfred Georg Mendelssohn, 7 January 1906–18 September 1980. Biographical Memoirs of Fellows of the Royal Society, 29:360–398, November 1983. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1983.0015.

Shoenberg:1985:PLK

[Sho85a] David Shoenberg. Piotr Leonidovich Kapitza, 9 July 1894–8 April 1984. Biographical Memoirs of Fellows of the Royal Society, 31:326– 374, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1985.0012.

Short:1985:ECA

[Sho85b] Roger Valentine Short. Emmanuel Ciprian Amoroso, 16 September 1901–30 October 1982. Biographical Memoirs of Fellows of the Royal Society, 31:2–30, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0001.

Siddiqi:2002:GPN

[Sid02] O. Siddiqi F.R.S. Guido Pontecorvo. 29 November 1907– 25 September 1999. Biographical Memoirs of Fellows of the Royal Society, 48:375–390, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2002.
0022.

Silverstein:2003:PGH

[Sil03] Arthur M. Silverstein. Philip George Houthem Gell. 20 October 1914–3 May 2001 Elected FRS 1969. Biographical Memoirs of Fellows of the Royal Society, 49:163–178, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0010.

Simmons:2018:SAF

[Sim18] Professor Robert Malcolm Simmons FRS. Sir Andrew Fielding Huxley OM. 22 November 1917–30 May 2012. Biographical Memoirs of Fellows of the Royal Society, 65:179–215, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0012.

Sinclair:2003:BEJ

[Sin03] Allan M. Sinclair. Barry Edward Johnson. 1 August 1937–5 May 2002. Biographical Memoirs of Fellows of the Royal Society, 49:249– 262, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2003.0014.

Slater:1966:TW

[SK66] William Kershaw Slater and H. G. H. Kearns. Thomas Wallace, 1891–1965. Biographical Memoirs of Fellows of the Royal Society, 12:503–519, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1966.
0025.

Smallman:2013:SAC

[SK13] R. E. Smallman CBE FRS FREng and J. F. Knott OBE FRS FREng. Sir Alan Cottrell FREng. 17 July 1919–15 February 2012. Biographical Memoirs of Fellows of the Royal Society, 59:93–124, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2012.0042.

Skempton:1970:AJS

[Ske70] Alec Westley Skempton. Alfred John Sutton Pippard, 1891–1969. Biographical Memoirs of Fellows of the Royal Society, 16:463–478, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1970.0019.

Skottsberg:1961:NES

[Sko61] Carl Johan Fredrik Skottsberg. Nils Eberhard Svedelius, 1873– 1960. Biographical Memoirs of Fellows of the Royal Society, 7:294– 312, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1961.0023.

Strang:1988:PYJ

[SL88] William John Strang and G. M. Lewis. Pierre Young, 12 June 1926–4 August 1985. Biographical Memoirs of Fellows of the Royal Society, 34:983–1001, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.
0029.

Suckling:1990:FLR

[SL90] Charles Walter Suckling and B. W. Langley. Francis Leslie Rose, 27 June 1909–3 March 1988. Biographical Memoirs of Fellows of the Royal Society, 36:489–524, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1990.0041.

Smith:1994:JLH

[SL94] David Cecil Smith and D. H. Lewis. John Laker Harley, 17 November 1911–13 December 1990. Biographical Memoirs of Fellows of the Royal Society, 39:157–175, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Seah:2018:ADH

[SL18] Martin P. Seah and Colin Lea. Anastasios Demetrios Hondros CMG FRS. 18 February 1930–13 September 2016. Biographical Memoirs of Fellows of the Royal Society, 64:231–248, June 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Sargent:2021:EMB

[SL21] Anneila I. Sargent and Malcolm S. Longair. Eleanor Margaret Burbidge. 12 August 1919–5 April 2020. Biographical Memoirs of Fellows of the Royal Society, 71:11–35, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0017.

Slater:1994:PDM

[Sla94] Edward Charles Slater. Peter Dennis Mitchell, 29 September 1920–10 April 1992. Biographical Memoirs of Fellows of the Royal Society, 40:283–305, November 1994. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.
0040.

Sleigh:1997:JAK

[Sle97] M. A. Sleigh. John Alwyne Kitching, O. B. E. 24 October 1908–1 April 1996. Biographical Memoirs of Fellows of the Royal Society, 43:269–284, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0015.

Siddle:2010:CNH

[SLO10] Kenneth Siddle, J. Paul Luzio, and Stephen O’Rahilly. Charles Nicholas Hales. 25 April 1935–15 September 2005. Biographical Memoirs of Fellows of the Royal Society, 56:105–130, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2009.0019.

Savill:2024:AAT

[SM24] A. Mark Savill and Ivan Marusic. Albert Alan Townsend. 22 January 1917–31 August 2010. Biographical Memoirs of Fellows of the Royal Society, 77:353–373, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Smallman:2010:CSR

[Sma10] R. E. Smallman. Correction for Sir Robert William Kerr Honeycombe. 2 May 1921–14 September 2007. Biographical Memoirs of Fellows of the Royal Society, 56:423, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2010.0006. See [CBE09].

Smith:1955:RNS

[Smi55] Kenneth Manley Smith. Redcliffe Nathan Salaman, 1874–1955. Biographical Memoirs of Fellows of the Royal Society, 1:238–245, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1955.0017.

Smith:1958:RWB

[Smi58] David Macleish Smith. Richard William Bailey, 1885–1957. Biographical Memoirs of Fellows of the Royal Society, 4:15–25, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1958.0002.

Smith:1963:HGC

[Smi63] James Eric Smith. Herbert Graham Cannon, 1897–1963. Biographical Memoirs of Fellows of the Royal Society, 9:55–68, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1963.0003.

Smith:1968:HMF

[Smi68] James Eric Smith. Harold Munro Fox, 1889–1967. Biographical Memoirs of Fellows of the Royal Society, 14:206–222, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0009.

Smith:1969:HHS

[Smi69] Kenneth Manley Smith. Harold Haydon Storey, 1894–1969. Biographical Memoirs of Fellows of the Royal Society, 15:238–246, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1969.0012.

Smith:1973:SFD

[Smi73] David Macleish Smith. Stanley Fabes Dorey, 1891–1972. Biographical Memoirs of Fellows of the Royal Society, 19:304–316, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1973.0012.

Smith:1982:RAS

[Smi82] Stanley Desmond Smith. Robert Allan Smith, 14 May 1909–16 May 1980. Biographical Memoirs of Fellows of the Royal Society, 28:479– 504, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1982.0019.

Smith:1998:SCW

[Smi98] K. C. A. Smith. Sir Charles William Oatley, O. B. E. 14 February 1904–11 March 1996. Biographical Memoirs of Fellows of the Royal Society, 44:331–347, January 1998. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1998.0022.

Smithies:1999:AGO

[Smi99] O. Smithies. Alexander George Ogston. 30 January 1911–29 June 1996. Biographical Memoirs of Fellows of the Royal Society, 45:349– 364, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0024.

Smiles:2005:JRP

[Smi05a] David Smiles. John Robert Philip. 18 January 1927–26 June 1999. Biographical Memoirs of Fellows of the Royal Society, 51:327–345, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2005.0021.

Smith:2005:CEM

[Smi05b] J. David Smith. Colin Eaborn. 15 March 1923–22 February 2004. Biographical Memoirs of Fellows of the Royal Society, 51:101–118, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2005.0007.

Smith:2015:GEP

[Smi15] A. F. M. Smith FRS. George Edward Pelham Box. 10 October 1919–28 March 2013. Biographical Memoirs of Fellows of the Royal Society, 61:23–37, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0015.

Smith:2016:KLJ

[Smi16] Roderick A. Smith FREng. Kenneth Langstreth Johnson. 19 March 1925–21 September 2015. Biographical Memoirs of Fellows of the Royal Society, 62:247–265, January 2016. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2016.0012.

Simberloff:2020:IAH

[SO20] Daniel Simberloff and Otso Ovaskainen. Ilkka Aulis Hanski. 14 February 1953–10 May 2016. Biographical Memoirs of Fellows of the Royal Society, 68:231–250, June 2020. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0033.

Skaer:2022:SHP

[SOD22] Helen Skaer, Michael J. O’Donnell, and Julian A. T. Dow FRSE. Simon Hugh Piper Maddrell. 11 December 1937–10 September 2020. Biographical Memoirs of Fellows of the Royal Society, 73:357–377, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2021.0039.

Sondheimer:1999:SAH

[Son99] E. H. Sondheimer. Sir Alan Herries Wilson. 2 July 1906– 30 September 1995. Biographical Memoirs of Fellows of the Royal Society, 45:547–562, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0036.

Southwood:1987:OWR

[Sou87] Sir Richard Southwood. Owain Westmacott Richards, 31 December 1901–10 November 1984. Biographical Memoirs of Fellows of the Royal Society, 33:537–571, December 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0019.

Southwood:2011:HEC

[Sou11] David Southwood. Harry Elliot CBE. 28 June 1920–5 July 2009. Biographical Memoirs of Fellows of the Royal Society, 57:97–127, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2010.0008.

Soward:2024:PHR

[Sow24] Andrew Michael Soward FRS. Paul Harry Roberts. 13 September 1929–17 November 2022. Biographical Memoirs of Fellows of the Royal Society, 76:365–387, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.

Stacey:1976:ELH

[SP76] Maurice Stacey and Elizabeth Percival. Edmund Langley Hirst, 21 July 1898–29 October 1975. Biographical Memoirs of Fellows of the Royal Society, 22:136–168, November 1976. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0006.

Sutcliffe:1979:PAS

[SP79] Reginald Cockcroft Sutcliffe and Frank Pasquill. Percival Albert Sheppard, 12 May 1907–22 December 1977. Biographical Memoirs of Fellows of the Royal Society, 25:534–553, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0019.

Spalding:2014:JHA

[Spa14] D. Brian Spalding FRS. John Hadji Argyris. 19 August 1913–2 April 2004. Biographical Memoirs of Fellows of the Royal Society, 60:23–37, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2013.
0003.

Spence:1970:OH

[Spe70] Robert Spence. Otto Hahn, 1879–1968. Biographical Memoirs of Fellows of the Royal Society, 16:279–313, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0010.

Spear:1994:PGL

[Spe94] Walter Eric Spear. Peter George Lecomber, 19 February 1941– 9 September 1992. Biographical Memoirs of Fellows of the Royal Society, 39:213–225, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Spiers:1991:WVM

[Spi91] F. W. Spiers. William Valentine Mayneord, 14 February 1902– 10 August 1988. Biographical Memoirs of Fellows of the Royal Society, 37:341–364, November 1991. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1991.
0017.

Spring:1995:JSJ

[Spr95] J. S. Spring. John Sutton, 8 July 1919–6 September 1992. Biographical Memoirs of Fellows of the Royal Society, 41:440–456, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1995.0026.

Stephens:1988:JAP

[SQ88] C. G. Stephens and James Patrick Quirk. James Arthur Prescott, 7 October 1890–6 February 1987. Biographical Memoirs of Fellows of the Royal Society, 34:639–668, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0021.

Schild:1976:HRI

[SR76] Heinz Otto Schild and Francis Leslie Rose. Harry Raymond Ing, 31 July–23 September 1974. Biographical Memoirs of Fellows of the Royal Society, 22:239–255, November 1976. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0010.

Sanders:2017:DHW

[SR17] Jeremy K. M. Sanders CBE FRS and Dame Carol V. Robinson DBE FMedSci FRS. Dudley Howard Williams. 25 May 1937–3 November 2010. Biographical Memoirs of Fellows of the Royal Society, 63:567–583, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2017.
0009.

Spanel:2024:DSN

[ŠR24] Patrik Španěl and Terence H. Risby. David Smith. 26 November 1935–15 February 2023. Biographical Memoirs of Fellows of the Royal Society, 77:335–351, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Self:2006:GPL

[SS06] S. Self and R. S. J. Sparks. George Patrick Leonard Walker. 2 March 1926–17 January 2005. Biographical Memoirs of Fellows of the Royal Society, 52:423–436, January 2006. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0029.

Stone:2023:ADM

[SSEF23] A. J. Stone, M. Stewart, D. S. Eisenberg, and R. A. Crowther FRS. Andrew David McLachlan. 25 January 1935–8 July 2022. Biographical Memoirs of Fellows of the Royal Society, 75:345–360, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2023.0014.

Sridharan:2023:REG

[SSGP23] Visuvanathar Sridharan, Paul J. Stevenson, H. Q. Nimal Gunaratne, and A. Prasanna de Silva. Ronald Ernest Grigg. 1 September 1935–10 January 2021. Biographical Memoirs of Fellows of the Royal Society, 75:175–197, December 2023. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0052.

Skehel:1999:HGP

[ST99] John Skehel and D. A. J. Tyrrell. Helio Gelli Pereira. 23 September 1918–16 August 1994. Biographical Memoirs of Fellows of the Royal Society, 45:379–396, January 1999. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0026.

Sullivan:2019:DCB

[ST19] Alice C. Sullivan FRSC and John Meurig Thomas FRS. Donald Charlton Bradley. 7 November 1924–20 December 2014. Biographical Memoirs of Fellows of the Royal Society, 67:119–141, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2019.0010.

Stanier:1958:HLG

[Sta58] William Arthur Stanier. Henry Lewis Guy, 1887–1956. Biographical Memoirs of Fellows of the Royal Society, 4:98–101, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0009.

Stavraky:1969:FRM

[Sta69] G. W. Stavraky. Frederick Robert Miller, 1881–1967. Biographical Memoirs of Fellows of the Royal Society, 15:172–184, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1969.0009.

Stacey:1979:JKN

[Sta79] Maurice Stacey. John Kenyon Netherton Jones, 28 January 1912–13 April 1977. Biographical Memoirs of Fellows of the Royal Society, 25:365–389, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1979.
0013.

Stafford:1986:JBA

[Sta86] Godfrey Harry Stafford. John Bertram Adams, 24 May 1920–3 March 1984. Biographical Memoirs of Fellows of the Royal Society, 32:1–34, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0001.

Stacey:1994:MHA

[Sta94] Maurice Stacey. Michael Heidelberger, 29 April 1888–25 June 1991. Biographical Memoirs of Fellows of the Royal Society, 39:177–197, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Steward:1967:PM

[Ste67] Frederick Campion Steward. Panchanan Maheshwari, 1904–1966. Biographical Memoirs of Fellows of the Royal Society, 13:256–266, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1967.0013.

Stevens:1984:KJF

[Ste84] Thomas Stevens Stevens, F.R.S. Karl Johann Freudenberg, 29 January 1886–3 April 1983. Biographical Memoirs of Fellows of the Royal Society, 30:168–189, November 1984. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1984.0006.

Stein:1986:JFD

[Ste86] W. D. Stein. James Frederic Danielli, 13 November 1911–22 April 1984. Biographical Memoirs of Fellows of the Royal Society, 32:115– 135, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1986.0005.

Stephenson:2020:EBJ

[Ste20] F. Anne Stephenson. Eric A. Barnard. 2 July 1927–23 May 2018. Biographical Memoirs of Fellows of the Royal Society, 69:37–61, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2020.0017.

Stiles:1965:DT

[Sti65] Walter Stiles. David Thoday, 1883–1964. Biographical Memoirs of Fellows of the Royal Society, 11:176–185, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1965.0012.

Stirling:2005:GLM

[Sti05] C. J. M. Stirling. Gordon Lowe. 31 May 1933–6 August 2003. Biographical Memoirs of Fellows of the Royal Society, 51:237–252, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2005.0015.

Stormer:1976:OHJ

[Stø76] L. Størmer. Olaf Holtedahl, 24 June 1885–26 August 1975. Biographical Memoirs of Fellows of the Royal Society, 22:193–205, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1976.0008.

Stoicheff:1994:HLW

[Sto94] Boris Peter Stoicheff. Harry Lambert Welsh, 23 March 1910–23 July 1984. Biographical Memoirs of Fellows of the Royal Society, 39:433–450, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Stoicheff:2003:GHP

[Sto03] Boris P. Stoicheff OC FRS. Gerhard Herzberg PC CC. 25 December 1904–3 March 1999. Biographical Memoirs of Fellows of the Royal Society, 49:179–195, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0011.

Stratton:1956:WSA

[Str56] Frederick John Marrian Stratton. Walter Sydney Adams, 1876– 1956. Biographical Memoirs of Fellows of the Royal Society, 2:1– 18, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1956.0001.

Stratton:1957:HNR

[Str57a] Frederick John Marrian Stratton. Henry Norris Russell, 1877–1957. Biographical Memoirs of Fellows of the Royal Society, 3:173–191, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1957.0012.

Stratton:1957:JE

[Str57b] Frederick John Marrian Stratton. John Evershed, 1864–1956. Biographical Memoirs of Fellows of the Royal Society, 3:40–51, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1957.0004.

Strang:1996:SAE

[Str96] William John Strang. Sir Archibald Edward Russell, C. B. E., 30 May 1904–29 May 1995. Biographical Memoirs of Fellows of the Royal Society, 42:381–384, November 1996. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0023.

Stubblefield:1965:EBB

[Stu65] Cyril James Stubblefield. Edward Battersby Bailey, 1881–1965. Biographical Memoirs of Fellows of the Royal Society, 11:1–21, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1965.0001.

Stubblefield:1966:RGS

[Stu66] Cyril James Stubblefield. Robert George Spencer Hudson, 1895– 1965. Biographical Memoirs of Fellows of the Royal Society, 12:320– 333, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1966.0015.

Stubblefield:1970:DNW

[Stu70] Cyril James Stubblefield. Darashaw Nosherwan Wadia, 1883–1969. Biographical Memoirs of Fellows of the Royal Society, 16:543–562, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1970.0023.

Stubblefield:1974:WSB

[Stu74] Cyril James Stubblefield. William Sawney Bisat, 1886–1973. Biographical Memoirs of Fellows of the Royal Society, 20:27–40, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1974.0002.

Stubblefield:1975:OMB

[Stu75] Cyril James Stubblefield. Oliver Meredith Boone Bulman, 20 May 1902–18 February 1974. Biographical Memoirs of Fellows of the Royal Society, 21:175–195, November 1975. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1975.0004.

Stubblefield:1977:GHM

[Stu77] Cyril James Stubblefield. George Hoole Mitchell, 31 December 1902–11 March 1976. Biographical Memoirs of Fellows of the Royal Society, 23:367–383, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.

Stubblefield:1985:EIW

[Stu85] Cyril James Stubblefield. Errol Ivor White, 30 June 1901– 11 January 1985. Biographical Memoirs of Fellows of the Royal Society, 31:633–651, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1985.
0022.

Stuart:1986:LRJ

[Stu86] John Trevor Stuart. Louis Rosenhead, 1 January 1906– 10 November 1984. Biographical Memoirs of Fellows of the Royal Society, 32:405–420, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0013.

Stuart:2009:LHO

[Stu09] J. T. Stuart FRS. Leslie Howarth OBE. 23 May 1911–22 September 2001. Biographical Memoirs of Fellows of the Royal Society, 55:107– 119, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2009.0013.

Stuart:2012:E

[Stu12] Trevor Stuart FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 58:1–2, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2012.
0040.

Stuart:2013:DWM

[Stu13a] J. T. Stuart FRS. Derek William Moore. 19 April 1931–15 July 2008. Biographical Memoirs of Fellows of the Royal Society, 59:241– 259, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2013.0014.

Stuart:2013:E

[Stu13b] Trevor Stuart FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 59:1–3, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2013.

0019. See correction [Ano13].

Stuart:2014:E

[Stu14] Trevor Stuart FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 60:1–4, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2014.
0022.

Stuart:2015:E

[Stu15] Trevor Stuart FRS. Editorial. Biographical Memoirs of Fellows of the Royal Society, 61:1–4, January 2015. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2015.
0027.

Sutton:1965:DB

[Sut65] Oliver Graham Sutton. David Brunt, 1886–1965. Biographical Memoirs of Fellows of the Royal Society, 11:41–52, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1965.0003.

Sutton:1970:HHR

[Sut70] John Sutton. Herbert Harold Read, 1889–1970. Biographical Memoirs of Fellows of the Royal Society, 16:479–497, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1970.0020.

Sutton:1980:WQK

[Sut80] John Sutton. William Quarrier Kennedy, 30 November 1903– 13 March 1979. Biographical Memoirs of Fellows of the Royal Society, 26:274–303, November 1980. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1980.
0008.

Schneider:1991:KFW

[SV91] William George Schneider and Z. Valenta. Karel František Weisner, 25 November 1919–28 November 1986. Biographical Memoirs of Fellows of the Royal Society, 37:461–490, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1991.0024.

Seitz:2000:EPW

[SVW00] Frederick Seitz, Erich Vogt, and Alvin M. Weinberg. Eugene Paul Wigner. 17 November 1902–1 January 1995. Biographical Memoirs of Fellows of the Royal Society, 46:577–592, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL http://rsbm.royalsocietypublishing.org/
content/46/577.abstract;   http://www.jstor.org/stable/
770417;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0102.

Synge:1990:ACC

[SW90] Richard Laurence Millington Synge and E. F. Williams. Albert Charles Chibnall, 28 January 1894–10 January 1988. Biographical Memoirs of Fellows of the Royal Society, 35:55–96, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Simpson:2021:JJM

[SW21a] Thomas J. Simpson FRS and Christine L. Willis. John (Jake) Macmillan. 13 September 1924–12 May 2014. Biographical Memoirs of Fellows of the Royal Society, 70:297–312, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0029.

Smith:2021:LMA

[SW21b] Robert Connon Smith and Nigel Weiss FRS. Leon Mestel. 5 August 1927–15 September 2017. Biographical Memoirs of Fellows of the Royal Society, 70:335–358, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0030.

Swaminathan:1996:BPP

[Swa96] Monkombu Sambasivan Swaminathan. Benjamin Peary Pal, 26 May 1906–14 September 1989. Biographical Memoirs of Fellows of the Royal Society, 42:267–274, November 1996. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0017.

Sykes:1973:NPA

[Syk73] Charles Sykes. Norman Percy Allen, 1903–1972. Biographical Memoirs of Fellows of the Royal Society, 19:1–18, December 1973.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1973.0001.

Synge:1967:HRM

[Syn67] Richard Laurence Millington Synge. Hedley Ralph Marston, 1900– 1965. Biographical Memoirs of Fellows of the Royal Society, 13:267– 293, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1967.0014.

Synge:1976:EVO

[Syn76] John Lighton Synge. Eamon De Valera, 14 October 1882–29 August 1975. Biographical Memoirs of Fellows of the Royal Society, 22:634– 653, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1976.0022.

Tabor:1969:FPB

[Tab69] David Tabor. Frank Philip Bowden, 1903–1968. Biographical Memoirs of Fellows of the Royal Society, 15:1–38, November 1969. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1969.0001.

Takhtajan:2022:LDF

[Tak22] Leon A. Takhtajan. Ludwig Dmitrievich Faddeev. 23 March 1934–26 February 2017. Biographical Memoirs of Fellows of the Royal Society, 73:251–275, December 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0003.

Tata:1990:RCO

[Tat90] Jamshed Rustom Tata. Robert Courrier, 6 October 1895–14 March 1986. Biographical Memoirs of Fellows of the Royal Society, 36:101– 123, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1990.0026.

Tata:1994:RVP

[Tat94] Jamshed Rustom Tata. Rosalind Venetia Pitt-Rivers, 4 March 1907–14 January 1990. Biographical Memoirs of Fellows of the Royal Society, 39:325–348, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Taylor:1958:IL

[Tay58] Hugh Stott Taylor. Irving Langmuir, 1881–1957. Biographical Memoirs of Fellows of the Royal Society, 4:167–184, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0015.

Taylor:1962:GTW

[Tay62] Geoffrey Ingram Taylor. Gilbert Thomas Walker, 1868–1958. Biographical Memoirs of Fellows of the Royal Society, 8:166–174, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1962.0013.

Taylor:1964:SWW

[Tay64] James Haward Taylor. Sidney William Wooldridge, 1900–1963. Biographical Memoirs of Fellows of the Royal Society, 10:371–387, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1964.0021.

Tayler:1996:SCO

[Tay96] Roger John Tayler. Subrahmanyan Chandrasekhar, 19 October 1910–21 August 1995. Biographical Memoirs of Fellows of the Royal Society, 42:80–94, November 1996. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1996.0006.

Taylor:2017:DJB

[Tay17] J. Roy Taylor. Daniel Joseph Bradley. 18 January 1928–7 February 2010. Biographical Memoirs of Fellows of the Royal Society, 63:23– 54, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2017.0012.

Thomson:2013:PJS

[TB13] Andrew J. Thomson OBE FRS and Laurence D. Barron FRS FRSE. Philip John Stephens. 9 October 1940–31 July 2012. Biographical Memoirs of Fellows of the Royal Society, 59:359–382, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2013.0013.

Todd:1976:RRS

[TC76] Alexander Robertus Todd and John Warcup Cornforth. Robert Robinson, 13 September 1886–8 February 1975. Biographical Memoirs of Fellows of the Royal Society, 22:414–527, November 1976. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1976.0018.

Todd:1981:RBW

[TC81] Alexander Robertus Todd and John Warcup Cornforth. Robert Burns Woodward, 10 April 1917–8 July 1979. Biographical Memoirs of Fellows of the Royal Society, 27:628–695, November 1981. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0025.

Thompson:1987:FDD

[TC87] Robert Henry Stewart Thompson and P. N. Campbell. Frank Dickens, 15 December 1899–25 June 1986. Biographical Memoirs of Fellows of the Royal Society, 33:187–210, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0007.

Turner:2018:NSM

[TC18] James J. Turner FRS and Michael A. Chesters. Norman Sheppard. 16 May 1921–10 April 2015. Biographical Memoirs of Fellows of the Royal Society, 65:357–380, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2017.0043.

Teich:2003:DMM

[Tei03] Mikuláš Teich. Dorothy Mary Moyle Needham. 22 September 1896–22 December 1987. Biographical Memoirs of Fellows of the Royal Society, 49:351–365, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0020.

Temple:1956:ETW

[Tem56] George Frederick James Temple. Edmund Taylor Whittaker, 1873– 1956. Biographical Memoirs of Fellows of the Royal Society, 2:299– 325, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1956.0021.

Temple:1965:LB

[Tem65] George Frederick James Temple. Leonard Bairstow, 1880–1963. Biographical Memoirs of Fellows of the Royal Society, 11:22–40, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1965.0002.

Terzian:2010:EES

[Ter10] Yervant Terzian. Edwin Ernest Salpeter. 3 December 1924– 26 November 2008. Biographical Memoirs of Fellows of the Royal Society, 56:391–399, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2010.
0005.

Temple:1967:WW

[TF67] George Frederick James Temple and H. T. Flint. William Wilson, 1875–1965. Biographical Memoirs of Fellows of the Royal Society, 13:386–391, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1967.
0020.

Tompkins:1971:EAG

[TG71] Frederick Clifford Tompkins and Charles Frederick Goodeve. Edward Armand Guggenheim, 1901–1970. Biographical Memoirs of Fellows of the Royal Society, 17:303–326, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1971.0012.

Trentham:2023:HFG

[TG23] David R. Trentham FRS and Michael A. Geeves. Herbert (Freddie) Gutfreund. 21 October 1921–21 March 2021. Biographical Memoirs of Fellows of the Royal Society, 74:211–235, April 2023.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0024.

Thomson:1971:WSF

[TH71] George Paget Thomson and A. A. Hall. William Scott Farren, 1892–1970. Biographical Memoirs of Fellows of the Royal Society, 17:215–241, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1971.
0009.

Talling:2016:JWG

[TH16] J. F. Talling FRS and S. I. Heaney. John Walter Guerrier Lund CBE. 27 November 1912–21 March 2015. Biographical Memoirs of Fellows of the Royal Society, 62:345–358, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2015.0025.

Thimann:1984:KAM

[Thi84] Kenneth Vivian Thimann. Kurt Albin Mothes, 3 November 1900–12 February 1983. Biographical Memoirs of Fellows of the Royal Society, 30:515–543, November 1984. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1984.0017.

Thomson:1955:GFC

[Tho55] George Paget Thomson. George Frederick Charles Searle, 1864– 1954. Biographical Memoirs of Fellows of the Royal Society, 1:246– 252, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1955.0018.

Thomson:1956:AOR

[Tho56] George Paget Thomson. Alexander Oliver Rankine, 1881–1956. Biographical Memoirs of Fellows of the Royal Society, 2:248–255, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1956.0017.

Thomson:1958:FAL

[Tho58] George Paget Thomson. Frederick Alexander Lindemann, Viscount Cherwell, 1886–1957. Biographical Memoirs of Fellows of the Royal Society, 4:45–71, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1958.
0005.

Thomas:1960:AA

[Tho60a] Hugh Hamshaw Thomas. Agnes Arber, 1879–1960. Biographical Memoirs of Fellows of the Royal Society, 6:1–11, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1960.0021.

Thompson:1960:GAK

[Tho60b] William Robin Thompson. Guy Anstruther Knox Marshall, 1871– 1959. Biographical Memoirs of Fellows of the Royal Society, 6:169– 181, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1960.0030.

Thompson:1961:HS

[Tho61] William Robin Thompson. Hugh Scott, 1885–1960. Biographical Memoirs of Fellows of the Royal Society, 7:229–242, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1961.0018.

Thomson:1963:CGD

[Tho63] George Paget Thomson. Charles Galton Darwin, 1887–1962. Biographical Memoirs of Fellows of the Royal Society, 9:69–85, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1963.0004.

Thomson:1965:HAW

[Tho65] George Paget Thomson. Harold Albert Wilson, 1874–1964. Biographical Memoirs of Fellows of the Royal Society, 11:186–201, November 1965. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1965.0013.

Thornton:1966:EJR

[Tho66] Henry Gerard Thornton. Edward John Russell, 1872–1965. Biographical Memoirs of Fellows of the Royal Society, 12:456–477, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0022.

Thompson:1973:CNH

[Tho73a] Harold Warris Thompson. Cyril Norman Hinshelwood, 1897–1967. Biographical Memoirs of Fellows of the Royal Society, 19:375–431, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0015.

Thorpe:1973:WRT

[Tho73b] William Homan Thorpe. William Robin Thompson, 1887–1972. Biographical Memoirs of Fellows of the Royal Society, 19:655–678, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0024.

Thorpe:1974:DLL

[Tho74] William Homan Thorpe. David Lambert Lack, 1910–1973. Biographical Memoirs of Fellows of the Royal Society, 20:271–293, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1974.0012.

Thorpe:1983:KFN

[Tho83] William Homan Thorpe. Karl Von Frisch, 20 November 1886–12 June 1982. Biographical Memoirs of Fellows of the Royal Society, 29:197–200, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1983.
0008.

Thompson:1991:RTG

[Tho91] Robert Henry Stewart Thompson. Ronald Thomson Grant, 5 November 1892–7 November 1989. Biographical Memoirs of Fellows of the Royal Society, 37:245–262, November 1991. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1991.0012.

Thomas:1998:SGB

[Tho98] L. Thomas. Sir Granville Beynon, C. B. E.. 24 May 1914–11 March 1996. Biographical Memoirs of Fellows of the Royal Society, 44:53– 62, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1998.0004.

Thompson:2013:AHC

[Tho13] J. Michael T. Thompson FRS. Amos Henry Chilver FREng, Baron Chilver of Cranfield. 30 October 1926–8 July 2012. Biographical Memoirs of Fellows of the Royal Society, 59:73–91, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2013.0010.

Thomas:2019:PNT

[Tho19] Hywel R. Thomas CBE FRS FREng FLSW MAE. Peter Neil Temple Wells CBE. 19 May 1936–22 April 2017. Biographical Memoirs of Fellows of the Royal Society, 66:463–477, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0022.

Thomas:2020:AHZ

[Tho20] Sir John Meurig Thomas FRS, (Hon) FRSE, (Hon) FREng. Ahmed Hassan Zewail. 26 February 1946–2 August 2016. Biographical Memoirs of Fellows of the Royal Society, 68:431–453, June 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0040.

Thorpe:2024:SRJ

[Tho24] Michael F. Thorpe. Sir Roger James Elliott. 8 December 1928–16 April 2018. Biographical Memoirs of Fellows of the Royal Society, 77:145–162, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Tilley:1957:NLB

[Til57] Cecil Edgar Tilley. Norman Levi Bowen, 1887–1956. Biographical Memoirs of Fellows of the Royal Society, 3:6–22, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1957.0002.

Tilley:1961:LJS

[Til61] Cecil Edgar Tilley. Leonard James Spencer, 1870–1959. Biographical Memoirs of Fellows of the Royal Society, 7:243–248, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1961.0019.

Tiselius:1960:KUL

[Tis60] Arne Wilhelm Kaurin Tiselius. Kaj Ulrik Linderstrøm-Lang, 1896– 1959. Biographical Memoirs of Fellows of the Royal Society, 6:157– 168, November 1960. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1960.0029.

Titchmarsh:1958:GBJ

[Tit58] Edward Charles Titchmarsh. George Barker Jeffery, 1891–1957. Biographical Memoirs of Fellows of the Royal Society, 4:128–137, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1958.0012.

Terentjev:2024:MWJ

[TJ24] Eugene Terentjev and Lisa Jardine-Wright OBE. Mark Warner. 26 January 1952–24 December 2021. Biographical Memoirs of Fellows of the Royal Society, 77:417–434, December 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Thomas:1974:SBG

[TK74] H. H. B. M. Thomas and Dietrich Küchemann. Sidney Barrington Gates, 1893–1973. Biographical Memoirs of Fellows of the Royal Society, 20:181–212, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1974.
0007.

Tyrrell:1990:AWD

[TM90] David Arthur John Tyrrell and K. McCarthy. Allan Watt Downie, 5 September 1901–26 January 1988. Biographical Memoirs of Fellows of the Royal Society, 35:97–112, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Thompson:1983:RAP

[TO83] Robert Henry Stewart Thompson and Alexander George Ogston. Rudolph Albert Peters, 13 April 1889–29 January 1982. Biographical Memoirs of Fellows of the Royal Society, 29:494–523, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1983.0018.

Turner:2012:OMP

[TO12] J. Stewart Turner FRS and Peter L. Olson. Owen Martin Phillips. 30 December 1930–13 October 2010. Biographical Memoirs of Fellows of the Royal Society, 58:229–244, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2012.0028.

Todd:1958:JHG

[Tod58] John Arthur Todd. John Hilton Grace, 1873–1958. Biographical Memoirs of Fellows of the Royal Society, 4:92–97, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0008.

Todd:1973:RAJ

[Tod73] Alexander Robertus Todd. Robert Arthur James Gascoyne-Cecil, 5th Marquess of Salisbury, 1893–1972. Biographical Memoirs of Fellows of the Royal Society, 19:621–627, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1973.0022.

Todd:1975:JKA

[Tod75] Alexander Robertus Todd. James Kenner, 13 April 1885–30 June 1974. Biographical Memoirs of Fellows of the Royal Society, 21:389– 405, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1975.0011.

Todd:1979:GWK

[Tod79] Alexander Robertus Todd. George Wallace Kenner, 16 November 1922–26 June 1978. Biographical Memoirs of Fellows of the Royal Society, 25:390–420, November 1979. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1979.
0014.

Todd:1988:FBF

[Tod88] Alexander Robertus Todd. Franz Bergel, 13 February 1900– 1 January 1987. Biographical Memoirs of Fellows of the Royal Society, 34:1–19, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1988.
0001.

Tolansky:1967:FZ

[Tol67] Samuel Tolansky. Frits Zernike, 1888–1966. Biographical Memoirs of Fellows of the Royal Society, 13:392–402, November 1967. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1967.0021.

Thomas:1975:CCI

[TP75] A. Rylands Thomas and Thomas Angus Lyall Paton. Claude Cavendish Inglis, 3 March 1883–29 August 1974. Biographical Memoirs of Fellows of the Royal Society, 21:366–388, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1975.0010.

Trentham:2011:SVP

[Tre11] David R. Trentham FRS. Samuel Victor Perry. 16 July 1918– 17 December 2009. Biographical Memoirs of Fellows of the Royal Society, 57:327–347, January 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2011.
0009.

Tickle:2024:LWO

[TS24] Cheryll Tickle FRS and Jim Smith FRS. Lewis Wolpert. 19 October 1929–28 January 2021. Biographical Memoirs of Fellows of the Royal Society, 76:477–499, June 2024. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0048.

Turnbull:1955:ARR

[Tur55a] Herbert Westren Turnbull. Archibald Read Richardson, 1881–1954. Biographical Memoirs of Fellows of the Royal Society, 1:223–237, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1955.0016.

Turner:1955:JTH

[Tur55b] Eustace Ebenezer Turner. John Theodore Hewitt, 1868–1954. Biographical Memoirs of Fellows of the Royal Society, 1:79–99, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1955.0007.

Turrill:1958:JCW

[Tur58] William Bertram Turrill. John Christopher Willis 1868–1958. Biographical Memoirs of Fellows of the Royal Society, 4:353–359, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1958.0027.

Turner:1962:JK

[Tur62] Eustace Ebenezer Turner. Joseph Kenyon, 1885–1961. Biographical Memoirs of Fellows of the Royal Society, 8:49–66, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1962.0005.

Turner:2023:PJL

[Tur23] Michael J. Turner. Peter John Lawrenson. 12 March 1933– 27 October 2017. Biographical Memoirs of Fellows of the Royal Society, 75:299–325, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0016.

Townsend:2021:JLH

[TW21] Colin R. Townsend and Andrew R. Watkinson. John Lander Harper. 27 May 1925–22 March 2009. Biographical Memoirs of Fellows of the Royal Society, 71:271–292, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0015.

Taylor:2022:JCP

[TW22] J. C. Taylor and D. A. Wilkinson. John Charlton Polkinghorne KBE. 16 October 1930–9 March 2021. Biographical Memoirs of Fellows of the Royal Society, 72:293–309, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0044.

Tyrrell:1987:JFE

[Tyr87] David Arthur John Tyrrell. John Franklin Enders, 10 February 1897–8 September 1985. Biographical Memoirs of Fellows of the Royal Society, 33:211–233, December 1987. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1987.0008.

Tyrrell:1991:CHA

[Tyr91] David Arthur John Tyrrell. Christopher Howard Andrewes, 7 June 1896–31 December 1987. Biographical Memoirs of Fellows of the Royal Society, 37:33–54, November 1991. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1991.0002.

Ubbelohde:1981:DMN

[Ubb81] Alfred Rene Jean Paul Ubbelohde. Dudley Maurice Newitt, 28 April 1894–14 March 1980. Biographical Memoirs of Fellows of the Royal Society, 27:365–378, November 1981. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0015.

Unsworth:2014:JLM

[Uns14] Michael H. Unsworth. John Lennox Monteith. 3 September 1929–20 July 2012. Biographical Memoirs of Fellows of the Royal Society, 60:299–329, January 2014. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2014.
0005.

vanDriel:2019:BPS

[van19] Henry M. van Driel FRSC. Boris Peter Stoicheff. 1 June 1924–15 April 2010. Biographical Memoirs of Fellows of the Royal Society, 66:403–422, June 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0026.

Vaughan:1987:HBF

[Vau87] Janet Maria Vaughan. Honor Bridget Fell, 22 May 1900–22 April 1986. Biographical Memoirs of Fellows of the Royal Society, 33:235– 259, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1987.0009.

vanCampen:1999:WTK

[vC99] D. H. van Campen. Warner Tjardus Koiter. 16 June 1914– 2 September 1997. Biographical Memoirs of Fellows of the Royal Society, 45:269–273, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0018.

Vinson:2018:JFT

[VC18] Gavin P. Vinson and John P. Coghlan. James Francis Tait. 1 December 1925–2 February 2014. Biographical Memoirs of Fellows of the Royal Society, 65:381–404, December 2018. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2018.0015.

vonEngel:1957:JSE

[vE57] A. von Engel. John Sealy Edward Townsend, 1868–1957. Biographical Memoirs of Fellows of the Royal Society, 3:256–272, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1957.0018.

vanEmden:2006:DML

[vEG06] Helmut F. van Emden and John Gurdon. Dame Miriam Louisa Rothschild. 5 August 1908–20 January 2005. Biographical Memoirs of Fellows of the Royal Society, 52:315–330, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0022.

Velupillai:2019:KJA

[Vel19] K. Vela Velupillai. Kenneth Joseph Arrow. 23 August 1921– 21 February 2017. Biographical Memoirs of Fellows of the Royal Society, 67:9–28, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0002.

Vincent:1998:SGM

[Vin98] E. A. Vincent. Sir George Malcolm Brown. 5 October 1925– 27 March 1997. Biographical Memoirs of Fellows of the Royal Society, 44:65–76, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0005.

Vincent:2009:RHO

[Vin09] Brian Vincent. Ronald Harry Ottewill OBE FRS. 8 February 1927–4 June 2008. Biographical Memoirs of Fellows of the Royal Society, 55:157–184, January 2009. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.
0010.

Vincent:2019:JND

[Vin19] Angela Vincent. John Newsom-Davis. 18 October 1932–24 August 2007. Biographical Memoirs of Fellows of the Royal Society, 67:327– 355, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2019.0028.

Vincent:2023:RPE

[Vin23] Angela Vincent FRS. Roger Philip Ekins. 22 September 1926–26 July 2016. Biographical Memoirs of Fellows of the Royal Society, 75:93–107, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0020.

Vijayan:2005:GNR

[VJ05] M. Vijayan and L. N. Johnson. Gopalasamudram Narayana Ramachandran. 8 October 1922–7 April 2001. Biographical Memoirs of Fellows of the Royal Society, 51:367–377, January 2005.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0024.

Varlamov:2024:AAA

[VL24] Andrey Varlamov and Peter Littlewood FRS. Alexei Alekseevich Abrikosov. 25 June 1928–29 March 2017. Biographical Memoirs of Fellows of the Royal Society, 76:9–26, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0030.

Vogt:1972:GWH

[Vog72] Marthe Louise Vogt. Geoffrey Wingfield Harris, 1913–1971. Biographical Memoirs of Fellows of the Royal Society, 18:309–329, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0010.

Vogel:2008:KSN

[Vog08] Steven Vogel. Knut Schmidt-Nielsen. 24 September 1915– 25 January 2007. Biographical Memoirs of Fellows of the Royal Society, 54:319–331, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2008.
0010.

Voneida:1997:RWS

[Von97] T. J. Voneida. Roger Wolcott Sperry. 20 August 1913–17 April 1994. Biographical Memoirs of Fellows of the Royal Society, 43:463– 470, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1997.0025.

Vinson:1988:JGP

[VT88] G. P. Vinson and James Francis Tait. John Guest Phillips, 13 June 1933–14 March 1987. Biographical Memoirs of Fellows of the Royal Society, 34:609–637, December 1988. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1988.0020.

Waterlow:1994:KLB

[WA94] John Conrad Waterlow and D. G. Armstrong. Kenneth Lyon Blaxter, 19 June 1919–18 April 1991. Biographical Memoirs of Fellows of the Royal Society, 39:35–58, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Waddington:1997:SRO

[Wad97] David Waddington. Sir Richard Oswald Chandler Norman, K. B. E. 27 April 1932–6 June 1993. Biographical Memoirs of Fellows of the Royal Society, 43:335–347, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1997.0019.

Wahle:2023:GAG

[Wah23] Klaus W. J. Wahle. George Alan Garton. 4 June 1922–13 May 2010. Biographical Memoirs of Fellows of the Royal Society, 75:109–128, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2023.0015.

Wain:1997:WRB

[Wai97] R. L. Wain. William Robert Boon. 20 March 1911–28 October 1994. Biographical Memoirs of Fellows of the Royal Society, 43:135–141, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1997.0008.

Walshe:1961:GJ

[Wal61] Francis Martin Rouse Walshe. Geoffrey Jefferson, 1886–1961. Biographical Memoirs of Fellows of the Royal Society, 7:127–135, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0010.

Walshe:1966:GMH

[Wal66] Francis Martin Rouse Walshe. Gordon Morgan Holmes, 1876–1965. Biographical Memoirs of Fellows of the Royal Society, 12:311–319, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0014.

Walsby:2006:GEF

[Wal06] Anthony E. Walsby. Gordon Elliott Fogg CBE. 26 April 1919– 30 January 2005. Biographical Memoirs of Fellows of the Royal Society, 52:97–116, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0008.

Waldegrave:2008:GPJ

[Wal08] William Waldegrave. George Patrick John Rushworth Jellicoe, 2nd Earl Jellicoe. 4 April 1918–22 February 2007. Biographical Memoirs of Fellows of the Royal Society, 54:169–174, January 2008. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2008.0004.

Warner:2017:SSE

[War17] Mark Warner FRS. Sir Sam Edwards. 1 February 1928–7 July 2015. Biographical Memoirs of Fellows of the Royal Society, 63:243–271, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2016.0028.

Wassle:2002:BBB

[Wäs02] Heinz Wässle. Brian Blundell Boycott. 10 December 1924–22 April 2000. Biographical Memoirs of Fellows of the Royal Society, 48: 51–68, January 2002. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2002.0004.

Watson:1955:JPH

[Wat55] David Meredith Seares Watson. James Peter Hill, 1873–1954. Biographical Memoirs of Fellows of the Royal Society, 1:101–117, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1955.0008.

Watson:1956:HGA

[Wat56] David Meredith Seares Watson. Henry George Albert Hickling, 1883–1954. Biographical Memoirs of Fellows of the Royal Society, 2:149–156, November 1956. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1956.
0011.

Watson:2006:RSP

[Wat06] Christopher Watson. Rendel Sebastian Pease. 2 November 1922–17 October 2004. Biographical Memoirs of Fellows of the Royal Society, 52:285–297, January 2006. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2006.
0020.

Watson:2018:JWC

[Wat18] Alan Watson FRS. James Watson Cronin. 29 September 1931–25 August 2016. Biographical Memoirs of Fellows of the Royal Society, 65:47–70, December 2018. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2018.
0021.

Watts:2019:MHP

[Wat19] Anthony Brian Watts. Martin Harold Phillips Bott. 12 July 1926–20 October 2018. Biographical Memoirs of Fellows of the Royal Society, 67:89–117, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2019.
0009.

Whittaker:1968:ACA

[WB68] John Macnaghten Whittaker and Maurice Stevenson Bartlett. Alexander Craig Aitken, 1895–1967. Biographical Memoirs of Fellows of the Royal Society, 14:1–14, November 1968. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1968.0001.

Willmer:1985:JRB

[WB85] Edward Nevill Willmer and P. C. J. Brunet. John Randal Baker, 23 October 1900–8 June 1984. Biographical Memoirs of Fellows of the Royal Society, 31:32–63, November 1985. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1985.0002.

Williams:2003:VMN

[WB03] Charles H. Williams and David P. Ballou. Vincent Massey 28 November 1926–26 August 2002 Elected FRS 1977. Biographical Memoirs of Fellows of the Royal Society, 49:335–350, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0019.

Watkins:2005:WTJ

[WB05] Winifred M. Watkins and Kenneth D. Bagshawe. Walter Thomas James Morgan CBE. 5 October 1900–10 February 2003. Biographical Memoirs of Fellows of the Royal Society, 51:291–302, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0018.

White:2020:SBJ

[WB20] John White FRS and Mark S. Bretscher FRS. Sydney Brenner. 13 January 1927–5 April 2019. Biographical Memoirs of Fellows of the Royal Society, 69:79–108, December 2020. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2020.0022.

Worsley:2023:JRL

[WBF23] Peter Worsley, Martin Bell, and Michael Fulford. John Robert Lawrence Allen. 25 October 1932–18 October 2020. Biographical Memoirs of Fellows of the Royal Society, 75:29–51, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2023.0009.

Wilson:2021:MMW

[WD21] Keith S. Wilson and Eleanor J. Dodson. Michael Mark Woolfson. 9 January 1927–23 December 2019. Biographical Memoirs of Fellows of the Royal Society, 71:529–553, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2021.0018.

Weatherall:2002:SCA

[Wea02] David Weatherall F.R.S. Sir Cyril Astley Clarke, C.B.E. 22 August 1907–21 November 2000. Biographical Memoirs of Fellows of the Royal Society, 48:69–85, January 2002. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0005.

Webster:2010:WGL

[Web10] Robert G. Webster. William Graeme Laver. 3 June 1929– 26 September 2008. Biographical Memoirs of Fellows of the Royal Society, 56:215–236, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2009.
0021.

Webb:2017:DEC

[Web17] David J. Webb. David Edgar Cartwright. 21 October 1926– 2 December 2015. Biographical Memoirs of Fellows of the Royal Society, 63:99–115, January 2017. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2017.
0001.

Weinberg:1990:ARJ

[Wei90] Felix Jiri Weinberg. Alfred Rene Jean Paul Ubbelohde, 14 December 1907–7 January 1988. Biographical Memoirs of Fellows of the Royal Society, 35:381–402, March 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Weiskrantz:1994:DEB

[Wei94] Lawrence Weiskrantz. Donald Eric Broadbent, 6 May 1926– 10 April 1993. Biographical Memoirs of Fellows of the Royal Society, 40:31–42, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.
0027.

Weinberg:2005:AGG

[Wei05] F. J. Weinberg. Alfred Gordon Gaydon. 26 September 1911– 16 April 2004. Biographical Memoirs of Fellows of the Royal Society, 51:169–184, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2005.

Weiss:2024:SMG

[Wei24] Robin A. Weiss FRS. Sir Michael George Parke Stoker. 4 July 1918–13 August 2013. Biographical Memoirs of Fellows of the Royal Society, 76:417–439, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.
0051.

Wells:1978:LTH

[Wel78] George Philip Wells. Lancelot Thomas Hogben, 9 December 1895–22 August 1975. Biographical Memoirs of Fellows of the Royal Society, 24:183–221, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1978.
0007.

Wells:2000:GRI

[Wel00] Alan A. Wells. George Rankin Irwin. 26 February 1907–9 October 1998. Biographical Memoirs of Fellows of the Royal Society, 46:269– 283, January 2000. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1999.0084.

Wells:2005:SGN

[Wel05] P. N. T. Wells. Sir Godfrey Newbold Hounsfield KT CBE. 28 August 1919–12 August 2004. Biographical Memoirs of Fellows of the Royal Society, 51:221–235, January 2005. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2005.0014.

Westergaard:1964:OW

[Wes64] M. Westergaard. Øjvind Winge, 1886–1964. Biographical Memoirs of Fellows of the Royal Society, 10:357–369, November 1964. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1964.0020.

West:1988:HGM

[Wes88] Richard Gilbert West. Harry Godwin, 9 May 1901–12 August 1985. Biographical Memoirs of Fellows of the Royal Society, 34:259–292, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1988.0010.

Westheimer:1995:FWC

[Wes95] Gerald Westheimer. Fergus William Campbell, 30 January 1924–3 May 1993. Biographical Memoirs of Fellows of the Royal Society, 41:104–116, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1995.
0007.

Waterston:2019:SJE

[WF19] Robert H. Waterston and Georgina Ferry. Sir John Edward Sulston CH. 27 March 1942–6 March 2018. Biographical Memoirs of Fellows of the Royal Society, 67:421–447, December 2019. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2019.0014.

Wigner:1977:MPM

[WH77] Eugene Paul Wigner and R. A. Hodgkin. Michael Polányi. 12 March 1891–22 February 1976. Biographical Memoirs of Fellows of the Royal Society, 23(2):413–448, November 1977. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL http://www.jstor.org/stable/769621;   https:
//royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1977.
0016.

Wheeler:1974:KGV

[Whe74] Robert Eric Mortimer Wheeler. King Gustaf VI Adolf of Sweden, K. G., 1882–1973. Biographical Memoirs of Fellows of the Royal Society, 20:213–216, December 1974. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1974.0008.

Wheeler:2001:SCS

[Whe01] Raymond L. Wheeler. Sir Christopher Sydney Cockerell, C.B.E., R.D.I. 4 June 1910–1 June 1999. Biographical Memoirs of Fellows of the Royal Society, 47:67–89, January 2001. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2001.0005.

Whittaker:1955:AE

[Whi55] Edmund Taylor Whittaker. Albert Einstein, 1879–1955. Biographical Memoirs of Fellows of the Royal Society, 1:37–67, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL http://www.jstor.org/stable/
769242;   https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0005.

Whiddington:1958:FH

[Whi58] Richard Whiddington. Frank Horton, 1878–1957. Biographical Memoirs of Fellows of the Royal Society, 4:117–127, November 1958. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1958.0011.

Whiddington:1962:JWR

[Whi62] Richard Whiddington. John Walter Ryde, 1898–1961. Biographical Memoirs of Fellows of the Royal Society, 8:104–117, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1962.0008.

White:1966:WDL

[Whi66a] Errol Ivor White. William Dickson Lang, 1878–1966. Biographical Memoirs of Fellows of the Royal Society, 12:366–386, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1966.0017.

Whittaker:1966:GNW

[Whi66b] John Macnaghten Whittaker. George Neville Watson, 1886–1965. Biographical Memoirs of Fellows of the Royal Society, 12:520–530, November 1966. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1966.0026.

White:1979:RGM

[Whi79] Frederick William George White. Robert Gordon Menzies, 20 December 1894–15 May 1978. Biographical Memoirs of Fellows of the Royal Society, 25:445–476, November 1979. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0016.

Whitteridge:1983:LGJ

[Whi83] David Whitteridge. Ludwig Guttmann, 3 July 1899–18 March 1980. Biographical Memoirs of Fellows of the Royal Society, 29:226–244, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1983.0010.

Whittington:1986:GGS

[Whi86] Harry Blackmore Whittington. George Gaylord Simpson, 16 June 1902–6 October 1984. Biographical Memoirs of Fellows of the Royal Society, 32:525–539, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.
0017.

Whitehouse:1990:RER

[Whi90] David J. Whitehouse. Richard Edmund Reason, 21 December 1903–20 March 1987. Biographical Memoirs of Fellows of the Royal Society, 36:435–462, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0039.

Whiffen:1994:LES

[Whi94a] David Hardy Whiffen. Leslie Ernest Sutton, 22 June 1906– 30 October 1992. Biographical Memoirs of Fellows of the Royal Society, 40:367–382, November 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1994.
0045.

Whitfield:1994:LHN

[Whi94b] M. Whitfield. Leslie Hugh Norman Cooper, 17 June 1905–19 September 1985. Biographical Memoirs of Fellows of the Royal Society, 39:111–128, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Whitby:1996:SFE

[Whi96] G. F. Whitby. Sir Frank Ewart Smith, 31 May 1897–14 June 1995. Biographical Memoirs of Fellows of the Royal Society, 42:421–431, November 1996. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1996.0026.

White:1999:DHM

[Whi99] Robert S. White. Drummond Hoyle Matthews. 5 February 1931–20 July 1997. Biographical Memoirs of Fellows of the Royal Society, 45:275–294, January 1999. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1999.
0019.

Whittington:2001:SCJ

[Whi01] H. B. Whittington. Sir (Cyril) James Stubblefield. 6 September 1901–23 October 1999. Biographical Memoirs of Fellows of the Royal Society, 47:453–464, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2001.
0027.

Whiting:2003:FEK

[Whi03] Donald Whiting. Frederick Ernest King. 2 May 1905–14 August 1999. Biographical Memoirs of Fellows of the Royal Society, 49:299– 314, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2003.0017.

Whittle:2004:MSB

[Whi04] Peter Whittle. Maurice Stevenson Bartlett. 18 June 1910– 8 January 2002. Biographical Memoirs of Fellows of the Royal Society, 50:15–33, January 2004. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2004.
0002.

White:2022:SJT

[Whi22] Robert S. White FRS. Sir John Theodore Houghton. 30 December 1931–15 April 2020. Biographical Memoirs of Fellows of the Royal Society, 73:291–314, December 2022. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2022.0014.

Widdowson:1995:RAM

[Wid95] Elsie May Widdowson. Robert Alexander McCance, 9 December 1898–5 March 1993. Biographical Memoirs of Fellows of the Royal Society, 41:262–280, November 1995. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1995.0016.

Wigglesworth:1956:PAB

[Wig56] Vincent Brian Wigglesworth. Patrick Alfred Buxton, 1892–1955. Biographical Memoirs of Fellows of the Royal Society, 2:69–84, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1956.0005.

Wigglesworth:1971:BPU

[Wig71] Vincent Brian Wigglesworth. Boris Petrovitch Uvarov, 1889–1970. Biographical Memoirs of Fellows of the Royal Society, 17:713–740, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1971.0029.

Wigglesworth:1982:CBW

[Wig82] Vincent Brian Wigglesworth. Carrington Bonsor Williams, 7 October 1889–12 July 1981. Biographical Memoirs of Fellows of the Royal Society, 28:666–684, November 1982. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0025.

Wigglesworth:1983:JWS

[Wig83] Vincent Brian Wigglesworth. John William Sutton Pringle, 22 July 1912–2 November 1982. Biographical Memoirs of Fellows of the Royal Society, 29:524–551, November 1983. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1983.0019.

Wilson:1955:HSA

[Wil55] William Wilson. Herbert Stanley Allen, 1873–1954. Biographical Memoirs of Fellows of the Royal Society, 1:5–10, November 1955. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1955.0002.

Wilson:1956:JWN

[Wil56] William Wilson. John William Nicholson, 1881–1955. Biographical Memoirs of Fellows of the Royal Society, 2:209–214, November 1956. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1956.0014.

Wilson:1960:OWR

[Wil60] William Wilson. Owen Willans Richardson, 1879–1959. Biographical Memoirs of Fellows of the Royal Society, 5:206–215, February 1960. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Wilson:1961:MDB

[Wil61] William Wilson. Maurice, Le Duc de Broglie, 1875–1960. Biographical Memoirs of Fellows of the Royal Society, 7:31–36, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1961.0003;   https://www.jstor.org/stable/
769394.

Willmer:1972:EFF

[Wil72] Edward Nevill Willmer. Emmanuel Fauré-Fremiet, 1883–1971. Biographical Memoirs of Fellows of the Royal Society, 18:187–221, November 1972. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1972.0006.

Williams:1975:WJP

[Wil75] Alwyn Williams. William John Pugh, 28 July 1892–18 March 1974. Biographical Memoirs of Fellows of the Royal Society, 21:485–495, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1975.0015.

Wilkins:1987:JTR

[Wil87] Maurice Hugh Frederick Wilkins. John Turton Randall, 23 March 1905–16 June 1984. Biographical Memoirs of Fellows of the Royal Society, 33:491–535, December 1987. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1987.
0018.

Williams:1994:FTB

[Wil94a] K. R. Williams. Francis Thomas Bacon, 21 December 1904–24 May 1992. Biographical Memoirs of Fellows of the Royal Society, 39:1– 18, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Willis:1994:ARC

[Wil94b] A. J. Willis. Arthur Roy Clapham, 24 May 1904–18 December 1990. Biographical Memoirs of Fellows of the Royal Society, 39:71– 90, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Williams:1995:BMF

[Wil95] Trevor I. Williams. Book review: Biographical Memoirs of Fellows of the Royal Society. Endeavour, 19(2):92, January 1995. CODEN ENDEAS. ISSN 0160-9327 (print), 1873-1929 (electronic).

Williams:2013:SHF

[Wil13] Gordon Williams FREng FRS. Sir Hugh Ford. 16 July 1913–28 May 2010. Biographical Memoirs of Fellows of the Royal Society, 59:145– 156, January 2013. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2012.0041.

Willis:2022:RBA

[Wil22] J. R. Willis. Ronald Bullough. 6 April 1931–20 November 2020. Biographical Memoirs of Fellows of the Royal Society, 72:101–115, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2021.0046.

Williams:2023:MFM

[Wil23a] David E. Williams. Martin Fleischmann. 29 March 1927–3 August 2012. Biographical Memoirs of Fellows of the Royal Society, 74: 161–184, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2022.0030.

Williams:2023:SMF

[Wil23b] Peter Williams FRS. Sir Martin Francis Wood. 19 April 1927– 23 November 2021. Biographical Memoirs of Fellows of the Royal Society, 75:479–501, December 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2023.

Wilson:2023:SIA

[Wil23c] R. J. M. (Iain) Wilson. Sir Ian Alexander McGregor. 26 August 1922–1 February 2007. Biographical Memoirs of Fellows of the Royal Society, 74:295–308, April 2023. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2022.
0041.

Winnifrith:1970:EEB

[Win70] John Winnifrith. Edward Ettingdean Bridges — Baron Bridges, 1892–1969. Biographical Memoirs of Fellows of the Royal Society, 16:36–56, November 1970. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1970.
0003.

Windle:2001:AKA

[Win01] Alan Windle. Andrew Keller. 22 August 1925–7 February 1999. Biographical Memoirs of Fellows of the Royal Society, 47:293–310, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2001.0017.

Wise:1983:DVG

[Wis83] M. J. Wise. David Victor Glass, 2 January 1912–23 September 1978. Biographical Memoirs of Fellows of the Royal Society, 29:201– 225, November 1983. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1983.0009.

Wilkes:2003:TKC

[WK03] Maurice Wilkes FRENG FRS and Hilary J. Kahn. Tom Kilburn CBE FREng. 11 August 1921–17 January 2001. Biographical Memoirs of Fellows of the Royal Society, 49:283–297, January 2003. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0016.

Westheimer:2002:KBJ

[WL02] Frank H. Westheimer For.Mem.R.S. and W. Lipscomb. Konrad Bloch. 21 January 1912–5 October 2000. Biographical Memoirs of Fellows of the Royal Society, 48:43–49, January 2002.

CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2002.0003.

Whitteridge:1984:CSH

[WM84] David Whitteridge and Patrick Anthony Merton. Charles Skinner Hallpike, 19 July 1900–26 September 1979. Biographical Memoirs of Fellows of the Royal Society, 30:282–295, November 1984. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1984.0010.

Wolfendale:1998:PHF

[Wol98] Arnold Wolfendale. Peter Howard Fowler. 27 February 1923– 8 November 1996. Biographical Memoirs of Fellows of the Royal Society, 44:177–189, January 1998. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1998.
0012.

Woledge:2001:DRW

[Wol01] Roger C. Woledge. Douglas Robert Wilkie. 2 October 1922–21 May 1998. Biographical Memoirs of Fellows of the Royal Society, 47:481– 495, January 2001. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.2001.0029.

Wolfendale:2003:GDR

[Wol03] Arnold Wolfendale FRS. George Dixon Rochester. 4 February 1908–26 December 2001. Biographical Memoirs of Fellows of the Royal Society, 49:415–429, January 2003. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2003.0024.

Woods:1957:AJK

[Woo57] Donald Devereux Woods. Albert Jan Kluyver, 1888–1956. Biographical Memoirs of Fellows of the Royal Society, 3:109–128, November 1957. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1957.0008.

Woolley:1961:HSJ

[Woo61] Richard Van Der Riet Woolley. Harold Spencer Jones, 1890–1960. Biographical Memoirs of Fellows of the Royal Society, 7:136–145, November 1961. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1961.0011.

Woolley:1971:CRD

[Woo71] Richard Van Der Riet Woolley. Charles Rundle Davidson, 1875– 1970. Biographical Memoirs of Fellows of the Royal Society, 17:193– 194, November 1971. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1971.0007.

Wood:1990:GMB

[Woo90] Alan Marshall Muir Wood. Geoffrey Morse Binnie, 13 November 1908–5 April 1989. Biographical Memoirs of Fellows of the Royal Society, 36:43–57, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1990.
0023.

Woolfson:1994:HSL

[Woo94] Michael Mark Woolfson. Henry Solomon Lipson, 11 March 1910–26 April 1991. Biographical Memoirs of Fellows of the Royal Society, 39:227–244, February 1994. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Wood:1995:JRJ

[Woo95a] Ronald Karslake Starr Wood. John Rishbeth, 10 July 1918–1 June 1991. Biographical Memoirs of Fellows of the Royal Society, 41:360– 376, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print),

1748-8494 (electronic). URL https://royalsocietypublishing.
org/doi/epdf/10.1098/rsbm.1995.0022.

Wooding:1995:CFK

[Woo95b] Norman Wooding. Christopher Frank Kearton, Baron Kearton, of Whitchurch, Bucks, Kt, O.B.E., 17 February 1911–2 July 1992. Biographical Memoirs of Fellows of the Royal Society, 41:220–241, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1995.0014.

Wood:1997:AMB

[Woo97a] Alan Muir Wood. Alfred Maurice Binnie, F. Eng. 6 February 1901–31 December 1986. Biographical Memoirs of Fellows of the Royal Society, 43:75–86, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0005.

Woolfson:1997:AJC

[Woo97b] M. M. Woolfson. Arthur James Cochran Wilson. 28 November 1914–1 July 1995. Biographical Memoirs of Fellows of the Royal Society, 43:523–535, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0029.

Woolfson:2005:WCJ

[Woo05] Michael Woolfson. William Cochran. 30 July 1922–28 August 2003. Biographical Memoirs of Fellows of the Royal Society, 51:67–85, January 2005. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2005.0005.

Wootton:2021:CPE

[Woo21] Robin Wootton. Charles Porter Ellington. 31 December 1952– 30 July 2019. Biographical Memoirs of Fellows of the Royal Society, 70:151–173, June 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2020.
0041.

Westoll:1975:ASR

[WP75] Thomas Stanley Westoll and Francis Rex Parrington. Alfred Sherwood Romer, 28 December 1894–5 November 1973. Biographical Memoirs of Fellows of the Royal Society, 21:496–516, November 1975. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1975.0016.

Walley:2021:JEF

[WPL21] Stephen M. Walley, William G. Proud, and Timothy G. Leighton FRS. John Edwin Field. 20 September 1936–21 October 2020. Biographical Memoirs of Fellows of the Royal Society, 71:145–170, December 2021. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2021.0013.

Wild:1995:JGB

[WR95] John Paul Wild and V. Radhakrishnan. John Gatenby Bolton, 5 June 1922–6 July 1993. Biographical Memoirs of Fellows of the Royal Society, 41:72–86, November 1995. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1995.
0005.

Wright:2024:CTP

[Wri24] Ian Wright. Colin Trevor Pillinger. 9 May 1943–7 May 2014. Biographical Memoirs of Fellows of the Royal Society, 76:315–339, June 2024. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.2023.0032.

Wunsch:1997:HMS

[Wun97] Carl Wunsch. Henry Melson Stommel. 27 September 1920– 17 January 1992. Biographical Memoirs of Fellows of the Royal Society, 43:493–502, January 1997. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1997.
0027.

Wurtz:2016:DHH

[Wur16] Robert H. Wurtz. David Hunter Hubel. 27 February 1926– 22 September 2013. Biographical Memoirs of Fellows of the Royal Society, 62:233–246, January 2016. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2016.
0022.

Wells:1986:RWM

[WW86] Alan Arthur Wells and E. G. West. Richard Weck, 5 March 1913–9 January 1986. Biographical Memoirs of Fellows of the Royal Society, 32:629–647, December 1986. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1986.

Watts:1999:GFM

[WW99] W. A. Watts and R. G. West. George Francis Mitchell. 15 October 1912–25 November 1997. Biographical Memoirs of Fellows of the Royal Society, 45:315–328, January 1999. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1999.0021.

Wilson:1979:JWM

[WZ79] Graham Selby Wilson and K. S. Zinnemann. James Walter McLeod, 2 January 1887–11 March 1978. Biographical Memoirs of Fellows of the Royal Society, 25:421–444, November 1979. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1979.0015.

Yates:2006:SFW

[Yat06] I. R. Yates. Sir Frederick William Page CBE FREng. 20 February 1917–29 May 2005. Biographical Memoirs of Fellows of the Royal Society, 52:231–250, January 2006. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0017.

Young:1974:BAH

[YF74] Frank George Young and V. G. Foglia. Bernardo Alberto Houssay, 1887–1971. Biographical Memoirs of Fellows of the Royal Society, 20:246–270, December 1974. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1974.
0011.

Young:1982:CHB

[YH82] Frank George Young and Charles Nicholas Hales. Charles Herbert Best, 27 February 1899–31 March 1978. Biographical Memoirs of Fellows of the Royal Society, 28:1–25, November 1982. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1982.0001.

Pearson:1986:SGH

[YHP^{+}86] Pierre Henry John Young, Lionel Haworth, H. Pearson, G. L. Wilde, and J. E. Ffowcs Williams. Stanley George Hooker, 30 September 1907–23 May 1984. Biographical Memoirs of Fellows of the Royal Society, 32:275–319, December 1986. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1986.0009.

Young:1992:PRO

[YL92] Alec David Young and Michael James Lighthill. Paul Robert Owen, 24 January 1920–11 November 1990. Biographical Memoirs of Fellows of the Royal Society, 38:267–285, November 1992. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1992.0014.

Yates:1963:RAF

[YM63] Frank Yates and Kenneth Mather. Ronald Aylmer Fisher, 1890–1962. Biographical Memoirs of Fellows of the Royal Society, 9:91–129, November 1963. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/10.1098/rsbm.1963.0006;
 https://www.jstor.org/stable/769423.

Yamamura:2006:YNJ

[YN06] Hirohei Yamamura and Shun-Ichi Nakamura. Yasutomi Nishizuka. 12 July 1932–4 November 2004. Biographical Memoirs of Fellows of the Royal Society, 52:219–230, January 2006. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.2006.0016.

Yonge:1962:TAS

[Yon62] Maurice Yonge. Thomas Alan Stephenson, 1898–1961. Biographical Memoirs of Fellows of the Royal Society, 8:136–148, November 1962. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL https://royalsocietypublishing.org/doi/
epdf/10.1098/rsbm.1962.0010.

Younger:2012:WTT

[You12] D. H. Younger. William Thomas Tutte. 14 May 1917–2 May 2002. Biographical Memoirs of Fellows of the Royal Society, 58:283–297, January 2012. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.2012.0036.

Young:1978:DWH

[YPS78] Alec David Young, R. C. Pankhurst, and D. L. Schultz. Douglas William Holder, 14 April 1923–18 April 1977. Biographical Memoirs of Fellows of the Royal Society, 24:223–244, November 1978. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1978.0008.

Yudkin:2010:JMN

[Yud10] Michael Yudkin. Joel Mandelstam. 13 November 1919–20 December 2008. Biographical Memoirs of Fellows of the Royal Society, 56:257–271, January 2010. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL https://

royalsocietypublishing.org/doi/epdf/10.1098/rsbm.2010.
0011.

Zeldovich:1990:EML

[ZK90] Yakov Borisovich Zeldovich and M. I. Kaganov. Evgenii Mikhailovich Lifshitz, 21 February 1915–29 October 1985. Biographical Memoirs of Fellows of the Royal Society, 36:335–357, December 1990. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1990.0035.

Zuckerman:1973:WEG

[Zuc73] Solly Zuckerman. Wilfrid Edward Le Gros Clark, 1895–1971. Biographical Memoirs of Fellows of the Royal Society, 19:216–233, December 1973. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-

8494 (electronic). URL https://royalsocietypublishing.org/
doi/epdf/10.1098/rsbm.1973.0008.

Zuckerman:1981:EMB

[Zuc81] Solly Zuckerman. Earl Mountbatten of Burma, K.G., O.M., 25 June 1900–27 August 1979. Biographical Memoirs of Fellows of the Royal Society, 27:355–364, November 1981. CO-DEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (elec-

tronic). URL https://royalsocietypublishing.org/doi/epdf/
10.1098/rsbm.1981.0014.

Zuckerman:1983:GWC

[Zuc83] Solly Zuckerman. George Washington Corner, 12 December 1889– 28 September 1981. Biographical Memoirs of Fellows of the Royal Society, 29:92–112, November 1983. CODEN BMFRA3.

ISSN 0080-4606 (print), 1748-8494 (electronic). URL https:/

/royalsocietypublishing.org/doi/epdf/10.1098/rsbm.1983.